### v2

# QOCO: A Quadratic Objective Conic Optimizer with Custom Solver Generation

Abstract
Second-order cone programs (SOCPs) with quadratic objective functions are common in optimal control and other fields. Most SOCP solvers which use interior-point methods are designed for linear objectives and convert quadratic objectives into linear ones via slack variables and extra constraints, despite the computational advantages of handling quadratic objectives directly. In applications like model-predictive control and online trajectory optimization, these SOCPs have known sparsity structures and require rapid solutions. When solving these problems, most solvers use sparse linear algebra routines, which introduce computational overhead and hinder performance. In contrast, custom linear algebra routines can exploit the known sparsity structure of problem data and be significantly faster. This work makes two key contributions: (1) the development of QOCO, an open-source C-based solver for quadratic objective SOCPs, and (2) the introduction of QOCOGEN, an open-source custom solver generator for quadratic objective SOCPs, which generates a solver written in C that leverages custom linear algebra. Both implement a primal-dual interior-point method with Mehrotra’s predictor-corrector. Our benchmarks show that QOCO is faster and more robust than many commonly used solvers, and solvers generated by QOCOGEN are significantly faster than QOCO and are free of dynamic memory allocation making them an attractive option for real-time optimization on resource-constrained embedded systems.
1 Introduction
We consider the quadratic objective second-order cone program (SOCP)
[…]
subject to $Gx\preceq_{\mathcal{K}}h$
$Ax=b,$
with optimization variable $x\in\mathbb{R}^{n}$. The cost is defined by the positive semidefinite matrix $P=P^{\top}\succeq 0$ and vector $c\in\mathbb{R}^{n}$. The constraints are defined by matrices $A\in\mathbb{R}^{p\times n}$ and $G\in\mathbb{R}^{m\times n}$ and vectors $b\in\mathbb{R}^{p}$ and $h\in\mathbb{R}^{m}$. The generalized inequality $\preceq_{\mathcal{K}}$ denotes membership in a closed, proper cone $\mathcal{K}$, i.e. $h-Gx\in\mathcal{K}$. We restrict $\mathcal{K}$ to be the Cartesian product
$\mathcal{K}=\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K},$
where $\mathcal{C}_{i}$ is either the non-negative orthant or a second-order cone. Each cone $\mathcal{C}_{i}$ corresponds to a subset of constraints, inducing a partition on $G$ and $h$, i.e., $h_{i}-G_{i}x\in\mathcal{C}_{i}$ for $i=1,\ldots,K$. Throughout this work, we assume that Problem 1 is feasible and has a bounded optimal objective.
Problem 1 appears in various applications, including model predictive control (MPC) [1], network flow optimization [2], portfolio optimization [3, 4], robust optimization [5, 6], and filter design [7], among others. Additionally, the solution of SOCPs is used as a subroutine in algorithms for nonconvex optimization such as sequential convex programming (SCP) [8, 9, 10, 11].
Problem 1 can be reformulated as a SOCP with a linear objective by introducing a slack variable $t$,
$\underset{x,t}{\text{minimize}}$ $t+c^{\top}x$
subject to $\left\|\begin{bmatrix}t-\frac{1}{2}\\ P^{1/2}x\end{bmatrix}\right\|_{2}\leq t+\frac{1}{2}$
$Gx\preceq_{\mathcal{K}}h$
$Ax=b.$
If $P$ is large (has many rows and columns), then computing $P^{1/2}$ can be prohibitively expensive, making this reformulation impractical. Solving this reformulation can also be slow if the matrix square root $P^{1/2}$ has significantly more nonzero elements than $P$. Thus it is advantageous for a SOCP solver to natively handle quadratic objective functions without resorting to the linear objective reformulation.
Some SOCP solvers that can solve Problem 1 directly include clarabel [12], gurobi [13], cosmo [14], and scs [15]. However, cosmo and scs implement operator-splitting methods. These methods are preferred for large-scale problems due to their low per-iteration cost but can struggle with problem scaling, conditioning, and achieving high-accuracy solutions [16]. For modestly sized SOCPs, an interior-point method (IPM) is preferred, due to its robustness to scaling and conditioning of the problem, and ability to converge to high accuracy solutions [17, Chapter 1]. One of the few open-source IPMs that can directly solve Problem 1 is clarabel. In addition to solving SOCPs, clarabel can also handle semidefinite programs, as well as problems involving exponential cones, power cones, and generalized power cones. Additionally, it supports infeasibility detection. However, clarabel is written in Rust. Although Rust is memory-safe due to its ownership system and can be used in embedded systems, its ecosystem is far less mature than that of C, and for legacy systems, such as aerospace software, C is still preferred.
Many engineering applications including linear MPC, nonlinear MPC [18], portfolio backtesting, sequential quadratic programming (SQP) [19, 20], and SCP, require solving SOCPs with a fixed, known sparsity structures in $P,A,G$ and a constant cone $\mathcal{K}$. These problems often arise in real-time settings, where computational efficiency is critical. For example, sequential convex programming (SCP) iteratively solves a sequence of SOCPs with quadratic objectives and identical sparsity patterns to find stationary points of nonconvex problems. SCP has been widely applied in aerospace trajectory optimization, where real-time performance is essential. Notable applications include powered-descent guidance [21, 22, 23], in-space rendezvous [24, 25], and hypersonic entry guidance [26, 27]. SCP is also the solution method used by NASA’s SPLICE program for lunar landing guidance [28, 29, 30, 31].
Most general-purpose SOCP solvers, such as clarabel [12], cosmo [14], ecos [32], gurobi [13], mosek [33], and scs rely on sparse linear algebra routines. Although sparse linear algebra allows the aforementioned solvers to solve SOCPs regardless of their sparsity structures, it can be quite slow due to the extra overhead of determining where the nonzero elements are before performing floating point operations, which leads to more CPU instructions being issued, more memory accesses, and more cache misses. However, when the sparsity structure of the problem matrices is known beforehand, it is possible to generate a custom solver that implements specialized linear algebra routines tailored to the sparsity structure of the problem. These custom solvers hardcode the exact floating point operations necessary, eliminating the overhead incurred by sparse linear algebra routines, resulting in significantly faster solves.
A custom solver generator is a tool that takes the sparsity structure of an optimization problem as an input and outputs a solver (typically in C) optimized for that specific sparsity structure without relying on sparse linear algebra. Two notable custom solver generators are cvxgen [34] and bsocp [35, 36]. cvxgen, which is academically licensed but not open-source, solves quadratic programs (QPs) rather than SOCPs (i.e. $\mathcal{K}$ in Problem 1 is the non-negative orthant) and bsocp, which is neither academically licensed nor open source, solves linear objective SOCPs (i.e. $P=0$ in Problem 1). Both have demonstrated substantial speed improvements over general-purpose solvers and have had a significant impact on aerospace trajectory optimization. For instance, after the development of lossless convexification [37, 38], bsocp was flight-tested on Masten’s Xombie vehicle to validate G-FOLD, a powered-descent guidance algorithm based on lossless convexification [39]. Later, cvxgen was used by SpaceX for landing their Falcon 9 boosters [40].
1.1 Contribution
This work makes two key contributions: (1) the development of qoco, an open-source C-based solver for quadratic objective SOCPs, and (2) the introduction of qocogen, an open-source custom solver generator for quadratic objective SOCPs which generates a custom solver written in C.
We demonstrate that qoco outperforms most commercial and open-source solvers in terms of both speed and robustness, and qococustom (the name of solvers generated by qocogen) is significantly faster than qoco. Both solvers are easy to use since qoco and qocogen can be called from CVXPY [41, 42] and CVXPYgen [43] respectively, allowing users to formulate optimization problems in a natural way following from math rather than manually converting the problem to the solver-required standard form. Additionally, qoco can be called from C/C++, Matlab, and Python, and qocogen can be called from Python.
Both implement a primal-dual interior point method with Mehrotra’s predictor-corrector [44]. However, they both use a variety of numerical enhancements to quickly and robustly solve the linear system that arises in the step direction computation. Specifically, we use an $LDL^{\top}$ factorization along with the Approximate Minimum Degree (AMD) [45, 46] heuristic to permute the coefficient matrix, minimizing fill-in for the factor $L$. We also apply static and dynamic regularization to the coefficient matrix to ensure that the matrix is invertible and the factorization succeeds. After solving the linear system, we apply iterative refinement to ensure that the original, unregularized system was solved to high accuracy, further enhancing numerical stability and robustness.
qoco and qocogen are available on GitHub at https://github.com/qoco-org.
1.2 Outline
Section 2 introduces the primal-dual interior point method implemented in qoco and qocogen. Section 3 explains the techniques used to ensure a stable factorization of the KKT matrix, which is essential to compute the step direction. Section 4 discusses the advantages of custom linear algebra over sparse linear algebra, and how qocogen generates custom solvers which exploit the known sparsity structure of problem data. Finally, Section 5 presents extensive numerical results comparing qoco and qococustom to existing solvers.
1.3 Notation
$(a,b)$ refers to the concatenation of $a\in\mathbb{R}^{a}$ and $b\in\mathbb{R}^{b}$ into a vector in $\mathbb{R}^{a+b}$
The $l$-dimensional non-negative orthant is given by +$\mathbb{R}^{l}_{+}=\{u\in\mathbb{R}^{l}\;|\;u_{i}\geq 0\}$
For a vector $u$ in the $l$-dimensional non-negative orthant, $u_{i}$ refers to the $i^{th}$ element of the vector
The $q$-dimensional second-order cone is given by $\mathcal{Q}^{q}=\{(u_{0},u_{1})\in\mathbb{R}\times\mathbb{R}^{q-1}\;|\;\|u_{1}\|_{2}\leq u_{0}\}$
For a vector $u$ in the $q$-dimensional second-order cone, we partition it as $u=(u_{0},u_{1})$, i.e. $u_{0}\in\mathbb{R}$, $u_{1}\in\mathbb{R}^{q-1}$ and $\|u_{1}\|_{2}\leq u_{0}$.
When $x\in\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K}$, $x_{i}$ refers to the $i^{th}$ subvector which is in cone $\mathcal{C}_{i}$
2 Primal-dual interior point method
In this section, we describe the primal-dual interior point algorithm we implement in qoco and qocogen, as outlined in Algorithm 1. This algorithm is equivalent to the coneqp algorithm outlined in [47] and follows a derivation similar to those presented in [48, Chapter 6], [49, 50].
If strong duality holds for Problem 1, the optimal primal-dual solution $(x^{*},s^{*},y^{*},z^{*})$ satisfies the Karush-Kuhn-Tucker (KKT) conditions [51, Chapter 5], which can be written as
$Px+c+A^{\top}y+G^{\top}z$ $=0$ (2a)
$Ax$ $=b$ (2b)
$Gx+s$ $=h$ (2c)
$s_{i}^{\top}z_{i}$ $=0\text{ for }i=1,\ldots K$ (2d)
$(s,z)$ $\in\mathcal{K}\times\mathcal{K}.$ (2e)
The primal-dual interior point method applies a modified Newton’s method to Equations (2a) - (2d), and a line search to satisfy Equation (2e). The modifications to Newton’s method correct for linearization errors in the Newton step and bias the search directions towards the interior of $\mathcal{K}$. This prevents the iterates from prematurely approaching the boundary of the cone $\mathcal{K}$, enabling longer steps without violating Equation (2e). It also ensures iterates do not converge to spurious solutions that satisfy Equations (2a) - (2d), but not Equation (2e) [50].
2.1 Central path derivation
To understand how the algorithm biases the search directions towards the interior of $\mathcal{K}$, we first introduce the concept of the central path.
Problem 1 can be equivalently rewritten as
[…]
subject to $Ax=b,$
where $\mathcal{I}_{\mathcal{K}}$, the indicator function of the cone $\mathcal{K}$, is defined as
$\mathcal{I}_{\mathcal{K}}(u)=\begin{cases}0,\text{ for }u\in\mathcal{K}\\ \infty,\text{ otherwise.}\end{cases}$
We then replace the indicator function, $\mathcal{I}_{\mathcal{K}}(u)$, with a smooth barrier function, $\phi_{\mathcal{K}}(u)$, which approaches $\infty$ as its argument approaches the boundary of the cone.
We use the barrier function
$\phi_{\mathcal{K}}(u)=\sum_{i=1}^{K}\phi_{i}(u_{i}),$
where $\phi_{i}$ is the barrier function for $\mathcal{C}_{i}$. The barrier function for the non-negative orthant and the second-order cone are
$\phi_{i}(u)=\begin{cases}-\sum_{j=1}^{l}\log u_{j},\;\;\;\;\;\;\quad\quad\text{ for }\mathcal{C}_{i}=\mathbb{R}^{l}_{+}\\ -(1/2)\log(u_{0}^{2}-u_{1}^{\top}u_{1}),\text{ for }\mathcal{C}_{i}=\mathcal{Q}^{q},\end{cases}$
where $\log$ is the natural logarithm.
Their gradients are
$\nabla\phi_{i}(u)=\begin{cases}(-1/u_{1},\ldots,-1/u_{l}),\text{ for }\mathcal{C}_{i}=\mathbb{R}^{l}_{+}\\ -(u^{\top}Ju)^{-1}Ju,\quad\;\;\;\;\text{ for }\mathcal{C}_{i}=\mathcal{Q}^{q},\end{cases}$ (3)
where
$J=\begin{bmatrix}1&0\\ 0&-I_{q-1}\end{bmatrix}.$
After replacing the indicator function with the barrier function, we obtain
[…]
subject to $Ax=b,$
where $\tau>0$ is a scalar. Denoting the optimal solution of Problem 4 as $x^{*}(\tau)$, it can be shown that as $\tau\to 0$, $x^{*}(\tau)\to x^{*}$.
The KKT condition for Problem 4 are
$Px+c+A^{\top}y-\tau\sum_{i=1}^{K}G_{i}^{\top}\nabla\phi_{i}(h_{i}-G_{i}x)$ $=0$
$Ax$ $=b$
$h-Gx$ $\in\mathcal{K}.$
If we define $s=h-Gx$ and $z_{i}=-\tau\nabla\phi_{i}(h_{i}-G_{i}x)$ for $i=1,\ldots K$, we can rewrite the KKT conditions as
$Px+c+A^{\top}y+G^{\top}z$ $=0$ (5a)
$Ax$ $=b$ (5b)
$Gx+s$ $=h$ (5c)
$z_{i}$ $=-\tau\nabla\phi_{i}(s_{i})\text{ for }i=1,\ldots K$ (5d)
$(s,z)$ $\in\mathcal{K}\times\mathcal{K},$ (5e)
where the condition $z\in\mathcal{K}$ arises because if $u\in\mathcal{K}$, then $-\nabla\phi(u)\in\mathcal{K}$ [47].
It is desirable to write Equation (5d), in a form where $s_{i}$ and $z_{i}$ appear symmetrically. To this end, we define the Jordan product [52, 49, 47], a commutative and linear operation, for the non-negative orthant and second-order cone as
$u\circ v=\begin{cases}(u_{1}v_{1},\ldots,u_{l}v_{l}),\quad\;\text{ for }\mathcal{C}_{i}=\mathbb{R}^{l}_{+}\\ (u^{\top}v,u_{0}v_{1}+v_{0}u_{1}),\text{ for }\mathcal{C}_{i}=\mathcal{Q}^{q},\end{cases}$
and the Jordan product for $\mathcal{K}$ as
$u\circ v=(u_{1}\circ v_{1},u_{2}\circ v_{2},\ldots,u_{K}\circ v_{K}),$ (6)
where $u_{i},v_{i}\in\mathcal{C}_{i}$.
The identity element $\boldsymbol{e}_{i}$ for cone $\mathcal{C}_{i}$ is defined as
$\boldsymbol{e}_{i}=\begin{cases}(1,1,\ldots,1),\text{ for }\mathcal{C}_{i}=\mathbb{R}^{l}_{+}\\ (1,0,\ldots,0),\text{ for }\mathcal{C}_{i}=\mathcal{Q}^{q},\end{cases}$
and the identity element for $\mathcal{K}$ is
$\boldsymbol{e}=(\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{K}).$ (7)
Taking the Jordan product with $s_{i}$ on both sides of Equation (5d) and substituting Equation (3), we obtain
$Px+c+A^{\top}y+G^{\top}z$ $=0$ (8a)
$Ax$ $=b$ (8b)
$Gx+s$ $=h$ (8c)
$s_{i}\circ z_{i}$ $=\tau\boldsymbol{e}_{i}\text{ for }i=1,\ldots K$ (8d)
$(s,z)$ $\in\mathcal{K}\times\mathcal{K}.$ (8e)
We then rewrite Equation (8d) by stacking $s_{i}$ and $z_{i}$ and using Equations (6) and (7) to obtain
$Px+c+A^{\top}y+G^{\top}z$ $=0$ (9a)
$Ax$ $=b$ (9b)
$Gx+s$ $=h$ (9c)
$s\circ z$ $=\tau\boldsymbol{e}$ (9d)
$(s,z)$ $\in\mathcal{K}\times\mathcal{K}.$ (9e)
The central path is defined as the trajectory of points parameterized by $\tau>0$, satisfying Equations (9a) - (9e). We denote the central path with the tuple $(x^{*}(\tau),s^{*}(\tau),y^{*}(\tau),z^{*}(\tau))$. From the above analysis, we see that the central path equations given by Equations (9a) - (9e) are equivalent to the optimality conditions for the barrier formulation of Problem 1 given by Problem 4. It can be shown that as $\tau\to 0$, $(x^{*}(\tau),s^{*}(\tau),y^{*}(\tau),z^{*}(\tau))\to(x^{*},s^{*},y^{*},z^{*})$.
The primal-dual interior-point method applies Newton’s method to move towards the central path rather than directly towards points satisfying Equation (2d). Since the central path lies within the cone, maintaining proximity to it allows the algorithm to take longer steps without violating Equation (2e) [50].
2.2 Nesterov-Todd scaling
If Newton’s method were applied directly to Equations (5a) - (5d), the coefficient matrix of the resulting linear system would lack symmetry (see Appendix A). As a result, it would be necessary to store the entire matrix, rather than only the upper or lower triangular portion. Additionally, solving this system would require matrix factorizations for non-symmetric matrices, which are generally more computationally expensive than those for symmetric matrices.
To derive a symmetric linear system, we introduce the following change of variables
$\tilde{s}=W^{-\top}s,\;\tilde{z}=Wz,$
where we choose $W$ such that the above transformation preserves cone membership and leaves the central path equations unchanged
$s\in\mathcal{K}\iff\tilde{s}\in\mathcal{K},\quad z\in\mathcal{K}\iff\tilde{z}\in\mathcal{K},\quad s\circ z=\tau\boldsymbol{e}\iff\tilde{s}\circ\tilde{z}=\tau\boldsymbol{e}.$ (10)
Using this transformation, the central path equations can be equivalently expressed as
$Px+c+A^{\top}y+G^{\top}z$ $=0$ (11a)
$Ax$ $=b$ (11b)
$Gx+s$ $=h$ (11c)
$(W^{-\top}s)\circ(Wz)$ $=\tau\boldsymbol{e}$ (11d)
$(s,z)$ $\in\mathcal{K}\times\mathcal{K}.$ (11e)
There are many matrices, $W$, which satisfy Equation (10). In particular, we use the Nesterov-Todd scaling matrix [53, 54]. This scaling matrix is determined based on the current iterates $s_{k}$ and $z_{k}$ and the unique point $w$ that satisfies
$\nabla^{2}\phi_{\mathcal{K}}(w)s_{k}=z_{k},$ (12)
where $\nabla^{2}\phi_{\mathcal{K}}(w)$ is the Hessian of the barrier function evaluated at $w$. Since the barrier function is strictly convex, $\nabla^{2}\phi_{\mathcal{K}}(w)$ is positive definite.
From this, we compute $W_{k}$ as
$\nabla^{2}\phi_{\mathcal{K}}(w)^{-1}=W_{k}^{\top}W_{k}.$ (13)
A key property that follows from Equations (12) and (13) is
$W_{k}^{-\top}s_{k}=W_{k}z_{k}.$ (14)
This property allows us to define
$\lambda_{k}=W_{k}^{-\top}s_{k}=W_{k}z_{k}.$ (15)
For more details on how to compute the scaling point $w$ and the scaling matrix $W_{k}$, see [47].
2.3 Computing search directions
Given the current iterate $(x_{k},s_{k},y_{k},z_{k})$ we define the duality measure as
$\mu_{k}=s_{k}^{\top}z_{k}/m,$ (16)
which represents the average violation of the complementary slackness condition given by Equation (2d). To compute a search direction for updating the current iterate, we apply Mehrotra’s predictor-corrector method [44]. This scheme computes a Newton step with three key objectives: reducing the duality measure, maintaining proximity to the central path, and correcting for the linearization error introduced during this step.
Applying Newton’s method to Equation (11) first requires a linearization of the central path equations around the current iterate $(x_{k},s_{k},y_{k},z_{k})$. This is done by substituting $(x,s,y,z)$ with $(x_{k}+\Delta x,s_{k}+\Delta s,y_{k}+\Delta y,z_{k}+\Delta z)$ and ignoring the higher order term $(W^{-\top}\Delta s)\circ(W\Delta z)$. This results in a linear system of the form
$P\Delta x+A^{\top}\Delta y+G^{\top}\Delta z$ $=-r_{x}$ (17a)
$A\Delta x$ $=-r_{y}$ (17b)
$G\Delta x+\Delta s$ $=-r_{z}$ (17c)
$(W_{k}^{-\top}s_{k})\circ(W_{k}\Delta z)+(W_{k}^{-\top}\Delta s)\circ(W_{k}z_{k})$ $=-r_{s}.$ (17d)
Here, $(r_{x},r_{y},r_{z},r_{s})$ are residual vectors that will be defined later.
Using Equation (15), we can rewrite the system as
$P\Delta x+A^{\top}\Delta y+G^{\top}\Delta z$ $=-r_{x}$ (18a)
$A\Delta x$ $=-r_{y}$ (18b)
$G\Delta x+\Delta s$ $=-r_{z}$ (18c)
$\lambda_{k}\circ(W_{k}\Delta z+W_{k}^{-\top}\Delta s)$ $=-r_{s}.$ (18d)
To obtain a symmetric system, we can eliminate $\Delta s$ and rewrite the equations in matrix form as
$\begin{bmatrix}P&A^{\top}&G^{\top}\\ A&0&0\\ G&0&-W_{k}^{\top}W_{k}\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}$ $=\begin{bmatrix}-r_{x}\\ -r_{y}\\ -r_{z}+W_{k}^{\top}(\lambda_{k}\backslash r_{s})\end{bmatrix}$ (19a)
$\Delta s$ $=-r_{z}-G\Delta x,$ (19b)
[…]
To account for the higher order term we ignored when forming Equation (17), we apply Mehrotra’s predictor-corrector [44]. This method consists of a predictor step, where a search direction is computed by taking a Newton step on Equation (11) with $\tau=0$, followed by a combined step. The combined step incorporates centering and correction terms that will help keep the next iterate close to the central path while compensating for the linearization error incurred by ignoring the $(W^{-\top}\Delta s)\circ(W\Delta z)$ term when linearizing Equation (11). Note that the predictor step is used to estimate this linearization error. Figure 1 illustrates Mehrotra’s predictor-corrector.
$(\boldsymbol{x_{k}},\boldsymbol{s_{k}},\boldsymbol{y_{k}},\boldsymbol{z_{k}})$$(\boldsymbol{x_{k+1}},\boldsymbol{s_{k+1}},\boldsymbol{y_{k+1}},\boldsymbol{z_{k+1}})$Combined StepPredictor StepFeasible SetCentral Path
Figure 1: Mehrotra’s predictor-corrector.
To compute the predictor step, also called the affine scaling direction, $(\Delta x_{a},\Delta s_{a},\Delta y_{a},\Delta z_{a})$, we solve Equation (19) with the residual vectors
$r_{x}$ $=Px_{k}+c+A^{\top}y_{k}+G^{\top}z_{k}$ (20a)
$r_{y}$ $=Ax_{k}-b$ (20b)
$r_{z}$ $=Gx_{k}+s_{k}-h$ (20c)
$r_{s}$ $=\lambda_{k}\circ\lambda_{k}.$ (20d)
Since taking a full step along the affine scaling direction can violate the constraint $(s,z)\in\mathcal{K}\times\mathcal{K}$, we compute the centering parameter $\sigma\in[0,1]$ as
$\alpha$ $=\sup\{\alpha\in[0,1]\;|\;(s_{k},z_{k})+\alpha(\Delta s_{a},\Delta z_{a})\in\mathcal{K}\times\mathcal{K}\}$
$\rho$ $=\frac{(s_{k}+\alpha\Delta s_{a})^{\top}(z_{k}+\alpha\Delta z_{a})}{s_{k}^{\top}z_{k}}$
$\sigma$ $=\max\{0,\min\{1,\rho\}^{3}\}.$
The centering parameter balances the need to reduce the duality measure while maintaining proximity to the central path, allowing for a longer step in the next iteration [50].
Finally, we compute the combined direction, which includes both predictor and corrector information. This direction, $(\Delta x,\Delta s,\Delta y,\Delta z)$ is obtained by solving Equation (19) with the residual vectors
$r_{x}$ $=Px_{k}+c+A^{\top}y_{k}+G^{\top}z_{k}$
$r_{y}$ $=Ax_{k}-b$
$r_{z}$ $=Gx_{k}+s_{k}-h$
$r_{s}$ $=\lambda_{k}\circ\lambda_{k}+(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})-\sigma\mu_{k}\boldsymbol{e}$
This direction moves the next iterate closer to $(x^{*}(\sigma\mu_{k}),s^{*}(\sigma\mu_{k}),y^{*}(\sigma\mu_{k}),z^{*}(\sigma\mu_{k}))$, a point on the central path where the duality measure is reduced by a factor of $\sigma$. Note that the residual vector $r_{s}$ has two additional terms when compared to Equation (20d): $(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})$ and $-\sigma\mu_{k}\boldsymbol{e}$. The former is to correct for linearization error in Newton’s method, and the latter is to maintain proximity to the central path and arises when applying Newton’s method to Equation (9d) with $\tau=\sigma\mu_{k}$.
We then compute the next iterate by performing a line search to ensure $(s_{k+1},z_{k+1})\in\mathcal{K}\times\mathcal{K}$. Specifically, we update the iterate as
$(x_{k+1},s_{k+1},y_{k+1},z_{k+1})=(x_{k},s_{k},y_{k},z_{k})+\alpha(\Delta x,\Delta s,\Delta y,\Delta z)$
where
$\alpha=\sup\left\{\alpha\in[0,1]\;|\;(s_{k},z_{k})+\frac{\alpha}{0.99}(\Delta s,\Delta z)\in\mathcal{K}\times\mathcal{K}\right\}$
The factor of $0.99$ ensures that $(s_{k+1},z_{k+1})$ remains strictly in the interior of $\mathcal{K}\times\mathcal{K}$.
2.4 Initialization
To initialize the primal-dual interior point method, we follow the approach outlined in [47]. The initialization procedure consists of two steps. First, we find $(x,s,y,z)$ that satisfy Equations (9a) - (9c). If $s$ and $z$ do not satisfy Equation (9e), a correction is applied to ensure they are in $\mathcal{K}$.
We begin by solving the optimization problem
[…]
subject to $Ax=b.$
Since this is an equality-constrained quadratic program, the KKT conditions correspond to the linear system
$\begin{bmatrix}P&A^{\top}&G^{\top}\\ A&0&0\\ G&0&-I\end{bmatrix}\begin{bmatrix}x\\ y\\ z\end{bmatrix}=\begin{bmatrix}-c\\ b\\ h\end{bmatrix}.$ (23)
We then set $s=-z$. This results in the tuple $(x,s,y,z)$ satisfying Equations (9a) - (9c), but $s$ and $z$ may not necessarily satisfy Equation (9e).
We then initialize $x_{0}$ and $y_{0}$ as $x_{0}=x$, $y_{0}=y$, and initialize $s_{0}$ as
$s_{0}=\begin{cases}s,\qquad\qquad\quad\;\;\text{ if }s\in\mathcal{K}\\ s+(1+\alpha_{s})\boldsymbol{e}\;\;\text{ otherwise},\end{cases}$
where $\alpha_{s}=\inf\{\alpha\;|\;s+\alpha\boldsymbol{e}\in\mathcal{K}\}$ and $\boldsymbol{e}$ is defined in Equation (7). We initialize $z_{0}$ as
$z_{0}=\begin{cases}z,\qquad\qquad\quad\;\;\text{ if }z\in\mathcal{K}\\ z+(1+\alpha_{z})\boldsymbol{e},\text{ otherwise}\end{cases}$
where $\alpha_{z}=\inf\{\alpha\;|\;z+\alpha\boldsymbol{e}\in\mathcal{K}\}$. Note that $\alpha_{s}$ and $\alpha_{z}$ are the minimum scalings of $\boldsymbol{e}$ that need to be added to $s$ and $z$ respectively to move them to the boundary of $\mathcal{K}$, but we scale $\boldsymbol{e}$ by $(1+\alpha_{s})$ and $(1+\alpha_{z})$ to ensure that $s_{0}$ and $z_{0}$ are strictly in the interior of $\mathcal{K}$.
Algorithm 1 Primal-Dual Interior Point Method
1:Optimal solution $(x^{*},s^{*},y^{*},z^{*})$
2:Initialization:
3: Solve KKT system (23) for $x,y,z$
4:$x_{0}\leftarrow x,\;y_{0}\leftarrow y$
5:$s\leftarrow-z$
6:if $s\in\mathcal{K}$ then
7:     $s_{0}\leftarrow s$
8:else
9:     $\alpha_{s}\leftarrow\inf\{\alpha\;|\;s+\alpha\boldsymbol{e}\in\mathcal{K}\}$
10:     $s_{0}\leftarrow s+(1+\alpha_{s})\boldsymbol{e}$
11:end if
12:if $z\in\mathcal{K}$ then
13:     $z_{0}\leftarrow z$
14:else
15:     $\alpha_{z}\leftarrow\inf\{\alpha\;|\;z+\alpha\boldsymbol{e}\in\mathcal{K}\}$
16:     $s_{0}\leftarrow z+(1+\alpha_{z})\boldsymbol{e}$
17:end if
18:repeat
19:     Compute duality measure:
20:      $\mu_{k}\leftarrow s_{k}^{\top}z_{k}/m$
21:     Compute Nesterov-Todd scaling:
22:      Construct $W_{k}$
23:      Calculate $\lambda_{k}=W_{k}^{-\top}s_{k}=W_{k}z_{k}$
24:     Compute affine direction $(\Delta x_{a},\Delta s_{a},\Delta y_{a},\Delta z_{a})$:
25:      Solve KKT system (19) with residuals (20)
26:     Calculate centering parameter:
27:      $\alpha\leftarrow\sup\{\alpha\in[0,1]\;|\;(s_{k},z_{k})+\alpha(\Delta s_{a},\Delta z_{a})\in\mathcal{K}\}$
28:      $\rho\leftarrow\frac{(s_{k}+\alpha\Delta s_{a})^{\top}(z_{k}+\alpha\Delta z_{a})}{s_{k}^{\top}z_{k}}$
29:      $\sigma\leftarrow\max\{0,\min\{1,\rho\}^{3}\}$
30:     Compute combined direction $(\Delta x,\Delta s,\Delta y,\Delta z)$:
31:      Solve (19) with residuals (22)
32:     Compute step-size:
33:      $\alpha\leftarrow\sup\left\{\alpha\in[0,1]\;|\;(s_{k},z_{k})+\frac{\alpha}{0.99}(\Delta s,\Delta z)\in\mathcal{K}\right\}$
34:     Update iterates:
35:      $x_{k+1}\leftarrow x_{k}+\alpha\Delta x$
36:      $s_{k+1}\leftarrow s_{k}+\alpha\Delta s$
37:      $y_{k+1}\leftarrow y_{k}+\alpha\Delta y$
38:      $z_{k+1}\leftarrow z_{k}+\alpha\Delta z$
39:until Stopping criteria (29) satisfied.
3 Algorithm Implementation
This section describes the implementation of the primal-dual interior point method in qoco and qocogen. We focus on three key aspects: the matrix factorization used to solve the system in (19), numerical techniques for ensuring a robust factorization and solution of (19), and the stopping criteria. These aspects apply to both qoco and qocogen, but while both use the same underlying matrix factorization, their implementations are different. These differences will be discussed in Section 4.1.
3.1 Linear System Solve
The coefficient matrix in (19), which we will refer to as the Karush-Kuhn-Tucker (KKT) matrix is given as
$K=\begin{bmatrix}P&A^{\top}&G^{\top}\\ A&0&0\\ G&0&-W_{k}^{\top}W_{k}\end{bmatrix}.$ (24)
Algorithm 1 requires solving the linear system $K\xi=r$ for two different residual vectors $r$. Since $K$ remains constant for both solves, we factorize it once and then apply two backsolves.
The KKT matrix (24) is symmetric but indefinite, meaning it has both positive and negative eigenvalues. A common approach for factoring such matrices is the Bunch-Parlett factorization [55], which factors $K$ as
$\Pi K\Pi^{\top}=LDL^{\top},$ (25)
where $\Pi$ is a permutation matrix, $L$ is a lower triangular matrix, and $D$ is a block diagonal matrix with $1\times 1$ and $2\times 2$ blocks. However, $\Pi$ must be selected during runtime with knowledge of the data in $K$, as the factorization may not exist for all permutations. Additionally, $K$ may not be invertible if, for example, $A$ is not full row rank.
Static Regularization
To ensure a numerically stable factorization that exists for all permutations $\Pi$ (allowing us to select $\Pi$ based solely on the sparsity pattern of $K$), we apply static regularization [56, 12, 34] to the KKT matrix
$\hat{K}=\left[\begin{array}[]{c|cc}P&A^{\top}&G^{\top}\\ \hline\cr A&0&0\\ G&0&-W_{k}^{\top}W_{k}\end{array}\right]+\left[\begin{array}[]{c|cc}\epsilon_{s}I&0&0\\ \hline\cr 0&-\epsilon_{s}I&0\\ 0&0&-\epsilon_{s}I\\ \end{array}\right].$ (26)
In qoco and qocogen, we use $\epsilon_{s}=10^{-8}$.
Now $\hat{K}$ is a quasidefinite matrix, which is a special case of a symmetric indefinite matrix where the $(1,1)$ block is positive definite and the $(2,2)$ block is negative definite [57].
For such matrices, the $LDL^{\top}$ factorization
$\Pi K\Pi^{\top}=LDL^{\top},$ (27)
exists for any permutation $\Pi$, and $D$ will be a diagonal matrix with known signs for its diagonal elements [57].
A judicious choice of $\Pi$ can minimize the fill-in for the factor $L$. Fill-in is the introduction of non-zero elements in positions of a matrix where they previously did not exist. Excessive fill-in can lead to several negative consequences, including increased storage requirements (since more non-zero elements must be stored in memory) and increased computational cost (due to more floating-point operations). Finding an optimal $\Pi$ to minimize fill-in is NP-complete [58], but heuristic methods such as the Approximate Minimum Degree (AMD) ordering can effectively compute near-optimal permutations [45]. In qoco and qocogen, we use Tim Davis’ implementation of the AMD ordering [46] to determine $\Pi$.
After computing $\Pi$, the $LDL^{\top}$ factorization consists of two phases: a symbolic and numeric factorization. The symbolic factorization determines the elimination tree, a data structure which encodes dependencies between elements in the matrix during the factorization process, and the sparsity pattern of the factor $L$. It only requires the sparsity pattern of the regularized KKT matrix $\hat{K}$, which remains unchanged throughout Algorithm 1. Therefore, the symbolic factorization needs to be computed only once. The numeric factorization calculates the numerical values of $L$ and $D$, which depend on the values of the nonzero elements of $\hat{K}$. Since these values change at each iteration of Algorithm 1, the numeric factorization must be computed at every step.
To compute the $LDL^{\top}$ factorization, we use a modified version of qdldl [56] with dynamic regularization for qoco, and a custom $LDL^{\top}$ routine for qocogen which we will discuss in Section 4.1.
Dynamic Regularization
Although the $LDL^{\top}$ factorization must theoretically exist for $\hat{K}$, due to the nature of floating-point arithmetic it is possible for diagonal elements $D_{ii}$ to be rounded to zero or be very close to zero leading to divide-by-zero errors and the factorization failing. To alleviate this and factor $\hat{K}$ in a stable manner, we apply dynamic regularization as given by Algorithm 2, which bounds the magnitude of $D_{ii}$ away from zero, by some $\epsilon_{d}>0$, and corrects for its sign if necessary [12, 34, 32]. In qoco and qocogen, we use $\epsilon_{d}=10^{-8}$.
Algorithm 2 Dynamic Regularization
1:Given: $\epsilon_{d}>0$
2:if $D_{ii}$ should be positive then
3:     if $D_{ii}<0$ then
4:         $D_{ii}=\epsilon_{d}$
5:     else
6:         $D_{ii}=D_{ii}+\epsilon_{d}$
7:     end if
8:else if $D_{ii}$ should be negative then
9:     if $D_{ii}>0$ then
10:         $D_{ii}=-\epsilon_{d}$
11:     else
12:         $D_{ii}=D_{ii}-\epsilon_{d}$
13:     end if
14:end if
Iterative Refinement
After static and dynamic regularization we end up solving the linear system $\hat{K}\xi=r$ rather than system $K\xi=r$. To get a solution to the latter system, we apply iterative refinement [59, Section 2.5.1]. This process corrects the solution by iteratively reducing the residual error, $\|r-K\xi\|$, ensuring higher numerical accuracy without requiring an additional factorization. It involves iteratively solving the system
$\hat{K}\Delta\xi^{k}=r-K\xi^{k},$ (28)
and updating $\xi^{k+1}=\xi^{k}+\Delta\xi^{k}$, where $\xi^{0}$ solves the first linear system (i.e. $\hat{K}\xi^{0}=r$). Note that since we have already computed the factorization of $\hat{K}$, one iteration of Equation (28) only requires backsolves and no matrix factorization.
3.2 Stopping Criterion
Algorithm 1 terminates when three conditions are met: primal feasibility (29a), stationarity (29b), and complementary slackness (29c). We use both absolute and relative tolerances to ensure robustness across different problem scales. Our stopping criteria is met if the following conditions hold:
$\left\|\begin{bmatrix}Ax_{k}-b\\ Gx_{k}+s_{k}-h\end{bmatrix}\right\|_{\infty}$ $\leq\epsilon_{\mathrm{abs}}+\epsilon_{\mathrm{rel}}\max\left\{\|Ax_{k}\|_{\infty},\|b\|_{\infty},\|Gx_{k}\|_{\infty},\|h\|_{\infty},\|s_{k}\|_{\infty}\right\}$ (29a)
$\|Px_{k}+A^{\top}y_{k}+G^{\top}z_{k}+c\|_{\infty}$ $\leq\epsilon_{\mathrm{abs}}+\epsilon_{\mathrm{rel}}\max\left\{\|Px_{k}\|_{\infty},\|A^{\top}y_{k}\|_{\infty},\|G^{\top}z_{k}\|_{\infty},\|c\|_{\infty}\right\}$ (29b)
$|s_{k}^{\top}z_{k}|$ $\leq\epsilon_{\mathrm{abs}}+\epsilon_{\mathrm{rel}}\max\left\{1,|p_{k}|,|d_{k}|\right\},$ (29c)
where the primal and dual objectives are
$p_{k}:=\frac{1}{2}x_{k}^{\top}Px_{k}+c^{\top}x_{k},\quad d_{k}:=-\frac{1}{2}x_{k}^{\top}Px_{k}-b^{\top}y_{k}-h^{\top}z_{k}.$
4 Sparsity Exploiting Custom Solver
This section discusses qocogen, which generates an implementation of Algorithm 1 called qococustom. Unlike generic solvers which use sparse linear algebra, qococustom uses a custom $LDL^{\top}$ factorization and other tailored linear algebra routines, which exploit the known sparsity structure of problem data, to solve Problem 1 significantly faster.
While qoco can solve instances of Problem 1 with any sparsity pattern for $P,A,$ and $G$ and cone $\mathcal{K}$, the use of custom linear algebra in qococustom restricts it to instances with the same sparsity structure for $P,A,$ and $G$ and optimizes over the same cone $\mathcal{K}$. This makes qococustom useful for applications that repeatedly solve optimization problems with identical sparsity structures, such as sequential convex programming [8].
Figure 2 illustrates the usage of qoco, qocogen, and qococustom. In the figure, the matrices $(P,A,G)$ contain two pieces of information: their sparsity patterns $\mathrm{str}(P,A,G)$ (the location of nonzero elements) and the values of the nonzero elements $\mathrm{data}(P,A,G)$. We see that qoco takes in $(P,A,G)$, which contains both the sparsity patterns and nonzero elements for the matrices, and the rest of the problem data and returns the optimal solution $(x^{*},s^{*},y^{*},z^{*})$. In contrast, qocogen only takes in the sparsity patterns of $(P,A,G)$ and cone $\mathcal{K}$, and then generates qococustom. This instance of qococustom can then solve optimization problems where the sparsity patterns of $(P,A,G)$ do not change, but the nonzero values for $(P,A,G)$ and the vector data $(c,b,h)$ can change.
Although qoco avoids dynamic memory allocation during the solution process, it still relies on setup and cleanup functions for dynamic memory allocation and deallocation before and after solving the problem. In contrast, qococustom exclusively uses static memory allocation, making it well-suited for resource-constrained embedded systems, where dynamic memory allocation can lead to non-deterministic behavior and memory fragmentation.
[…]
Figure 2: Usage of qoco, qocogen, and qococustom.
4.1 Custom Linear Algebra
A key difference between qococustom and qoco is that qococustom uses custom linear algebra routines rather than sparse linear algebra. The main motivation for custom routines is to speed up the $LDL^{\top}$ factorization of the regularized KKT matrix, which is the most computationally expensive step of Algorithm 1. Typically, the KKT matrix is factored using a sparse linear algebra routine that stores the matrix in either compressed sparse column (CSC) or compressed sparse row (CSR) format. However, these sparse routines involve more than just the floating-point operations required for factorization. They also incur additional overhead to locate nonzero elements and their positions in the matrix. For example, in a CSC matrix, accessing data first requires indexing into the column pointer and row index arrays. This results in more CPU instructions, increased memory accesses, and a higher likelihood of cache misses. We provide empirical evidence supporting these claims in Appendix B.
When the sparsity structure of the KKT matrix is known a priori it is possible to write a custom $LDL^{\top}$ factorization routine which only contains code to perform the necessary floating-point operations and data accesses, eliminating the additional overhead that typical sparse linear algebra routines have. This makes the custom $LDL^{\top}$ factorization significantly faster than a sparse $LDL^{\top}$ factorization. We also note that since the AMD ordering we employ to compute the permutation matrix $\Pi$ only depends on the sparsity pattern of the KKT matrix, we can determine $\Pi$ at the time of code generation, so the generated code does not include the code to compute the AMD ordering.
A concrete example of a custom linear algebra routine for matrix-vector multiplication is given by Listing 1, where the sparsity pattern of $A$ is fixed and defined in Equation (30). Note that the custom routine only includes the necessary floating-point operations. Although Listing 1 depicts Python code, the custom linear algebra routines in qococustom are in C.
$\begin{bmatrix}y_{0}\\ y_{1}\\ y_{2}\end{bmatrix}=\underbrace{\begin{bmatrix}a_{00}&0&a_{02}\\ 0&0&0\\ a_{20}&0&0\end{bmatrix}}_{A}\begin{bmatrix}x_{0}\\ x_{1}\\ x_{2}\end{bmatrix}$ (30)
1 # Adata stores the nonzeros of A in column-major format, i.e.
2 # Adata[0] = a00
3 # Adata[1] = a20
4 # Adata[2] = a02
5 def custom_matvec(y, Adata, x):
6 y[0] = Adata[0]*x[0] + Adata[2]*x[2]
7 y[1] = 0
8 y[2] = Adata[1]*x[0]
Listing 1: Custom matrix-vector multiplication.
A drawback of these custom routines is that the amount of code generated increases with the problem size. As a result, the time to generate and compile qococustom can become burdensome as the problem sizes get very large.
4.2 Code Generation
qocogen is written is Python and example code for generating qococustom for a problem family defined by Equation (31) is given by Listing 2. The generated solver qococustom is customized to the sparsity pattern of the matrices $P,A,G$ which are passed to the function qocogen.generate_solver() as sparse scipy matrices.
[…]
The file tree for qococustom is given by Figure 3. The file qoco_custom.c contains the qoco_custom_solve() function which implements Algorithm 1, ldl.c contains the custom $LDL^{\top}$ factorization of the regularized KKT matrix, CMakeLists.txt is the CMake configuration file which automates the build process [60], runtest.c gives an example of calling qococustom with default data and times the resulting solve, and the remaining files contain various helper functions needed to implement Algorithm 1.
1 import qocogen
2 import numpy as np
3 from scipy import sparse
4
5 # Define problem data
6 P = sparse.diags([2, 2, 2, 0], 0).tocsc()
7
8 c = np.array([0, 0, 0, 1])
9 G = -sparse.identity(4).tocsc()
10 h = np.zeros(4)
11 A = sparse.csc_matrix([[1, 1, 0, 0], [0, 1, 1, 0]])
12 b = np.array([1, 1])
13
14 l = 1
15 n = 4
16 m = 4
17 p = 2
18 nsoc = 1
19 q = np.array([3])
20
21 # Generate custom solver.
22 qocogen.generate_solver(n, m, p, P, c, A, b, G, h, l, nsoc, q)
Listing 2: Generating a custom solver with qocogen.
[…]
Figure 3: qococustom file structure.
Listing 3 provides an example of calling qococustom from C, updating the problem data, and solving it again.
1 #include "qoco_custom.h"
2
3 int main() {
4 // Instantiate workspace.
5 Workspace work;
6
7 // Set default settings, but settings can be modified with
8 // work.settings.setting_name = ...
9 set_default_settings(&work);
10
11 // Loads the default data for the problem
12 // (i.e. the data passed to the qocogen.generate() call).
13 load_data(&work);
14
15 // Solve original problem.
16 qoco_custom_solve(&work);
[…]
18
19 // Can modify non-zero elements of P,A,G and c, b, h.
20 update_P(&work, Pnew);
21 update_A(&work, Anew);
22 update_G(&work, Gnew);
23 update_c(&work, cnew);
24 update_b(&work, bnew);
25 update_h(&work, hnew);
26
27 // Solve updated problem.
28 qoco_custom_solve(&work);
[…]
30 }
Listing 3: Calling qococustom from C/C++.
5 Numerical Results
We benchmarked qoco and qococustom against several conic interior-point solvers: the open-source solvers ecos [32] and clarabel [12], the commercial solvers gurobi [13] and mosek [33], and against the academically licensed custom solver generator cvxgen which generates a primal-dual interior point method for quadratic programs. We did not test against the custom solver generator bsocp since it is not academically licensed. For all solvers, we used their default settings but set the tolerances $\epsilon_{abs}=\epsilon_{rel}=10^{-7}$, and disabled the presolve features of gurobi and mosek to ensure all solvers are given the same problem formulation.
We evaluated our solvers on benchmark problems spanning multiple fields and a range of problem sizes. In our experiments, we define problem size as the total number of nonzero elements in $A$, $G$, and in the upper half of $P$. For all benchmark problems, we tested qoco, but only tested qococustom for the smaller problems, because code generation and compile times become prohibitively long for larger problem instances. We also evaluated qoco on the difficult Maros–Mészáros problems, a repository of 138 challenging quadratic programs (QPs) [61]. To evaluate the solvers’ performance, we used performance profiles [62] and the shifted geometric mean, both of which are frequently used to assess solver performance [12, 32, 56, 14, 63].
All numerical results were generated on a desktop computer with an AMD Ryzen 9 7950X3D processor running at 4.2 GHz and with 128 GB of RAM. Our benchmarks are written in Python and we used CVXPY [41, 42] to interface with the solvers. Our benchmarks are publicly available [64].
We provide detailed numeric results in Appendix D, which includes iteration counts and solve times for each solver on all problems we ran. Because Gurobi does not report iteration counts via the CVXPY interface for SOCPs, these values are omitted from our results.
Shifted Geometric Mean
We use the normalized shifted geometric mean to assign a scalar value to the performance of a solver on a test set of problems. The shifted geometric mean of $N$ runtimes for solver $s$ is computed as
$g_{s}=\left[\prod_{p=1}^{N}(t_{s,p}+k)\right]^{1/N}-k,$
where $t_{s,p}$ is the runtime of solver $s$ on problem $p$ and $k=1$ is the shift. If the solver does not find an optimal solution, then the runtime is set to the maximum allowable runtime (10 seconds for benchmark problems and 1200 seconds for the Maros–Mészáros problems).
We then define the normalized shifted geometric mean for solver $s$ as
$r_{s}=g_{s}/\min_{s}g_{s},$
so the solver with the lowest shifted geometric mean will have a normalized shifted geometric mean of $1.00$.
Performance Profiles
We also plot the relative and absolute performance profiles to compare solvers. We first define the relative performance ratio of solver $s$ on problem $p$ as
$u_{s,p}=t_{s,p}/\min_{s}t_{s,p}.$
The relative performance profile then plots $f^{r}_{s}(\tau)$ where
$f^{r}_{s}(\tau)=\frac{1}{N}\sum_{p=1}^{N}\mathcal{I}_{\leq\tau}(u_{s,p}),$
and $\mathcal{I}_{\leq\tau}(z)=1$ if $z\leq\tau$ and $\mathcal{I}_{\leq\tau}(z)=0$ otherwise. If the relative performance profile passes through the point $(x,y)$ for solver $s$, this means that solver $s$ solves a fraction $y$ of the problems within a factor of $x$ of the fastest solver.
The absolute performance profile plots $f^{a}_{s}(\tau)$ where
$f^{a}_{s}(\tau)=\frac{1}{N}\sum_{p=1}^{N}\mathcal{I}_{\leq\tau}(t_{s,p}).$
If the absolute performance profile passes through the point $(x,y)$ for solver $s$, this means that solver $s$ solves a fraction $y$ of the problems within $x$ time. For both relative and absolute profiles, the profile of the highest-performing solver will lie above the profiles of the other solvers.
5.1 Benchmark Problems
We solved optimization problems from five problem classes: robust Kalman filtering, group lasso regression [65], the losslessly convexified powered-descent guidance problem [37], Markowitz portfolio optimization [3], and the oscillating masses control problem [66]. The first three of these are SOCPs and the last two are QPs. For each of the five classes, we considered 10 different problem sizes, generating 20 unique problem instances per size, for a total of 1000 distinct optimization problems. Details on the problems we solved can be found in Appendix C. For these problems, we plotted runtime as a function of problem size in addition to reporting performance profiles and shifted geometric means.
For each solver, we solved each optimization problem 100 times and recorded the minimum runtime, which includes both setup and solve time. Since the computer runs various background processes that can artificially inflate execution time, we conducted 100 runs and reported the minimum execution time, thereby providing a more accurate measure of the solver’s performance.
Although the oscillating masses and portfolio optimization problems are QPs and could be solved with cvxgen, we limited testing to the two smaller oscillating masses problems. For larger instances, cvxgen was unable to generate code due to the size of the problems. For the portfolio optimization problems, we opted not to test cvxgen due to the sparsity pattern of the factor matrix $F$ in Problem 34. To specify the sparsity pattern of $F$ in cvxgen, we must manually provide the locations of nonzero elements into the web interface, which would have been prohibitively expensive, since $F$ has hundreds of nonzero elements. In contrast, to generate code with qocogen, we pass in problem data as sparse scipy matrices rather than manually specifying sparsity patterns. We could have specified $F$ as a dense matrix and only populated the nonzero elements for cvxgen, this would severely hinder its performance and result in an unfair comparison.
When computing normalized shifted geometric mean, relative performance profiles, and absolute performance profiles, all solvers must be tested on the same set of problems. Since we test qococustom on only half of the problems we provide two versions of the aforementioned metrics. The first version, given by Figure 5, includes clarabel, ecos, gurobi, mosek, and qoco evaluated on all problems. The second version, given by Figure 6, also includes qococustom, but the solvers are evaluated on the half of the benchmark problems which qococustom is tested on.
Figures 4, 5, and 6 show that qoco is the fastest generic solver on our benchmark problems. Additionally, qococustom is the fastest solver by a significant margin and is as performant as cvxgen on the $40$ smallest oscillating mass problems. This latter result is unsurprising, as qococustom uses an identical algorithm to cvxgen when solving QPs and both utilize custom linear algebra routines.
Figure 4: Solvetime in seconds vs problem size for benchmark problems
(a) Relative performance profile
(b) Absolute performance profile
QOCO Clarabel ECOS Gurobi Mosek
Shifted GM 1.0 1.5 2.4 14.4 3.8
Failure Rate (%) 0.0 0.0 0.0 0.0 0.0
(c) Shifted geometric means and failure rates
Figure 5: Performance profiles for all benchmark problems
(a) Relative performance profile
(b) Absolute performance profile
QOCO Custom QOCO Clarabel ECOS Gurobi Mosek
Shifted GM 1.0 2.1 3.5 6.9 30.4 15.1
Failure Rate (%) 0.0 0.0 0.0 0.0 0.0 0.0
(c) Shifted geometric means and failure rates
Figure 6: Performance profiles for benchmark problems with custom solver
5.2 Maros–Mészáros problems
The Maros–Mészáros problems are a set of 138 QPs that include a wide range of problem sizes (from $3$ to around $300,000$ variables) and contain very difficult problems that have ill-conditioning and poor scaling [61]. These properties make the Maros–Mészáros problems a good test set to assess the robustness of solvers. We test qoco but not qococustom on these problems since many problems are far too large to generate code for. Additionally, we ran each problem once rather than the 100 runs we did for the benchmark problems since many of the Maros–Mészáros problems are extremely large, and running them 100 times per solver would be extremely cumbersome.
Figure 7 shows that the proprietary solver Gurobi successfully solved the largest fraction of the problems followed by qoco, clarabel, mosek, and ecos. We can also see that for the majority of the problems, qoco was the fastest solver, whereas mosek and ecos tended to be the slowest. The latter result is unsurprising as mosek and ecos can only handle linear objectives and the reformulation of the quadratic objective into a second-order cone likely introduced significant fill-in, slowing these solvers down. We also observed that for many of the largest problems, Gurobi was the fastest solver, due to multithreading in its matrix factorization.
(a) Relative performance profile
(b) Absolute performance profile
QOCO Clarabel ECOS Gurobi Mosek
Shifted GM 2.0 3.0 27.4 1.0 3.8
Failure Rate (%) 6.5 8.7 31.9 2.9 9.4
(c) Shifted geometric means and failure rates
Figure 7: Performance profiles for Maros–Mészáros problems
6 Conclusion
We have presented qoco, a C-based solver, and qocogen, a custom solver generator for quadratic objective second-order cone programs (SOCPs). Both implement primal-dual interior-point methods, with qocogen generating custom solvers written in C, called qococustom, that exploit the sparsity pattern of problems to gain a computational advantage. We demonstrate that qoco is robust and faster than most of its competitors and qococustom is significantly faster than qoco, making it a useful tool for real-time applications, such as model-predictive control, trajectory optimization, and other domains where computational efficiency is critical.
Since both qoco and qocogen are open-source, they are accessible to users in both academia and industry. qoco and qocogen are integrated with CVXPY and CVXPYgen respectively, allowing users to formulate optimization problems in a natural way following from math rather than manually converting the problem to the solver-required standard form. This integration enhances usability of our solvers.
Documentation and installation instructions for qoco and qocogen are available at https://qoco-org.github.io/qoco/index.html.
Acknowledgments
This research was supported by ONR grant N00014-20-1-2288, AFOSR grant FA9550-20-1-0053, and Blue Origin, LLC. Government sponsorship is acknowledged. The authors would like to thank Danylo Malyuta for his review of this paper.
Appendix A Nonsymmetric Newton System
In this section we will show that Newton steps applied to Equations (5a), (5b), (5c), (5d) result in a nonsymmetric linear system which is less efficient to store and factor than the symmetric system presented in Equation (19).
We must first introduce the arrow matrix, a matrix that will allow us to rewrite the Jordan product, defined in Equation (6), as a matrix-vector multiplication. The arrow matrix can be written for vectors in the non-negative orthant and second-order cone as
$\mathrm{Arw}(x)=\begin{cases}\mathrm{diag}(x)\qquad\;\;\text{if }x\in\mathbb{R}^{l}_{+}\\ \begin{bmatrix}x_{0}&x_{1}^{\top}\\ x_{1}&x_{0}I\end{bmatrix}\quad\text{if }x\in\mathcal{Q}^{q}.\end{cases}$
For a vector in cone $\mathcal{K}$ where
$\mathcal{K}=\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K},$
and $\mathcal{C}_{i}$ is the non-negative orthant or a second-order cone, we can write the arrow matrix as
$\mathrm{Arw}(x)=\mathrm{blkdiag}(\mathrm{Arw}(x_{1}),\ldots,\mathrm{Arw}(x_{K}))\quad\text{where }x_{i}\in\mathcal{C}_{i},$
Note that if $x$ is in the interior of $\mathcal{K}$, then $\mathrm{Arw}(x)$ is positive definite and thus is invertible.
We can then write the Jordan product for two vectors $u,v\in\mathcal{K}$ as
$x\circ y=\mathrm{Arw}(x)y,$
We now linearize Equations (9a) - (9e) about the current iterate, $(x_{k},s_{k},y_{k},z_{k})$, write the Jordan products using the arrow matrix and get the linear system
$\begin{bmatrix}P&0&A^{\top}&G^{\top}\\ 0&\mathrm{Arw}(z_{k})&0&\mathrm{Arw}(s_{k})\\ A&0&0&0\\ G&I&0&0\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y\\ \Delta z\end{bmatrix}=\begin{bmatrix}-r_{x}\\ -r_{s}\\ -r_{y}\\ -r_{z}\end{bmatrix}$
We can attempt to symmetrize the coefficient matrix by multiplying the second equation by $\mathrm{Arw}(s_{k})^{-1}$ to get
$\begin{bmatrix}P&0&A^{\top}&G^{\top}\\ 0&\mathrm{Arw}(s_{k})^{-1}\mathrm{Arw}(z_{k})&0&I\\ A&0&0&0\\ G&I&0&0\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y\\ \Delta z\end{bmatrix}=\begin{bmatrix}-r_{x}\\ -\mathrm{Arw}(s_{k})^{-1}r_{s}\\ -r_{y}\\ -r_{z}\end{bmatrix}.$ (32)
However, the coefficient matrix above will only be symmetric if $\mathrm{Arw}(s_{k})^{-1}\mathrm{Arw}(z_{k})$ is symmetric. This can only be ensured if $\mathcal{K}$ only consisted of the non-negative orthant and no second-order cones. In other words, the system in (32) would only be symmetric if Problem 1 was a quadratic program rather than a second-order cone program. Even if we were to eliminate $\Delta s$ from Equation (32), we would still not have a symmetric linear system.
Appendix B Custom LDL Factorization Performance
This section provides empirical evidence supporting our claims in Sections 1 and 4.1. Specifically, we show that sparse linear algebra incurs extra overhead from identifying and locating nonzero elements, leading to more CPU instructions, memory accesses and cache misses. In contrast, a custom implementation that hardcodes the exact memory accesses and floating point operations reduces these inefficiencies. We specifically compare qdldl (used in qoco) against our custom $LDL^{\top}$ factorization (used in qococustom), since the most expensive operation in a generic implementation of Algorithm 1 is the $LDL^{\top}$ factorization. To compare the two matrix factorization implementations, we use Linux perf, a profiling tool that provides CPU performance metrics, which will allow us to quantify the computational overhead of sparse versus custom factorizations.
As a representative example of a matrix we would factor, we consider the KKT matrix associated with the Linear Quadratic Regulator (LQR). The LQR problem is
$\begin{split}\underset{x,u}{\mathrm{minimize}}\quad&\frac{1}{2}\left(\sum_{k=1}^{T}x_{k}^{\top}Q_{k}x_{k}+\sum_{k=1}^{T-1}u_{k}^{\top}R_{k}u_{k}\right)\\ \mathrm{subject\;to}\quad&x_{k+1}=A_{k}x_{k}+B_{k}u_{k}\;\quad\forall k\in[1,T-1]\\ &x_{1}=x_{\mathrm{init}},\\ \end{split}$
and its associated KKT matrix is
$K=\begin{bmatrix}P&H^{\top}\\ H&-\epsilon I\end{bmatrix},$
where
$\begin{split}P&=\mathrm{blkdiag}(Q_{1},Q_{2},\ldots,Q_{T},R_{1},R_{2},\ldots,R_{T-1})\\ H&=\begin{bmatrix}A_{1}&-I&&&&&B_{1}\\ &A_{2}&-I&&&&&B_{2}\\ &&&\ddots&&&&&\ddots\\ &&&&A_{N-1}&-I&&&&B_{N-1}\\ I&&&&&&&&&\\ \end{bmatrix}.\end{split}$
Notice that we apply static regularization to ensure that the $(2,2)$ block of K is negative definite, making $K$ quasidefinite. The sparsity pattern of $K$ is depicted by Figure 8.
Figure 8: Sparsity pattern of LQR KKT matrix for T=5
For simplicity, we assume time-invariant system and cost matrices, i.e., $Q_{k}=R_{k}=I$ and fixed matrices $A\in\mathbb{R}^{6\times 6}$, $B\in\mathbb{R}^{6\times 3}$. We choose the elements of $A$ and $B$ randomly and consider time horizons $T$: $\{5,15,50,75,100\}$. The actual elements of these matrices do not matter much, as our goal with these experiments is to understand the CPU behavior which (to first order) is unaffected by the value of elements in the matrices. To get accurate profiling results, we run each matrix factorization 1 million times to get the average runtime, number of instructions, number of L1 data cache loads, number of L1 cache misses, number of branches, and number of branch mispredictions.
From Table LABEL:table:perf, we can see that the custom $LDL^{\top}$ factorization is about four times faster than the sparse factorization because it does much less work: issuing fewer instructions, performing fewer memory accesses, and incurring fewer L1 cache misses. Although qdldl has a far more branches than our custom $LDL^{\top}$ implementation, the branches are quite predictable, and pipeline stalls are avoided.
By hardcoding the exact memory accesses and floating point operations, our custom $LDL^{\top}$ factorization eliminates much of the overhead associated with sparse matrix operations. This substantial reduction in memory accesses and instructions leads to a significantly faster factorization.
Table 1: Perf results for qdldl and custom $LDL^{\top}$ factorization
Runtime (us) Instructions L1 cache loads L1 cache misses Branches Branch misses
Size qdldl custom qdldl custom qdldl custom qdldl custom qdldl custom qdldl custom
5 2.103 0.435 36650 4858 14049 1851 0 0 5419 5 1 0
15 7.507 1.777 132796 18053 50550 6442 572 27 19124 8 4 0
50 27.077 6.117 469289 62046 179709 21468 2539 1204 67081 17 11 1
75 38.848 8.485 709731 93912 279033 32211 3735 1828 101347 24 14 2
100 49.577 12.894 950101 125747 365981 43559 5567 2459 135597 35 18 5
Appendix C Problem Classes
In this section, we describe the five problem classes considered in our numerical experiments, along with various problem sizes and instances within each problem class. We use the notation $\mathcal{U}(a,b)$ to denote a uniform distribution on the interval $[a,b]$, $\mathcal{N}(\mu,\sigma)$ for a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$, $0_{m\times n}$ for the matrix of all zeros in $\mathbb{R}^{m\times n}$, and $I_{n}$ for the identity matrix in $\mathbb{R}^{n\times n}$.
C.1 Robust Kalman Filter
We consider the robust Kalman filtering problem for vehicle tracking, as outlined in [67]. This problem is a quadratic objective second-order cone program (SOCP) and takes the form
$\begin{split}\underset{x_{k},w_{k},v_{k}}{\text{minimize}}\quad&\sum_{k=0}^{N-1}\left(\|w_{k}\|_{2}^{2}+\tau\phi_{\rho}(v_{k})\right)\\ \text{subject to}\quad&x_{k+1}=Ax_{k}+Bw_{k}\quad\forall k\in[0,N-1]\\ \quad&y_{k}=Cx_{k}+v_{k}\quad\forall k\in[0,N-1],\\ \end{split}$ (33)
where $\phi_{\rho}$ is the Huber loss function
$\phi_{\rho}(z)=\begin{cases}\|z\|_{2}^{2}&\|z\|_{2}\leq\rho\\ 2\rho\|z\|_{2}-\rho^{2}&\|z\|_{2}>\rho.\\ \end{cases}$
In this problem, $x_{k}\in\mathbb{R}^{4}$ represents the state of the vehicle at timestep $k$, $w_{k}\in\mathbb{R}^{2}$ is an unknown force vector acting on the vehicle at timestep $k$, $v_{k}\in\mathbb{R}^{2}$ is the measurement noise vector at timestep $k$, and $y_{k}\in\mathbb{R}^{2}$ represents a noisy measurement of the vehicle’s state at timestep $k$. The matrices $A$ and $B$ are known dynamics matrices for the system, and $C$ is the known observation matrix. To reconstruct the vehicle’s state trajectory, we solve Problem 33.
We use the system matrices
$A$ $=\begin{bmatrix}1&0&\left(1-\frac{\gamma}{2}\Delta t\right)\Delta t&0\\ 0&1&0&\left(1-\frac{\gamma}{2}\Delta t\right)\Delta t\\ 0&0&1-\gamma\Delta t&0\\ 0&0&0&1-\gamma\Delta t\end{bmatrix}$
$\vspace{10mm}B$ $=\begin{bmatrix}\frac{1}{2}\Delta t^{2}&0&\\ 0&\frac{1}{2}\Delta t^{2}\\ \Delta t&0\\ 0&\Delta t\end{bmatrix}$
$C$ $=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ \end{bmatrix},$
which correspond to a two-dimensional double integrator model with linear velocity drag and unit mass, where the input is a force and we observe the position of the vehicle. We use $\gamma=0.05$ for the velocity drag parameter, $\Delta t$ for the discretization time, and $\rho=2$ and $\tau=2$ for the Huber loss function.
Problem Sizes
To generate different problem sizes, we vary the number of timesteps $N$. We consider the following values for $N$: $\{25,50,75,125,175,225,300,375,450,500\}$. For each of these problem sizes, we set $\Delta t=T/(N-1)$, where $T=50$ is the time horizon.
Problem Instances
[…]
C.2 Lossless Convexification
The losslessly convexified powered-descent guidance problem is an optimal control problem that aims to compute a soft landing trajectory for a rocket while respecting relevant constraints. One important constraint is the lower thrust bound on the engine, which is nonconvex. However, it has been shown that this constraint admits a lossless convex relaxation [37, 38]. This means that the solution to the relaxed convex problem guarantees a solution to the original nonconvex problem. The losslessly convexified problem can be written as an SOCP
$\begin{split}\underset{x,z,u,\sigma}{\text{minimize}}\quad&-z_{T}\\ \text{subject to}\quad&x_{k+1}=Ax_{k}+Bu_{k}+g\quad\forall k\in[0,T-1]\\ \quad&z_{k+1}=z_{k}-\alpha\sigma_{k}\Delta t\quad\forall k\in[0,T-1]\\ \quad&\|u_{k}\|_{2}\leq\sigma_{k}\quad\forall k\in[0,T-1]\\ \quad&\log(m_{\mathrm{wet}}-\alpha\rho_{2}k\Delta t)\leq z_{k}\leq\log(m_{\mathrm{wet}}-\alpha\rho_{1}k\Delta t)\quad\forall k\in[0,T-1]\\ \quad&\mu_{1,k}\left[1-[z_{k}-z_{0,k}]+\frac{[z_{k}-z_{0,k}]^{2}}{2}\right]\leq\sigma_{k}\leq\mu_{2,k}[1-(z_{k}-z_{0,k})]\quad\forall k\in[0,T-1]\\ \quad&e_{3}^{\top}u_{k}\geq\sigma_{k}\cos(\theta_{\mathrm{max}})\quad\forall k\in[0,T-1]\\ \quad&x_{0}=x_{\mathrm{init}},\;z_{0}=\log(m_{\mathrm{wet}}),\;z_{T}\geq\log(m_{\mathrm{dry}}),\end{split}$
where $x_{k}\in\mathbb{R}^{6}$ is the position and velocity of the rocket at timestep $k$, $u_{k}\in\mathbb{R}^{3}$ is the thrust per unit mass at timestep $k$, $z_{k}\in\mathbb{R}$ is the natural logarithm of the rocket’s mass at timestep $k$, and $\sigma_{k}\in\mathbb{R}$ is a slack variable at timestep $k$. Further we define $z_{0,k}\coloneqq\log(m_{\mathrm{wet}}-\alpha\rho_{2}k\Delta t)$, $\mu_{1,k}\coloneqq\rho_{1}{\rm e}^{-z_{0}}$, and $\mu_{2,k}\coloneqq\rho_{2}{\rm e}^{-z_{0}}$. The parameter $\rho_{1}\in\mathbb{R}$ is the minimum thrust of the engine, $\rho_{2}\in\mathbb{R}$ is the maximum thrust of the engine, $\alpha\in\mathbb{R}$ is a mass depletion parameter that describes how efficient the engine is, $m_{\mathrm{wet}}\in\mathbb{R}$ is the sum of the dry mass of the rocket and the initial fuel mass, $\theta_{\mathrm{max}}\in\mathbb{R}$ is the maximum angle from vertical that the thrust vector can make, $x_{\mathrm{init}}\in\mathbb{R}^{6}$ is the initial state of the rocket, $\Delta t\in\mathbb{R}$ is the discretization time, and $e_{3}$ is the third canonical basis vector in $\mathbb{R}^{3}$. The dynamics matrices $A$ and $B$, and the vector $g$, are defined as
$A$ $=\begin{bmatrix}I_{3}&\Delta t(I_{3})\\ 0_{3\times 3}&I_{3}\end{bmatrix}$
$\vspace{10mm}B$ $=\begin{bmatrix}\frac{1}{2}\Delta t^{2}(I_{3})\\ \Delta t(I_{3})\end{bmatrix}$
[…]
For this problem class we use the parameters
$g_{0}=9.807,\;\rho_{1}=100,\;\rho_{2}=500,\;m_{\mathrm{dry}}=25,\;m_{\mathrm{wet}}=35,\;\theta_{\mathrm{max}}=\pi/4,\;\alpha=0.001.\$
Problem Sizes
To generate different problem sizes, we vary the number of timesteps $T$. The discretization time is computed as $\Delta t=t_{f}/(T-1)$, where $t_{f}=20$. We consider the following values for $T$: $\{15,50,75,100,125,150,200,250,300,350\}$.
Problem Instances
For each problem size (i.e., for each number of timesteps), we generate 20 random problem instances, which involves generating a random initial state $x_{\mathrm{init}}$. We generate $x_{\mathrm{init}}$ as
$x_{\mathrm{init}}\sim\begin{bmatrix}\mathcal{U}(-10,10)\\ \mathcal{U}(-10,10)\\ \mathcal{U}(200,400)\\ 0\\ 0\\ 0\end{bmatrix}.$
C.3 Group Lasso
The group lasso problem is a quadratic objective SOCP. It is a variation of the least-squares regression problem, where the regression variables are partitioned into disjoint groups, and a regularization term is added to encourage group sparsity. Let $x=[x^{(1)},x^{(2)},\ldots,x^{(N)}]$ represent the partitioning of the regression variables, where each $x^{(i)}$ is a group of variables. We can write the problem as
[…]
This can be written as an SOCP by introducing a slack variable $t\in\mathbb{R}^{N}$, and an auxiliary variable $y\in\mathbb{R}^{m}$ to avoid computing $A^{\top}A$ and maintain sparsity in the Hessian of the objective function. The problem is now
$\begin{split}\underset{x,t}{\text{minimize}}\quad&\|y\|_{2}^{2}+\lambda\sum_{i=1}^{N}t_{i}\\ \text{subject to}\quad&\|x^{(i)}\|_{2}\leq t_{i}\quad\forall i\in[1,N]\\ \quad&y=Ax-b,\end{split}$
where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$.
Problem Sizes
To generate different problem sizes, we vary the number of groups $N$. We consider the following values for $N$: $\{1,2,3,4,5,8,10,12,14,16\}$.
Problem Instances
For each problem size (i.e., for each number of groups) we generate 20 random problem instances, corresponding to the generation of the matrix $A$ and the vector $b$. We set $A$ to have $m=250N$ rows and $n=10N$ columns, where $x$ is partitioned into $N$ groups, each containing ten regression variables. The matrix $A$ is sparse, with 10% of its elements nonzero, and the nonzero elements are drawn from the uniform distribution $\mathcal{U}(0,1)$.
To generate $b$, we first generate $\hat{x}\in\mathbb{R}^{n}$ and partition it into $N$ groups of ten variables. Half of the groups are set to zero, and the remaining elements are drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$. Next, we generate an error vector $e\in\mathbb{R}^{m}$ with elements drawn from $\mathcal{N}(0,1/n)$. Finally, we compute $b=A\hat{x}+e$. Note that each problem instance will have a different sparsity pattern for $A$, which requires us to generate a new custom solver for each instance.
C.4 Portfolio Optimization
We consider the Markowitz portfolio optimization problem with a no-short-selling constraint [3, 51]. This problem is the QP
[…]
where $x\in\mathbb{R}^{n}$ is the allocation vector, with $x_{i}$ representing the percentage of resources invested in asset $i$, $\mu\in\mathbb{R}^{n}$ is the vector of expected returns, and $\Sigma\in\mathbb{S}^{n}_{+}$ is the covariance matrix of asset returns.
Typically, the return covariance matrix $\Sigma$ is approximated as the sum of a low-rank matrix and a diagonal matrix. Thus, we write $\Sigma=FF^{\top}+D$, where $F\in\mathbb{R}^{n\times k}$ is a factor matrix of rank $k$. We then add the auxiliary variable $y$ and the constraint $y=F^{\top}x$ to avoid computing $FF^{\top}$ and maintain sparsity in the Hessian of the objective function. The final problem can be rewritten as
$\begin{split}\underset{x,y}{\text{minimize}}\quad&x^{\top}Dx+y^{\top}y-\gamma^{-1}\mu^{\top}x\\ \text{subject to}\quad&y=F^{\top}x\\ \quad&\sum_{i=1}^{n}x_{i}=1\\ \quad&x_{i}\geq 0\quad\forall i\in[1,n].\\ \end{split}$ (34)
Problem Sizes
To generate different problem sizes, we vary the number of factors, $k$. We consider the following values for $k$: $\{2,4,6,8,10,15,20,25,30,35\}$.
Problem Instances
For each problem size (i.e., each number of factors), we generate 20 random problem instances. This corresponds to generating the matrices $D$, $F$, and the vector $\mu$. We set the number of assets to $n=100k$, and for all instances, we take $\gamma=1$. The diagonal elements of $D$ are drawn from $\mathcal{N}(0,1)$, and the elements of $\mu$ are drawn from $\mathcal{N}(0,\sqrt{k})$. The matrix $F$ is a sparse matrix, with 50% of its elements being nonzero, and its nonzero elements are drawn from $\mathcal{N}(0,1)$. Note that each problem instance will have a different sparsity pattern for $F$, which requires us to generate a new custom solver for each instance.
C.5 Oscillating Masses
The oscillating masses problem is an optimal control problem involving a system of $N$ masses in one dimension, where each mass is connected to its neighboring mass by a spring, and the outermost two masses are attached to a fixed wall. The goal is to apply forces to each mass to bring the system to equilibrium [66]. This problem can be formulated as a QP
$\begin{split}\underset{x,u}{\mathrm{minimize}}\quad&\frac{1}{2}\left(\sum_{k=0}^{T}x_{k}^{\top}Qx_{k}+\sum_{k=0}^{T-1}u_{k}^{\top}Ru_{k}\right)\\ \mathrm{subject\;to}\quad&x_{k+1}=Ax_{k}+Bu_{k}\;\quad\forall k\in[0,T-1]\\ &x_{0}=x_{\mathrm{init}}\\ &\|x_{k}\|_{\infty}\leq x_{\mathrm{max}}\quad\forall k\in[0,T]\\ &\|u_{k}\|_{\infty}\leq u_{\mathrm{max}}\quad\forall k\in[0,T-1].\end{split}$
where $x_{k}\in\mathbb{R}^{2N}$ is the position and velocity of the masses at timestep $k$, $u_{k}\in\mathbb{R}^{N}$ is force applied to each mass at timestep $k$, and $Q\in\mathbb{S}^{2N}_{+}$ and $R\in\mathbb{S}^{N}_{+}$ penalize deviation from equilibrium and control effort respectively.
The dynamics matrices $A$ and $B$ are derived from an exact discretization of the continuous-time linear dynamics, using a zero-order hold on control input
$A$ $={\rm e}^{A_{c}\Delta t}$
$\vspace{10mm}B$ $=A_{c}^{-1}(A-I_{2N})B_{c},$
where
$A_{c}$ $=\begin{bmatrix}0_{N\times N}&I_{N}\\ L_{N}&0_{N\times N}\end{bmatrix}$
$\vspace{10mm}B_{c}$ $=\begin{bmatrix}0_{N\times N}\\ I_{N}\end{bmatrix}.$
Here, $L_{N}\in\mathbb{R}^{N\times N}$ is a symmetric tridiagonal matrix with $-2$ on the main diagonal and $1$ on the subdiagonal and superdiagonal.
For this problem class, we use the parameters
$N=4,\;\Delta t=0.25,\;x_{\mathrm{max}}=2,\;u_{\mathrm{max}}=5.$
Problem Sizes
To generate different problem sizes, we vary the number of timesteps $T$. We consider the following values for $T$: $\{8,20,32,44,56,76,96,116,136,156\}$.
Problem Instances
For each problem size (i.e., each number of timesteps), we generate 20 random problem instances. This involves generating a random initial state $x_{\mathrm{init}}$, as well as the matrices $Q$ and $R$. We draw each element of $x_{\mathrm{init}}$ from $\mathcal{N}(0,1)$, then clip the elements such that they lie within the interval $[-0.9x_{\mathrm{max}},0.9x_{\mathrm{max}}]$. This ensures that the initial state satisfies the state constraints, making the resulting problem feasible. The matrices $Q$ and $R$ are generated as diagonal matrices, with each diagonal element drawn from the uniform distribution $\mathcal{U}(0,10)$.
Appendix D Detailed Numeric Results
Table 2: Iterations and solver runtimes for Maros–Mészáros problems
Iterations Solver Runtime (s)
Problem Size qoco clarabel ecos gurobi mosek qoco clarabel ecos gurobi mosek
LASER 9216 10 10 - 21 24 0.00258 0.00414 - 0.02421 0.03150
QPCBOEI1 4743 21 15 - 21 33 0.00613 0.00632 - 0.01534 0.03697
HUESTIS 39968 7 8 - 14 - 0.01261 0.02198 - 0.03369 -
QBANDM 3007 21 20 32 29 22 0.00345 0.00502 0.01198 0.01520 0.02451
QSC205 775 15 18 17 30 15 0.00111 0.00192 0.00183 0.00950 0.00823
LISWET11 40002 - - - 40 31 - - - 0.15115 0.26106
LISWET8 40002 - - - 417 70 - - - 1.83018 0.90131
AUG2DC 60200 0 0 - 2 15 0.01392 0.02377 - 0.02266 0.46025
QSEBA 6548 30 29 - 27 30 0.00717 0.01452 - 0.02000 0.05188
AUG2D 59800 0 0 - 2 11 0.01088 0.02133 - 0.01970 0.30137
QGROW15 7365 18 20 17 20 28 0.00616 0.01114 0.01225 0.02356 0.06942
CONT-200 319196 13 10 30 12 28 3.41491 3.91965 8.99180 0.75313 2.27627
LISWET1 40002 - - - 27 52 - - - 0.09510 0.55497
ZECEVIC2 9 6 7 8 8 7 0.00002 0.00005 0.00004 0.00275 0.00102
QFFFFF80 8997 95 25 45 48 22 0.08216 0.03295 0.49989 0.06666 0.10850
AUG2DCQP 80400 11 12 - 17 - 0.08341 0.14402 - 0.09956 -
QPILOTNO 16258 39 31 43 56 27 0.09448 0.11113 0.21933 0.12265 0.15589
CVXQP2_M 6733 13 9 30 15 17 0.03255 0.02544 11.03589 0.06452 1.07815
POWELL20 30000 38 - - 60 - 0.05451 - - 0.21773 -
STADAT2 15997 17 44 - 18 21 0.00939 0.04388 - 0.02200 0.04019
QPCBOEI2 1623 40 19 - 26 48 0.00298 0.00274 - 0.00741 0.01278
PRIMAL3 22292 11 8 18 17 8 0.01240 0.01680 0.03449 0.01265 0.04159
QSHIP08S 21178 14 15 26 24 18 0.02043 0.02884 1.58040 0.04072 0.17699
DUALC2 1645 18 10 13 11 16 0.00054 0.00061 0.00091 0.00860 0.00116
LISWET4 40002 10 22 - 13 19 0.01575 0.05955 - 0.05271 0.13555
QPCSTAIR 4872 25 22 - 23 39 0.00897 0.01026 - 0.02300 0.05212
PRIMALC2 2076 24 15 30 11 16 0.00130 0.00142 0.00353 0.00385 0.00828
LISWET6 40002 9 17 - 13 20 0.01906 0.05229 - 0.05445 0.14808
QSHIP04L 8506 14 15 26 21 28 0.00769 0.00918 0.02240 0.02517 0.06894
QE226 3824 18 24 22 25 19 0.00466 0.00709 0.02897 0.01216 0.01492
PRIMAL4 17520 9 9 16 16 11 0.01266 0.01348 0.02799 0.00947 0.04455
QSCSD8 13844 11 10 15 15 12 0.00735 0.01020 0.05332 0.02534 0.06593
QSCORPIO 1824 15 11 13 24 15 0.00128 0.00169 0.00258 0.00776 0.03019
CONT-201 290393 12 10 - 14 - 4.17722 5.71965 - 0.82300 -
GOULDQP2 3142 9 12 20 13 10 0.00188 0.00313 0.58865 0.00962 0.05237
LISWET10 40002 - - - 125 20 - - - 0.44295 0.14301
HS268 40 5 10 15 12 15 0.00003 0.00007 0.00009 0.00356 0.00057
QBORE3D 1834 18 28 24 26 20 0.00202 0.00442 0.00487 0.00950 0.01198
QSHIP12L 83825 18 15 34 23 39 0.21193 0.19337 79.39230 0.10314 4.08685
QSHIP04S 5866 14 14 24 19 21 0.00373 0.00592 0.01427 0.01438 0.04390
QSHARE1B 1415 28 28 46 30 31 0.00200 0.00343 0.00512 0.00929 0.02247
QGFRDXPN 3889 24 20 - 43 31 0.00589 0.00682 - 0.03540 0.08180
CONT-050 19796 10 9 24 10 18 0.03964 0.04369 0.10372 0.02808 0.08417
MOSARQP1 8467 9 10 24 15 11 0.00520 0.00843 0.03596 0.01658 0.02250
BOYD2 517049 68 - - 78 - 2.99925 - - 6.10176 -
QSCAGR25 2182 17 19 34 24 - 0.00160 0.00345 0.00844 0.01023 -
HUES-MOD 39968 13 12 - 13 23 0.02296 0.02834 - 0.02918 0.36701
YAO 8007 - - - 25 36 - - - 0.01411 0.05955
CONT-101 72693 10 10 - 13 25 0.30755 0.44878 - 0.09657 0.47646
HS53 24 5 6 9 8 7 0.00003 0.00004 0.00005 0.00273 0.00030
CVXQP1_M 7482 13 10 - 16 21 0.04223 0.03895 - 0.09931 0.43868
STCQP2 48135 7 8 25 12 14 0.08184 0.10365 0.53351 0.21152 0.16697
QGROW22 10930 22 25 20 25 31 0.00986 0.01891 0.02126 0.03386 0.10940
HS35 11 5 7 10 9 9 0.00002 0.00005 0.00004 0.00366 0.00077
STCQP1 48135 7 7 28 - 24 0.03555 0.04303 0.32154 - 0.24021
STADAT1 15997 20 15 - 185 - 0.01086 0.01243 - 0.17253 -
PRIMAL1 6140 12 9 17 18 8 0.00325 0.00402 0.00986 0.00567 0.00878
DUAL3 6441 9 11 13 11 7 0.00313 0.00378 0.00485 0.01875 0.00723
HS21 8 9 8 8 8 10 0.00002 0.00003 0.00003 0.00255 0.00027
QBEACONF 3664 12 - 29 18 22 0.00182 - 0.00717 0.00796 0.01373
QISRAEL 3109 26 25 31 23 23 0.00366 0.00576 0.01015 0.01732 0.00747
LISWET2 40002 9 17 - 15 20 0.01800 0.04961 - 0.05521 0.14195
DTOC3 49994 4 6 - 8 26 0.01081 0.02869 - 0.02627 0.38133
GOULDQP3 3840 6 7 25 9 17 0.00109 0.00237 3.86470 0.01007 0.17719
PRIMAL2 8691 11 8 16 17 8 0.00431 0.00527 0.01403 0.00586 0.01458
TAME 7 4 5 5 6 4 0.00002 0.00004 0.00003 0.00368 0.00031
PRIMALC5 2860 20 13 14 11 14 0.00141 0.00149 0.00241 0.00348 0.00938
AUG3DC 10419 0 0 14 2 10 0.00227 0.00270 0.03786 0.00950 0.04388
AUG3DCQP 14292 9 10 29 16 22 0.01343 0.01937 0.08650 0.02952 0.16977
QSCFXM2 7228 32 30 36 56 - 0.01022 0.01737 0.04520 0.05266 -
VALUES 4428 12 13 - - - 0.00149 0.00242 - - -
QAFIRO 121 15 13 14 15 13 0.00017 0.00023 0.00020 0.00347 0.00100
QETAMACR 7761 26 19 25 28 27 0.05831 0.04680 0.91543 0.04582 0.10767
QSCSD6 7070 14 13 16 17 13 0.00407 0.00613 0.02588 0.01711 0.05499
QSTAIR 5423 23 33 32 41 40 0.01108 0.01966 0.03227 0.03192 0.05891
PRIMALC1 2514 25 16 27 10 20 0.00237 0.00153 0.00347 0.00359 0.01178
QSHELL 40488 28 36 - 33 27 0.10774 0.12657 - 0.07681 0.18787
CVXQP2_L 67483 22 9 - 16 31 15.34297 9.37783 - 2.75816 750.19697
QSCFXM1 3779 26 25 35 49 31 0.00448 0.00762 0.02215 0.02710 0.08131
DUAL1 3813 10 12 14 12 12 0.00182 0.00219 0.00283 0.01026 0.00219
UBH1 66033 7 17 - 16 8 0.08430 0.08050 - 0.05857 0.15917
DUALC5 2276 10 10 18 14 12 0.00043 0.00084 0.00126 0.01239 0.00120
LISWET9 40002 - - - - 62 - - - - 0.75226
Q25FV47 71470 27 25 34 37 53 0.18642 0.18053 2.34299 0.16627 0.52865
PRIMALC8 5182 20 12 14 10 22 0.00247 0.00322 0.00401 0.00440 0.02708
QPCBLEND 657 16 16 17 20 17 0.00060 0.00095 0.00097 0.00530 0.00192
LOTSCHD 72 7 8 18 11 15 0.00004 0.00008 0.00011 0.00273 0.00067
QSCAGR7 585 15 15 29 23 32 0.00044 0.00084 0.00174 0.00461 0.00756
QSCRS8 4472 23 31 33 35 30 0.00722 0.01372 0.01325 0.02854 0.06120
CVXQP1_S 734 8 8 12 13 12 0.00054 0.00085 0.01330 0.00830 0.00374
QADLITTL 567 11 13 17 16 14 0.00041 0.00076 0.00140 0.00485 0.00326
CVXQP2_S 660 8 9 12 14 11 0.00045 0.00073 0.01244 0.00789 0.00310
QRECIPE 988 18 16 19 20 15 0.00110 0.00153 0.00195 0.00703 0.00469
QSCSD1 3893 11 10 12 16 12 0.00267 0.00363 0.00896 0.01347 0.01324
GENHS28 43 0 0 9 1 5 0.00002 0.00004 0.00008 0.00576 0.00039
QSCTAP1 2325 15 10 10 17 13 0.00176 0.00185 0.00272 0.00821 0.02044
QBRANDY 2462 19 18 28 40 28 0.00330 0.00420 0.00662 0.01947 0.02632
HS52 14 0 0 7 2 8 0.00002 0.00003 0.00005 0.00333 0.00039
HS51 14 0 0 7 2 7 0.00001 0.00002 0.00004 0.00242 0.00025
CVXQP3_S 808 9 10 13 14 14 0.00050 0.00089 0.01110 0.00681 0.00338
CVXQP3_L 82481 - 10 - 14 42 - 21.86824 - 6.19038 187.95385
STADAT3 31997 17 54 28 27 22 0.02096 0.10802 0.08724 0.06329 0.07766
S268 40 5 10 15 12 15 0.00002 0.00006 0.00007 0.00266 0.00044
QSTANDAT 5030 24 20 11 47 22 0.00536 0.00792 0.04844 0.03088 0.05748
EXDATA 1137750 16 21 19 14 13 4.82029 6.01972 36.06845 1.83909 0.83910
QGROW7 3550 19 21 16 20 27 0.00310 0.00556 0.00557 0.01416 0.03009
DUAL4 3024 8 11 13 10 13 0.00134 0.00224 0.00236 0.00856 0.00247
QFORPLAN 5591 28 21 - 56 - 0.00617 0.00799 - 0.02855 -
QSCTAP3 12401 16 11 11 19 12 0.01999 0.01361 0.07613 0.03330 0.08330
QSHARE2B 828 19 15 24 21 20 0.00065 0.00092 0.00141 0.00396 0.00258
HS35MOD 12 9 11 11 15 13 0.00002 0.00003 0.00005 0.00227 0.00032
DUAL2 4796 8 10 13 11 7 0.00168 0.00263 0.00429 0.00878 0.00594
MOSARQP2 4775 8 9 27 16 13 0.00445 0.00614 0.02766 0.01492 0.01722
QPTEST 10 5 7 9 10 10 0.00001 0.00003 0.00003 0.00225 0.00021
AUG3D 9219 0 0 13 2 6 0.00191 0.00252 0.03937 0.00711 0.04820
QSIERRA 11557 22 26 38 24 21 0.01371 0.02585 0.05065 0.03655 0.12325
LISWET5 40002 8 9 - 13 39 0.01421 0.03237 - 0.05394 0.40437
LISWET3 40002 9 16 - 13 20 0.01809 0.05045 - 0.05186 0.14603
LISWET12 40002 - - - - 60 - - - - 0.72638
CVXQP3_M 8231 23 11 - 18 21 0.08985 0.05016 - 0.14954 0.46277
HS118 108 11 11 9 13 10 0.00007 0.00012 0.00011 0.00229 0.00048
QSHIP12S 28344 18 15 31 24 26 0.02525 0.03026 7.90644 0.04829 0.79477
DUALC1 1998 23 11 29 15 14 0.00076 0.00074 0.00162 0.01125 0.00095
CVXQP1_L 74982 - 10 - 15 37 - 19.33309 - 4.44083 157.56906
LISWET7 40002 11 - - 58 44 0.01843 - - 0.26488 0.49234
CONT-100 79596 11 9 22 11 19 0.32731 0.38091 0.80000 0.08902 0.27225
QSCFXM3 10369 31 32 34 58 31 0.02114 0.02997 0.07188 0.08477 0.09320
DPKLO1 1652 0 0 13 2 4 0.00018 0.00027 0.00204 0.00283 0.00136
KSIP 18871 14 14 21 17 11 0.00666 0.00982 0.01018 0.01295 0.00380
DUALC8 4076 14 9 27 14 22 0.00111 0.00136 0.00493 0.02466 0.00201
QSHIP08L 52050 14 15 28 22 20 0.11742 0.14559 13.08404 0.08743 0.45644
AUG3DQP 13092 11 11 31 17 13 0.01592 0.02040 0.08446 0.03143 0.07889
AUG2DQP 80000 11 13 - 17 - 0.07453 0.14683 - 0.10080 -
HS76 20 6 6 10 10 7 0.00002 0.00004 0.00005 0.00300 0.00038
QSCTAP2 9371 14 10 10 18 11 0.00958 0.00923 0.03774 0.02440 0.07304
BOYD1 745507 37 - - 30 - 0.90575 - - 1.08528 -
QCAPRI 3147 40 31 - 33 39 0.00557 0.00889 - 0.02008 0.06086
CONT-300 653093 12 10 - 14 - 15.35224 15.73776 - 2.06562 -
Table 3: Iterations and solver runtimes for robust Kalman filter problems
Iterations Solver Runtime (s)
Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
robust_kalman_filter_N_25_i_0 775 12 15 - 12 9 9 0.00091 0.00167 0.00801 0.00420 0.00051 0.00020
robust_kalman_filter_N_25_i_1 775 10 15 - 11 9 9 0.00081 0.00170 0.00895 0.00415 0.00051 0.00020
robust_kalman_filter_N_25_i_2 775 11 18 - 12 10 10 0.00087 0.00197 0.00998 0.00416 0.00056 0.00021
robust_kalman_filter_N_25_i_3 775 10 17 - 12 9 9 0.00082 0.00170 0.00776 0.00398 0.00044 0.00020
robust_kalman_filter_N_25_i_4 775 11 15 - 12 9 9 0.00087 0.00166 0.00933 0.00478 0.00051 0.00019
robust_kalman_filter_N_25_i_5 775 10 14 - 10 8 8 0.00080 0.00160 0.00887 0.00400 0.00047 0.00018
robust_kalman_filter_N_25_i_6 775 11 15 - 12 9 9 0.00086 0.00169 0.01004 0.00413 0.00052 0.00020
robust_kalman_filter_N_25_i_7 775 10 17 - 11 9 9 0.00080 0.00193 0.00975 0.00389 0.00051 0.00020
robust_kalman_filter_N_25_i_8 775 10 15 - 11 9 9 0.00083 0.00163 0.00885 0.00390 0.00052 0.00020
robust_kalman_filter_N_25_i_9 775 9 18 - 11 9 9 0.00075 0.00176 0.00724 0.00356 0.00046 0.00018
robust_kalman_filter_N_25_i_10 775 11 16 - 13 10 10 0.00075 0.00153 0.00809 0.00419 0.00048 0.00020
robust_kalman_filter_N_25_i_11 775 12 19 - 12 10 10 0.00080 0.00177 0.00831 0.00378 0.00049 0.00020
robust_kalman_filter_N_25_i_12 775 10 17 - 12 9 9 0.00068 0.00159 0.00728 0.00372 0.00044 0.00018
robust_kalman_filter_N_25_i_13 775 8 18 - 13 8 8 0.00059 0.00175 0.00799 0.00413 0.00041 0.00016
robust_kalman_filter_N_25_i_14 775 10 17 - 12 9 9 0.00068 0.00162 0.00760 0.00381 0.00045 0.00019
robust_kalman_filter_N_25_i_15 775 10 16 - 11 8 8 0.00070 0.00157 0.00805 0.00351 0.00042 0.00016
robust_kalman_filter_N_25_i_16 775 10 17 - 12 9 9 0.00070 0.00161 0.00762 0.00386 0.00044 0.00018
robust_kalman_filter_N_25_i_17 775 9 17 - 11 9 9 0.00064 0.00180 0.01127 0.00367 0.00045 0.00018
robust_kalman_filter_N_25_i_18 775 11 16 - 12 9 9 0.00080 0.00156 0.00795 0.00396 0.00048 0.00018
robust_kalman_filter_N_25_i_19 775 11 19 - 11 10 10 0.00075 0.00196 0.00793 0.00370 0.00051 0.00021
robust_kalman_filter_N_50_i_0 1550 13 18 - 13 12 12 0.00167 0.00415 0.01567 0.00701 0.00127 0.00053
robust_kalman_filter_N_50_i_1 1550 14 19 - 14 12 12 0.00209 0.00410 0.01851 0.00789 0.00126 0.00054
robust_kalman_filter_N_50_i_2 1550 11 20 - 12 10 10 0.00167 0.00441 0.01895 0.00712 0.00108 0.00044
robust_kalman_filter_N_50_i_3 1550 11 19 - 12 10 10 0.00169 0.00426 0.02149 0.00714 0.00108 0.00040
robust_kalman_filter_N_50_i_4 1550 11 20 - 11 10 10 0.00166 0.00426 0.01637 0.00651 0.00108 0.00046
robust_kalman_filter_N_50_i_5 1550 11 19 - 11 10 10 0.00168 0.00386 0.01841 0.00680 0.00101 0.00045
robust_kalman_filter_N_50_i_6 1550 12 19 - 12 11 11 0.00157 0.00363 0.01517 0.00638 0.00103 0.00044
robust_kalman_filter_N_50_i_7 1550 10 17 - 10 9 9 0.00135 0.00338 0.01494 0.00546 0.00089 0.00036
robust_kalman_filter_N_50_i_8 1550 11 20 - 10 10 10 0.00145 0.00374 0.01444 0.00548 0.00096 0.00041
robust_kalman_filter_N_50_i_9 1550 12 24 - 12 10 10 0.00152 0.00454 0.01599 0.00647 0.00097 0.00041
robust_kalman_filter_N_50_i_10 1550 11 21 - 11 9 9 0.00145 0.00406 0.01574 0.00576 0.00093 0.00037
robust_kalman_filter_N_50_i_11 1550 11 19 - 12 10 10 0.00146 0.00374 0.01475 0.00612 0.00099 0.00040
robust_kalman_filter_N_50_i_12 1550 12 18 - 11 11 11 0.00155 0.00402 0.01724 0.00593 0.00117 0.00050
robust_kalman_filter_N_50_i_13 1550 13 21 - 14 12 12 0.00199 0.00425 0.01969 0.00778 0.00127 0.00048
robust_kalman_filter_N_50_i_14 1550 13 19 - 13 12 12 0.00167 0.00421 0.01539 0.00650 0.00109 0.00053
robust_kalman_filter_N_50_i_15 1550 11 19 - 12 10 10 0.00172 0.00364 0.01478 0.00624 0.00096 0.00039
robust_kalman_filter_N_50_i_16 1550 12 21 - 11 10 10 0.00155 0.00401 0.01620 0.00600 0.00094 0.00044
robust_kalman_filter_N_50_i_17 1550 13 21 - 12 11 11 0.00165 0.00394 0.01572 0.00610 0.00102 0.00043
robust_kalman_filter_N_50_i_18 1550 11 18 - 11 10 10 0.00144 0.00345 0.01480 0.00610 0.00102 0.00040
robust_kalman_filter_N_50_i_19 1550 13 19 - 13 12 12 0.00170 0.00380 0.01551 0.00675 0.00116 0.00053
robust_kalman_filter_N_75_i_0 2325 13 18 - 12 11 11 0.00250 0.00548 0.02104 0.00899 0.00153 0.00067
robust_kalman_filter_N_75_i_1 2325 12 19 - 11 10 10 0.00229 0.00552 0.02102 0.00838 0.00147 0.00062
robust_kalman_filter_N_75_i_2 2325 11 21 - 11 10 10 0.00221 0.00629 0.02039 0.00836 0.00145 0.00062
robust_kalman_filter_N_75_i_3 2325 13 21 - 14 11 11 0.00247 0.00600 0.02210 0.01016 0.00161 0.00067
robust_kalman_filter_N_75_i_4 2325 13 22 - 14 12 12 0.00251 0.00646 0.02187 0.01024 0.00163 0.00073
robust_kalman_filter_N_75_i_5 2325 12 23 - 13 11 11 0.00232 0.00661 0.02040 0.00925 0.00154 0.00068
robust_kalman_filter_N_75_i_6 2325 12 21 - 12 10 10 0.00230 0.00634 0.02142 0.00912 0.00143 0.00063
robust_kalman_filter_N_75_i_7 2325 12 24 - 12 10 10 0.00230 0.00727 0.02007 0.00882 0.00146 0.00069
robust_kalman_filter_N_75_i_8 2325 12 21 - 12 10 10 0.00229 0.00603 0.01992 0.00915 0.00142 0.00062
robust_kalman_filter_N_75_i_9 2325 12 20 - 11 10 10 0.00232 0.00575 0.02182 0.00839 0.00143 0.00063
robust_kalman_filter_N_75_i_10 2325 14 23 - 14 12 12 0.00264 0.00648 0.02124 0.01016 0.00168 0.00072
robust_kalman_filter_N_75_i_11 2325 13 21 - 12 11 11 0.00246 0.00622 0.01754 0.00891 0.00154 0.00067
robust_kalman_filter_N_75_i_12 2325 12 22 - 12 11 11 0.00234 0.00624 0.02128 0.00904 0.00154 0.00069
robust_kalman_filter_N_75_i_13 2325 13 21 - 12 11 11 0.00245 0.00611 0.02185 0.00887 0.00154 0.00068
robust_kalman_filter_N_75_i_14 2325 13 19 - 12 11 11 0.00244 0.00558 0.02053 0.00882 0.00160 0.00065
robust_kalman_filter_N_75_i_15 2325 14 22 - 15 12 12 0.00267 0.00634 0.02140 0.01082 0.00165 0.00075
robust_kalman_filter_N_75_i_16 2325 13 22 - 12 11 11 0.00246 0.00630 0.01927 0.00899 0.00154 0.00069
robust_kalman_filter_N_75_i_17 2325 12 21 - 12 10 10 0.00231 0.00611 0.02065 0.00906 0.00142 0.00062
robust_kalman_filter_N_75_i_18 2325 13 20 - 14 11 11 0.00250 0.00578 0.02298 0.01035 0.00154 0.00069
robust_kalman_filter_N_75_i_19 2325 12 20 - 13 10 10 0.00229 0.00568 0.02107 0.00992 0.00144 0.00062
robust_kalman_filter_N_125_i_0 3875 13 22 - 14 11 11 0.00418 0.01056 0.03251 0.01433 0.00257 0.00117
robust_kalman_filter_N_125_i_1 3875 14 23 - 13 12 12 0.00447 0.01152 0.03139 0.01300 0.00282 0.00140
robust_kalman_filter_N_125_i_2 3875 13 20 - 12 11 11 0.00420 0.01122 0.03640 0.01222 0.00300 0.00117
robust_kalman_filter_N_125_i_3 3875 13 21 - 13 10 10 0.00488 0.01034 0.03417 0.01317 0.00240 0.00120
robust_kalman_filter_N_125_i_4 3875 13 20 - 12 10 10 0.00416 0.00991 0.03691 0.01212 0.00247 0.00106
robust_kalman_filter_N_125_i_5 3875 13 26 - 14 12 12 0.00418 0.01285 0.03163 0.01427 0.00275 0.00125
robust_kalman_filter_N_125_i_6 3875 14 21 - 13 12 12 0.00443 0.01017 0.03108 0.01283 0.00276 0.00126
robust_kalman_filter_N_125_i_7 3875 13 21 - 14 12 12 0.00416 0.01198 0.03990 0.01367 0.00319 0.00141
robust_kalman_filter_N_125_i_8 3875 11 23 - 12 10 10 0.00421 0.01136 0.03093 0.01260 0.00252 0.00119
robust_kalman_filter_N_125_i_9 3875 12 23 - 12 10 10 0.00454 0.01325 0.04026 0.01384 0.00275 0.00120
robust_kalman_filter_N_125_i_10 3875 15 24 - 15 12 12 0.00554 0.01366 0.04077 0.01713 0.00317 0.00142
robust_kalman_filter_N_125_i_11 3875 13 24 - 13 11 11 0.00496 0.01335 0.04015 0.01551 0.00300 0.00130
robust_kalman_filter_N_125_i_12 3875 13 23 - 15 11 11 0.00493 0.01106 0.03885 0.01553 0.00258 0.00115
robust_kalman_filter_N_125_i_13 3875 11 21 - 11 10 10 0.00365 0.01220 0.03033 0.01155 0.00277 0.00105
robust_kalman_filter_N_125_i_14 3875 14 23 - 13 11 11 0.00519 0.01256 0.03178 0.01334 0.00298 0.00129
robust_kalman_filter_N_125_i_15 3875 13 21 - 13 11 11 0.00485 0.01181 0.03807 0.01472 0.00300 0.00116
robust_kalman_filter_N_125_i_16 3875 12 21 - 13 11 11 0.00459 0.01058 0.03455 0.01461 0.00260 0.00129
robust_kalman_filter_N_125_i_17 3875 13 23 - 13 12 12 0.00419 0.01117 0.03539 0.01349 0.00280 0.00124
robust_kalman_filter_N_125_i_18 3875 13 21 - 12 11 11 0.00416 0.01051 0.03393 0.01216 0.00258 0.00119
robust_kalman_filter_N_125_i_19 3875 14 21 - 13 12 12 0.00440 0.01028 0.03846 0.01306 0.00281 0.00122
robust_kalman_filter_N_175_i_0 5425 13 23 - 14 11 11 0.00576 0.01609 0.04946 0.02003 0.00372 0.00182
robust_kalman_filter_N_175_i_1 5425 13 24 - 14 11 11 0.00586 0.01695 0.04905 0.01895 0.00369 0.00167
robust_kalman_filter_N_175_i_2 5425 13 24 - 14 11 11 0.00581 0.01677 0.04803 0.01918 0.00385 0.00162
robust_kalman_filter_N_175_i_3 5425 13 22 - 13 11 11 0.00583 0.01531 0.04494 0.01763 0.00383 0.00159
robust_kalman_filter_N_175_i_4 5425 13 22 - 14 11 11 0.00675 0.01538 0.04717 0.01978 0.00374 0.00166
robust_kalman_filter_N_175_i_5 5425 14 26 - 14 11 11 0.00622 0.01812 0.04802 0.01980 0.00373 0.00162
robust_kalman_filter_N_175_i_6 5425 15 22 - 14 12 12 0.00657 0.01523 0.04868 0.01976 0.00405 0.00197
robust_kalman_filter_N_175_i_7 5425 14 25 - 14 11 11 0.00651 0.01728 0.04697 0.01966 0.00377 0.00163
robust_kalman_filter_N_175_i_8 5425 14 24 - 15 11 11 0.00620 0.01666 0.04773 0.02065 0.00381 0.00183
robust_kalman_filter_N_175_i_9 5425 14 24 - 15 12 12 0.00727 0.01647 0.04878 0.02087 0.00401 0.00198
robust_kalman_filter_N_175_i_10 5425 13 24 - 14 11 11 0.00584 0.01975 0.04389 0.01975 0.00431 0.00169
robust_kalman_filter_N_175_i_11 5425 13 24 - 15 12 12 0.00679 0.01684 0.04417 0.02087 0.00403 0.00175
robust_kalman_filter_N_175_i_12 5425 14 23 - 16 12 12 0.00621 0.01603 0.04882 0.02320 0.00398 0.00170
robust_kalman_filter_N_175_i_13 5425 12 26 - 14 10 10 0.00543 0.02103 0.04248 0.01993 0.00403 0.00168
robust_kalman_filter_N_175_i_14 5425 13 22 - 13 10 10 0.00674 0.01584 0.05128 0.01833 0.00350 0.00145
robust_kalman_filter_N_175_i_15 5425 14 26 - 15 13 13 0.00618 0.01838 0.04897 0.02098 0.00427 0.00187
robust_kalman_filter_N_175_i_16 5425 12 22 - 13 11 11 0.00553 0.01566 0.04850 0.01898 0.00372 0.00162
robust_kalman_filter_N_175_i_17 5425 14 25 - 15 11 11 0.00623 0.01752 0.04675 0.02127 0.00372 0.00160
robust_kalman_filter_N_175_i_18 5425 13 25 - 14 12 12 0.00594 0.01770 0.04420 0.01985 0.00397 0.00170
robust_kalman_filter_N_175_i_19 5425 13 24 - 15 11 11 0.00584 0.01683 0.04848 0.02111 0.00374 0.00161
robust_kalman_filter_N_225_i_0 6975 13 24 - 14 11 - 0.00757 0.02219 0.05807 0.02588 0.00496 -
robust_kalman_filter_N_225_i_1 6975 13 24 - 13 11 - 0.00748 0.02199 0.05756 0.02292 0.00512 -
robust_kalman_filter_N_225_i_2 6975 13 26 - 15 12 - 0.00754 0.02367 0.06344 0.02702 0.00524 -
robust_kalman_filter_N_225_i_3 6975 12 23 - 13 10 - 0.00708 0.02084 0.06611 0.02305 0.00457 -
robust_kalman_filter_N_225_i_4 6975 13 25 - 14 11 - 0.00757 0.02393 0.06104 0.02447 0.00496 -
robust_kalman_filter_N_225_i_5 6975 13 25 - 13 11 - 0.00874 0.02696 0.05785 0.02351 0.00492 -
robust_kalman_filter_N_225_i_6 6975 13 26 - 15 11 - 0.00880 0.02462 0.06758 0.02683 0.00490 -
robust_kalman_filter_N_225_i_7 6975 13 24 - 16 11 - 0.00755 0.02537 0.05830 0.02795 0.00567 -
robust_kalman_filter_N_225_i_8 6975 13 24 - 13 11 - 0.00877 0.02552 0.06356 0.02308 0.00569 -
robust_kalman_filter_N_225_i_9 6975 12 25 - 14 11 - 0.00816 0.02334 0.05694 0.02564 0.00492 -
robust_kalman_filter_N_225_i_10 6975 14 26 - 15 12 - 0.00796 0.02748 0.06755 0.02696 0.00605 -
robust_kalman_filter_N_225_i_11 6975 14 25 - 15 12 - 0.00928 0.02284 0.06145 0.02675 0.00528 -
robust_kalman_filter_N_225_i_12 6975 12 25 - 12 10 - 0.00697 0.02360 0.06125 0.02157 0.00459 -
robust_kalman_filter_N_225_i_13 6975 13 24 - 14 11 - 0.00759 0.02204 0.08955 0.02526 0.00563 -
robust_kalman_filter_N_225_i_14 6975 13 24 - 15 11 - 0.00754 0.02213 0.05645 0.02666 0.00491 -
robust_kalman_filter_N_225_i_15 6975 14 24 - 13 10 - 0.00797 0.02213 0.06271 0.02311 0.00459 -
robust_kalman_filter_N_225_i_16 6975 13 24 - 13 11 - 0.00750 0.02191 0.06950 0.02277 0.00486 -
robust_kalman_filter_N_225_i_17 6975 14 23 - 13 11 - 0.00794 0.02353 0.06576 0.02390 0.00566 -
robust_kalman_filter_N_225_i_18 6975 13 25 - 15 11 - 0.00877 0.02643 0.06683 0.03059 0.00569 -
robust_kalman_filter_N_225_i_19 6975 13 25 - 15 12 - 0.00878 0.02639 0.06711 0.03064 0.00606 -
robust_kalman_filter_N_300_i_0 9300 14 23 - 15 11 - 0.01230 0.02920 0.07393 0.03522 0.00675 -
robust_kalman_filter_N_300_i_1 9300 13 25 - 16 11 - 0.00996 0.03078 0.07924 0.03700 0.00673 -
robust_kalman_filter_N_300_i_2 9300 14 25 - 13 10 - 0.01053 0.03125 0.08683 0.03018 0.00626 -
robust_kalman_filter_N_300_i_3 9300 13 25 - 14 11 - 0.01003 0.03218 0.08651 0.03377 0.00679 -
robust_kalman_filter_N_300_i_4 9300 13 23 - 13 11 - 0.01006 0.02924 0.08816 0.03026 0.00672 -
robust_kalman_filter_N_300_i_5 9300 13 25 - 15 11 - 0.01000 0.03138 0.08341 0.03518 0.00674 -
robust_kalman_filter_N_300_i_6 9300 13 24 - 14 11 - 0.00996 0.03027 0.08081 0.03316 0.00685 -
robust_kalman_filter_N_300_i_7 9300 14 25 - 16 11 - 0.01068 0.03071 0.08734 0.03687 0.00688 -
robust_kalman_filter_N_300_i_8 9300 13 25 - 15 10 - 0.00999 0.03179 0.08249 0.03536 0.00639 -
robust_kalman_filter_N_300_i_9 9300 14 26 - 13 11 - 0.01063 0.03282 0.08182 0.03022 0.00677 -
robust_kalman_filter_N_300_i_10 9300 13 24 - 13 11 - 0.01002 0.03008 0.08660 0.03015 0.00678 -
robust_kalman_filter_N_300_i_11 9300 13 25 - 16 11 - 0.01003 0.03607 0.08646 0.03671 0.00776 -
robust_kalman_filter_N_300_i_12 9300 14 26 - 15 11 - 0.01245 0.03253 0.07619 0.04085 0.00691 -
robust_kalman_filter_N_300_i_13 9300 13 25 - 15 11 - 0.01001 0.03161 0.08319 0.03523 0.00678 -
robust_kalman_filter_N_300_i_14 9300 14 25 - 15 11 - 0.01066 0.03163 0.07509 0.03532 0.00786 -
robust_kalman_filter_N_300_i_15 9300 13 27 - 15 11 - 0.00997 0.03292 0.08569 0.03499 0.00687 -
robust_kalman_filter_N_300_i_16 9300 16 26 - 16 12 - 0.01194 0.03260 0.08105 0.03728 0.00725 -
robust_kalman_filter_N_300_i_17 9300 13 23 - 13 10 - 0.00987 0.02901 0.09474 0.03012 0.00634 -
robust_kalman_filter_N_300_i_18 9300 13 24 - 13 10 - 0.00986 0.03074 0.08284 0.03015 0.00637 -
robust_kalman_filter_N_300_i_19 9300 13 25 - 19 11 - 0.01003 0.03526 0.08672 0.05072 0.00690 -
robust_kalman_filter_N_375_i_0 11625 13 26 - 13 11 - 0.01465 0.04878 0.10385 0.03771 0.01025 -
robust_kalman_filter_N_375_i_1 11625 13 26 - 16 10 - 0.01451 0.04146 0.10711 0.04625 0.00820 -
robust_kalman_filter_N_375_i_2 11625 12 26 - 14 11 - 0.01180 0.04283 0.10340 0.04159 0.00882 -
robust_kalman_filter_N_375_i_3 11625 13 26 - 14 11 - 0.01253 0.04203 0.09593 0.03980 0.00894 -
robust_kalman_filter_N_375_i_4 11625 13 24 - 16 11 - 0.01258 0.03876 0.09791 0.04590 0.00876 -
robust_kalman_filter_N_375_i_5 11625 12 23 - 13 10 - 0.01174 0.03728 0.11823 0.03755 0.00820 -
robust_kalman_filter_N_375_i_6 11625 14 25 - 16 12 - 0.01331 0.04077 0.13839 0.05476 0.00940 -
robust_kalman_filter_N_375_i_7 11625 14 25 - 15 11 - 0.01339 0.04071 0.11648 0.04368 0.00874 -
robust_kalman_filter_N_375_i_8 11625 13 26 - 16 11 - 0.01259 0.04182 0.10793 0.04589 0.01014 -
robust_kalman_filter_N_375_i_9 11625 14 27 - 15 11 - 0.01346 0.04374 0.13064 0.04391 0.00891 -
robust_kalman_filter_N_375_i_10 11625 14 25 - 14 11 - 0.01348 0.04713 0.11812 0.04127 0.01033 -
robust_kalman_filter_N_375_i_11 11625 12 26 - 15 10 - 0.01367 0.04176 0.10039 0.04392 0.00828 -
robust_kalman_filter_N_375_i_12 11625 13 24 - 16 11 - 0.01250 0.04469 0.11300 0.04598 0.01010 -
robust_kalman_filter_N_375_i_13 11625 13 27 - 14 11 - 0.01467 0.04935 0.11771 0.04580 0.01018 -
robust_kalman_filter_N_375_i_14 11625 14 26 - 15 11 - 0.01557 0.04223 0.13709 0.04383 0.00888 -
robust_kalman_filter_N_375_i_15 11625 14 26 - 15 11 - 0.01346 0.04199 0.10943 0.04412 0.00887 -
robust_kalman_filter_N_375_i_16 11625 13 27 - 16 12 - 0.01264 0.04334 0.10027 0.04589 0.01077 -
robust_kalman_filter_N_375_i_17 11625 13 26 - 16 11 - 0.01258 0.04814 0.10561 0.04790 0.01013 -
robust_kalman_filter_N_375_i_18 11625 13 24 - 15 11 - 0.01459 0.03878 0.11506 0.05058 0.01005 -
robust_kalman_filter_N_375_i_19 11625 14 25 - 16 11 - 0.01344 0.04601 0.11903 0.04579 0.01016 -
robust_kalman_filter_N_450_i_0 13950 13 26 - 16 11 - 0.01755 0.05943 0.13787 0.06379 0.01251 -
robust_kalman_filter_N_450_i_1 13950 13 27 - 15 11 - 0.01761 0.06068 0.13805 0.06043 0.01265 -
robust_kalman_filter_N_450_i_2 13950 13 26 - 15 11 - 0.01755 0.05289 0.15038 0.06417 0.01084 -
robust_kalman_filter_N_450_i_3 13950 13 28 - 15 10 - 0.01499 0.06268 0.13392 0.05238 0.01051 -
robust_kalman_filter_N_450_i_4 13950 13 31 - 15 11 - 0.01753 0.05969 0.12347 0.06045 0.01090 -
robust_kalman_filter_N_450_i_5 13950 14 26 - 15 11 - 0.01596 0.05177 0.13747 0.05258 0.01252 -
robust_kalman_filter_N_450_i_6 13950 14 29 - 16 11 - 0.01600 0.05609 0.13698 0.05515 0.01088 -
robust_kalman_filter_N_450_i_7 13950 13 27 - 16 11 - 0.01501 0.05145 0.14600 0.05486 0.01091 -
robust_kalman_filter_N_450_i_8 13950 14 28 - 16 12 - 0.01609 0.05432 0.13131 0.05507 0.01152 -
robust_kalman_filter_N_450_i_9 13950 13 29 - 14 10 - 0.01497 0.05663 0.14099 0.04749 0.01039 -
robust_kalman_filter_N_450_i_10 13950 14 27 - 15 10 - 0.01595 0.06025 0.14367 0.05227 0.01175 -
robust_kalman_filter_N_450_i_11 13950 12 25 - 15 10 - 0.01635 0.05020 0.13812 0.06005 0.01029 -
robust_kalman_filter_N_450_i_12 13950 15 26 - 16 12 - 0.01697 0.05084 0.13032 0.05577 0.01160 -
robust_kalman_filter_N_450_i_13 13950 13 26 - 16 11 - 0.01504 0.05123 0.12504 0.05499 0.01091 -
robust_kalman_filter_N_450_i_14 13950 13 26 - 14 10 - 0.01502 0.05176 0.13912 0.04784 0.01034 -
robust_kalman_filter_N_450_i_15 13950 14 28 - 16 10 - 0.01606 0.05606 0.12310 0.05692 0.01182 -
robust_kalman_filter_N_450_i_16 13950 14 29 - 16 11 - 0.01590 0.05792 0.13368 0.05491 0.01091 -
robust_kalman_filter_N_450_i_17 13950 13 27 - 14 11 - 0.01747 0.05376 0.15784 0.05465 0.01091 -
robust_kalman_filter_N_450_i_18 13950 14 26 - 15 13 - 0.01610 0.05996 0.13933 0.05233 0.01228 -
robust_kalman_filter_N_450_i_19 13950 13 25 - 17 11 - 0.01744 0.04958 0.14799 0.06067 0.01084 -
robust_kalman_filter_N_500_i_0 15500 14 25 - 14 11 - 0.01781 0.05705 0.18759 0.05262 0.01240 -
robust_kalman_filter_N_500_i_1 15500 13 25 - 16 11 - 0.01664 0.05575 0.14857 0.06107 0.01249 -
robust_kalman_filter_N_500_i_2 15500 13 25 - 15 10 - 0.01664 0.05711 0.15082 0.05883 0.01156 -
robust_kalman_filter_N_500_i_3 15500 13 26 - 14 11 - 0.01672 0.06066 0.16043 0.05296 0.01227 -
robust_kalman_filter_N_500_i_4 15500 13 29 - 16 11 - 0.01941 0.07490 0.15324 0.06134 0.01428 -
robust_kalman_filter_N_500_i_5 15500 13 25 - 15 10 - 0.01652 0.06453 0.14032 0.05833 0.01341 -
robust_kalman_filter_N_500_i_6 15500 14 29 - 16 11 - 0.02072 0.06508 0.17977 0.06136 0.01437 -
robust_kalman_filter_N_500_i_7 15500 13 25 - 14 11 - 0.01673 0.06524 0.16417 0.05285 0.01429 -
robust_kalman_filter_N_500_i_8 15500 14 27 - 16 11 - 0.02074 0.07022 0.17136 0.06185 0.01437 -
robust_kalman_filter_N_500_i_9 15500 13 30 - 16 10 - 0.01941 0.06671 0.17077 0.06119 0.01201 -
robust_kalman_filter_N_500_i_10 15500 14 26 - 15 11 - 0.01779 0.05815 0.14614 0.05873 0.01234 -
robust_kalman_filter_N_500_i_11 15500 13 28 - 15 11 - 0.01657 0.06165 0.14662 0.05840 0.01236 -
robust_kalman_filter_N_500_i_12 15500 13 26 - 16 10 - 0.01666 0.05863 0.15932 0.06092 0.01163 -
robust_kalman_filter_N_500_i_13 15500 12 25 - 15 10 - 0.01554 0.05651 0.14555 0.05786 0.01232 -
robust_kalman_filter_N_500_i_14 15500 13 26 - 16 11 - 0.01957 0.05692 0.16304 0.06113 0.01237 -
robust_kalman_filter_N_500_i_15 15500 14 27 - 16 11 - 0.01773 0.05933 0.19075 0.06103 0.01254 -
robust_kalman_filter_N_500_i_16 15500 13 26 - 16 11 - 0.01670 0.05790 0.16259 0.06289 0.01237 -
robust_kalman_filter_N_500_i_17 15500 13 27 - 16 11 - 0.01671 0.06048 0.15740 0.06095 0.01250 -
robust_kalman_filter_N_500_i_18 15500 13 26 - 15 11 - 0.01664 0.05763 0.16208 0.05853 0.01254 -
robust_kalman_filter_N_500_i_19 15500 13 28 - 16 11 - 0.01677 0.07028 0.16353 0.06141 0.01430 -
Table 4: Iterations and solver runtimes for lossless convexification problems
Iterations Solver Runtime (s)
Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
lcvx_N_15_i_0 644 8 7 - 7 9 9 0.00084 0.00067 0.00505 0.00198 0.00055 0.00026
lcvx_N_15_i_1 644 8 7 - 7 9 9 0.00082 0.00065 0.00515 0.00193 0.00056 0.00024
lcvx_N_15_i_2 644 10 7 - 7 10 10 0.00097 0.00068 0.00535 0.00194 0.00060 0.00028
lcvx_N_15_i_3 644 8 7 - 8 10 10 0.00085 0.00064 0.00505 0.00240 0.00060 0.00028
lcvx_N_15_i_4 644 9 7 - 7 9 9 0.00090 0.00059 0.00443 0.00195 0.00047 0.00023
lcvx_N_15_i_5 644 10 7 - 7 9 9 0.00082 0.00059 0.00432 0.00162 0.00051 0.00023
lcvx_N_15_i_6 644 7 7 - 7 10 10 0.00065 0.00058 0.00440 0.00162 0.00051 0.00026
lcvx_N_15_i_7 644 9 7 - 7 10 10 0.00076 0.00058 0.00454 0.00161 0.00051 0.00026
lcvx_N_15_i_8 644 8 7 - 7 9 9 0.00071 0.00058 0.00430 0.00168 0.00047 0.00023
lcvx_N_15_i_9 644 8 7 - 8 10 10 0.00072 0.00056 0.00457 0.00215 0.00051 0.00025
lcvx_N_15_i_10 644 8 7 - 7 9 9 0.00071 0.00059 0.00456 0.00173 0.00048 0.00023
lcvx_N_15_i_11 644 7 7 - 7 9 9 0.00065 0.00058 0.00430 0.00163 0.00047 0.00023
lcvx_N_15_i_12 644 9 7 - 8 10 10 0.00083 0.00055 0.00444 0.00195 0.00051 0.00025
lcvx_N_15_i_13 644 7 7 - 7 9 9 0.00064 0.00058 0.00446 0.00163 0.00047 0.00023
lcvx_N_15_i_14 644 8 7 - 7 10 10 0.00071 0.00056 0.00443 0.00174 0.00053 0.00026
lcvx_N_15_i_15 644 8 7 - 7 10 10 0.00077 0.00056 0.00445 0.00163 0.00051 0.00025
lcvx_N_15_i_16 644 7 7 - 7 9 9 0.00065 0.00058 0.00448 0.00160 0.00047 0.00023
lcvx_N_15_i_17 644 8 7 - 7 10 10 0.00075 0.00059 0.00459 0.00163 0.00052 0.00025
lcvx_N_15_i_18 644 8 7 - 7 9 9 0.00071 0.00059 0.00444 0.00161 0.00048 0.00023
lcvx_N_15_i_19 644 9 7 - 8 10 10 0.00078 0.00059 0.00461 0.00197 0.00051 0.00025
lcvx_N_50_i_0 2114 9 8 - 10 10 10 0.00278 0.00205 0.01024 0.00657 0.00166 0.00093
lcvx_N_50_i_1 2114 9 7 - 11 10 10 0.00261 0.00185 0.01066 0.00742 0.00167 0.00094
lcvx_N_50_i_2 2114 9 7 - 10 10 10 0.00260 0.00212 0.01193 0.00662 0.00192 0.00101
lcvx_N_50_i_3 2114 10 9 - 13 10 10 0.00339 0.00269 0.01270 0.00951 0.00194 0.00100
lcvx_N_50_i_4 2114 8 7 - 10 10 10 0.00281 0.00209 0.01227 0.00737 0.00191 0.00103
lcvx_N_50_i_5 2114 8 7 - 11 10 10 0.00244 0.00185 0.01040 0.00765 0.00166 0.00092
lcvx_N_50_i_6 2114 9 7 - 11 10 10 0.00259 0.00182 0.01134 0.00764 0.00165 0.00093
lcvx_N_50_i_7 2114 8 8 - 10 10 10 0.00239 0.00218 0.01072 0.00672 0.00167 0.00099
lcvx_N_50_i_8 2114 10 8 - 10 10 10 0.00274 0.00213 0.01097 0.00670 0.00165 0.00090
lcvx_N_50_i_9 2114 10 8 - 10 10 10 0.00277 0.00214 0.01102 0.00672 0.00167 0.00087
lcvx_N_50_i_10 2114 10 8 - 11 10 10 0.00278 0.00212 0.01095 0.00784 0.00168 0.00089
lcvx_N_50_i_11 2114 10 8 - 11 10 10 0.00291 0.00203 0.01069 0.00713 0.00166 0.00092
lcvx_N_50_i_12 2114 9 7 - 10 10 10 0.00257 0.00183 0.01087 0.00660 0.00166 0.00092
lcvx_N_50_i_13 2114 9 8 - 11 10 10 0.00272 0.00204 0.01022 0.00728 0.00167 0.00100
lcvx_N_50_i_14 2114 10 7 - 10 10 10 0.00289 0.00210 0.01576 0.00823 0.00189 0.00102
lcvx_N_50_i_15 2114 9 8 - 11 10 10 0.00300 0.00207 0.01070 0.00796 0.00167 0.00092
lcvx_N_50_i_16 2114 9 7 - 10 10 10 0.00255 0.00176 0.01094 0.00739 0.00167 0.00091
lcvx_N_50_i_17 2114 10 9 - 11 10 10 0.00290 0.00269 0.01292 0.00770 0.00192 0.00101
lcvx_N_50_i_18 2114 9 8 - 11 10 10 0.00310 0.00238 0.01204 0.00853 0.00192 0.00102
lcvx_N_50_i_19 2114 12 9 - 12 10 10 0.00382 0.00267 0.01244 0.00827 0.00191 0.00101
lcvx_N_75_i_0 3164 10 8 - 10 10 10 0.00483 0.00367 0.01726 0.01255 0.00281 0.00157
lcvx_N_75_i_1 3164 10 9 - 12 10 10 0.00503 0.00395 0.01880 0.01296 0.00284 0.00157
lcvx_N_75_i_2 3164 10 8 - 11 10 10 0.00520 0.00350 0.01749 0.01170 0.00284 0.00160
lcvx_N_75_i_3 3164 8 8 - 10 10 10 0.00420 0.00326 0.01437 0.01028 0.00242 0.00139
lcvx_N_75_i_4 3164 10 9 - 12 10 10 0.00426 0.00398 0.01629 0.01286 0.00282 0.00159
lcvx_N_75_i_5 3164 9 8 - 10 10 10 0.00456 0.00373 0.01735 0.01230 0.00288 0.00159
lcvx_N_75_i_6 3164 8 8 - 10 10 10 0.00421 0.00314 0.01460 0.01023 0.00243 0.00137
lcvx_N_75_i_7 3164 11 9 - 11 10 10 0.00464 0.00343 0.01502 0.01178 0.00242 0.00141
lcvx_N_75_i_8 3164 10 8 - 11 10 10 0.00437 0.00302 0.01486 0.01073 0.00244 0.00142
lcvx_N_75_i_9 3164 11 9 - 11 10 10 0.00460 0.00341 0.01494 0.01078 0.00244 0.00139
lcvx_N_75_i_10 3164 10 8 - 10 10 10 0.00404 0.00311 0.01478 0.01152 0.00243 0.00141
lcvx_N_75_i_11 3164 8 8 - 10 10 10 0.00363 0.00314 0.01447 0.01010 0.00244 0.00141
lcvx_N_75_i_12 3164 11 9 - 11 10 10 0.00470 0.00344 0.01473 0.01054 0.00243 0.00142
lcvx_N_75_i_13 3164 9 8 - 11 10 10 0.00410 0.00300 0.01479 0.01059 0.00244 0.00158
lcvx_N_75_i_14 3164 10 9 - 11 10 10 0.00442 0.00348 0.01499 0.01232 0.00246 0.00144
lcvx_N_75_i_15 3164 10 9 - 11 10 10 0.00438 0.00346 0.01438 0.01071 0.00243 0.00140
lcvx_N_75_i_16 3164 10 8 - 11 10 10 0.00440 0.00300 0.01516 0.01072 0.00247 0.00145
lcvx_N_75_i_17 3164 11 9 - 11 10 10 0.00467 0.00345 0.01504 0.01073 0.00247 0.00142
lcvx_N_75_i_18 3164 10 8 - 11 10 10 0.00407 0.00326 0.01829 0.01182 0.00244 0.00140
lcvx_N_75_i_19 3164 9 8 - 10 10 10 0.00383 0.00301 0.01468 0.01010 0.00243 0.00145
lcvx_N_100_i_0 4214 9 8 - 10 10 10 0.00531 0.00413 0.01846 0.01415 0.00332 0.00202
lcvx_N_100_i_1 4214 10 9 - 11 10 10 0.00587 0.00528 0.02145 0.01364 0.00380 0.00219
lcvx_N_100_i_2 4214 12 9 - 12 11 11 0.00772 0.00505 0.02538 0.01594 0.00419 0.00245
lcvx_N_100_i_3 4214 10 9 - 11 10 10 0.00713 0.00530 0.02058 0.01506 0.00382 0.00222
lcvx_N_100_i_4 4214 8 8 - 10 10 10 0.00522 0.00469 0.02155 0.01427 0.00382 0.00224
lcvx_N_100_i_5 4214 11 9 - 11 10 10 0.00756 0.00530 0.02170 0.01503 0.00379 0.00221
lcvx_N_100_i_6 4214 10 9 - 11 10 10 0.00692 0.00461 0.01840 0.01704 0.00329 0.00188
lcvx_N_100_i_7 4214 11 9 - 12 11 11 0.00616 0.00435 0.01749 0.01617 0.00362 0.00216
lcvx_N_100_i_8 4214 9 8 - 10 10 10 0.00523 0.00412 0.01727 0.01262 0.00332 0.00196
lcvx_N_100_i_9 4214 9 8 - 10 10 10 0.00474 0.00404 0.01726 0.01276 0.00331 0.00191
lcvx_N_100_i_10 4214 10 8 - 11 10 10 0.00560 0.00395 0.01779 0.01347 0.00331 0.00189
lcvx_N_100_i_11 4214 10 8 - 11 10 10 0.00574 0.00388 0.01737 0.01600 0.00331 0.00190
lcvx_N_100_i_12 4214 10 9 - 12 10 10 0.00592 0.00456 0.01834 0.01565 0.00331 0.00194
lcvx_N_100_i_13 4214 10 9 - 12 11 11 0.00578 0.00507 0.01823 0.01437 0.00418 0.00242
lcvx_N_100_i_14 4214 11 9 - 12 11 11 0.00722 0.00435 0.02542 0.01744 0.00388 0.00205
lcvx_N_100_i_15 4214 10 9 - 11 10 10 0.00565 0.00536 0.02161 0.01337 0.00387 0.00223
lcvx_N_100_i_16 4214 10 8 - 10 10 10 0.00666 0.00479 0.02177 0.01598 0.00380 0.00220
lcvx_N_100_i_17 4214 12 9 - 12 10 10 0.00783 0.00533 0.02160 0.01689 0.00381 0.00219
lcvx_N_100_i_18 4214 10 8 - 10 10 10 0.00668 0.00410 0.01870 0.01297 0.00334 0.00221
lcvx_N_100_i_19 4214 11 9 - 11 10 10 0.00760 0.00529 0.02164 0.01657 0.00380 0.00222
lcvx_N_125_i_0 5264 11 9 - 12 11 11 0.00889 0.00639 0.03142 0.02272 0.00506 0.00297
lcvx_N_125_i_1 5264 11 9 - 11 10 10 0.00876 0.00562 0.03142 0.01875 0.00406 0.00235
lcvx_N_125_i_2 5264 10 9 - 12 10 10 0.00742 0.00549 0.02604 0.02137 0.00407 0.00233
lcvx_N_125_i_3 5264 11 9 - 11 10 10 0.00759 0.00566 0.02148 0.01894 0.00401 0.00237
lcvx_N_125_i_4 5264 9 8 - 10 10 10 0.00647 0.00494 0.02034 0.01619 0.00407 0.00232
lcvx_N_125_i_5 5264 11 9 - 12 10 10 0.00770 0.00540 0.02805 0.01804 0.00406 0.00233
lcvx_N_125_i_6 5264 11 9 - 11 10 10 0.00745 0.00566 0.02949 0.01920 0.00406 0.00234
lcvx_N_125_i_7 5264 11 8 - 11 10 10 0.00727 0.00513 0.02713 0.01875 0.00404 0.00275
lcvx_N_125_i_8 5264 9 8 - 10 10 10 0.00589 0.00582 0.02469 0.01816 0.00489 0.00276
lcvx_N_125_i_9 5264 10 9 - 11 10 10 0.00749 0.00572 0.02529 0.01747 0.00409 0.00272
lcvx_N_125_i_10 5264 9 8 - 11 10 10 0.00597 0.00513 0.01997 0.02011 0.00405 0.00234
lcvx_N_125_i_11 5264 8 8 - 11 10 10 0.00561 0.00499 0.02039 0.01770 0.00411 0.00237
lcvx_N_125_i_12 5264 9 8 - 11 10 10 0.00714 0.00491 0.02246 0.01729 0.00410 0.00233
lcvx_N_125_i_13 5264 11 8 - 10 10 10 0.00701 0.00595 0.03027 0.01666 0.00471 0.00275
lcvx_N_125_i_14 5264 9 8 - 10 10 10 0.00709 0.00590 0.03029 0.01800 0.00471 0.00278
lcvx_N_125_i_15 5264 11 9 - 12 10 10 0.00911 0.00544 0.02933 0.02052 0.00405 0.00234
lcvx_N_125_i_16 5264 11 9 - 11 10 10 0.00753 0.00575 0.02619 0.01939 0.00402 0.00245
lcvx_N_125_i_17 5264 11 9 - 12 11 11 0.00774 0.00546 0.02632 0.02054 0.00441 0.00256
lcvx_N_125_i_18 5264 11 9 - 12 10 10 0.00783 0.00549 0.02611 0.02030 0.00404 0.00274
lcvx_N_125_i_19 5264 11 9 - 11 10 10 0.00806 0.00575 0.02910 0.01742 0.00408 0.00232
lcvx_N_150_i_0 6314 10 8 - 11 10 - 0.00786 0.00612 0.02926 0.02194 $0$
lcvx_N_150_i_1 6314 11 9 - 12 11 - 0.00957 0.00654 0.03172 0.02258 0.00538 -
lcvx_N_150_i_2 6314 9 8 - 11 10 - 0.00803 0.00585 0.02330 0.02031 0.00490 -
lcvx_N_150_i_3 6314 10 9 - 12 10 - 0.00922 0.00777 0.02501 0.02377 0.00494 -
lcvx_N_150_i_4 6314 9 8 - 11 10 - 0.00948 0.00680 0.02824 0.02464 0.00571 -
lcvx_N_150_i_5 6314 11 10 - 12 11 - 0.00982 0.00724 0.02544 0.02362 0.00539 -
lcvx_N_150_i_6 6314 8 8 - 11 10 - 0.00673 0.00608 0.02354 0.02240 0.00495 -
lcvx_N_150_i_7 6314 9 8 - 11 10 - 0.00718 0.00586 0.02917 0.02252 0.00494 -
lcvx_N_150_i_8 6314 11 9 - 11 10 - 0.00915 0.00702 0.03529 0.02027 0.00497 -
lcvx_N_150_i_9 6314 12 11 - 12 11 - 0.01013 0.00809 0.03704 0.02204 0.00533 -
lcvx_N_150_i_10 6314 11 9 - 11 10 - 0.00926 0.00695 0.03020 0.01989 0.00495 -
lcvx_N_150_i_11 6314 10 8 - 11 10 - 0.00868 0.00591 0.03679 0.02231 0.00496 -
lcvx_N_150_i_12 6314 10 8 - 11 10 - 0.00866 0.00586 0.02538 0.02202 0.00493 -
lcvx_N_150_i_13 6314 12 9 - 13 11 - 0.01014 0.00660 0.02530 0.02384 0.00540 -
lcvx_N_150_i_14 6314 9 8 - 11 10 - 0.00776 0.00591 0.02464 0.02160 0.00496 -
lcvx_N_150_i_15 6314 10 8 - 11 10 - 0.00869 0.00580 0.03040 0.02198 0.00492 -
lcvx_N_150_i_16 6314 11 9 - 12 10 - 0.00948 0.00658 0.03178 0.02267 0.00492 -
lcvx_N_150_i_17 6314 9 8 - 11 10 - 0.00766 0.00680 0.03032 0.02280 0.00571 -
lcvx_N_150_i_18 6314 9 8 - 11 10 - 0.00852 0.00605 0.02975 0.02419 0.00494 -
lcvx_N_150_i_19 6314 11 9 - 11 10 - 0.00920 0.00709 0.03412 0.01958 0.00491 -
lcvx_N_200_i_0 8414 10 9 - 12 11 - 0.01143 0.00892 0.03671 0.02733 0.00717 -
lcvx_N_200_i_1 8414 10 9 - 12 11 - 0.01170 0.00886 0.03590 0.02720 0.00720 -
lcvx_N_200_i_2 8414 9 9 - 12 11 - 0.01059 0.00905 0.03362 0.02715 0.00720 -
lcvx_N_200_i_3 8414 12 9 - 12 11 - 0.01327 0.00889 0.04336 0.02782 0.00710 -
lcvx_N_200_i_4 8414 12 9 - 12 11 - 0.01386 0.00874 0.03456 0.03067 0.00713 -
lcvx_N_200_i_5 8414 11 9 - 11 10 - 0.01195 0.00924 0.04326 0.02603 0.00664 -
lcvx_N_200_i_6 8414 11 8 - 11 10 - 0.01121 0.00793 0.03630 0.02847 0.00658 -
lcvx_N_200_i_7 8414 11 9 - 12 11 - 0.01286 0.00881 0.05245 0.03026 0.00714 -
lcvx_N_200_i_8 8414 10 8 - 12 10 - 0.01117 0.00798 0.03576 0.03125 0.00659 -
lcvx_N_200_i_9 8414 9 8 - 11 10 - 0.00971 0.00796 0.03288 0.02907 0.00663 -
lcvx_N_200_i_10 8414 11 12 - 12 11 - 0.01259 0.01169 0.04709 0.02880 0.00715 -
lcvx_N_200_i_11 8414 9 8 - 13 10 - 0.01061 0.00797 0.03540 0.02946 0.00659 -
lcvx_N_200_i_12 8414 10 8 - 11 10 - 0.01031 0.00781 0.03495 0.02987 0.00659 -
lcvx_N_200_i_13 8414 10 8 - 11 10 - 0.01042 0.00799 0.03464 0.02826 0.00668 -
lcvx_N_200_i_14 8414 11 11 - 12 11 - 0.01269 0.01097 0.04928 0.03008 0.00718 -
lcvx_N_200_i_15 8414 11 9 - 12 10 - 0.01204 0.00929 0.04334 0.03131 0.00658 -
lcvx_N_200_i_16 8414 9 8 - 11 10 - 0.00971 0.00813 0.04864 0.02568 0.00668 -
lcvx_N_200_i_17 8414 12 9 - 12 11 - 0.01312 0.00888 0.04630 0.03020 0.00717 -
lcvx_N_200_i_18 8414 9 8 - 11 10 - 0.00966 0.00914 0.03348 0.02590 0.00764 -
lcvx_N_200_i_19 8414 12 10 - 12 11 - 0.01595 0.01134 0.04035 0.03376 0.00833 -
lcvx_N_250_i_0 10514 10 9 - 12 11 - 0.01613 0.01238 0.04357 0.03788 0.01035 -
lcvx_N_250_i_1 10514 11 10 - 12 11 - 0.01734 0.01417 0.04455 0.03854 0.01039 -
lcvx_N_250_i_2 10514 11 9 - 12 11 - 0.01491 0.01096 0.06412 0.03377 0.00894 -
lcvx_N_250_i_3 10514 9 8 - 11 10 - 0.01209 0.00976 0.04371 0.03196 0.00824 -
lcvx_N_250_i_4 10514 10 8 - 12 10 - 0.01297 0.00965 0.05525 0.03753 0.00829 -
lcvx_N_250_i_5 10514 11 8 - 11 10 - 0.01413 0.00983 0.05159 0.03588 0.00827 -
lcvx_N_250_i_6 10514 9 8 - 12 10 - 0.01215 0.00974 0.04114 0.03851 0.00829 -
lcvx_N_250_i_7 10514 11 8 - 11 10 - 0.01415 0.00984 0.07292 0.03271 0.00830 -
lcvx_N_250_i_8 10514 12 12 - 12 11 - 0.01583 0.01423 0.07551 0.03456 $0$
lcvx_N_250_i_9 10514 9 8 - 11 10 - 0.01214 0.01133 0.04417 0.03677 0.00958 -
lcvx_N_250_i_10 10514 12 9 - 11 10 - 0.01802 0.01307 0.06299 0.03559 0.00956 -
lcvx_N_250_i_11 10514 11 9 - 11 10 - 0.01673 0.01162 0.06861 0.03580 0.00821 -
lcvx_N_250_i_12 10514 12 9 - 13 11 - 0.01621 0.01095 0.06591 0.03763 0.00890 -
lcvx_N_250_i_13 10514 11 8 - 11 10 - 0.01414 0.01006 0.06107 0.03165 0.00960 -
lcvx_N_250_i_14 10514 10 9 - 12 11 - 0.01417 0.01097 0.05955 0.03464 0.00902 -
lcvx_N_250_i_15 10514 9 8 - 11 10 - 0.01203 0.00981 0.06903 0.03250 0.00830 -
lcvx_N_250_i_16 10514 10 8 - 11 10 - 0.01306 0.00984 0.05888 0.03269 0.00824 -
lcvx_N_250_i_17 10514 10 8 - 11 10 - 0.01316 0.01139 0.04364 0.03279 0.00822 -
lcvx_N_250_i_18 10514 13 9 - 12 11 - 0.02033 0.01271 0.06754 0.03774 0.01034 -
lcvx_N_250_i_19 10514 10 9 - 11 10 - 0.01540 0.01307 0.06816 0.03144 0.00956 -
lcvx_N_300_i_0 12614 11 11 - 13 11 - 0.01840 0.01611 0.06711 0.04499 0.01068 -
lcvx_N_300_i_1 12614 11 9 - 12 10 - 0.01711 0.01391 0.07451 0.04436 0.00989 -
lcvx_N_300_i_2 12614 10 11 - 13 11 - 0.01689 0.01585 0.05271 0.04656 0.01069 -
lcvx_N_300_i_3 12614 12 10 - 13 12 - 0.01890 0.01467 0.07362 0.04415 0.01154 -
lcvx_N_300_i_4 12614 10 8 - 12 10 - 0.01585 0.01191 0.07343 0.04025 0.00994 -
lcvx_N_300_i_5 12614 12 10 - 14 11 - 0.01901 0.01476 0.08295 0.04839 0.01077 -
lcvx_N_300_i_6 12614 10 8 - 13 10 - 0.01555 0.01262 0.05116 0.04782 0.01023 -
lcvx_N_300_i_7 12614 9 8 - 12 10 - 0.01461 0.01368 0.07382 0.04529 0.01144 -
lcvx_N_300_i_8 12614 10 9 - 13 11 - 0.01857 0.01335 0.05193 0.04822 0.01078 -
lcvx_N_300_i_9 12614 11 9 - 13 11 - 0.01695 0.01322 0.06888 0.04389 0.01076 -
lcvx_N_300_i_10 12614 13 10 - 13 11 - 0.02362 0.01457 0.08371 0.04780 0.01241 -
lcvx_N_300_i_11 12614 10 9 - 13 11 - 0.01585 0.01350 0.05985 0.04982 0.01250 -
lcvx_N_300_i_12 12614 11 9 - 13 11 - 0.01714 0.01542 0.07645 0.04790 0.01251 -
lcvx_N_300_i_13 12614 9 9 - 13 11 - 0.01711 0.01361 0.04904 0.04500 0.01081 -
lcvx_N_300_i_14 12614 10 9 - 12 10 - 0.01572 0.01389 0.08113 0.04145 0.00987 -
lcvx_N_300_i_15 12614 11 8 - 12 10 - 0.01689 0.01197 0.06695 0.04455 0.00989 -
lcvx_N_300_i_16 12614 9 8 - 12 10 - 0.01464 0.01368 0.05260 0.04061 0.01152 -
lcvx_N_300_i_17 12614 10 8 - 12 10 - 0.01862 0.01196 0.08553 0.04934 0.00985 -
lcvx_N_300_i_18 12614 9 8 - 12 10 - 0.01446 0.01191 0.04953 0.04100 0.00994 -
lcvx_N_300_i_19 12614 10 10 - 13 11 - 0.01679 0.01469 0.05266 0.04568 0.01084 -
lcvx_N_350_i_0 14714 9 9 - 13 11 - 0.02004 0.01545 0.07680 0.05598 0.01249 -
lcvx_N_350_i_1 14714 11 10 - 13 11 - 0.02061 0.01728 0.06032 0.05012 0.01260 -
lcvx_N_350_i_2 14714 11 9 - 12 11 - 0.01994 0.01560 0.06840 0.05095 0.01260 -
lcvx_N_350_i_3 14714 10 8 - 12 10 - 0.01839 0.01380 0.07929 0.05342 0.01161 -
lcvx_N_350_i_4 14714 9 8 - 12 10 - 0.01700 0.01389 0.05833 0.04700 0.01158 -
lcvx_N_350_i_5 14714 11 9 - 12 11 - 0.01983 0.01550 0.10310 0.05331 0.01261 -
lcvx_N_350_i_6 14714 10 10 - 12 11 - 0.01853 0.01738 0.06233 0.04658 0.01265 -
lcvx_N_350_i_7 14714 11 9 - 12 11 - 0.01990 0.01556 0.10463 0.05172 0.01256 -
lcvx_N_350_i_8 14714 10 8 - 12 10 - 0.01841 0.01379 0.05820 0.04862 0.01168 -
lcvx_N_350_i_9 14714 11 9 - 12 11 - 0.01967 0.01550 0.07686 0.04768 0.01260 -
lcvx_N_350_i_10 14714 9 8 - 12 11 - 0.01696 0.01403 0.06066 0.05296 0.01266 -
lcvx_N_350_i_11 14714 12 10 - 14 12 - 0.02212 0.01970 0.08369 0.05846 0.01574 -
lcvx_N_350_i_12 14714 12 10 - 14 12 - 0.02615 0.01714 0.07824 0.05940 0.01350 -
lcvx_N_350_i_13 14714 11 10 - 14 11 - 0.02099 0.01721 0.08116 0.05812 0.01254 -
lcvx_N_350_i_14 14714 10 11 - 13 11 - 0.01949 0.01833 0.07396 0.05615 0.01257 -
lcvx_N_350_i_15 14714 11 9 - 14 11 - 0.01968 0.01787 0.07639 0.06019 0.01456 -
lcvx_N_350_i_16 14714 11 9 - 12 11 - 0.02357 0.01550 0.07644 0.05193 0.01264 -
lcvx_N_350_i_17 14714 8 8 - 12 10 - 0.01551 0.01374 0.05710 0.05051 0.01343 -
lcvx_N_350_i_18 14714 9 11 - 13 11 - 0.01701 0.01860 0.06107 0.05173 0.01263 -
lcvx_N_350_i_19 14714 10 8 - 12 10 - 0.01837 0.01601 0.05973 0.05333 0.01342 -
Table 5: Iterations and solver runtimes for group lasso problems
Iterations Solver Runtime (s)
Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
group_lasso_N_1_i_0 761 4 11 - 5 6 6 0.00025 0.00109 0.01260 0.00126 0.00019 0.00006
group_lasso_N_1_i_1 761 4 11 - 7 6 6 0.00026 0.00108 0.01325 0.00153 0.00020 0.00006
group_lasso_N_1_i_2 761 4 10 - 7 6 6 0.00027 0.00097 0.01365 0.00152 0.00019 0.00006
group_lasso_N_1_i_3 761 4 12 - 7 6 6 0.00026 0.00119 0.01350 0.00156 0.00020 0.00007
group_lasso_N_1_i_4 761 4 11 - 7 6 6 0.00026 0.00110 0.01364 0.00151 0.00020 0.00007
group_lasso_N_1_i_5 761 4 11 - 7 6 6 0.00026 0.00112 0.01389 0.00151 0.00020 0.00006
group_lasso_N_1_i_6 761 4 11 - 6 6 6 0.00026 0.00107 0.01362 0.00139 0.00019 0.00006
group_lasso_N_1_i_7 761 5 10 - 7 6 6 0.00030 0.00097 0.01332 0.00150 0.00019 0.00006
group_lasso_N_1_i_8 761 4 12 - 7 6 6 0.00026 0.00114 0.01329 0.00159 0.00019 0.00006
group_lasso_N_1_i_9 761 5 10 - 6 6 6 0.00030 0.00109 0.01349 0.00148 0.00023 0.00007
group_lasso_N_1_i_10 761 4 9 - 7 6 6 0.00029 0.00101 0.01752 0.00167 0.00022 0.00007
group_lasso_N_1_i_11 761 4 11 - 5 6 6 0.00030 0.00123 0.01574 0.00139 0.00022 0.00007
group_lasso_N_1_i_12 761 5 9 - 6 6 6 0.00033 0.00100 0.01571 0.00153 0.00022 0.00007
group_lasso_N_1_i_13 761 4 11 - 7 6 6 0.00029 0.00122 0.01684 0.00166 0.00022 0.00007
group_lasso_N_1_i_14 761 4 12 - 6 6 6 0.00029 0.00129 0.01562 0.00154 0.00022 0.00007
group_lasso_N_1_i_15 761 4 12 - 7 6 6 0.00030 0.00121 0.01704 0.00169 0.00022 0.00007
group_lasso_N_1_i_16 761 4 11 - 7 6 6 0.00030 0.00121 0.01793 0.00166 0.00022 0.00007
group_lasso_N_1_i_17 761 4 12 - 7 6 6 0.00029 0.00120 0.01522 0.00164 0.00019 0.00006
group_lasso_N_1_i_18 761 5 11 - 7 6 6 0.00029 0.00110 0.01455 0.00152 0.00019 0.00007
group_lasso_N_1_i_19 761 5 10 - 5 6 6 0.00030 0.00098 0.01329 0.00126 0.00019 0.00007
group_lasso_N_2_i_0 2022 6 11 - 7 7 7 0.00080 0.00246 0.03452 0.00311 0.00054 0.00021
group_lasso_N_2_i_1 2022 7 11 - 7 7 7 0.00087 0.00248 0.04259 0.00309 0.00053 0.00022
group_lasso_N_2_i_2 2022 5 11 - 7 7 7 0.00073 0.00292 0.03949 0.00314 0.00057 0.00022
group_lasso_N_2_i_3 2022 7 12 - 8 8 8 0.00102 0.00316 0.04699 0.00371 0.00072 0.00024
group_lasso_N_2_i_4 2022 6 11 - 7 7 7 0.00092 0.00293 0.04221 0.00346 0.00062 0.00022
group_lasso_N_2_i_5 2022 7 11 - 9 8 8 0.00103 0.00243 0.04056 0.00357 0.00057 0.00021
group_lasso_N_2_i_6 2022 6 11 - 7 7 7 0.00079 0.00245 0.03730 0.00311 0.00052 0.00020
group_lasso_N_2_i_7 2022 6 10 - 7 7 7 0.00078 0.00230 0.03744 0.00310 0.00054 0.00020
group_lasso_N_2_i_8 2022 6 10 - 8 7 7 0.00079 0.00234 0.03938 0.00333 0.00053 0.00020
group_lasso_N_2_i_9 2022 6 11 - 8 7 7 0.00079 0.00235 0.03837 0.00336 0.00053 0.00022
group_lasso_N_2_i_10 2022 6 11 - 7 6 6 0.00080 0.00289 0.04676 0.00311 0.00056 0.00019
group_lasso_N_2_i_11 2022 7 11 - 7 7 7 0.00102 0.00287 0.04623 0.00345 0.00061 0.00019
group_lasso_N_2_i_12 2022 6 10 - 7 7 7 0.00091 0.00270 0.05127 0.00349 0.00061 0.00022
group_lasso_N_2_i_13 2022 7 13 - 7 7 7 0.00101 0.00347 0.04492 0.00356 0.00062 0.00022
group_lasso_N_2_i_14 2022 6 12 - 7 7 7 0.00094 0.00259 0.03859 0.00344 0.00055 0.00020
group_lasso_N_2_i_15 2022 6 11 - 8 7 7 0.00079 0.00243 0.03694 0.00336 0.00052 0.00022
group_lasso_N_2_i_16 2022 6 11 - 7 7 7 0.00081 0.00240 0.04211 0.00311 0.00052 0.00022
group_lasso_N_2_i_17 2022 6 11 - 7 7 7 0.00080 0.00251 0.03627 0.00307 0.00052 0.00020
group_lasso_N_2_i_18 2022 7 11 - 8 7 7 0.00087 0.00286 0.04580 0.00335 0.00062 0.00022
group_lasso_N_2_i_19 2022 6 10 - 8 7 7 0.00092 0.00235 0.03820 0.00342 0.00052 0.00022
group_lasso_N_3_i_0 3783 7 12 - 7 8 8 0.00179 0.00632 0.09303 0.00640 0.00136 0.00045
group_lasso_N_3_i_1 3783 7 12 - 8 8 8 0.00181 0.00663 0.10263 0.00699 0.00131 0.00045
group_lasso_N_3_i_2 3783 6 11 - 8 7 7 0.00167 0.00665 0.06413 0.00738 0.00134 0.00036
group_lasso_N_3_i_3 3783 6 11 - 8 7 7 0.00184 0.00611 0.09396 0.00745 0.00122 0.00040
group_lasso_N_3_i_4 3783 7 12 - 9 8 8 0.00179 0.00652 0.09673 0.00747 0.00129 0.00045
group_lasso_N_3_i_5 3783 6 12 - 8 7 7 0.00169 0.00627 0.09973 0.00699 0.00123 0.00040
group_lasso_N_3_i_6 3783 6 12 - 8 7 7 0.00167 0.00674 0.09808 0.00601 0.00123 0.00040
group_lasso_N_3_i_7 3783 7 12 - 9 8 8 0.00178 0.00665 0.10586 0.00756 0.00131 0.00045
group_lasso_N_3_i_8 3783 8 12 - 9 8 8 0.00195 0.00654 0.10374 0.00751 0.00131 0.00045
group_lasso_N_3_i_9 3783 7 12 - 9 7 7 0.00178 0.00671 0.09236 0.00744 0.00124 0.00040
group_lasso_N_3_i_10 3783 7 12 - 8 8 8 0.00179 0.00694 0.10241 0.00586 0.00145 0.00041
group_lasso_N_3_i_11 3783 8 12 - 10 8 8 0.00221 0.00650 0.09390 0.00712 0.00134 0.00046
group_lasso_N_3_i_12 3783 8 12 - 10 8 8 0.00194 0.00719 0.09329 0.00809 0.00146 0.00041
group_lasso_N_3_i_13 3783 7 12 - 8 7 7 0.00203 0.00664 0.09344 0.00617 0.00122 0.00041
group_lasso_N_3_i_14 3783 8 12 - 9 8 8 0.00198 0.00643 0.05164 0.00744 0.00136 0.00045
group_lasso_N_3_i_15 3783 8 12 - 9 8 8 0.00192 0.00626 0.09667 0.00624 0.00132 0.00045
group_lasso_N_3_i_16 3783 8 12 - 9 8 8 0.00198 0.00616 0.06490 0.00748 0.00137 0.00045
group_lasso_N_3_i_17 3783 7 11 - 9 7 7 0.00184 0.00646 0.09337 0.00752 0.00123 0.00036
group_lasso_N_3_i_18 3783 7 12 - 8 7 7 0.00200 0.00671 0.08989 0.00693 0.00124 0.00040
group_lasso_N_3_i_19 3783 7 12 - 8 8 8 0.00179 0.00707 0.09407 0.00697 0.00140 0.00039
group_lasso_N_4_i_0 6044 7 11 - 9 8 8 0.00372 0.00825 0.08675 0.00929 0.00254 0.00068
group_lasso_N_4_i_1 6044 7 11 - 9 7 7 0.00331 0.01034 0.07636 0.01163 0.00241 0.00060
group_lasso_N_4_i_2 6044 7 11 - 9 8 8 0.00332 0.00936 0.09203 0.01130 0.00233 0.00074
group_lasso_N_4_i_3 6044 6 11 - 7 7 7 0.00279 0.01065 0.07408 0.00972 0.00239 0.00060
group_lasso_N_4_i_4 6044 6 10 - 8 7 7 0.00309 0.00910 0.07713 0.01098 0.00240 0.00060
group_lasso_N_4_i_5 6044 7 11 - 9 8 8 0.00332 0.00932 0.07537 0.01131 0.00234 0.00075
group_lasso_N_4_i_6 6044 8 11 - 9 9 9 0.00319 0.01044 0.07465 0.01045 0.00272 0.00075
group_lasso_N_4_i_7 6044 7 11 - 9 8 8 0.00332 0.01038 0.08235 0.01195 0.00257 0.00067
group_lasso_N_4_i_8 6044 7 12 - 8 7 7 0.00332 0.01096 0.09033 0.01048 0.00239 0.00060
group_lasso_N_4_i_9 6044 8 12 - 9 8 8 0.00364 0.01153 0.09164 0.01165 0.00254 0.00067
group_lasso_N_4_i_10 6044 7 12 - 9 7 7 0.00338 0.01116 0.08548 0.00989 0.00240 0.00060
group_lasso_N_4_i_11 6044 7 11 - 8 7 7 0.00350 0.01017 0.08882 0.01088 0.00240 0.00061
group_lasso_N_4_i_12 6044 8 12 - 10 8 8 0.00373 0.01022 0.07332 0.01159 0.00231 0.00074
group_lasso_N_4_i_13 6044 8 12 - 10 8 8 0.00325 0.01025 0.08315 0.00932 0.00229 0.00074
group_lasso_N_4_i_14 6044 7 11 - 9 8 8 0.00303 0.01005 0.08752 0.01085 0.00255 0.00067
group_lasso_N_4_i_15 6044 7 11 - 8 7 7 0.00332 0.00827 0.08699 0.00823 0.00240 0.00060
group_lasso_N_4_i_16 6044 7 12 - 8 8 8 0.00328 0.01145 0.08421 0.01026 0.00248 0.00067
group_lasso_N_4_i_17 6044 7 12 - 8 7 7 0.00326 0.01063 0.07738 0.00827 0.00220 0.00066
group_lasso_N_4_i_18 6044 7 12 - 9 8 8 0.00307 0.01160 0.07562 0.01173 0.00255 0.00068
group_lasso_N_4_i_19 6044 7 11 - 9 8 8 0.00336 0.01011 0.08114 0.01054 0.00258 0.00067
group_lasso_N_5_i_0 8805 7 12 - 9 8 8 0.00539 0.01336 0.15490 0.01425 0.00376 0.00113
group_lasso_N_5_i_1 8805 7 12 - 8 8 8 0.00478 0.01432 0.12789 0.01328 0.00416 0.00099
group_lasso_N_5_i_2 8805 7 12 - 9 8 8 0.00527 0.01689 0.13607 0.01258 0.00412 0.00103
group_lasso_N_5_i_3 8805 7 12 - 9 8 8 0.00520 0.01569 0.12261 0.01514 0.00412 0.00101
group_lasso_N_5_i_4 8805 7 12 - 8 8 8 0.00522 0.01640 0.13366 0.01215 0.00409 0.00102
group_lasso_N_5_i_5 8805 8 12 - 9 8 8 0.00557 0.01594 0.13463 0.01448 0.00412 0.00103
group_lasso_N_5_i_6 8805 8 13 - 11 9 9 0.00557 0.01799 0.14506 0.01703 0.00430 0.00111
group_lasso_N_5_i_7 8805 8 13 - 11 9 9 0.00557 0.01593 0.13196 0.01626 0.00433 0.00114
group_lasso_N_5_i_8 8805 7 12 - 9 8 8 0.00578 0.01632 0.14206 0.01525 0.00414 0.00100
group_lasso_N_5_i_9 8805 8 12 - 10 9 9 0.00563 0.01265 0.10661 0.01594 0.00396 0.00125
group_lasso_N_5_i_10 8805 8 12 - 10 8 8 0.00507 0.01444 0.12481 0.01522 0.00386 0.00112
group_lasso_N_5_i_11 8805 7 12 - 9 8 8 0.00474 0.01315 0.13159 0.01476 0.00375 0.00113
group_lasso_N_5_i_12 8805 7 12 - 9 8 8 0.00471 0.01321 0.11506 0.01415 0.00376 0.00104
group_lasso_N_5_i_13 8805 7 12 - 9 8 8 0.00480 0.01528 0.12522 0.01420 0.00379 0.00113
group_lasso_N_5_i_14 8805 8 13 - 10 8 8 0.00504 0.01657 0.13945 0.01488 0.00412 0.00103
group_lasso_N_5_i_15 8805 6 11 - 9 7 7 0.00489 0.01440 0.15750 0.01530 0.00392 0.00091
group_lasso_N_5_i_16 8805 8 13 - 10 8 8 0.00556 0.01707 0.15121 0.01627 0.00411 0.00103
group_lasso_N_5_i_17 8805 7 12 - 8 8 8 0.00530 0.01363 0.12424 0.01428 0.00412 0.00098
group_lasso_N_5_i_18 8805 8 12 - 9 8 8 0.00552 0.01342 0.13743 0.01483 0.00376 0.00114
group_lasso_N_5_i_19 8805 7 12 - 8 8 8 0.00477 0.01616 0.15250 0.01364 0.00410 0.00103
group_lasso_N_8_i_0 20088 8 12 - 11 8 - 0.01663 0.03164 0.42595 0.03193 0.01264 -
group_lasso_N_8_i_1 20088 8 12 - 11 9 - 0.01533 0.03202 0.43969 0.03056 0.01308 -
group_lasso_N_8_i_2 20088 8 11 - 10 8 - 0.01536 0.03429 0.52114 0.03041 0.01263 -
group_lasso_N_8_i_3 20088 8 12 - 11 9 - 0.01680 0.03238 0.53397 0.03206 0.01312 -
group_lasso_N_8_i_4 20088 8 13 - 11 9 - 0.01536 0.03464 0.43105 0.03163 0.01310 -
group_lasso_N_8_i_5 20088 8 12 - 11 9 - 0.01529 0.03548 0.44647 0.03194 0.01441 -
group_lasso_N_8_i_6 20088 7 11 - 9 8 - 0.01607 0.03011 0.43955 0.03064 0.01260 -
group_lasso_N_8_i_7 20088 8 12 - 10 8 - 0.01534 0.03510 0.37856 0.03001 0.01385 -
group_lasso_N_8_i_8 20088 8 13 - 11 9 - 0.01667 0.03326 0.44272 0.03138 0.01322 -
group_lasso_N_8_i_9 20088 8 12 - 10 8 - 0.01531 0.03289 0.36489 0.02823 0.01272 -
group_lasso_N_8_i_10 20088 8 12 - 10 8 - 0.01546 0.02967 0.44515 0.03007 0.01382 -
group_lasso_N_8_i_11 20088 8 12 - 10 8 - 0.01682 0.03171 0.36262 0.02546 0.01268 -
group_lasso_N_8_i_12 20088 7 11 - 9 8 - 0.01448 0.03253 0.39282 0.02447 0.01382 -
group_lasso_N_8_i_13 20088 8 12 - 10 8 - 0.01684 0.03112 0.38244 0.03053 0.01271 -
group_lasso_N_8_i_14 20088 7 12 - 10 8 - 0.01587 0.03158 0.43929 0.02986 0.01263 -
group_lasso_N_8_i_15 20088 8 12 - 10 8 - 0.01534 0.03216 0.43672 0.02979 0.01272 -
group_lasso_N_8_i_16 20088 8 12 - 10 8 - 0.01541 0.03232 0.45250 0.03021 0.01268 -
group_lasso_N_8_i_17 20088 7 12 - 9 8 - 0.01487 0.03449 0.39412 0.02670 0.01387 -
group_lasso_N_8_i_18 20088 7 12 - 9 8 - 0.01594 0.03084 0.45118 0.02881 0.01258 -
group_lasso_N_8_i_19 20088 8 11 - 10 8 - 0.01543 0.03438 0.48799 0.03068 0.01379 -
group_lasso_N_10_i_0 30110 8 12 - 10 8 - 0.02990 0.04260 0.81705 0.03441 0.02297 -
group_lasso_N_10_i_1 30110 8 13 - 10 8 - 0.02640 0.04393 0.71839 0.03325 0.02295 -
group_lasso_N_10_i_2 30110 8 13 - 10 8 - 0.02634 0.04325 0.71089 0.03299 0.02309 -
group_lasso_N_10_i_3 30110 8 12 - 10 8 - 0.02629 0.04314 0.82863 0.03320 0.02288 -
group_lasso_N_10_i_4 30110 8 12 - 10 9 - 0.02683 0.04325 0.74742 0.03280 0.02406 -
group_lasso_N_10_i_5 30110 8 12 - 10 8 - 0.02697 0.04647 0.71730 0.03281 0.02506 -
group_lasso_N_10_i_6 30110 8 13 - 11 9 - 0.02953 0.04871 0.72293 0.03796 0.02569 -
group_lasso_N_10_i_7 30110 8 12 - 11 8 - 0.02942 0.04018 0.78965 0.03761 0.02297 -
group_lasso_N_10_i_8 30110 8 12 - 10 8 - 0.02650 0.04055 0.73732 0.03352 0.02290 -
group_lasso_N_10_i_9 30110 7 12 - 10 8 - 0.02551 0.04065 0.70643 0.03675 $0$
group_lasso_N_10_i_10 30110 8 13 - 10 9 - 0.02617 0.04443 0.71020 0.03381 0.02367 -
group_lasso_N_10_i_11 30110 7 12 - 10 8 - 0.02521 0.04700 0.85205 0.03292 0.02512 -
group_lasso_N_10_i_12 30110 8 12 - 11 8 - 0.02932 0.03969 0.84706 0.03805 0.02333 -
group_lasso_N_10_i_13 30110 8 12 - 10 8 - 0.02654 0.04085 0.71497 0.03355 0.02305 -
group_lasso_N_10_i_14 30110 8 12 - 11 8 - 0.02646 0.04156 0.71650 0.03545 0.02291 -
group_lasso_N_10_i_15 30110 8 13 - 11 8 - 0.02719 0.04412 0.70731 0.03440 0.02288 -
group_lasso_N_10_i_16 30110 8 12 - 10 8 - 0.02660 0.04692 0.72722 0.03345 0.02508 -
group_lasso_N_10_i_17 30110 8 12 - 10 8 - 0.02956 0.04225 0.81926 0.03716 0.02288 -
group_lasso_N_10_i_18 30110 8 12 - 10 8 - 0.02757 0.04193 0.71638 0.03325 0.02331 -
group_lasso_N_10_i_19 30110 7 11 - 9 8 - 0.02621 0.03705 0.57778 0.03114 0.02293 -
group_lasso_N_12_i_0 42132 8 12 - 10 8 - 0.04176 0.05646 1.13886 0.04270 0.03342 -
group_lasso_N_12_i_1 42132 8 12 - 11 8 - 0.04718 0.06379 1.21747 0.04359 0.03531 -
group_lasso_N_12_i_2 42132 7 12 - 10 7 - 0.04523 0.06090 1.29220 0.04656 0.03292 -
group_lasso_N_12_i_3 42132 8 13 - 11 8 - 0.04103 0.06254 1.14120 0.04422 0.03333 -
group_lasso_N_12_i_4 42132 8 12 - 11 9 - 0.04334 0.05760 1.33111 0.04989 0.03406 -
group_lasso_N_12_i_5 42132 8 12 - 10 8 - 0.04179 0.05618 1.13627 0.04234 0.03340 -
group_lasso_N_12_i_6 42132 8 12 - 10 8 - 0.04743 0.06710 1.14156 0.04154 0.03561 -
group_lasso_N_12_i_7 42132 8 12 - 10 8 - 0.04925 0.06160 1.40540 0.04672 0.03352 -
group_lasso_N_12_i_8 42132 8 12 - 10 8 - 0.04718 0.05998 1.35485 0.04598 0.03311 -
group_lasso_N_12_i_9 42132 8 12 - 10 8 - 0.04905 0.05656 1.21177 0.04653 0.03330 -
group_lasso_N_12_i_10 42132 8 12 - 10 8 - 0.04734 0.06416 1.16632 0.04149 0.03473 -
group_lasso_N_12_i_11 42132 8 12 - 10 8 - 0.04929 0.05646 1.16314 0.05081 0.03312 -
group_lasso_N_12_i_12 42132 8 12 - 11 8 - 0.04759 0.06398 1.17661 0.05339 0.03473 -
group_lasso_N_12_i_13 42132 8 12 - 10 8 - 0.04926 0.05970 1.15317 0.05015 0.03347 -
group_lasso_N_12_i_14 42132 8 12 - 10 8 - 0.04236 0.05699 1.15607 0.04171 0.03296 -
group_lasso_N_12_i_15 42132 8 12 - 11 8 - 0.04935 0.06750 1.37147 0.04819 0.03532 -
group_lasso_N_12_i_16 42132 8 12 - 11 8 - 0.05005 0.06500 1.13286 0.04832 0.03555 -
group_lasso_N_12_i_17 42132 8 12 - 10 8 - 0.04763 0.06239 1.19049 0.04723 0.03542 -
group_lasso_N_12_i_18 42132 8 12 - 11 8 - 0.04341 0.06100 1.56013 0.04476 0.03346 -
group_lasso_N_12_i_19 42132 8 12 - 11 8 - 0.04741 0.06446 1.36957 0.05266 $0$
group_lasso_N_14_i_0 56154 7 11 - 10 8 - 0.06732 0.08825 1.84934 0.06011 0.04971 -
group_lasso_N_14_i_1 56154 8 12 - 10 8 - 0.07072 0.09272 1.74199 0.06349 0.04997 -
group_lasso_N_14_i_2 56154 8 12 - 11 8 - 0.07052 0.08224 1.75298 0.06603 0.04582 -
group_lasso_N_14_i_3 56154 8 12 - 10 8 - 0.06288 0.08092 1.81541 0.05672 0.04587 -
group_lasso_N_14_i_4 56154 8 12 - 11 8 - 0.06408 0.08039 1.77375 0.06373 0.04576 -
group_lasso_N_14_i_5 56154 7 11 - 10 8 - 0.05957 0.08723 1.73803 0.05546 0.05051 -
group_lasso_N_14_i_6 56154 8 13 - 11 9 - 0.07036 0.08742 1.78430 0.06126 0.05179 -
group_lasso_N_14_i_7 56154 8 13 - 11 9 - 0.06455 0.08512 1.75818 0.05494 0.04697 -
group_lasso_N_14_i_8 56154 8 12 - 11 8 - 0.07086 0.08050 1.71973 0.06242 0.04588 -
group_lasso_N_14_i_9 56154 7 11 - 10 8 - 0.06089 0.07734 1.76336 0.05817 0.04569 -
group_lasso_N_14_i_10 56154 7 11 - 10 8 - 0.06136 0.07749 2.37906 0.06483 0.04582 -
group_lasso_N_14_i_11 56154 8 12 - 11 8 - 0.06271 0.08033 1.74188 0.05966 0.04585 -
group_lasso_N_14_i_12 56154 7 12 - 11 8 - 0.05962 0.09189 1.82984 0.05886 0.05023 -
group_lasso_N_14_i_13 56154 8 12 - 10 8 - 0.07108 0.08138 1.73734 0.05703 0.04582 -
group_lasso_N_14_i_14 56154 7 12 - 10 8 - 0.06805 0.09148 1.83138 0.05970 0.05013 -
group_lasso_N_14_i_15 56154 7 12 - 9 8 - 0.06833 0.08114 1.76118 0.06040 0.04593 -
group_lasso_N_14_i_16 56154 8 12 - 11 8 - 0.07149 0.08162 1.74090 0.05902 0.04518 -
group_lasso_N_14_i_17 56154 8 13 - 11 8 - 0.06544 0.08407 1.72602 0.05831 0.04578 -
group_lasso_N_14_i_18 56154 7 12 - 10 8 - 0.05882 0.08192 1.85183 0.06374 0.04648 -
group_lasso_N_14_i_19 56154 8 12 - 10 8 - 0.07138 0.08001 1.72838 0.05553 0.04590 -
group_lasso_N_16_i_0 72176 7 12 - 9 8 - 0.08672 0.10810 2.73494 0.06387 0.06441 -
group_lasso_N_16_i_1 72176 8 12 - 9 8 - 0.09175 0.10956 2.57650 0.06641 0.06458 -
group_lasso_N_16_i_2 72176 7 13 - 8 8 - 0.08420 0.11815 2.47541 0.06230 0.06381 -
group_lasso_N_16_i_3 72176 7 12 - 8 8 - 0.09670 0.11035 2.61814 0.06587 0.06468 -
group_lasso_N_16_i_4 72176 7 12 - 9 8 - 0.08805 0.11655 2.53132 0.06833 0.06994 -
group_lasso_N_16_i_5 72176 7 12 - 8 8 - 0.08591 0.11024 2.56714 0.06273 0.06479 -
group_lasso_N_16_i_6 72176 8 12 - 9 8 - 0.09397 0.12189 2.57184 0.06543 0.07089 -
group_lasso_N_16_i_7 72176 7 12 - 8 8 - 0.09662 0.11715 2.62896 0.07091 0.07108 -
group_lasso_N_16_i_8 72176 7 12 - 8 8 - 0.09697 0.10666 2.52536 0.06546 0.06463 -
group_lasso_N_16_i_9 72176 8 12 - 9 8 - 0.10105 0.11045 2.59268 0.07594 0.06850 -
group_lasso_N_16_i_10 72176 8 12 - 8 8 - 0.08739 0.12383 2.55628 0.06493 0.07137 -
group_lasso_N_16_i_11 72176 8 12 - 8 8 - 0.09985 0.11001 2.27946 0.07167 0.06446 -
group_lasso_N_16_i_12 72176 8 12 - 9 8 - 0.09288 0.12516 2.65303 0.06511 0.07138 -
group_lasso_N_16_i_13 72176 8 12 - 9 8 - 0.10170 0.12524 2.70107 0.06619 0.07118 -
group_lasso_N_16_i_14 72176 8 12 - 8 8 - 0.10075 0.11424 2.69967 0.06278 0.06494 -
group_lasso_N_16_i_15 72176 8 12 - 9 8 - 0.10100 0.12544 2.59960 0.06730 0.07098 -
group_lasso_N_16_i_16 72176 8 12 - 9 8 - 0.10077 0.11113 2.91541 0.07759 0.06459 -
group_lasso_N_16_i_17 72176 7 12 - 8 8 - 0.08758 0.12504 2.53327 0.07146 0.07131 -
group_lasso_N_16_i_18 72176 8 12 - 8 8 - 0.10206 0.11034 2.64789 0.07121 0.06484 -
group_lasso_N_16_i_19 72176 7 12 - 8 8 - 0.08468 0.11729 2.65421 0.07299 0.07011 -
Table 6: Iterations and solver runtimes for portfolio optimization problems
Iterations Solver Runtime (s)
Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
portfolio_N_2_i_0 804 9 12 14 12 9 9 0.00046 0.00099 0.00263 0.00485 0.00028 0.00013
portfolio_N_2_i_1 804 9 14 13 12 8 8 0.00045 0.00108 0.00255 0.00486 0.00027 0.00011
portfolio_N_2_i_2 804 8 12 11 11 8 8 0.00041 0.00102 0.00251 0.00461 0.00026 0.00010
portfolio_N_2_i_3 804 8 15 12 12 8 8 0.00041 0.00120 0.00252 0.00486 0.00025 0.00011
portfolio_N_2_i_4 804 9 14 13 11 8 8 0.00044 0.00113 0.00259 0.00461 0.00026 0.00010
portfolio_N_2_i_5 804 10 14 13 14 9 9 0.00049 0.00112 0.00256 0.00535 0.00027 0.00012
portfolio_N_2_i_6 804 10 18 12 12 10 10 0.00049 0.00151 0.00255 0.00485 0.00030 0.00014
portfolio_N_2_i_7 804 8 10 11 11 7 7 0.00041 0.00090 0.00248 0.00447 0.00024 0.00010
portfolio_N_2_i_8 804 9 17 11 10 9 9 0.00045 0.00150 0.00251 0.00423 0.00027 0.00012
portfolio_N_2_i_9 804 16 13 14 13 9 9 0.00070 0.00101 0.00258 0.00493 0.00028 0.00012
portfolio_N_2_i_10 804 10 13 15 13 9 9 0.00050 0.00100 0.00262 0.00490 0.00027 0.00011
portfolio_N_2_i_11 804 10 12 13 13 10 10 0.00044 0.00087 0.00229 0.00451 0.00028 0.00013
portfolio_N_2_i_12 804 10 11 13 12 10 10 0.00043 0.00079 0.00225 0.00426 0.00026 0.00012
portfolio_N_2_i_13 804 9 12 11 11 8 8 0.00040 0.00085 0.00220 0.00402 0.00023 0.00010
portfolio_N_2_i_14 804 9 14 13 13 9 9 0.00040 0.00095 0.00228 0.00444 0.00025 0.00011
portfolio_N_2_i_15 804 9 11 12 12 8 8 0.00040 0.00077 0.00224 0.00424 0.00022 0.00010
portfolio_N_2_i_16 804 10 14 11 12 8 8 0.00043 0.00095 0.00224 0.00425 0.00022 0.00010
portfolio_N_2_i_17 804 8 11 10 9 8 8 0.00037 0.00081 0.00219 0.00359 0.00023 0.00010
portfolio_N_2_i_18 804 9 13 11 11 9 9 0.00040 0.00092 0.00223 0.00400 0.00024 0.00011
portfolio_N_2_i_19 804 10 11 12 11 7 7 0.00044 0.00081 0.00224 0.00401 0.00021 0.00009
portfolio_N_4_i_0 2008 10 13 13 12 9 9 0.00107 0.00191 0.00293 0.00703 0.00054 0.00026
portfolio_N_4_i_1 2008 10 14 15 12 11 11 0.00107 0.00243 0.00310 0.00709 0.00062 0.00033
portfolio_N_4_i_2 2008 11 15 12 12 10 10 0.00131 0.00256 0.00287 0.00670 0.00058 0.00030
portfolio_N_4_i_3 2008 11 16 14 14 9 9 0.00113 0.00220 0.00299 0.00807 0.00056 0.00027
portfolio_N_4_i_4 2008 9 17 13 12 8 8 0.00098 0.00239 0.00284 0.00705 0.00049 0.00024
portfolio_N_4_i_5 2008 11 13 12 14 9 9 0.00113 0.00193 0.00279 0.00805 0.00053 0.00028
portfolio_N_4_i_6 2008 10 14 13 11 9 9 0.00107 0.00204 0.00290 0.00650 0.00054 0.00030
portfolio_N_4_i_7 2008 11 15 13 12 9 9 0.00110 0.00211 0.00293 0.00689 0.00053 0.00028
portfolio_N_4_i_8 2008 11 13 12 12 9 9 0.00109 0.00189 0.00292 0.00708 0.00053 0.00027
portfolio_N_4_i_9 2008 11 16 13 13 10 10 0.00111 0.00240 0.00295 0.00742 0.00057 0.00030
portfolio_N_4_i_10 2008 11 18 15 14 10 10 0.00112 0.00246 0.00311 0.00797 0.00057 0.00031
portfolio_N_4_i_11 2008 9 15 14 12 9 9 0.00097 0.00223 0.00291 0.00702 0.00056 0.00027
portfolio_N_4_i_12 2008 12 14 12 12 9 9 0.00120 0.00207 0.00278 0.00698 0.00054 0.00027
portfolio_N_4_i_13 2008 10 15 14 11 9 9 0.00102 0.00230 0.00303 0.00666 0.00053 0.00027
portfolio_N_4_i_14 2008 10 17 15 14 9 9 0.00106 0.00233 0.00305 0.00750 0.00054 0.00027
portfolio_N_4_i_15 2008 11 12 12 13 10 10 0.00115 0.00185 0.00289 0.00746 0.00058 0.00030
portfolio_N_4_i_16 2008 11 17 14 12 10 10 0.00114 0.00270 0.00292 0.00698 0.00058 0.00030
portfolio_N_4_i_17 2008 9 13 14 13 9 9 0.00098 0.00187 0.00292 0.00748 0.00052 0.00028
portfolio_N_4_i_18 2008 11 14 13 14 9 9 0.00112 0.00195 0.00292 0.00776 0.00052 0.00027
portfolio_N_4_i_19 2008 9 19 12 12 8 8 0.00096 0.00286 0.00285 0.00692 0.00049 0.00027
portfolio_N_6_i_0 3612 13 16 14 14 11 11 0.00215 0.00403 0.00395 0.01172 0.00112 0.00056
portfolio_N_6_i_1 3612 14 16 14 12 8 8 0.00224 0.00401 0.00387 0.01036 0.00090 0.00041
portfolio_N_6_i_2 3612 12 14 14 13 10 10 0.00210 0.00408 0.00452 0.01138 0.00124 0.00057
portfolio_N_6_i_3 3612 15 18 13 14 10 10 0.00279 0.00505 0.00437 0.01193 0.00121 0.00057
portfolio_N_6_i_4 3612 12 14 13 15 11 11 0.00235 0.00424 0.00442 0.01279 0.00129 0.00063
portfolio_N_6_i_5 3612 11 13 12 12 9 9 0.00221 0.00385 0.00422 0.01174 0.00112 0.00052
portfolio_N_6_i_6 3612 10 15 14 14 9 9 0.00208 0.00457 0.00453 0.01331 0.00118 0.00052
portfolio_N_6_i_7 3612 12 14 14 13 9 9 0.00237 0.00393 0.00451 0.01255 0.00114 0.00051
portfolio_N_6_i_8 3612 12 16 14 14 10 10 0.00235 0.00405 0.00454 0.01340 0.00124 0.00053
portfolio_N_6_i_9 3612 13 15 15 14 10 10 0.00217 0.00389 0.00398 0.01178 0.00102 0.00057
portfolio_N_6_i_10 3612 11 15 14 13 10 10 0.00198 0.00365 0.00387 0.01253 0.00104 0.00050
portfolio_N_6_i_11 3612 11 14 13 11 9 9 0.00194 0.00415 0.00434 0.00978 0.00114 0.00051
portfolio_N_6_i_12 3612 12 15 16 13 11 11 0.00236 0.00388 0.00410 0.01128 0.00117 0.00056
portfolio_N_6_i_13 3612 11 15 13 13 10 10 0.00197 0.00428 0.00435 0.01116 0.00122 0.00053
portfolio_N_6_i_14 3612 13 16 14 14 10 10 0.00252 0.00461 0.00452 0.01329 0.00122 0.00058
portfolio_N_6_i_15 3612 12 17 13 14 10 10 0.00238 0.00484 0.00438 0.01335 0.00120 0.00058
portfolio_N_6_i_16 3612 11 17 14 12 9 9 0.00225 0.00431 0.00386 0.01178 0.00103 0.00047
portfolio_N_6_i_17 3612 13 15 15 13 10 10 0.00250 0.00414 0.00485 0.01259 0.00121 0.00057
portfolio_N_6_i_18 3612 12 20 15 14 10 10 0.00244 0.00614 0.00461 0.01338 0.00123 0.00057
portfolio_N_6_i_19 3612 13 13 12 14 9 9 0.00248 0.00333 0.00375 0.01202 0.00096 0.00046
portfolio_N_8_i_0 5616 11 15 14 14 11 11 0.00327 0.00665 0.00574 0.01784 0.00183 0.00101
portfolio_N_8_i_1 5616 12 15 13 13 11 11 0.00396 0.00685 0.00584 0.01858 0.00189 0.00097
portfolio_N_8_i_2 5616 12 16 13 13 11 11 0.00392 0.00691 0.00569 0.01836 0.00181 0.00090
portfolio_N_8_i_3 5616 13 15 16 15 10 10 0.00418 0.00551 0.00537 0.02087 0.00136 0.00087
portfolio_N_8_i_4 5616 13 16 13 12 11 11 0.00357 0.00635 0.00478 0.01551 0.00159 0.00100
portfolio_N_8_i_5 5616 12 15 13 15 10 10 0.00340 0.00580 0.00476 0.01890 0.00138 0.00083
portfolio_N_8_i_6 5616 14 17 14 14 10 10 0.00377 0.00628 0.00500 0.01778 0.00137 0.00085
portfolio_N_8_i_7 5616 12 14 15 15 9 9 0.00340 0.00544 0.00526 0.01896 0.00133 0.00084
portfolio_N_8_i_8 5616 13 17 14 15 11 11 0.00363 0.00637 0.00525 0.01838 0.00150 0.00088
portfolio_N_8_i_9 5616 14 18 16 15 10 10 0.00379 0.00621 0.00547 0.01805 0.00144 0.00084
portfolio_N_8_i_10 5616 13 18 15 16 10 10 0.00360 0.00625 0.00516 0.01937 0.00135 0.00087
portfolio_N_8_i_11 5616 12 18 15 14 10 10 0.00338 0.00660 0.00522 0.01747 0.00138 0.00084
portfolio_N_8_i_12 5616 11 15 13 13 9 9 0.00322 0.00561 0.00483 0.01652 0.00138 0.00074
portfolio_N_8_i_13 5616 14 17 15 14 10 10 0.00376 0.00652 0.00520 0.01770 0.00143 0.00083
portfolio_N_8_i_14 5616 12 17 15 15 10 10 0.00342 0.00622 0.00521 0.01825 0.00136 0.00086
portfolio_N_8_i_15 5616 14 18 15 15 10 10 0.00379 0.00653 0.00515 0.01854 0.00143 0.00084
portfolio_N_8_i_16 5616 14 17 15 15 12 12 0.00379 0.00609 0.00514 0.01866 0.00158 0.00098
portfolio_N_8_i_17 5616 12 15 15 13 9 9 0.00337 0.00571 0.00525 0.01621 0.00133 0.00075
portfolio_N_8_i_18 5616 12 14 14 15 10 10 0.00345 0.00547 0.00507 0.01890 0.00136 0.00083
portfolio_N_8_i_19 5616 12 16 15 15 10 10 0.00342 0.00583 0.00514 0.01860 0.00137 0.00083
portfolio_N_10_i_0 8020 13 18 13 14 9 9 0.00567 0.01012 0.00598 0.02258 0.00187 0.00144
portfolio_N_10_i_1 8020 12 19 14 13 10 10 0.00543 0.01160 0.00641 0.02101 0.00202 0.00157
portfolio_N_10_i_2 8020 14 18 15 14 10 10 0.00595 0.00946 0.00654 0.02224 0.00202 0.00158
portfolio_N_10_i_3 8020 14 16 15 15 10 10 0.00590 0.00945 0.00747 0.02371 0.00237 0.00178
portfolio_N_10_i_4 8020 12 16 15 14 9 9 0.00619 0.00860 0.00655 0.02196 0.00188 0.00143
portfolio_N_10_i_5 8020 11 15 15 14 11 11 0.00517 0.00979 0.00648 0.02214 0.00222 0.00195
portfolio_N_10_i_6 8020 14 17 17 15 11 11 0.00587 0.01039 0.00793 0.02337 0.00256 0.00198
portfolio_N_10_i_7 8020 14 15 14 14 11 11 0.00677 0.01008 0.00753 0.02500 0.00258 0.00195
portfolio_N_10_i_8 8020 13 16 16 13 11 11 0.00647 0.00867 0.00691 0.02054 0.00224 0.00172
portfolio_N_10_i_9 8020 12 15 15 13 11 11 0.00541 0.00941 0.00759 0.02041 0.00249 0.00195
portfolio_N_10_i_10 8020 13 19 15 14 11 11 0.00653 0.01125 0.00770 0.02481 0.00257 0.00196
portfolio_N_10_i_11 8020 14 18 17 15 10 10 0.00681 0.01118 0.00832 0.02592 0.00240 0.00161
portfolio_N_10_i_12 8020 13 18 14 13 12 12 0.00645 0.00939 0.00615 0.02070 0.00242 0.00185
portfolio_N_10_i_13 8020 13 13 14 14 9 9 0.00562 0.00730 0.00627 0.02257 0.00187 0.00143
portfolio_N_10_i_14 8020 13 18 16 15 10 10 0.00577 0.00979 0.00696 0.02354 0.00206 0.00161
portfolio_N_10_i_15 8020 12 16 15 13 11 11 0.00540 0.00855 0.00646 0.02049 0.00221 0.00174
portfolio_N_10_i_16 8020 13 18 15 14 12 12 0.00570 0.00988 0.00650 0.02208 0.00235 0.00188
portfolio_N_10_i_17 8020 13 17 16 15 10 10 0.00572 0.00930 0.00695 0.02366 0.00207 0.00157
portfolio_N_10_i_18 8020 13 18 15 13 10 10 0.00565 0.01098 0.00751 0.02122 0.00236 0.00179
portfolio_N_10_i_19 8020 13 15 14 14 12 12 0.00643 0.00842 0.00650 0.02185 0.00245 0.00186
portfolio_N_15_i_0 15780 13 18 14 15 11 - 0.01573 0.01625 0.01104 0.03435 0.00476 -
portfolio_N_15_i_1 15780 13 18 15 16 11 - 0.01555 0.01698 0.01102 0.03537 0.00480 -
portfolio_N_15_i_2 15780 13 16 17 16 11 - 0.01575 0.01525 0.01227 0.03390 0.00488 -
portfolio_N_15_i_3 15780 12 18 15 15 12 - 0.01520 0.02108 0.01112 0.03209 0.00529 -
portfolio_N_15_i_4 15780 13 18 15 14 10 - 0.01562 0.01648 0.01098 0.03003 0.00447 -
portfolio_N_15_i_5 15780 13 15 15 15 10 - 0.01560 0.01402 0.01105 0.03340 0.00443 -
portfolio_N_15_i_6 15780 14 15 17 15 11 - 0.01616 0.01385 0.01165 0.03171 0.00479 -
portfolio_N_15_i_7 15780 13 16 16 15 10 - 0.01560 0.01443 0.01140 0.03200 0.00441 -
portfolio_N_15_i_8 15780 14 19 17 15 11 - 0.01606 0.01592 0.01228 0.03191 0.00484 -
portfolio_N_15_i_9 15780 13 14 16 15 10 - 0.01554 0.01470 0.01153 0.03193 0.00446 -
portfolio_N_15_i_10 15780 14 17 15 16 10 - 0.01614 0.01479 0.01128 0.03533 0.00447 -
portfolio_N_15_i_11 15780 13 17 14 15 11 - 0.01566 0.01509 0.01094 0.03390 0.00480 -
portfolio_N_15_i_12 15780 14 21 17 16 11 - 0.01614 0.01952 0.01213 0.03332 0.00481 -
portfolio_N_15_i_13 15780 13 16 15 14 12 - 0.01543 0.01474 0.01108 0.02999 0.00519 -
portfolio_N_15_i_14 15780 12 17 16 16 12 - 0.01511 0.01508 0.01151 0.03332 0.00521 -
portfolio_N_15_i_15 15780 15 17 18 16 13 - 0.01665 0.01487 0.01277 0.03564 0.00569 -
portfolio_N_15_i_16 15780 12 16 15 15 11 - 0.01519 0.01448 0.01143 0.03405 0.00487 -
portfolio_N_15_i_17 15780 13 15 15 14 11 - 0.01557 0.01546 0.01118 0.03211 0.00489 -
portfolio_N_15_i_18 15780 13 17 15 15 10 - 0.01562 0.01611 0.01145 0.03200 0.00438 -
portfolio_N_15_i_19 15780 12 17 16 14 10 - 0.01538 0.01628 0.01147 0.03246 0.00449 -
portfolio_N_20_i_0 26040 13 15 16 14 11 - 0.01579 0.02319 0.01691 0.04306 0.00867 -
portfolio_N_20_i_1 26040 13 19 16 15 10 - 0.01596 0.02887 0.01702 0.04652 $0$
portfolio_N_20_i_2 26040 13 17 16 15 13 - 0.01584 0.02901 0.01678 0.04624 0.00982 -
portfolio_N_20_i_3 26040 13 16 16 14 11 - 0.01462 0.02260 0.01738 0.04299 0.00871 -
portfolio_N_20_i_4 26040 15 19 16 16 13 - 0.01763 0.03265 0.01731 0.04858 0.01017 -
portfolio_N_20_i_5 26040 13 16 16 15 10 - 0.01851 0.02140 0.01698 0.04604 0.00802 -
portfolio_N_20_i_6 26040 14 20 16 15 11 - 0.01711 0.02815 0.01691 0.04646 0.00873 -
portfolio_N_20_i_7 26040 15 17 17 15 11 - 0.01803 0.03054 0.02014 0.04598 0.00995 -
portfolio_N_20_i_8 26040 13 18 15 15 11 - 0.01829 0.02487 0.01659 0.04978 0.00834 -
portfolio_N_20_i_9 26040 14 17 15 15 11 - 0.01674 0.02765 0.01654 0.04574 0.00868 -
portfolio_N_20_i_10 26040 14 18 17 14 11 - 0.01678 0.02426 0.01734 0.04360 0.00856 -
portfolio_N_20_i_11 26040 13 16 17 15 13 - 0.01566 0.02243 0.01722 0.04618 0.01007 -
portfolio_N_20_i_12 26040 15 17 16 16 11 - 0.01770 0.02716 0.01693 0.04894 0.00866 -
portfolio_N_20_i_13 26040 13 15 14 13 10 - 0.01550 0.02817 0.01549 0.04090 0.00795 -
portfolio_N_20_i_14 26040 13 18 17 14 10 - 0.01559 0.02457 0.01755 0.04368 0.00788 -
portfolio_N_20_i_15 26040 13 16 16 14 11 - 0.01849 0.03036 0.01962 0.05036 0.01012 -
portfolio_N_20_i_16 26040 13 17 15 15 12 - 0.01848 0.02328 0.01900 0.05309 0.01077 -
portfolio_N_20_i_17 26040 13 18 17 16 12 - 0.01567 0.02475 0.01766 0.04937 0.00933 -
portfolio_N_20_i_18 26040 13 18 17 15 10 - 0.01576 0.02845 0.01770 0.04576 0.00794 -
portfolio_N_20_i_19 26040 14 16 16 16 11 - 0.01681 0.02650 0.01691 0.04792 0.00869 -
portfolio_N_25_i_0 38800 15 18 18 15 12 - 0.02811 0.03697 0.02796 0.06077 0.01479 -
portfolio_N_25_i_1 38800 13 19 17 15 12 - 0.02489 0.03911 0.02666 0.06077 0.01491 -
portfolio_N_25_i_2 38800 14 15 17 16 11 - 0.02627 0.03189 0.02655 0.06440 0.01396 -
portfolio_N_25_i_3 38800 15 20 16 15 12 - 0.02832 0.03838 0.02565 0.06078 0.01489 -
portfolio_N_25_i_4 38800 13 16 16 15 11 - 0.02451 0.03297 0.02543 0.06389 0.01387 -
portfolio_N_25_i_5 38800 14 15 16 15 12 - 0.02645 0.03208 0.02487 0.06108 0.01474 -
portfolio_N_25_i_6 38800 14 18 16 15 11 - 0.02653 0.03622 0.02487 0.06103 0.01371 -
portfolio_N_25_i_7 38800 14 19 17 14 13 - 0.02668 0.04121 0.03103 0.05905 0.01871 -
portfolio_N_25_i_8 38800 14 17 16 15 11 - 0.03083 0.03273 0.02523 0.06138 0.01385 -
portfolio_N_25_i_9 38800 14 19 17 16 12 - 0.02653 0.03570 0.02635 0.06449 0.01502 -
portfolio_N_25_i_10 38800 14 18 15 15 11 - 0.03111 0.03593 0.02820 0.07083 0.01615 -
portfolio_N_25_i_11 38800 15 21 17 16 11 - 0.02816 0.03953 0.02618 0.06469 0.01383 -
portfolio_N_25_i_12 38800 15 18 16 15 12 - 0.02831 0.03316 0.02426 0.06071 0.01493 -
portfolio_N_25_i_13 38800 16 19 17 16 11 - 0.02977 0.03784 0.02601 0.06431 0.01380 -
portfolio_N_25_i_14 38800 14 16 17 15 10 - 0.02643 0.03204 0.02684 0.06072 0.01276 -
portfolio_N_25_i_15 38800 13 16 17 14 11 - 0.02484 0.03209 0.02575 0.05748 0.01383 -
portfolio_N_25_i_16 38800 14 19 16 14 12 - 0.02663 0.06186 0.02918 0.05802 0.01725 -
portfolio_N_25_i_17 38800 14 15 17 14 11 - 0.03080 0.03863 0.03059 0.06712 0.01616 -
portfolio_N_25_i_18 38800 15 18 16 16 12 - 0.03304 0.03557 0.02505 0.06457 0.01495 -
portfolio_N_25_i_19 38800 14 17 17 14 12 - 0.02603 0.03617 0.02648 0.05771 0.01506 -
portfolio_N_30_i_0 54060 15 17 17 15 11 - 0.04078 0.04232 0.03562 0.07782 0.01979 -
portfolio_N_30_i_1 54060 15 18 19 14 11 - 0.04038 0.04706 0.03848 0.07431 0.01976 -
portfolio_N_30_i_2 54060 14 20 16 14 14 - 0.03791 0.04927 0.03356 0.07436 0.02453 -
portfolio_N_30_i_3 54060 14 20 17 16 12 - 0.03790 0.05631 0.03541 0.08273 0.02143 -
portfolio_N_30_i_4 54060 14 17 17 15 13 - 0.03847 0.04522 0.03540 0.07840 0.02313 -
portfolio_N_30_i_5 54060 14 18 18 15 11 - 0.03849 0.04513 0.03699 0.08154 0.01995 -
portfolio_N_30_i_6 54060 12 19 16 14 11 - 0.03361 0.04782 0.03416 0.07452 0.01964 -
portfolio_N_30_i_7 54060 14 17 17 16 11 - 0.03834 0.04254 0.03533 0.08245 0.01989 -
portfolio_N_30_i_8 54060 15 17 16 15 12 - 0.04104 0.04122 0.03423 0.07857 0.02149 -
portfolio_N_30_i_9 54060 14 18 17 14 11 - 0.03878 0.05691 0.03565 0.07451 0.02303 -
portfolio_N_30_i_10 54060 13 18 16 14 11 - 0.04241 0.04240 0.03351 0.07457 0.02012 -
portfolio_N_30_i_11 54060 14 18 17 15 12 - 0.03828 0.05314 0.03513 0.07874 0.02485 -
portfolio_N_30_i_12 54060 14 18 16 16 11 - 0.04465 0.04759 0.03386 0.08244 0.01961 -
portfolio_N_30_i_13 54060 15 21 16 15 12 - 0.04048 0.05437 0.03384 0.07851 0.02141 -
portfolio_N_30_i_14 54060 16 19 17 15 11 - 0.04310 0.05265 0.03532 0.07844 0.01982 -
portfolio_N_30_i_15 54060 14 19 17 15 12 - 0.03834 0.04615 0.03562 0.07836 0.02172 -
portfolio_N_30_i_16 54060 15 17 17 16 14 - 0.04096 0.05073 0.04097 0.08245 0.02876 -
portfolio_N_30_i_17 54060 16 16 15 14 11 - 0.05060 0.04024 0.03210 0.07422 0.02012 -
portfolio_N_30_i_18 54060 15 19 16 15 12 - 0.04045 0.04823 0.03355 0.07831 0.02184 -
portfolio_N_30_i_19 54060 16 18 16 15 11 - 0.04294 0.04589 0.03384 0.07850 0.02013 -
portfolio_N_35_i_0 71820 15 19 18 16 12 - 0.05584 0.05739 0.04426 0.10356 0.02967 -
portfolio_N_35_i_1 71820 15 20 16 15 12 - 0.05662 0.06042 0.04109 0.09836 0.02957 -
portfolio_N_35_i_2 71820 14 18 16 14 11 - 0.05282 0.07135 0.04395 0.10664 0.03179 -
portfolio_N_35_i_3 71820 15 20 17 16 13 - 0.06657 0.07905 0.04376 0.10612 0.03325 -
portfolio_N_35_i_4 71820 16 19 16 15 13 - 0.07068 0.07492 0.04398 0.11254 0.03697 -
portfolio_N_35_i_5 71820 14 19 17 16 13 - 0.06208 0.06215 0.04377 0.11941 0.03192 -
portfolio_N_35_i_6 71820 15 18 16 15 12 - 0.05657 0.05702 0.03915 0.09702 0.02968 -
portfolio_N_35_i_7 71820 17 17 17 16 12 - 0.06331 0.05526 0.04299 0.10258 0.02950 -
portfolio_N_35_i_8 71820 15 20 16 16 13 - 0.05627 0.06575 0.04163 0.10279 0.03188 -
portfolio_N_35_i_9 71820 14 17 18 16 12 - 0.05318 0.06848 0.04473 0.10844 0.03426 -
portfolio_N_35_i_10 71820 15 18 17 17 11 - 0.06658 0.06619 0.04656 0.12444 0.03171 -
portfolio_N_35_i_11 71820 15 20 15 17 12 - 0.06651 0.06644 0.04067 0.10984 0.02958 -
portfolio_N_35_i_12 71820 15 18 16 15 11 - 0.05663 0.06125 0.04135 0.09925 0.02731 -
portfolio_N_35_i_13 71820 15 20 17 15 12 - 0.05644 0.06690 0.04293 0.09958 0.02954 -
portfolio_N_35_i_14 71820 16 20 19 16 13 - 0.05986 0.07504 0.04919 0.10278 0.03688 -
portfolio_N_35_i_15 71820 15 19 16 15 11 - 0.06674 0.06148 0.04025 0.11378 0.02719 -
portfolio_N_35_i_16 71820 14 18 18 15 13 - 0.05262 0.06579 0.04489 0.09922 0.03192 -
portfolio_N_35_i_17 71820 13 17 16 16 13 - 0.04921 0.05822 0.04275 0.10370 0.03170 -
portfolio_N_35_i_18 71820 16 19 18 16 13 - 0.05986 0.06595 0.04472 0.10778 0.03192 -
portfolio_N_35_i_19 71820 16 17 17 15 12 - 0.05994 0.05646 0.04273 0.09732 0.02933 -
Table 7: Iterations and solver runtimes for oscillating masses problems
Iterations Solver Runtime (s)
Problem Size clarabel ecos gurobi mosek qoco qoco_custom cvxgen clarabel ecos gurobi mosek qoco qoco_custom cvxgen
oscillating_masses_N_8_i_0 1344 5 13 9 10 5 5 7 0.00055 0.00128 0.00327 0.00375 0.00029 0.00012 0.00016
oscillating_masses_N_8_i_1 1344 5 14 9 10 5 5 6 0.00052 0.00132 0.00326 0.00377 0.00029 0.00011 0.00013
oscillating_masses_N_8_i_2 1344 5 17 9 11 5 5 6 0.00052 0.00156 0.00329 0.00427 0.00030 0.00012 0.00014
oscillating_masses_N_8_i_3 1344 5 14 9 11 5 5 6 0.00052 0.00130 0.00327 0.00422 0.00030 0.00011 0.00013
oscillating_masses_N_8_i_4 1344 5 12 10 11 5 5 7 0.00053 0.00114 0.00338 0.00434 0.00030 0.00011 0.00015
oscillating_masses_N_8_i_5 1344 5 12 9 10 5 5 6 0.00052 0.00113 0.00326 0.00381 0.00030 0.00012 0.00014
oscillating_masses_N_8_i_6 1344 10 13 10 12 7 7 10 0.00093 0.00105 0.00304 0.00426 0.00034 0.00016 0.00022
oscillating_masses_N_8_i_7 1344 5 13 12 11 5 5 6 0.00047 0.00119 0.00357 0.00400 0.00030 0.00012 0.00014
oscillating_masses_N_8_i_8 1344 5 14 8 11 4 4 6 0.00052 0.00125 0.00316 0.00404 0.00026 0.00010 0.00013
oscillating_masses_N_8_i_9 1344 5 15 9 11 5 5 6 0.00052 0.00133 0.00326 0.00407 0.00030 0.00011 0.00013
oscillating_masses_N_8_i_10 1344 5 14 8 13 5 5 6 0.00052 0.00125 0.00317 0.00472 0.00030 0.00012 0.00014
oscillating_masses_N_8_i_11 1344 5 13 10 12 5 5 6 0.00052 0.00119 0.00337 0.00426 0.00030 0.00011 0.00013
oscillating_masses_N_8_i_12 1344 5 14 9 12 5 5 7 0.00052 0.00130 0.00332 0.00453 0.00030 0.00012 0.00016
oscillating_masses_N_8_i_13 1344 5 12 9 10 5 5 6 0.00052 0.00113 0.00330 0.00375 0.00030 0.00012 0.00013
oscillating_masses_N_8_i_14 1344 5 12 9 10 5 5 6 0.00052 0.00095 0.00284 0.00375 0.00026 0.00011 0.00012
oscillating_masses_N_8_i_15 1344 5 13 9 10 5 5 7 0.00047 0.00105 0.00284 0.00353 0.00025 0.00011 0.00015
oscillating_masses_N_8_i_16 1344 6 14 10 11 6 6 8 0.00055 0.00115 0.00293 0.00379 0.00029 0.00013 0.00017
oscillating_masses_N_8_i_17 1344 5 13 8 11 5 5 6 0.00047 0.00102 0.00276 0.00365 0.00026 0.00012 0.00013
oscillating_masses_N_8_i_18 1344 5 12 9 10 4 4 6 0.00047 0.00096 0.00286 0.00334 0.00023 0.00009 0.00012
oscillating_masses_N_8_i_19 1344 5 12 9 11 5 5 6 0.00048 0.00097 0.00288 0.00379 0.00026 0.00011 0.00012
oscillating_masses_N_20_i_0 3312 5 15 9 12 5 5 7 0.00111 0.00287 0.00466 0.01414 0.00060 0.00028 0.00040
oscillating_masses_N_20_i_1 3312 5 14 10 13 5 5 7 0.00113 0.00263 0.00488 0.01447 0.00060 0.00028 0.00038
oscillating_masses_N_20_i_2 3312 5 15 10 13 5 5 7 0.00112 0.00280 0.00497 0.01484 0.00061 0.00029 0.00040
oscillating_masses_N_20_i_3 3312 7 14 12 13 7 7 8 0.00141 0.00265 0.00529 0.01475 0.00077 0.00038 0.00045
oscillating_masses_N_20_i_4 3312 7 14 9 14 7 7 9 0.00146 0.00266 0.00462 0.01572 0.00077 0.00039 0.00051
oscillating_masses_N_20_i_5 3312 5 14 9 12 5 5 6 0.00109 0.00262 0.00462 0.01343 0.00061 0.00028 0.00034
oscillating_masses_N_20_i_6 3312 5 16 10 12 5 5 7 0.00110 0.00296 0.00492 0.01387 0.00061 0.00028 0.00040
oscillating_masses_N_20_i_7 3312 5 14 9 13 5 5 6 0.00109 0.00270 0.00463 0.01458 0.00060 0.00029 0.00034
oscillating_masses_N_20_i_8 3312 5 13 9 13 5 5 7 0.00111 0.00255 0.00467 0.01429 0.00062 0.00031 0.00039
oscillating_masses_N_20_i_9 3312 5 16 10 13 5 5 6 0.00109 0.00297 0.00492 0.01376 0.00064 0.00027 0.00034
oscillating_masses_N_20_i_10 3312 5 14 9 13 5 5 6 0.00110 0.00269 0.00463 0.01413 0.00062 0.00028 0.00035
oscillating_masses_N_20_i_11 3312 5 14 9 13 5 5 7 0.00111 0.00268 0.00467 0.01373 0.00061 0.00031 0.00040
oscillating_masses_N_20_i_12 3312 5 15 9 13 5 5 7 0.00112 0.00276 0.00465 0.01485 0.00062 0.00031 0.00040
oscillating_masses_N_20_i_13 3312 5 14 9 12 5 5 6 0.00108 0.00262 0.00464 0.01416 0.00062 0.00027 0.00035
oscillating_masses_N_20_i_14 3312 5 15 9 13 5 5 6 0.00109 0.00280 0.00465 0.01435 0.00061 0.00029 0.00038
oscillating_masses_N_20_i_15 3312 5 13 9 12 5 5 6 0.00110 0.00253 0.00469 0.01373 0.00062 0.00029 0.00034
oscillating_masses_N_20_i_16 3312 5 14 9 14 5 5 6 0.00109 0.00270 0.00460 0.01496 0.00061 0.00029 0.00035
oscillating_masses_N_20_i_17 3312 5 15 10 13 5 5 7 0.00116 0.00286 0.00488 0.01485 0.00061 0.00029 0.00040
oscillating_masses_N_20_i_18 3312 5 15 9 14 5 5 6 0.00111 0.00279 0.00467 0.01549 0.00061 0.00029 0.00035
oscillating_masses_N_20_i_19 3312 5 15 9 13 5 5 7 0.00110 0.00288 0.00464 0.01489 0.00063 0.00028 0.00040
oscillating_masses_N_32_i_0 5280 5 14 10 14 5 5 - 0.00172 0.00422 0.00671 0.02045 0.00097 0.00053 -
oscillating_masses_N_32_i_1 5280 5 14 9 13 5 5 - 0.00181 0.00420 0.00629 0.01951 0.00096 0.00048 -
oscillating_masses_N_32_i_2 5280 5 15 9 12 5 5 - 0.00173 0.00442 0.00628 0.01808 0.00101 0.00047 -
oscillating_masses_N_32_i_3 5280 7 13 11 12 6 6 - 0.00219 0.00385 0.00695 0.01738 0.00109 0.00056 -
oscillating_masses_N_32_i_4 5280 9 15 10 14 8 8 - 0.00281 0.00461 0.00681 0.02077 0.00137 0.00073 -
oscillating_masses_N_32_i_5 5280 5 15 9 11 5 5 - 0.00172 0.00446 0.00631 0.01756 0.00098 0.00048 -
oscillating_masses_N_32_i_6 5280 6 14 10 11 6 6 - 0.00194 0.00415 0.00677 0.01658 0.00110 0.00056 -
oscillating_masses_N_32_i_7 5280 6 15 11 12 6 6 - 0.00196 0.00443 0.00702 0.01703 0.00110 0.00058 -
oscillating_masses_N_32_i_8 5280 5 16 9 13 5 5 - 0.00176 0.00477 0.00631 0.01762 0.00099 0.00048 -
oscillating_masses_N_32_i_9 5280 5 14 9 12 5 5 - 0.00171 0.00414 0.00630 0.01706 0.00099 0.00048 -
oscillating_masses_N_32_i_10 5280 6 15 10 12 5 5 - 0.00193 0.00448 0.00678 0.01870 0.00097 0.00048 -
oscillating_masses_N_32_i_11 5280 5 13 9 12 5 5 - 0.00172 0.00391 0.00622 0.01769 0.00102 0.00048 -
oscillating_masses_N_32_i_12 5280 5 15 9 11 5 5 - 0.00170 0.00438 0.00631 0.01586 0.00099 0.00049 -
oscillating_masses_N_32_i_13 5280 5 15 9 14 5 5 - 0.00173 0.00454 0.00632 0.01974 0.00097 0.00048 -
oscillating_masses_N_32_i_14 5280 6 16 10 11 6 6 - 0.00196 0.00470 0.00678 0.01605 0.00111 0.00057 -
oscillating_masses_N_32_i_15 5280 5 16 9 13 5 5 - 0.00173 0.00480 0.00629 0.01822 0.00098 0.00053 -
oscillating_masses_N_32_i_16 5280 5 15 9 13 5 5 - 0.00174 0.00444 0.00622 0.01941 0.00096 0.00048 -
oscillating_masses_N_32_i_17 5280 5 15 10 12 5 5 - 0.00171 0.00446 0.00665 0.01877 0.00097 0.00048 -
oscillating_masses_N_32_i_18 5280 5 15 9 13 5 5 - 0.00170 0.00459 0.00731 0.01839 0.00115 0.00048 -
oscillating_masses_N_32_i_19 5280 5 15 10 13 5 5 - 0.00173 0.00454 0.00670 0.01873 0.00100 0.00048 -
oscillating_masses_N_44_i_0 7248 5 15 9 12 5 5 - 0.00230 0.00611 0.00789 0.02372 0.00134 0.00065 -
oscillating_masses_N_44_i_1 7248 6 16 10 12 6 6 - 0.00266 0.00640 0.00840 0.02427 0.00153 0.00080 -
oscillating_masses_N_44_i_2 7248 6 15 11 13 6 6 - 0.00272 0.00604 0.00884 0.02677 0.00156 0.00080 -
oscillating_masses_N_44_i_3 7248 5 14 9 13 5 5 - 0.00240 0.00599 0.00783 0.02612 0.00134 0.00068 -
oscillating_masses_N_44_i_4 7248 5 16 9 13 5 5 - 0.00235 0.00668 0.00783 0.02650 0.00137 0.00069 -
oscillating_masses_N_44_i_5 7248 5 14 9 12 5 5 - 0.00235 0.00573 0.00780 0.02321 0.00136 0.00068 -
oscillating_masses_N_44_i_6 7248 6 15 11 13 6 6 - 0.00278 0.00623 0.00880 0.02700 0.00154 0.00080 -
oscillating_masses_N_44_i_7 7248 5 15 9 12 5 5 - 0.00235 0.00624 0.00781 0.02492 0.00132 0.00066 -
oscillating_masses_N_44_i_8 7248 5 15 9 13 5 5 - 0.00238 0.00622 0.00785 0.02579 0.00133 0.00069 -
oscillating_masses_N_44_i_9 7248 5 15 9 12 5 5 - 0.00239 0.00604 0.00778 0.02401 0.00134 0.00069 -
oscillating_masses_N_44_i_10 7248 5 15 9 13 5 5 - 0.00246 0.00608 0.00775 0.02662 0.00134 0.00069 -
oscillating_masses_N_44_i_11 7248 5 14 9 12 5 5 - 0.00236 0.00586 0.00777 0.02545 0.00133 0.00068 -
oscillating_masses_N_44_i_12 7248 5 16 9 13 5 5 - 0.00234 0.00642 0.00786 0.02667 0.00134 0.00068 -
oscillating_masses_N_44_i_13 7248 5 14 9 12 5 5 - 0.00235 0.00589 0.00779 0.02546 0.00133 0.00076 -
oscillating_masses_N_44_i_14 7248 5 16 9 13 5 5 - 0.00240 0.00668 0.00790 0.02632 0.00134 0.00068 -
oscillating_masses_N_44_i_15 7248 5 14 9 14 5 5 - 0.00237 0.00578 0.00781 0.02633 0.00135 0.00069 -
oscillating_masses_N_44_i_16 7248 5 16 9 13 5 5 - 0.00237 0.00647 0.00781 0.02624 0.00135 0.00068 -
oscillating_masses_N_44_i_17 7248 5 15 9 13 5 5 - 0.00237 0.00618 0.00781 0.02594 0.00135 0.00070 -
oscillating_masses_N_44_i_18 7248 5 14 9 13 5 5 - 0.00238 0.00576 0.00776 0.02651 0.00133 0.00069 -
oscillating_masses_N_44_i_19 7248 5 14 9 13 5 5 - 0.00238 0.00584 0.00783 0.02656 0.00134 0.00068 -
oscillating_masses_N_56_i_0 9216 5 15 9 14 5 5 - 0.00304 0.00893 0.00976 0.03563 0.00197 0.00101 -
oscillating_masses_N_56_i_1 9216 5 16 9 12 5 5 - 0.00340 0.00968 0.00980 0.02983 0.00201 0.00101 -
oscillating_masses_N_56_i_2 9216 5 16 9 12 5 5 - 0.00341 0.00834 0.00839 0.03087 0.00171 0.00092 -
oscillating_masses_N_56_i_3 9216 5 15 9 13 5 5 - 0.00306 0.00785 0.00836 0.03244 0.00178 0.00092 -
oscillating_masses_N_56_i_4 9216 5 15 9 13 5 5 - 0.00351 0.00900 0.00832 0.03240 0.00198 0.00101 -
oscillating_masses_N_56_i_5 9216 7 16 11 13 6 6 - 0.00401 0.00826 0.00938 0.03206 0.00194 0.00107 -
oscillating_masses_N_56_i_6 9216 5 14 9 12 5 5 - 0.00304 0.00734 0.00846 0.03137 0.00173 0.00090 -
oscillating_masses_N_56_i_7 9216 5 19 10 14 5 5 - 0.00298 0.00972 0.00894 0.03386 0.00170 0.00091 -
oscillating_masses_N_56_i_8 9216 5 14 9 12 5 5 - 0.00300 0.00740 0.00838 0.02979 0.00175 0.00091 -
oscillating_masses_N_56_i_9 9216 6 16 9 13 6 6 - 0.00340 0.00822 0.00859 0.03321 0.00194 0.00105 -
oscillating_masses_N_56_i_10 9216 5 19 9 12 5 5 - 0.00301 0.00978 0.00830 0.03279 0.00171 0.00091 -
oscillating_masses_N_56_i_11 9216 5 14 9 12 5 5 - 0.00301 0.00738 0.00836 0.03012 0.00172 0.00089 -
oscillating_masses_N_56_i_12 9216 5 15 9 12 5 5 - 0.00299 0.00775 0.00846 0.03024 0.00172 0.00092 -
oscillating_masses_N_56_i_13 9216 5 14 9 12 5 5 - 0.00299 0.00730 0.00832 0.03158 0.00170 0.00090 -
oscillating_masses_N_56_i_14 9216 5 15 9 13 5 5 - 0.00298 0.00762 0.00838 0.03245 0.00172 0.00090 -
oscillating_masses_N_56_i_15 9216 6 14 11 13 6 6 - 0.00351 0.00844 0.01098 0.03534 0.00228 0.00119 -
oscillating_masses_N_56_i_16 9216 5 16 9 13 5 5 - 0.00353 0.00958 0.00973 0.03549 0.00200 0.00101 -
oscillating_masses_N_56_i_17 9216 6 15 10 14 6 6 - 0.00393 0.00903 0.01064 0.03770 0.00225 0.00120 -
oscillating_masses_N_56_i_18 9216 5 16 9 13 5 5 - 0.00351 0.00846 0.00840 0.03598 0.00170 0.00090 -
oscillating_masses_N_56_i_19 9216 7 16 10 16 7 7 - 0.00404 0.00836 0.00929 0.03900 0.00224 0.00121 -
oscillating_masses_N_76_i_0 12496 5 16 9 12 5 - - 0.00404 0.01138 0.01050 0.04013 0.00236 - -
oscillating_masses_N_76_i_1 12496 6 15 10 13 5 - - 0.00479 0.01055 0.01141 0.04462 0.00233 - -
oscillating_masses_N_76_i_2 12496 5 15 9 13 5 - - 0.00409 0.01073 0.01062 0.04632 0.00236 - -
oscillating_masses_N_76_i_3 12496 5 14 9 13 6 - - 0.00409 0.00973 0.01075 0.04351 0.00268 - -
oscillating_masses_N_76_i_4 12496 5 14 9 12 5 - - 0.00406 0.01169 0.01070 0.04310 0.00235 - -
oscillating_masses_N_76_i_5 12496 5 15 9 13 5 - - 0.00476 0.01237 0.01246 0.04689 0.00272 - -
oscillating_masses_N_76_i_6 12496 5 15 9 13 5 - - 0.00472 0.01268 0.01276 0.04681 0.00270 - -
oscillating_masses_N_76_i_7 12496 5 15 9 14 5 - - 0.00471 0.01057 0.01069 0.04750 0.00236 - -
oscillating_masses_N_76_i_8 12496 6 14 10 12 6 - - 0.00469 0.00984 0.01144 0.04088 0.00266 - -
oscillating_masses_N_76_i_9 12496 5 14 10 14 5 - - 0.00415 0.00987 0.01140 0.04779 0.00235 - -
oscillating_masses_N_76_i_10 12496 6 15 10 12 6 - - 0.00454 0.01066 0.01137 0.04373 0.00266 - -
oscillating_masses_N_76_i_11 12496 5 14 9 12 5 - - 0.00412 0.00997 0.01064 0.04083 0.00234 - -
oscillating_masses_N_76_i_12 12496 7 15 10 14 7 - - 0.00533 0.01259 0.01348 0.04971 0.00346 - -
oscillating_masses_N_76_i_13 12496 7 17 12 14 6 - - 0.00618 0.01371 0.01473 0.04861 0.00310 - -
oscillating_masses_N_76_i_14 12496 5 15 9 14 5 - - 0.00477 0.01248 0.01232 0.05148 0.00271 - -
oscillating_masses_N_76_i_15 12496 5 16 9 13 5 - - 0.00476 0.01134 0.01066 0.04689 0.00232 - -
oscillating_masses_N_76_i_16 12496 5 15 9 11 5 - - 0.00433 0.01051 0.01066 0.03930 0.00235 - -
oscillating_masses_N_76_i_17 12496 5 15 9 13 5 - - 0.00411 0.01072 0.01072 0.04533 0.00236 - -
oscillating_masses_N_76_i_18 12496 5 16 9 13 5 - - 0.00406 0.01292 0.01258 0.04323 0.00273 - -
oscillating_masses_N_76_i_19 12496 5 15 9 12 5 - - 0.00471 0.01225 0.01249 0.04638 0.00272 - -
oscillating_masses_N_96_i_0 15776 5 15 9 13 5 - - 0.00602 0.01550 0.01450 0.05848 0.00348 - -
oscillating_masses_N_96_i_1 15776 6 16 11 13 6 - - 0.00698 0.01420 0.01402 0.05978 0.00339 - -
oscillating_masses_N_96_i_2 15776 5 16 9 13 5 - - 0.00511 0.01426 0.01238 0.05680 $0$
oscillating_masses_N_96_i_3 15776 5 17 9 13 5 - - 0.00512 0.01756 0.01263 0.05612 0.00347 - -
oscillating_masses_N_96_i_4 15776 5 13 13 12 5 - - 0.00581 0.01338 0.01845 0.05332 0.00346 - -
oscillating_masses_N_96_i_5 15776 5 15 9 15 5 - - 0.00603 0.01337 0.01244 0.07083 0.00305 - -
oscillating_masses_N_96_i_6 15776 5 14 9 13 5 - - 0.00512 0.01271 0.01272 0.05773 0.00301 - -
oscillating_masses_N_96_i_7 15776 5 15 9 13 5 - - 0.00520 0.01367 0.01246 0.05746 0.00308 - -
oscillating_masses_N_96_i_8 15776 5 16 9 13 5 - - 0.00523 0.01456 0.01244 0.05736 $0$
oscillating_masses_N_96_i_9 15776 5 15 9 11 5 - - 0.00513 0.01334 0.01273 0.05200 $0$
oscillating_masses_N_96_i_10 15776 6 14 10 14 6 - - 0.00571 0.01237 0.01338 0.05674 0.00340 - -
oscillating_masses_N_96_i_11 15776 6 15 10 12 6 - - 0.00590 0.01362 0.01320 0.05358 0.00341 - -
oscillating_masses_N_96_i_12 15776 5 17 9 13 5 - - 0.00512 0.01509 0.01248 0.05466 0.00298 - -
oscillating_masses_N_96_i_13 15776 5 15 9 14 5 - - 0.00520 0.01339 0.01243 0.05965 $0$
oscillating_masses_N_96_i_14 15776 5 17 9 13 5 - - 0.00518 0.01539 0.01246 0.05609 0.00302 - -
oscillating_masses_N_96_i_15 15776 6 15 11 12 6 - - 0.00584 0.01363 0.01427 0.05236 0.00339 - -
oscillating_masses_N_96_i_16 15776 5 14 10 14 5 - - 0.00528 0.01269 0.01350 0.05957 0.00299 - -
oscillating_masses_N_96_i_17 15776 5 16 9 15 5 - - 0.00510 0.01405 0.01259 0.06245 0.00307 - -
oscillating_masses_N_96_i_18 15776 5 14 9 12 5 - - 0.00504 0.01257 0.01469 0.05114 0.00347 - -
oscillating_masses_N_96_i_19 15776 5 16 9 13 5 - - 0.00514 0.01434 0.01248 0.05771 0.00298 - -
oscillating_masses_N_116_i_0 19056 5 15 9 13 5 - - 0.00627 0.01601 0.01484 0.06574 0.00368 - -
oscillating_masses_N_116_i_1 19056 6 14 10 12 6 - - 0.00695 0.01539 0.01599 0.06194 0.00419 - -
oscillating_masses_N_116_i_2 19056 5 15 9 12 5 - - 0.00637 0.01643 0.01469 0.06273 0.00367 - -
oscillating_masses_N_116_i_3 19056 5 16 9 13 5 - - 0.00622 0.01964 0.01696 0.06735 0.00422 - -
oscillating_masses_N_116_i_4 19056 5 15 10 12 5 - - 0.00715 0.01912 0.01840 0.06859 0.00423 - -
oscillating_masses_N_116_i_5 19056 6 16 10 12 6 - - 0.00811 0.01714 0.01554 0.07024 0.00412 - -
oscillating_masses_N_116_i_6 19056 6 14 10 10 6 - - 0.00646 0.01515 0.01561 0.05609 0.00416 - -
oscillating_masses_N_116_i_7 19056 5 15 9 13 5 - - 0.00638 0.01621 0.01461 0.06767 0.00369 - -
oscillating_masses_N_116_i_8 19056 7 17 11 14 7 - - 0.00841 0.01864 0.01665 0.07346 0.00465 - -
oscillating_masses_N_116_i_9 19056 5 15 9 14 5 - - 0.00626 0.01642 0.01470 0.07328 0.00373 - -
oscillating_masses_N_116_i_10 19056 5 14 10 13 5 - - 0.00623 0.01727 0.01818 0.06714 0.00425 - -
oscillating_masses_N_116_i_11 19056 5 16 9 13 5 - - 0.00700 0.01690 0.01750 0.06795 0.00426 - -
oscillating_masses_N_116_i_12 19056 5 16 9 12 5 - - 0.00608 0.01737 0.01494 0.06339 0.00366 - -
oscillating_masses_N_116_i_13 19056 5 15 9 13 6 - - 0.00619 0.01623 0.01471 0.07116 0.00419 - -
oscillating_masses_N_116_i_14 19056 6 15 10 13 6 - - 0.00718 0.01647 0.01562 0.07111 0.00415 - -
oscillating_masses_N_116_i_15 19056 5 16 9 13 5 - - 0.00619 0.01733 0.01465 0.07095 0.00366 - -
oscillating_masses_N_116_i_16 19056 5 14 9 14 5 - - 0.00620 0.01531 0.01466 0.07134 0.00363 - -
oscillating_masses_N_116_i_17 19056 5 16 9 14 5 - - 0.00627 0.02028 0.01707 0.07243 0.00424 - -
oscillating_masses_N_116_i_18 19056 5 15 10 12 6 - - 0.00719 0.01678 0.01573 0.07083 0.00417 - -
oscillating_masses_N_116_i_19 19056 5 16 9 14 5 - - 0.00635 0.01737 0.01461 0.06908 0.00362 - -
oscillating_masses_N_136_i_0 22336 5 16 9 13 5 - - 0.00721 0.02032 0.01688 0.08071 0.00432 - -
oscillating_masses_N_136_i_1 22336 5 15 10 12 5 - - 0.00733 0.01903 0.01791 0.07587 0.00430 - -
oscillating_masses_N_136_i_2 22336 5 14 9 12 5 - - 0.00743 0.01827 0.01725 0.07667 0.00436 - -
oscillating_masses_N_136_i_3 22336 5 16 10 13 5 - - 0.00734 0.02020 0.01796 0.08116 0.00437 - -
oscillating_masses_N_136_i_4 22336 5 15 9 13 5 - - 0.00729 0.01963 0.01699 0.07803 0.00438 - -
oscillating_masses_N_136_i_5 22336 5 16 10 12 5 - - 0.00737 0.02324 0.02100 0.07585 0.00503 - -
oscillating_masses_N_136_i_6 22336 5 15 10 14 5 - - 0.00860 0.02233 0.02066 0.09022 0.00504 - -
oscillating_masses_N_136_i_7 22336 5 15 9 14 5 - - 0.00839 0.01910 0.01967 0.09066 0.00509 - -
oscillating_masses_N_136_i_8 22336 5 15 9 13 5 - - 0.00723 0.01913 0.01691 0.07890 0.00433 - -
oscillating_masses_N_136_i_9 22336 5 14 9 12 5 - - 0.00725 0.01794 0.01715 0.07905 0.00436 - -
oscillating_masses_N_136_i_10 22336 6 14 10 13 6 - - 0.00821 0.01806 0.01861 0.07904 $0$
oscillating_masses_N_136_i_11 22336 6 15 10 13 6 - - 0.00817 0.01932 0.01792 0.07561 0.00504 - -
oscillating_masses_N_136_i_12 22336 5 15 9 13 5 - - 0.00727 0.01911 0.01712 0.07888 0.00435 - -
oscillating_masses_N_136_i_13 22336 5 15 9 14 5 - - 0.00751 0.01968 0.01682 0.08226 0.00433 - -
oscillating_masses_N_136_i_14 22336 5 16 9 14 5 - - 0.00726 0.02032 0.01678 0.08191 0.00440 - -
oscillating_masses_N_136_i_15 22336 6 15 9 12 6 - - 0.00809 0.01898 0.01684 0.07451 0.00493 - -
oscillating_masses_N_136_i_16 22336 6 16 10 12 6 - - 0.00810 0.01970 0.01790 0.07366 0.00493 - -
oscillating_masses_N_136_i_17 22336 6 16 10 12 5 - - 0.00806 0.01983 0.01801 0.07859 0.00436 - -
oscillating_masses_N_136_i_18 22336 5 17 9 13 5 - - 0.00729 0.02159 0.01689 0.08000 0.00432 - -
oscillating_masses_N_136_i_19 22336 6 16 9 12 6 - - 0.00810 0.01987 0.01685 0.07215 0.00495 - -
oscillating_masses_N_156_i_0 25616 6 15 10 14 7 - - 0.00975 0.02223 0.02161 0.09490 0.00636 - -
oscillating_masses_N_156_i_1 25616 6 14 10 13 6 - - 0.00928 0.02040 0.02191 0.09083 0.00581 - -
oscillating_masses_N_156_i_2 25616 7 20 12 14 5 - - 0.01080 0.02879 0.02462 0.09866 0.00515 - -
oscillating_masses_N_156_i_3 25616 5 16 9 14 5 - - 0.00833 0.02689 0.02182 0.09493 0.00589 - -
oscillating_masses_N_156_i_4 25616 6 16 10 13 6 - - 0.01073 0.02285 0.02305 0.09158 0.00571 - -
oscillating_masses_N_156_i_5 25616 5 16 9 12 5 - - 0.00830 0.02309 0.02070 0.08317 0.00504 - -
oscillating_masses_N_156_i_6 25616 5 15 9 14 5 - - 0.00851 0.02211 0.02133 0.09482 $0$
oscillating_masses_N_156_i_7 25616 5 16 9 13 5 - - 0.00815 0.02310 0.02083 0.08544 0.00506 - -
oscillating_masses_N_156_i_8 25616 5 15 9 13 5 - - 0.00841 0.02230 0.02131 0.09227 0.00514 - -
oscillating_masses_N_156_i_9 25616 5 15 9 14 5 - - 0.00839 0.02164 0.02064 0.09737 0.00506 - -
oscillating_masses_N_156_i_10 25616 6 15 10 13 6 - - 0.00960 0.02565 0.02176 0.09410 0.00671 - -
oscillating_masses_N_156_i_11 25616 5 17 9 13 5 - - 0.00818 0.02474 0.02085 0.08873 0.00504 - -
oscillating_masses_N_156_i_12 25616 5 15 9 13 5 - - 0.00840 0.02249 0.02040 0.09105 0.00503 - -
oscillating_masses_N_156_i_13 25616 5 14 9 12 5 - - 0.00841 0.02069 0.02099 0.08307 0.00514 - -
oscillating_masses_N_156_i_14 25616 6 16 9 12 6 - - 0.00932 0.02232 0.02112 0.08811 0.00664 - -
oscillating_masses_N_156_i_15 25616 5 14 9 14 5 - - 0.00832 0.02064 0.02225 0.09329 0.00589 - -
oscillating_masses_N_156_i_16 25616 6 13 11 11 6 - - 0.00927 0.01946 0.02327 0.07696 0.00576 - -
oscillating_masses_N_156_i_17 25616 5 17 9 13 5 - - 0.00838 0.02463 0.02070 0.09187 0.00505 - -
oscillating_masses_N_156_i_18 25616 7 16 13 13 7 - - 0.01111 0.02653 0.02615 0.08826 0.00740 - -
oscillating_masses_N_156_i_19 25616 5 15 9 13 5 - - 0.00985 0.02505 0.02235 0.09991 0.00586 - -

### v3

@@
 We consider the quadratic objective second-order cone program (SOCP)
-[…]
+$\underset{x}{\text{minimize}}$ $\frac{1}{2}x^{\top}Px+c^{\top}x$ (1)
 subject to $Gx\preceq_{\mathcal{K}}h$
@@
 We restrict $\mathcal{K}$ to be the Cartesian product
-$\mathcal{K}=\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K},$
+$\mathcal{K}:=\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K},$
 where $\mathcal{C}_{i}$ is either the non-negative orthant or a second-order cone.
@@
 Although sparse linear algebra allows the aforementioned solvers to solve SOCPs regardless of their sparsity structures, it can be quite slow due to the extra overhead of determining where the nonzero elements are before performing floating point operations, which leads to more CPU instructions being issued, more memory accesses, and more cache misses.
-However, when the sparsity structure of the problem matrices is known beforehand, it is possible to generate a custom solver that implements specialized linear algebra routines tailored to the sparsity structure of the problem.
-These custom solvers hardcode the exact floating point operations necessary, eliminating the overhead incurred by sparse linear algebra routines, resulting in significantly faster solves.
+However, when the sparsity structure of the problem matrices is known beforehand, it is possible to generate a custom solver that implements customized linear algebra routines tailored to the sparsity structure of the problem.
+Specifically, in sparse linear algebra matrices are stored as three arrays: a row/column index array, column/row pointer array, and data array.
+Before performing a floating point operation on a nonzero element in the matrix, sparse linear algebra routines must first index into the pointer and index arrays to determine the location of the element in the data array.
+In customized linear algebra, the index and pointer arrays do not exist, and we directly index into the data array, eliminating the overhead of sparse linear algebra, resulting in significantly faster operations.
+We discuss this further in Section 4.1.
 A custom solver generator is a tool that takes the sparsity structure of an optimization problem as an input and outputs a solver (typically in C) optimized for that specific sparsity structure without relying on sparse linear algebra.
@@
 Later, cvxgen was used by SpaceX for landing their Falcon 9 boosters [40].
+A seemingly related but distinct tool is cvxpygen [41].
+It leverages cvxpy to parse the optimization problem to the standard form for solvers and generates C code for updating problem data and retrieving solutions.
+cvxpygen then wraps a backend solver without (or with very little) modification to its source code.
+Because it customizes only the parsing and solution retrieval while relying entirely on an existing solver, we do not classify cvxpygen as a custom solver generator.
 1.1 Contribution
-This work makes two key contributions: (1) the development of qoco, an open-source C-based solver for quadratic objective SOCPs, and (2) the introduction of qocogen, an open-source custom solver generator for quadratic objective SOCPs which generates a custom solver written in C.
+This work makes two key contributions: (1) the development of qoco, an open-source C-based solver for quadratic objective SOCPs, and (2) the introduction of qocogen, an open-source custom solver generator for quadratic objective SOCPs which generates a custom solver written in C  [^3] Both are available on GitHub at https://github.com/qoco-org.
 We demonstrate that qoco outperforms most commercial and open-source solvers in terms of both speed and robustness, and qococustom (the name of solvers generated by qocogen) is significantly faster than qoco.
@@
 After solving the linear system, we apply iterative refinement to ensure that the original, unregularized system was solved to high accuracy, further enhancing numerical stability and robustness.
-qoco and qocogen are available on GitHub at https://github.com/qoco-org.
 1.2 Outline
@@
 1.3 Notation
-$(a,b)$ refers to the concatenation of $a\in\mathbb{R}^{a}$ and $b\in\mathbb{R}^{b}$ into a vector in $\mathbb{R}^{a+b}$
-The $l$-dimensional non-negative orthant is given by +$\mathbb{R}^{l}_{+}=\{u\in\mathbb{R}^{l}\;|\;u_{i}\geq 0\}$
-For a vector $u$ in the $l$-dimensional non-negative orthant, $u_{i}$ refers to the $i^{th}$ element of the vector
-The $q$-dimensional second-order cone is given by $\mathcal{Q}^{q}=\{(u_{0},u_{1})\in\mathbb{R}\times\mathbb{R}^{q-1}\;|\;\|u_{1}\|_{2}\leq u_{0}\}$
-For a vector $u$ in the $q$-dimensional second-order cone, we partition it as $u=(u_{0},u_{1})$, i.e. $u_{0}\in\mathbb{R}$, $u_{1}\in\mathbb{R}^{q-1}$ and $\|u_{1}\|_{2}\leq u_{0}$.
-When $x\in\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K}$, $x_{i}$ refers to the $i^{th}$ subvector which is in cone $\mathcal{C}_{i}$
+We use $(a,b)$ to denote the concatenation of vectors $a\in\mathbb{R}^{a}$ and $b\in\mathbb{R}^{b}$, resulting in a vector in $\mathbb{R}^{a+b}$.
+The non-negative orthant in $\mathbb{R}^{l}$ is denoted by $\mathbb{R}^{l}_{+}=\{u\in\mathbb{R}^{l}\mid u_{i}\geq 0\}$, and for any $u\in\mathbb{R}^{l}_{+}$, we use $u_{i}$ to refer to its $i^{\text{th}}$ element.
+The $q$-dimensional second-order cone is defined as $\mathcal{Q}^{q}=\{(u_{0},u_{1})\in\mathbb{R}\times\mathbb{R}^{q-1}\mid\|u_{1}\|_{2}\leq u_{0}\}$, where any $u\in\mathcal{Q}^{q}$ is written as $u=(u_{0},u_{1})$ with $u_{0}\in\mathbb{R}$ and $u_{1}\in\mathbb{R}^{q-1}$.
+Finally, when a vector $x$ lies in the Cartesian product of cones, i.e., $x\in\mathcal{C}_{1}\times\mathcal{C}_{2}\times\cdots\times\mathcal{C}_{K}$, we denote $x_{i}$ as its $i^{\text{th}}$ subvector belonging to cone $\mathcal{C}_{i}$.
 2 Primal-dual interior point method
@@
 Problem 1 can be equivalently rewritten as
-[…]
+$\underset{x}{\text{minimize}}$ $\frac{1}{2}x^{\top}Px+c^{\top}x+\mathcal{I}_{\mathcal{K}}(h-Gx)$
 subject to $Ax=b,$
@@
 After replacing the indicator function with the barrier function, we obtain
-[…]
+$\underset{x}{\text{minimize}}$ $\frac{1}{2}x^{\top}Px+c^{\top}x+\tau\sum_{i=1}^{K}\phi_{i}(h_{i}-G_{i}x)$ (4)
 subject to $Ax=b,$
@@
 $h-Gx$ $\in\mathcal{K}.$
-If we define $s=h-Gx$ and $z_{i}=-\tau\nabla\phi_{i}(h_{i}-G_{i}x)$ for $i=1,\ldots K$, we can rewrite the KKT conditions as
+If we define s=hGxs=h-Gx and $z_{i}=-\tau\nabla\phi_{i}(h_{i}-G_{i}x)$ for $i=1,\ldots K$, we can rewrite the KKT conditions as
 $Px+c+A^{\top}y+G^{\top}z$ $=0$ (5a)
@@
 $(s,z)$ $\in\mathcal{K}\times\mathcal{K},$ (5e)
-where the condition $z\in\mathcal{K}$ arises because if $u\in\mathcal{K}$, then $-\nabla\phi(u)\in\mathcal{K}$ [47].
+where the condition $z\in\mathcal{K}$ arises because if $u\in\mathcal{K}$, then $\nabla\phi(u)\in\mathcal{K}$ [47].
 It is desirable to write Equation (5d), in a form where $s_{i}$ and $z_{i}$ appear symmetrically.
@@
 $\Delta s$ $=-r_{z}-G\Delta x,$ (19b)
-[…]
+where the operator $\backslash$ represents the inverse of the Jordan product $\circ$, meaning that $x\backslash(x\circ y)=y$.
 To account for the higher order term we ignored when forming Equation (17), we apply Mehrotra’s predictor-corrector [44].
@@
 Since taking a full step along the affine scaling direction can violate the constraint $(s,z)\in\mathcal{K}\times\mathcal{K}$, we compute the centering parameter $\sigma\in[0,1]$ as
-$\alpha$ $=\sup\{\alpha\in[0,1]\;|\;(s_{k},z_{k})+\alpha(\Delta s_{a},\Delta z_{a})\in\mathcal{K}\times\mathcal{K}\}$
-$\rho$ $=\frac{(s_{k}+\alpha\Delta s_{a})^{\top}(z_{k}+\alpha\Delta z_{a})}{s_{k}^{\top}z_{k}}$
-$\sigma$ $=\max\{0,\min\{1,\rho\}^{3}\}.$
+$\begin{split}\alpha&=\sup\{\alpha\in[0,1]\;|\;(s_{k},z_{k})+\alpha(\Delta s_{a},\Delta z_{a})\in\mathcal{K}\times\mathcal{K}\}\\ \rho&=\frac{(s_{k}+\alpha\Delta s_{a})^{\top}(z_{k}+\alpha\Delta z_{a})}{s_{k}^{\top}z_{k}}\\ \sigma&=\max\{0,\min\{1,\rho\}^{3}\}.\end{split}$
 The centering parameter balances the need to reduce the duality measure while maintaining proximity to the central path, allowing for a longer step in the next iteration [50].
@@
 This direction, $(\Delta x,\Delta s,\Delta y,\Delta z)$ is obtained by solving Equation (19) with the residual vectors
-$r_{x}$ $=Px_{k}+c+A^{\top}y_{k}+G^{\top}z_{k}$
-$r_{y}$ $=Ax_{k}-b$
-$r_{z}$ $=Gx_{k}+s_{k}-h$
-$r_{s}$ $=\lambda_{k}\circ\lambda_{k}+(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})-\sigma\mu_{k}\boldsymbol{e}$
+$r_{x}$ $=Px_{k}+c+A^{\top}y_{k}+G^{\top}z_{k}$ (21a)
+$r_{y}$ $=Ax_{k}-b$ (21b)
+$r_{z}$ $=Gx_{k}+s_{k}-h$ (21c)
+$r_{s}$ $=\lambda_{k}\circ\lambda_{k}+(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})-\sigma\mu_{k}\boldsymbol{e}$ (21d)
 This direction moves the next iterate closer to $(x^{*}(\sigma\mu_{k}),s^{*}(\sigma\mu_{k}),y^{*}(\sigma\mu_{k}),z^{*}(\sigma\mu_{k}))$, a point on the central path where the duality measure is reduced by a factor of $\sigma$.
-Note that the residual vector $r_{s}$ has two additional terms when compared to Equation (20d): $(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})$ and $-\sigma\mu_{k}\boldsymbol{e}$.
+Note that the residual vector $r_{s}$ has two additional terms when compared to Equation (20d): $(W_{k}^{-\top}\Delta s_{a})\circ(W_{k}\Delta z_{a})$ and $\sigma\mu_{k}\boldsymbol{e}$.
 The former is to correct for linearization error in Newton’s method, and the latter is to maintain proximity to the central path and arises when applying Newton’s method to Equation (9d) with $\tau=\sigma\mu_{k}$.
@@
 We begin by solving the optimization problem
-[…]
+$\underset{x}{\text{minimize}}$ $\frac{1}{2}x^{\top}Px+c^{\top}x+\|Gx-h\|_{2}^{2}$
 subject to $Ax=b.$
@@
 $\begin{bmatrix}P&A^{\top}&G^{\top}\\ A&0&0\\ G&0&-I\end{bmatrix}\begin{bmatrix}x\\ y\\ z\end{bmatrix}=\begin{bmatrix}-c\\ b\\ h\end{bmatrix}$. (22)
-We then set $s=-z$. This results in the tuple $(x,s,y,z)$ satisfying Equations (9a) - (9c), but $s$ and $z$ may not necessarily satisfy Equation (9e).
+We then set s=zs=-z. This results in the tuple $(x,s,y,z)$ satisfying Equations (9a) - (9c), but $s$ and $z$ may not necessarily satisfy Equation (9e).
 We then initialize $x_{0}$ and $y_{0}$ as $x_{0}=x$, $y_{0}=y$, and initialize $s_{0}$ as
@@
 3: Solve KKT system (22) for $x,y,z$
-4:$x_{0}\leftarrow x,\;y_{0}\leftarrow y$
-5:$s\leftarrow-z$
+[…]
+[…]
 6:if $s\in\mathcal{K}$ then
@@
 4.2 Code generation
-qocogen is written is Python and example code for generating qococustom for a problem family defined by Equation (31) is given by Listing 2.
+qocogen is written in Python and example code for generating qococustom for a problem family defined by Equation (30) is given by Listing 2.
 The generated solver qococustom is customized to the sparsity pattern of the matrices $P,A,G$ which are passed to the function qocogen.generate_solver() as sparse scipy matrices.
-[…]
+$\begin{array}[]{ll}\underset{x}{\text{minimize}}&x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}\\ \mbox{subject to}&x_{1}+x_{2}=1\\ &x_{2}+x_{3}=1\\ &x_{1}\geq 0\\ &\sqrt{x_{3}^{2}+x_{4}^{2}}\leq x_{2}\end{array}$ (30)
 The file tree for qococustom is given by Figure 3.
@@
 5 Numerical results
-We benchmarked qoco and qococustom against several conic interior-point solvers: the open-source solvers ecos [32] and clarabel [12], the commercial solvers gurobi [13] and mosek [33], and against the academically licensed custom solver generator cvxgen which generates a primal-dual interior point method for quadratic programs.
-We did not test against the custom solver generator bsocp since it is not academically licensed.
-For all solvers, we used their default settings but set the tolerances $\epsilon_{abs}=\epsilon_{rel}=10^{-7}$, and disabled the presolve features of gurobi and mosek to ensure all solvers are given the same problem formulation.
-We evaluated our solvers on benchmark problems spanning multiple fields and a range of problem sizes.
+We benchmark qoco and qococustom against several conic interior-point solvers  [^4] Our benchmarks are publicly available at https://github.com/qoco-org/qoco-benchmarks: the open-source solvers ecos [32] and clarabel [12], the commercial solvers gurobi [13] and mosek [33], and against the academically licensed custom solver generator cvxgen which generates a primal-dual interior point method for quadratic programs.
+We do not test against the custom solver generator bsocp since it is not academically licensed.
+For all solvers, we use their default settings but set the tolerances $\epsilon_{abs}=\epsilon_{rel}=10^{-7}$.
 In our experiments, we define problem size as the total number of nonzero elements in $A$, $G$, and in the upper half of $P$.
-For all benchmark problems, we tested qoco, but only tested qococustom for the smaller problems, because code generation and compile times become prohibitively long for larger problem instances.
-We also evaluated qoco on the difficult Maros–Mészáros problems, a repository of 138 challenging quadratic programs (QPs) [61].
-To evaluate the solvers’ performance, we used performance profiles [62] and the shifted geometric mean, both of which are frequently used to assess solver performance [12, 32, 56, 14, 63].
+We evaluate our solvers on a set of benchmark problems we developed, a set of model-predictive control problems from [61], the Maros–Mészáros problems [62], and least-squares problems derived from the SuiteSparse repository [63].
+To evaluate the solvers’ performance, we use performance profiles [64] and the shifted geometric mean, both of which are frequently used to assess solver performance [12, 32, 56, 14, 65].
 All numerical results were generated on a desktop computer with an AMD Ryzen 9 7950X3D processor running at 4.2 GHz and with 128 GB of RAM.
-Our benchmarks are written in Python and we used CVXPY [41, 42] to interface with the solvers.
-Our benchmarks are publicly available [64].
+Our benchmarks are written in Python and we use cvxpy [42, 43] to interface with the solvers.
 We provide detailed numeric results in Appendix D, which includes iteration counts and solve times for each solver on all problems we ran.
@@
 where $t_{s,p}$ is the runtime of solver $s$ on problem $p$ and $k=1$ is the shift.
-If the solver does not find an optimal solution, then the runtime is set to the maximum allowable runtime (10 seconds for benchmark problems and 1200 seconds for the Maros–Mészáros problems).
+If the solver does not find an optimal solution, then the runtime is set to 10 seconds for benchmark problems, 0.5 seconds for the MPC problems, 1200 seconds for the Maros–Mészáros problems, and 600 seconds for the SuiteSparse least-squares problems.
+These times were chosen to be roughly an order of magnitude higher than the slowest successful solve on the respective problem set.
 We then define the normalized shifted geometric mean for solver $s$ as
@@
 5.1 Benchmark problems
-We solved optimization problems from five problem classes: robust Kalman filtering, group lasso regression [65], the losslessly convexified powered-descent guidance problem [37], Markowitz portfolio optimization [3], and the oscillating masses control problem [66].
+We solve optimization problems from five problem classes: robust Kalman filtering, group lasso regression [66], the losslessly convexified powered-descent guidance problem [37], Markowitz portfolio optimization [3], and the oscillating masses control problem [67].
 The first three of these are SOCPs and the last two are QPs.
-For each of the five classes, we considered 10 different problem sizes, generating 20 unique problem instances per size, for a total of 1000 distinct optimization problems.
-Details on the problems we solved can be found in Appendix C.
-For these problems, we plotted runtime as a function of problem size in addition to reporting performance profiles and shifted geometric means.
-For each solver, we solved each optimization problem 100 times and recorded the minimum runtime, which includes both setup and solve time.
-Since the computer runs various background processes that can artificially inflate execution time, we conducted 100 runs and reported the minimum execution time, thereby providing a more accurate measure of the solver’s performance.
-Although the oscillating masses and portfolio optimization problems are QPs and could be solved with cvxgen, we limited testing to the two smaller oscillating masses problems.
+For each of the five classes, we consider 10 different problem sizes, generating 20 unique problem instances per size, for a total of 1000 distinct optimization problems.
+Details on the problems we solve can be found in Appendix C.
+For these problems, we plot runtime as a function of problem size in addition to reporting performance profiles and shifted geometric means.
+We test qoco on all problems, but only test qococustom on the smaller problems, as code generation and compile times become prohibitively long for the larger instances.
+For each solver, we solve each optimization problem 100 times and record the minimum runtime, which includes both setup and solve time.
+Since the computer runs various background processes that can artificially inflate execution time, we conduct 100 runs and report the minimum execution time, thereby providing a more accurate measure of the solver’s performance.
+Although the oscillating masses and portfolio optimization problems are QPs and can be solved with cvxgen, we limit testing to the two smaller oscillating masses problems.
 For larger instances, cvxgen was unable to generate code due to the size of the problems.
-For the portfolio optimization problems, we opted not to test cvxgen due to the sparsity pattern of the factor matrix $F$ in Problem 34.
+For the portfolio optimization problems, we do not test cvxgen due to the sparsity pattern of the factor matrix $F$ in Problem 33.
 To specify the sparsity pattern of $F$ in cvxgen, we must manually provide the locations of nonzero elements into the web interface, which would have been prohibitively expensive, since $F$ has hundreds of nonzero elements.
 In contrast, to generate code with qocogen, we pass in problem data as sparse scipy matrices rather than manually specifying sparsity patterns.
-We could have specified $F$ as a dense matrix and only populated the nonzero elements for cvxgen, this would severely hinder its performance and result in an unfair comparison.
+Although we can specify $F$ as a dense matrix and only populate the nonzero elements for cvxgen, this would severely hinder its performance and result in an unfair comparison.
 When computing normalized shifted geometric mean, relative performance profiles, and absolute performance profiles, all solvers must be tested on the same set of problems.
@@
 The second version, given by Figure 6, also includes qococustom, but the solvers are evaluated on the half of the benchmark problems which qococustom is tested on.
-Figures 4, 5, and 6 show that qoco is the fastest generic solver on our benchmark problems.
-Additionally, qococustom is the fastest solver by a significant margin and is as performant as cvxgen on the $40$ smallest oscillating mass problems.
+Figures 4, 5, and 6 show that qococustom is the fastest solver on our benchmark problems by a significant margin and is as performant as cvxgen on the $40$ smallest oscillating mass problems.
 This latter result is unsurprising, as qococustom uses an identical algorithm to cvxgen when solving QPs and both utilize custom linear algebra routines.
+Additionally, qoco is the fastest open-source generic solver on all of the problems, although it is slower than Mosek on the group lasso problems and slower than Gurobi on some of the largest portfolio optimization problems.
+However, we notice that if we disable presolve for Mosek and Gurobi, qoco is faster on these problems as well.
 Figure 4: Solvetime in seconds vs problem size for benchmark problems
@@
 QOCO Clarabel ECOS Gurobi Mosek
-Shifted GM 1.0 1.5 2.4 14.4 3.8
+Shifted GM 1.0 1.6 2.5 10.2 1.8
 Failure Rate (%) 0.0 0.0 0.0 0.0 0.0
@@
 QOCO Custom QOCO Clarabel ECOS Gurobi Mosek
-Shifted GM 1.0 2.1 3.5 6.9 30.4 15.1
+Shifted GM 1.0 2.0 3.5 6.9 23.9 7.0
 Failure Rate (%) 0.0 0.0 0.0 0.0 0.0 0.0
@@
 Figure 6: Performance profiles for benchmark problems with custom solver
+5.2 Model-predictive control problems
+We consider a set of 64 model-predictive control problems taken from [61] and also used by [12].
+Note that we exclude the eight “nonlinear chain” problems as they were too large to generate custom solvers for.
+Similarly to the benchmark problems, we solve each MPC problem 100 times and record the minimum runtime.
+These problems take the following form
+[…]
+Figure 7 shows that clarabel, gurobi, and mosek solve all problems successfully, qoco and qococustom each fail on only one, and ecos fails on seven.
+Among the 63 problems solved by both qoco and qococustom, qococustom is the fastest solver by a wide margin, followed by qoco.
+(a) Relative performance profile
+(b) Absolute performance profile
+QOCO Custom QOCO Clarabel ECOS Gurobi Mosek
+Shifted GM 4.4 4.5 1.0 31.4 1.7 2.5
+Failure Rate (%) 1.6 1.6 0.0 10.9 0.0 0.0
+(c) Shifted geometric means and failure rates
+Figure 7: Performance profiles for model-predictive control problems
 5.3 Maros–Mészáros problems
-The Maros–Mészáros problems are a set of 138 QPs that include a wide range of problem sizes (from $3$ to around $300,000$ variables) and contain very difficult problems that have ill-conditioning and poor scaling [61].
+The Maros–Mészáros problems are a set of 138 challenging QPs that include a wide range of problem sizes (from $3$ to around $300,000$ variables) and contain very difficult problems that have ill-conditioning and poor scaling [62].
 These properties make the Maros–Mészáros problems a good test set to assess the robustness of solvers.
 We test qoco but not qococustom on these problems since many problems are far too large to generate code for.
-Additionally, we ran each problem once rather than the 100 runs we did for the benchmark problems since many of the Maros–Mészáros problems are extremely large, and running them 100 times per solver would be extremely cumbersome.
+Additionally, we run each problem once rather than the 100 runs we do for the benchmark and MPC problems since many of the Maros–Mészáros problems are extremely large, and running them 100 times per solver is extremely cumbersome.
 Figure 8 shows that the proprietary solver Gurobi successfully solved the largest fraction of the problems followed by qoco, clarabel, mosek, and ecos.
 We can also see that for the majority of the problems, qoco was the fastest solver, whereas mosek and ecos tended to be the slowest.
-The latter result is unsurprising as mosek and ecos can only handle linear objectives and the reformulation of the quadratic objective into a second-order cone likely introduced significant fill-in, slowing these solvers down.
-We also observed that for many of the largest problems, Gurobi was the fastest solver, due to multithreading in its matrix factorization.
+The latter result is unsurprising as mosek and ecos can only handle linear objectives and the reformulation of the quadratic objective into a second-order cone likely introduces significant fill-in, slowing these solvers down.
+We also observe that for many of the largest problems, Gurobi is the fastest solver, due to multithreading in its matrix factorization.
 (a) Relative performance profile
@@
 QOCO Clarabel ECOS Gurobi Mosek
-Shifted GM 2.0 3.0 27.4 1.0 3.8
-Failure Rate (%) 6.5 8.7 31.9 2.9 9.4
+Shifted GM 2.4 3.6 33.9 1.0 4.1
+Failure Rate (%) 6.5 8.7 31.9 2.9 8.7
 (c) Shifted geometric means and failure rates
 Figure 8: Performance profiles for Maros–Mészáros problems
+5.4 SuiteSparse least-squares problems
+Finally, we consider a set of 23 least-squares problems $Ax\approx b$, where the matrix $A\in\mathbb{R}^{m\times n}$ is drawn from the SuiteSparse Matrix Collection [63].
+Following [56] and [12], we get an approximate solution to each least-squares problem by solving a Huber regression and lasso regression problem, which can both be formulated as constrained quadratic programs [51].
+These optimization problems are quite large with some having over a million optimization variables, so we do not test qococustom due to the prohibitively long code-generation time.
+Additionally, we run each problem once rather than 100 times.
+The Huber regression problem is given by
+[…]
+where $a_{i}^{\top}$ is the $i^{th}$ row of $A$ and $\phi:\mathbb{R}\to\mathbb{R}$ is defined as
+$\phi(z)=\begin{cases}z^{2}&\text{if }|z|\leq\delta\\ \delta(2|z|-\delta)&\text{if }|z|>\delta\end{cases},$
+where we use $\delta=1$.
+The lasso regression problem is given by
+[…]
+where we set $\lambda=\|A^{\top}b\|_{\infty}$.
+Figure 9 shows that qoco and clarabel successfully solve all problems, with qoco being the fastest on most instances.
+As with the Maros–Mészáros problems, we observe that Gurobi is the fastest on some of the largest problems, likely due to it multithreaded matrix factorization.
+(a) Relative performance profile
+(b) Absolute performance profile
+QOCO Clarabel ECOS Gurobi Mosek
+Shifted GM 1.0 1.2 7.6 1.6 1.2
+Failure Rate (%) 0.0 0.0 23.9 8.7 2.2
+(c) Shifted geometric means and failure rates
+Figure 9: Performance profiles for SuiteSparse least-squares problems
 6 Conclusion
@@
 We can attempt to symmetrize the coefficient matrix by multiplying the second equation by $\mathrm{Arw}(s_{k})^{-1}$ to get
-$\begin{bmatrix}P&0&A^{\top}&G^{\top}\\ 0&\mathrm{Arw}(s_{k})^{-1}\mathrm{Arw}(z_{k})&0&I\\ A&0&0&0\\ G&I&0&0\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y\\ \Delta z\end{bmatrix}=\begin{bmatrix}-r_{x}\\ -\mathrm{Arw}(s_{k})^{-1}r_{s}\\ -r_{y}\\ -r_{z}\end{bmatrix}.$ (32)
+$\begin{bmatrix}P&0&A^{\top}&G^{\top}\\ 0&\mathrm{Arw}(s_{k})^{-1}\mathrm{Arw}(z_{k})&0&I\\ A&0&0&0\\ G&I&0&0\end{bmatrix}\begin{bmatrix}\Delta x\\ \Delta s\\ \Delta y\\ \Delta z\end{bmatrix}=\begin{bmatrix}-r_{x}\\ -\mathrm{Arw}(s_{k})^{-1}r_{s}\\ -r_{y}\\ -r_{z}\end{bmatrix}$.
+(31)
 However, the coefficient matrix above will only be symmetric if $\mathrm{Arw}(s_{k})^{-1}\mathrm{Arw}(z_{k})$ is symmetric.
@@
 As a representative example of a matrix we would factor, we consider the KKT matrix associated with the Linear Quadratic Regulator (LQR). The LQR problem is
-$\begin{split}\underset{x,u}{\mathrm{minimize}}\quad&\frac{1}{2}\left(\sum_{k=1}^{T}x_{k}^{\top}Q_{k}x_{k}+\sum_{k=1}^{T-1}u_{k}^{\top}R_{k}u_{k}\right)\\ \mathrm{subject\;to}\quad&x_{k+1}=A_{k}x_{k}+B_{k}u_{k}\;\quad\forall k\in[1,T-1]\\ &x_{1}=x_{\mathrm{init}},\\ \end{split}$
+[…]
 and its associated KKT matrix is
@@
 where
-$\begin{split}P&=\mathrm{blkdiag}(Q_{1},Q_{2},\ldots,Q_{T},R_{1},R_{2},\ldots,R_{T-1})\\ H&=\begin{bmatrix}A_{1}&-I&&&&&B_{1}\\ &A_{2}&-I&&&&&B_{2}\\ &&&\ddots&&&&&\ddots\\ &&&&A_{N-1}&-I&&&&B_{N-1}\\ I&&&&&&&&&\\ \end{bmatrix}.\end{split}$
+[…]
 Notice that we apply static regularization to ensure that the $(2,2)$ block of K is negative definite, making $K$ quasidefinite. The sparsity pattern of $K$ is depicted by Figure 10.
@@
 We consider the robust Kalman filtering problem for vehicle tracking, as outlined in [68]. This problem is a quadratic objective second-order cone program (SOCP) and takes the form
-$\begin{split}\underset{x_{k},w_{k},v_{k}}{\text{minimize}}\quad&\sum_{k=0}^{N-1}\left(\|w_{k}\|_{2}^{2}+\tau\phi_{\rho}(v_{k})\right)\\ \text{subject to}\quad&x_{k+1}=Ax_{k}+Bw_{k}\quad\forall k\in[0,N-1]\\ \quad&y_{k}=Cx_{k}+v_{k}\quad\forall k\in[0,N-1],\\ \end{split}$ (33)
+[…]
 where $\phi_{\rho}$ is the Huber loss function
-$\phi_{\rho}(z)=\begin{cases}\|z\|_{2}^{2}&\|z\|_{2}\leq\rho\\ 2\rho\|z\|_{2}-\rho^{2}&\|z\|_{2}>\rho.\\ \end{cases}$
+[…]
 In this problem, $x_{k}\in\mathbb{R}^{4}$ represents the state of the vehicle at timestep $k$, $w_{k}\in\mathbb{R}^{2}$ is an unknown force vector acting on the vehicle at timestep $k$, $v_{k}\in\mathbb{R}^{2}$ is the measurement noise vector at timestep $k$, and $y_{k}\in\mathbb{R}^{2}$ represents a noisy measurement of the vehicle’s state at timestep $k$.
 The matrices $A$ and $B$ are known dynamics matrices for the system, and $C$ is the known observation matrix.
-To reconstruct the vehicle’s state trajectory, we solve Problem 33.
+To reconstruct the vehicle’s state trajectory, we solve Problem 32.
 We use the system matrices
 $A$ $=\begin{bmatrix}1&0&\left(1-\frac{\gamma}{2}\Delta t\right)\Delta t&0\\ 0&1&0&\left(1-\frac{\gamma}{2}\Delta t\right)\Delta t\\ 0&0&1-\gamma\Delta t&0\\ 0&0&0&1-\gamma\Delta t\end{bmatrix}$
-$\vspace{10mm}B$ $=\begin{bmatrix}\frac{1}{2}\Delta t^{2}&0&\\ 0&\frac{1}{2}\Delta t^{2}\\ \Delta t&0\\ 0&\Delta t\end{bmatrix}$
+B
+[…]
+$=\begin{bmatrix}\frac{1}{2}\Delta t^{2}&0&\\ 0&\frac{1}{2}\Delta t^{2}\\ \Delta t&0\\ 0&\Delta t\end{bmatrix}$
 $C$ $=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ \end{bmatrix}$,
@@
 Problem instances
-[…]
+For each problem size, we generate 20 random problem instances, corresponding to 20 different observation trajectories $y_{t}$.
+To do this, we first randomly generate the force and noise vectors $w_{t}$ and $v_{t}$, respectively, for $t\in[0,N-1]$.
+Then, we forward-simulate an initial state of $x_{0}=[0\;0\;0\;0]$ through the dynamics defined in the constraints of Problem 32.
+Each element of $w_{t}$ and $v_{t}$ is drawn from a standard Gaussian distribution $\mathcal{N}(0,1)$.
+Additionally, for 20 % of the elements of $v_{t}$, we introduce outlier noise by drawing elements from $\mathcal{N}(0,20)$, to simulate noise with a larger magnitude.
 C.2 Lossless convexification
@@
 $A$ $=\begin{bmatrix}I_{3}&\Delta t(I_{3})\\ 0_{3\times 3}&I_{3}\end{bmatrix}$
-$\vspace{10mm}B$ $=\begin{bmatrix}\frac{1}{2}\Delta t^{2}(I_{3})\\ \Delta t(I_{3})\end{bmatrix}$
+B
 […]
+$=\begin{bmatrix}\frac{1}{2}\Delta t^{2}(I_{3})\\ \Delta t(I_{3})\end{bmatrix}$
+$g$ $=[0\;\;0\;\;-0.5g_{0}\Delta t^{2}\;\;0\;\;0\;\;-g_{0}\Delta t].$
 For this problem class we use the parameters
-$g_{0}=9.807,\;\rho_{1}=100,\;\rho_{2}=500,\;m_{\mathrm{dry}}=25,\;m_{\mathrm{wet}}=35,\;\theta_{\mathrm{max}}=\pi/4,\;\alpha=0.001.\$
+$g_{0}=9.807,\;\rho_{1}=100,\;\rho_{2}=500,\;m_{\mathrm{dry}}=25,\;m_{\mathrm{wet}}=35,\;\theta_{\mathrm{max}}=\pi/4,\;\alpha=0.001.\\$
 Problem sizes
@@
 The final problem can be rewritten as
-$\begin{split}\underset{x,y}{\text{minimize}}\quad&x^{\top}Dx+y^{\top}y-\gamma^{-1}\mu^{\top}x\\ \text{subject to}\quad&y=F^{\top}x\\ \quad&\sum_{i=1}^{n}x_{i}=1\\ \quad&x_{i}\geq 0\quad\forall i\in[1,n].\\ \end{split}$ (34)
+[…]
 Problem sizes
@@
 $A$ $={\rm e}^{A_{c}\Delta t}$
-$\vspace{10mm}B$ $=A_{c}^{-1}(A-I_{2N})B_{c},$
+B
+[…]
+$=A_{c}^{-1}(A-I_{2N})B_{c},$
 where
 $A_{c}$ $=\begin{bmatrix}0_{N\times N}&I_{N}\\ L_{N}&0_{N\times N}\end{bmatrix}$
-$\vspace{10mm}B_{c}$ $=\begin{bmatrix}0_{N\times N}\\ I_{N}\end{bmatrix}.$
-Here, $L_{N}\in\mathbb{R}^{N\times N}$ is a symmetric tridiagonal matrix with $-2$ on the main diagonal and $1$ on the subdiagonal and superdiagonal.
+Bc
+[…]
+$=\begin{bmatrix}0_{N\times N}\\ I_{N}\end{bmatrix}$.
+Here, $L_{N}\in\mathbb{R}^{N\times N}$ is a symmetric tridiagonal matrix with 2-2 on the main diagonal and $1$ on the subdiagonal and superdiagonal.
 For this problem class, we use the parameters
@@
 This involves generating a random initial state $x_{\mathrm{init}}$, as well as the matrices $Q$ and $R$.
-We draw each element of $x_{\mathrm{init}}$ from $\mathcal{N}(0,1)$, then clip the elements such that they lie within the interval $[-0.9x_{\mathrm{max}},0.9x_{\mathrm{max}}]$.
+We draw each element of $x_{\mathrm{init}}$ from $\mathcal{N}(0,1)$, then clip the elements such that they lie within the interval $0.9x_{\mathrm{max}},0.9x_{\mathrm{max}}]$.
 This ensures that the initial state satisfies the state constraints, making the resulting problem feasible.
@@
 Appendix D Detailed numeric results
+Table 2: Iterations and solver runtimes for robust Kalman filter problems
+Iterations Solver Runtime (s)
+Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
+robust_kalman_filter_N_25_i_0 775 12 15 - 9 9 9 0.00082 0.00148 0.00618 0.00123 0.00043 0.00018
+robust_kalman_filter_N_25_i_1 775 10 15 - 10 9 9 0.00071 0.00154 0.00652 0.00129 0.00046 0.00018
+robust_kalman_filter_N_25_i_2 775 11 18 - 11 10 10 0.00077 0.00175 0.00701 0.00134 0.00047 0.00020
+robust_kalman_filter_N_25_i_3 775 10 17 - 10 9 9 0.00072 0.00171 0.00616 0.00129 0.00045 0.00018
+robust_kalman_filter_N_25_i_4 775 11 15 - 11 9 9 0.00077 0.00148 0.00630 0.00137 0.00043 0.00018
+robust_kalman_filter_N_25_i_5 775 10 14 - 9 8 8 0.00071 0.00145 0.00659 0.00124 0.00041 0.00016
+robust_kalman_filter_N_25_i_6 775 11 15 - 10 9 9 0.00076 0.00152 0.00656 0.00130 0.00045 0.00018
+robust_kalman_filter_N_25_i_7 775 10 17 - 11 9 9 0.00071 0.00168 0.00728 0.00136 0.00045 0.00018
+robust_kalman_filter_N_25_i_8 775 10 15 - 9 9 9 0.00071 0.00147 0.00614 0.00123 0.00044 0.00018
+robust_kalman_filter_N_25_i_9 775 9 18 - 11 9 9 0.00066 0.00182 0.00633 0.00133 0.00045 0.00018
+robust_kalman_filter_N_25_i_10 775 11 16 - 11 10 10 0.00077 0.00161 0.00668 0.00135 0.00048 0.00020
+robust_kalman_filter_N_25_i_11 775 12 19 - 11 10 10 0.00084 0.00185 0.00647 0.00134 0.00048 0.00020
+robust_kalman_filter_N_25_i_12 775 10 17 - 10 9 9 0.00071 0.00166 0.00576 0.00128 0.00045 0.00018
+robust_kalman_filter_N_25_i_13 775 8 18 - 10 8 8 0.00061 0.00182 0.00687 0.00129 0.00040 0.00016
+robust_kalman_filter_N_25_i_14 775 10 17 - 10 9 9 0.00071 0.00166 0.00582 0.00129 0.00044 0.00018
+robust_kalman_filter_N_25_i_15 775 10 16 - 9 8 8 0.00069 0.00162 0.00623 0.00123 0.00039 0.00016
+robust_kalman_filter_N_25_i_16 775 10 17 - 9 9 9 0.00072 0.00164 0.00665 0.00124 0.00043 0.00017
+robust_kalman_filter_N_25_i_17 775 9 17 - 10 9 9 0.00065 0.00175 0.00625 0.00130 0.00043 0.00018
+robust_kalman_filter_N_25_i_18 775 11 16 - 10 9 9 0.00076 0.00155 0.00579 0.00130 0.00044 0.00018
+robust_kalman_filter_N_25_i_19 775 11 19 - 10 10 10 0.00077 0.00186 0.00706 0.00129 0.00047 0.00020
+robust_kalman_filter_N_50_i_0 1550 13 18 - 11 12 12 0.00174 0.00364 0.01170 0.00272 0.00108 0.00048
+robust_kalman_filter_N_50_i_1 1550 14 19 - 11 12 12 0.00185 0.00362 0.01123 0.00275 0.00110 0.00048
+robust_kalman_filter_N_50_i_2 1550 11 20 - 10 10 10 0.00146 0.00393 0.01189 0.00268 0.00095 0.00041
+robust_kalman_filter_N_50_i_3 1550 11 19 - 9 10 10 0.00148 0.00375 0.01069 0.00246 0.00092 0.00040
+robust_kalman_filter_N_50_i_4 1550 11 20 - 9 10 10 0.00149 0.00381 0.01077 0.00245 0.00092 0.00040
+robust_kalman_filter_N_50_i_5 1550 11 19 - 9 10 10 0.00149 0.00377 0.01188 0.00243 0.00092 0.00041
+robust_kalman_filter_N_50_i_6 1550 12 19 - 10 11 11 0.00164 0.00373 0.01144 0.00260 0.00098 0.00045
+robust_kalman_filter_N_50_i_7 1550 10 17 - 8 9 9 0.00137 0.00347 0.01058 0.00228 0.00086 0.00037
+robust_kalman_filter_N_50_i_8 1550 11 20 - 9 10 10 0.00150 0.00386 0.01149 0.00257 0.00093 0.00040
+robust_kalman_filter_N_50_i_9 1550 12 24 - 9 10 10 0.00159 0.00463 0.01121 0.00246 0.00093 0.00040
+robust_kalman_filter_N_50_i_10 1550 11 21 - 9 9 9 0.00150 0.00410 0.01281 0.00245 0.00087 0.00037
+robust_kalman_filter_N_50_i_11 1550 11 19 - 9 10 10 0.00152 0.00377 0.01052 0.00246 0.00098 0.00040
+robust_kalman_filter_N_50_i_12 1550 12 18 - 9 11 11 0.00161 0.00358 0.01095 0.00244 0.00099 0.00044
+robust_kalman_filter_N_50_i_13 1550 13 21 - 12 12 12 0.00173 0.00411 0.01202 0.00287 0.00107 0.00048
+robust_kalman_filter_N_50_i_14 1550 13 19 - 12 12 12 0.00176 0.00377 0.01116 0.00286 0.00110 0.00049
+robust_kalman_filter_N_50_i_15 1550 11 19 - 10 10 10 0.00150 0.00379 0.01168 0.00259 0.00096 0.00041
+robust_kalman_filter_N_50_i_16 1550 12 21 - 9 10 10 0.00162 0.00414 0.01058 0.00243 0.00092 0.00040
+robust_kalman_filter_N_50_i_17 1550 13 21 - 10 11 11 0.00172 0.00398 0.01104 0.00255 0.00101 0.00044
+robust_kalman_filter_N_50_i_18 1550 11 18 - 9 10 10 0.00152 0.00356 0.01133 0.00243 0.00094 0.00041
+robust_kalman_filter_N_50_i_19 1550 13 19 - 11 12 12 0.00175 0.00372 0.01152 0.00270 0.00107 0.00048
+robust_kalman_filter_N_75_i_0 2325 13 18 - 12 11 11 0.00257 0.00556 0.01736 0.00523 0.00147 0.00068
+robust_kalman_filter_N_75_i_1 2325 12 19 - 10 10 10 0.00236 0.00565 0.01971 0.00477 0.00135 0.00062
+robust_kalman_filter_N_75_i_2 2325 11 21 - 11 10 10 0.00227 0.00648 0.01652 0.00496 0.00140 0.00061
+robust_kalman_filter_N_75_i_3 2325 13 21 - 13 11 11 0.00258 0.00616 0.01908 0.00540 0.00146 0.00068
+robust_kalman_filter_N_75_i_4 2325 13 22 - 12 12 12 0.00261 0.00666 0.01768 0.00523 0.00158 0.00073
+robust_kalman_filter_N_75_i_5 2325 12 23 - 12 11 11 0.00240 0.00680 0.01752 0.00521 0.00149 0.00068
+robust_kalman_filter_N_75_i_6 2325 12 21 - 10 10 10 0.00241 0.00646 0.01624 0.00471 0.00142 0.00061
+robust_kalman_filter_N_75_i_7 2325 12 24 - 11 10 10 0.00236 0.00723 0.01604 0.00488 0.00140 0.00063
+robust_kalman_filter_N_75_i_8 2325 12 21 - 11 10 10 0.00235 0.00619 0.01583 0.00499 0.00135 0.00063
+robust_kalman_filter_N_75_i_9 2325 12 20 - 11 10 10 0.00241 0.00601 0.01868 0.00502 0.00136 0.00062
+robust_kalman_filter_N_75_i_10 2325 14 23 - 13 12 12 0.00276 0.00668 0.01850 0.00548 0.00159 0.00074
+robust_kalman_filter_N_75_i_11 2325 13 21 - 11 11 11 0.00254 0.00644 0.01732 0.00497 0.00147 0.00068
+robust_kalman_filter_N_75_i_12 2325 12 22 - 12 11 11 0.00246 0.00644 0.01831 0.00520 0.00149 0.00067
+robust_kalman_filter_N_75_i_13 2325 13 21 - 11 11 11 0.00256 0.00626 0.01602 0.00493 0.00144 0.00067
+robust_kalman_filter_N_75_i_14 2325 13 19 - 11 11 11 0.00258 0.00577 0.01731 0.00497 0.00150 0.00068
+robust_kalman_filter_N_75_i_15 2325 14 22 - 13 12 12 0.00275 0.00643 0.01746 0.00544 0.00161 0.00074
+robust_kalman_filter_N_75_i_16 2325 13 22 - 11 11 11 0.00256 0.00649 0.01592 0.00489 0.00147 0.00068
+robust_kalman_filter_N_75_i_17 2325 12 21 - 11 10 10 0.00237 0.00628 0.01567 0.00495 0.00137 0.00061
+robust_kalman_filter_N_75_i_18 2325 13 20 - 12 11 11 0.00256 0.00593 0.01666 0.00523 0.00147 0.00067
+robust_kalman_filter_N_75_i_19 2325 12 20 - 11 10 10 0.00236 0.00578 0.01685 0.00494 0.00141 0.00062
+robust_kalman_filter_N_125_i_0 3875 13 22 - 14 11 11 0.00427 0.01094 0.03237 0.00975 0.00244 0.00116
+robust_kalman_filter_N_125_i_1 3875 14 23 - 13 12 12 0.00459 0.01191 0.03259 0.00893 0.00264 0.00127
+robust_kalman_filter_N_125_i_2 3875 13 20 - 12 11 11 0.00432 0.01002 0.03394 0.00857 0.00247 0.00117
+robust_kalman_filter_N_125_i_3 3875 13 21 - 12 10 10 0.00424 0.01075 0.03132 0.00857 0.00225 0.00109
+robust_kalman_filter_N_125_i_4 3875 13 20 - 11 10 10 0.00428 0.01025 0.03491 0.00816 0.00226 0.00107
+robust_kalman_filter_N_125_i_5 3875 13 26 - 12 12 12 0.00430 0.01322 0.03353 0.00852 0.00264 0.00127
+robust_kalman_filter_N_125_i_6 3875 14 21 - 13 12 12 0.00458 0.01053 0.03500 0.00932 0.00266 0.00129
+robust_kalman_filter_N_125_i_7 3875 13 21 - 13 12 12 0.00436 0.01074 0.03184 0.00902 0.00267 0.00127
+robust_kalman_filter_N_125_i_8 3875 11 23 - 11 10 10 0.00369 0.01136 0.03438 0.00813 0.00226 0.00107
+robust_kalman_filter_N_125_i_9 3875 12 23 - 11 10 10 0.00402 0.01188 0.03560 0.00808 0.00227 0.00107
+robust_kalman_filter_N_125_i_10 3875 15 24 - 13 12 12 0.00484 0.01217 0.03564 0.00894 0.00264 0.00124
+robust_kalman_filter_N_125_i_11 3875 13 24 - 11 11 11 0.00436 0.01190 0.03224 0.00818 0.00245 0.00116
+robust_kalman_filter_N_125_i_12 3875 13 23 - 13 11 11 0.00436 0.01151 0.03240 0.00891 0.00246 0.00117
+robust_kalman_filter_N_125_i_13 3875 11 21 - 11 10 10 0.00375 0.01088 0.03264 0.00812 0.00225 0.00106
+robust_kalman_filter_N_125_i_14 3875 14 23 - 12 11 11 0.00460 0.01119 0.03209 0.00857 0.00247 0.00119
+robust_kalman_filter_N_125_i_15 3875 13 21 - 12 11 11 0.00426 0.01063 0.03336 0.00861 0.00249 0.00117
+robust_kalman_filter_N_125_i_16 3875 12 21 - 12 11 11 0.00404 0.01096 0.03084 0.00858 0.00249 0.00117
+robust_kalman_filter_N_125_i_17 3875 13 23 - 12 12 12 0.00425 0.01165 0.03524 0.00862 0.00269 0.00125
+robust_kalman_filter_N_125_i_18 3875 13 21 - 12 11 11 0.00430 0.01095 0.03311 0.00846 0.00249 0.00116
+robust_kalman_filter_N_125_i_19 3875 14 21 - 12 12 12 0.00452 0.01066 0.03371 0.00850 0.00265 0.00126
+robust_kalman_filter_N_175_i_0 5425 13 23 - 13 11 11 0.00593 0.01662 0.03626 0.01213 0.00344 0.00164
+robust_kalman_filter_N_175_i_1 5425 13 24 - 13 11 11 0.00608 0.01747 0.03183 0.01210 0.00348 0.00163
+robust_kalman_filter_N_175_i_2 5425 13 24 - 14 11 11 0.00597 0.01745 0.03231 0.01318 0.00341 0.00162
+robust_kalman_filter_N_175_i_3 5425 13 22 - 12 11 11 0.00595 0.01584 0.03097 0.01173 0.00340 0.00163
+robust_kalman_filter_N_175_i_4 5425 13 22 - 12 11 11 0.00597 0.01597 0.03450 0.01165 0.00344 0.00164
+robust_kalman_filter_N_175_i_5 5425 14 26 - 13 11 11 0.00642 0.01877 0.03103 0.01214 0.00344 0.00166
+robust_kalman_filter_N_175_i_6 5425 15 22 - 13 12 12 0.00680 0.01587 0.03132 0.01206 0.00372 0.00180
+robust_kalman_filter_N_175_i_7 5425 14 25 - 13 11 11 0.00642 0.01788 0.03193 0.01218 0.00341 0.00166
+robust_kalman_filter_N_175_i_8 5425 14 24 - 13 11 11 0.00636 0.01728 0.03399 0.01199 0.00339 0.00166
+robust_kalman_filter_N_175_i_9 5425 14 24 - 14 12 12 0.00639 0.01703 0.03370 0.01262 0.00367 0.00179
+robust_kalman_filter_N_175_i_10 5425 13 24 - 12 11 11 0.00604 0.01771 0.03132 0.01146 0.00337 0.00168
+robust_kalman_filter_N_175_i_11 5425 13 24 - 14 12 12 0.00600 0.01756 0.03318 0.01306 0.00368 0.00177
+robust_kalman_filter_N_175_i_12 5425 14 23 - 14 12 12 0.00636 0.01663 0.03131 0.01259 0.00388 0.00176
+robust_kalman_filter_N_175_i_13 5425 12 26 - 13 10 10 0.00559 0.01881 0.03124 0.01259 0.00313 0.00150
+robust_kalman_filter_N_175_i_14 5425 13 22 - 12 10 10 0.00599 0.01625 0.02931 0.01162 0.00319 0.00153
+robust_kalman_filter_N_175_i_15 5425 14 26 - 13 13 13 0.00641 0.01890 0.03223 0.01215 0.00396 0.00194
+robust_kalman_filter_N_175_i_16 5425 12 22 - 12 11 11 0.00566 0.01615 0.03234 0.01155 0.00344 0.00165
+robust_kalman_filter_N_175_i_17 5425 14 25 - 13 11 11 0.00644 0.01818 0.03067 0.01220 0.00341 0.00167
+robust_kalman_filter_N_175_i_18 5425 13 25 - 12 12 12 0.00610 0.01829 0.03079 0.01211 0.00365 0.00178
+robust_kalman_filter_N_175_i_19 5425 13 24 - 15 11 11 0.00604 0.01730 0.03026 0.01383 0.00338 0.00164
+robust_kalman_filter_N_225_i_0 6975 13 24 - 12 11 - 0.00769 0.02292 0.03904 0.01513 0.00441 -
+robust_kalman_filter_N_225_i_1 6975 13 24 - 12 11 - 0.00768 0.02281 0.03951 0.01509 0.00436 -
+robust_kalman_filter_N_225_i_2 6975 13 26 - 13 12 - 0.00775 0.02437 0.03929 0.01586 0.00472 -
+robust_kalman_filter_N_225_i_3 6975 12 23 - 12 10 - 0.00729 0.02152 0.04337 0.01506 0.00401 -
+robust_kalman_filter_N_225_i_4 6975 13 25 - 13 11 - 0.00771 0.02463 0.04028 0.01566 0.00462 -
+robust_kalman_filter_N_225_i_5 6975 13 25 - 12 11 - 0.00769 0.02433 0.04326 0.01502 0.00453 -
+robust_kalman_filter_N_225_i_6 6975 13 26 - 13 11 - 0.00777 0.02557 0.04165 0.01581 0.00439 -
+robust_kalman_filter_N_225_i_7 6975 13 24 - 14 11 - 0.00774 0.02269 0.04116 0.01626 0.00455 -
+robust_kalman_filter_N_225_i_8 6975 13 24 - 12 11 - 0.00770 0.02280 0.03837 0.01508 0.00461 -
+robust_kalman_filter_N_225_i_9 6975 12 25 - 13 11 - 0.00718 0.02402 0.03773 0.01582 0.00451 -
+robust_kalman_filter_N_225_i_10 6975 14 26 - 13 12 - 0.00819 0.02443 0.04265 0.01576 0.00476 -
+robust_kalman_filter_N_225_i_11 6975 14 25 - 14 12 - 0.00824 0.02341 0.04128 0.01714 0.00497 -
+robust_kalman_filter_N_225_i_12 6975 12 25 - 11 10 - 0.00712 0.02437 0.04371 0.01434 0.00404 -
+robust_kalman_filter_N_225_i_13 6975 13 24 - 13 11 - 0.00762 0.02290 0.04400 0.01571 0.00451 -
+robust_kalman_filter_N_225_i_14 6975 13 24 - 15 11 - 0.00778 0.02273 0.04209 0.01791 0.00471 -
+robust_kalman_filter_N_225_i_15 6975 14 24 - 12 10 - 0.00818 0.02301 0.04042 0.01514 0.00451 -
+robust_kalman_filter_N_225_i_16 6975 13 24 - 12 11 - 0.00771 0.02254 0.04902 0.01508 0.00452 -
+robust_kalman_filter_N_225_i_17 6975 14 23 - 12 11 - 0.00817 0.02096 0.04042 0.01522 0.00446 -
+robust_kalman_filter_N_225_i_18 6975 13 25 - 13 11 - 0.00774 0.02364 0.04153 0.01579 0.00436 -
+robust_kalman_filter_N_225_i_19 6975 13 25 - 13 12 - 0.00775 0.02357 0.04061 0.01564 0.00471 -
+robust_kalman_filter_N_300_i_0 9300 14 23 - 13 11 - 0.01091 0.03031 0.05073 0.02093 0.00585 -
+robust_kalman_filter_N_300_i_1 9300 13 25 - 13 11 - 0.01028 0.03182 0.05255 0.02074 0.00591 -
+robust_kalman_filter_N_300_i_2 9300 14 25 - 12 10 - 0.01087 0.03232 0.05360 0.01995 0.00539 -
+robust_kalman_filter_N_300_i_3 9300 13 25 - 12 11 - 0.01032 0.03302 0.05028 0.01985 0.00587 -
+robust_kalman_filter_N_300_i_4 9300 13 23 - 13 11 - 0.01033 0.03029 0.05402 0.02151 0.00584 -
+robust_kalman_filter_N_300_i_5 9300 13 25 - 13 11 - 0.01028 0.03236 0.05103 0.02079 0.00621 -
+robust_kalman_filter_N_300_i_6 9300 13 24 - 12 11 - 0.01024 0.03124 0.05269 0.01998 0.00599 -
+robust_kalman_filter_N_300_i_7 9300 14 25 - 13 11 - 0.01096 0.03146 0.05782 0.02094 0.00612 -
+robust_kalman_filter_N_300_i_8 9300 13 25 - 13 10 - 0.01023 0.03290 0.05050 0.02072 0.00536 -
+robust_kalman_filter_N_300_i_9 9300 14 26 - 12 11 - 0.01093 0.03401 0.05052 0.01978 0.00599 -
+robust_kalman_filter_N_300_i_10 9300 13 24 - 12 11 - 0.01030 0.03104 0.04965 0.02001 0.00583 -
+robust_kalman_filter_N_300_i_11 9300 13 25 - 14 11 - 0.01033 0.03223 0.04701 0.02167 0.00578 -
+robust_kalman_filter_N_300_i_12 9300 14 26 - 13 11 - 0.01110 0.03346 0.05214 0.02155 0.00584 -
+robust_kalman_filter_N_300_i_13 9300 13 25 - 14 11 - 0.01037 0.03264 0.05290 0.02281 0.00585 -
+robust_kalman_filter_N_300_i_14 9300 14 25 - 13 11 - 0.01104 0.03283 0.05176 0.02094 0.00585 -
+robust_kalman_filter_N_300_i_15 9300 13 27 - 13 11 - 0.01018 0.03389 0.05460 0.02070 0.00577 -
+robust_kalman_filter_N_300_i_16 9300 16 26 - 14 12 - 0.01230 0.03374 0.05565 0.02153 0.00629 -
+robust_kalman_filter_N_300_i_17 9300 13 23 - 12 10 - 0.01023 0.02993 0.05101 0.01985 0.00539 -
+robust_kalman_filter_N_300_i_18 9300 13 24 - 12 10 - 0.01019 0.03166 0.05298 0.01995 0.00538 -
+robust_kalman_filter_N_300_i_19 9300 13 25 - 14 11 - 0.01033 0.03166 0.05133 0.02138 0.00581 -
+robust_kalman_filter_N_375_i_0 11625 13 26 - 12 11 - 0.01295 0.04382 0.06739 0.02676 0.00735 -
+robust_kalman_filter_N_375_i_1 11625 13 26 - 14 10 - 0.01291 0.04279 0.06874 0.02889 0.00674 -
+robust_kalman_filter_N_375_i_2 11625 12 26 - 12 11 - 0.01215 0.04399 0.06708 0.02649 0.00731 -
+robust_kalman_filter_N_375_i_3 11625 13 26 - 13 11 - 0.01293 0.04319 0.06396 0.02766 0.00730 -
+robust_kalman_filter_N_375_i_4 11625 13 24 - 15 11 - 0.01297 0.04012 0.06814 0.03156 0.00728 -
+robust_kalman_filter_N_375_i_5 11625 12 23 - 12 10 - 0.01214 0.03820 0.06889 0.02659 0.00704 -
+robust_kalman_filter_N_375_i_6 11625 14 25 - 13 12 - 0.01376 0.04189 0.06935 0.02769 0.00789 -
+robust_kalman_filter_N_375_i_7 11625 14 25 - 13 11 - 0.01381 0.04206 0.07491 0.02782 0.00727 -
+robust_kalman_filter_N_375_i_8 11625 13 26 - 14 11 - 0.01294 0.04346 0.07001 0.02881 0.00724 -
+robust_kalman_filter_N_375_i_9 11625 14 27 - 13 11 - 0.01389 0.04595 0.06740 0.02766 0.00733 -
+robust_kalman_filter_N_375_i_10 11625 14 25 - 12 11 - 0.01390 0.04195 0.07056 0.02656 0.00729 -
+robust_kalman_filter_N_375_i_11 11625 12 26 - 13 10 - 0.01212 0.04352 0.06386 0.02762 0.00665 -
+robust_kalman_filter_N_375_i_12 11625 13 24 - 14 11 - 0.01294 0.03993 0.06879 0.02891 0.00729 -
+robust_kalman_filter_N_375_i_13 11625 13 27 - 12 11 - 0.01298 0.04403 0.06671 0.02637 0.00735 -
+robust_kalman_filter_N_375_i_14 11625 14 26 - 14 11 - 0.01387 0.04313 0.06759 0.02902 0.00731 -
+robust_kalman_filter_N_375_i_15 11625 14 26 - 13 11 - 0.01382 0.04325 0.07478 0.02744 0.00736 -
+robust_kalman_filter_N_375_i_16 11625 13 27 - 13 12 - 0.01301 0.04496 0.06681 0.02751 0.00791 -
+robust_kalman_filter_N_375_i_17 11625 13 26 - 15 11 - 0.01300 0.04334 0.06621 0.03148 0.00726 -
+robust_kalman_filter_N_375_i_18 11625 13 24 - 13 11 - 0.01283 0.03984 0.06962 0.02743 0.00733 -
+robust_kalman_filter_N_375_i_19 11625 14 25 - 14 11 - 0.01394 0.04152 0.07587 0.02906 0.00733 -
+robust_kalman_filter_N_450_i_0 13950 13 26 - 14 11 - 0.01558 0.05342 0.08442 0.03470 0.00877 -
+robust_kalman_filter_N_450_i_1 13950 13 27 - 13 11 - 0.01559 0.05502 0.08003 0.03476 0.00872 -
+robust_kalman_filter_N_450_i_2 13950 13 26 - 13 11 - 0.01556 0.05468 0.08639 0.03301 0.00879 -
+robust_kalman_filter_N_450_i_3 13950 13 28 - 14 10 - 0.01539 0.05612 0.09875 0.03590 0.00811 -
+robust_kalman_filter_N_450_i_4 13950 13 31 - 13 11 - 0.01551 0.06150 0.08927 0.03265 0.00875 -
+robust_kalman_filter_N_450_i_5 13950 14 26 - 14 11 - 0.01651 0.05324 0.09395 0.03425 0.00880 -
+robust_kalman_filter_N_450_i_6 13950 14 29 - 14 11 - 0.01651 0.05747 0.08958 0.03447 0.00876 -
+robust_kalman_filter_N_450_i_7 13950 13 27 - 15 11 - 0.01555 0.05310 0.08318 0.03585 0.00877 -
+robust_kalman_filter_N_450_i_8 13950 14 28 - 14 12 - 0.01658 0.05595 0.08977 0.03436 0.00967 -
+robust_kalman_filter_N_450_i_9 13950 13 29 - 13 10 - 0.01544 0.05851 0.08958 0.03259 0.00811 -
+robust_kalman_filter_N_450_i_10 13950 14 27 - 14 10 - 0.01656 0.05439 0.10099 0.03602 0.00807 -
+robust_kalman_filter_N_450_i_11 13950 12 25 - 13 10 - 0.01446 0.05188 0.09591 0.03300 0.00833 -
+robust_kalman_filter_N_450_i_12 13950 15 26 - 13 12 - 0.01738 0.05263 0.09965 0.03307 0.00944 -
+robust_kalman_filter_N_450_i_13 13950 13 26 - 14 11 - 0.01559 0.05297 0.09955 0.03434 0.00880 -
+robust_kalman_filter_N_450_i_14 13950 13 26 - 13 10 - 0.01553 0.05320 0.07826 0.03279 0.00811 -
+robust_kalman_filter_N_450_i_15 13950 14 28 - 13 10 - 0.01669 0.05794 0.08767 0.03325 0.00813 -
+robust_kalman_filter_N_450_i_16 13950 14 29 - 14 11 - 0.01654 0.05936 0.08538 0.03584 0.00873 -
+robust_kalman_filter_N_450_i_17 13950 13 27 - 13 11 - 0.01547 0.05590 0.08448 0.03308 0.00878 -
+robust_kalman_filter_N_450_i_18 13950 14 26 - 13 13 - 0.01660 0.05390 0.08806 0.03296 0.01010 -
+robust_kalman_filter_N_450_i_19 13950 13 25 - 14 11 - 0.01542 0.05115 0.09033 0.03430 0.00879 -
+robust_kalman_filter_N_500_i_0 15500 14 25 - 12 11 - 0.01843 0.05906 0.10041 0.03531 0.00977 -
+robust_kalman_filter_N_500_i_1 15500 13 25 - 14 11 - 0.01730 0.05758 0.09165 0.03863 0.00972 -
+robust_kalman_filter_N_500_i_2 15500 13 25 - 12 10 - 0.01713 0.05923 0.09384 0.03503 0.00894 -
+robust_kalman_filter_N_500_i_3 15500 13 26 - 12 11 - 0.01729 0.06229 0.08355 0.03555 0.00964 -
+robust_kalman_filter_N_500_i_4 15500 13 29 - 14 11 - 0.01718 0.06716 0.08791 0.03857 0.00981 -
+robust_kalman_filter_N_500_i_5 15500 13 25 - 13 10 - 0.01721 0.05794 0.08527 0.03686 0.00897 -
+robust_kalman_filter_N_500_i_6 15500 14 29 - 14 11 - 0.01838 0.06719 0.08700 0.03852 0.00976 -
+robust_kalman_filter_N_500_i_7 15500 13 25 - 13 11 - 0.01732 0.05852 0.09098 0.03673 0.00975 -
+robust_kalman_filter_N_500_i_8 15500 14 27 - 14 11 - 0.01850 0.06306 0.09047 0.03830 0.00987 -
+robust_kalman_filter_N_500_i_9 15500 13 30 - 13 10 - 0.01717 0.06862 0.08998 0.03689 0.00901 -
+robust_kalman_filter_N_500_i_10 15500 14 26 - 13 11 - 0.01861 0.05992 0.09209 0.03657 0.00973 -
+robust_kalman_filter_N_500_i_11 15500 13 28 - 13 11 - 0.01721 0.06351 0.08248 0.03656 0.00975 -
+robust_kalman_filter_N_500_i_12 15500 13 26 - 14 10 - 0.01726 0.06046 0.08330 0.03811 0.00894 -
+robust_kalman_filter_N_500_i_13 15500 12 25 - 13 10 - 0.01613 0.05756 0.08654 0.03681 0.00897 -
+robust_kalman_filter_N_500_i_14 15500 13 26 - 14 11 - 0.01735 0.05895 0.09669 0.04051 0.00978 -
+robust_kalman_filter_N_500_i_15 15500 14 27 - 15 11 - 0.01839 0.06165 0.09525 0.04201 0.00972 -
+robust_kalman_filter_N_500_i_16 15500 13 26 - 14 11 - 0.01729 0.05945 0.09317 0.03859 0.00977 -
+robust_kalman_filter_N_500_i_17 15500 13 27 - 14 11 - 0.01714 0.06225 0.09272 0.03811 0.00970 -
+robust_kalman_filter_N_500_i_18 15500 13 26 - 14 11 - 0.01720 0.05936 0.08589 0.03816 0.00980 -
+robust_kalman_filter_N_500_i_19 15500 13 28 - 14 11 - 0.01728 0.06322 0.09300 0.03831 0.00972 -
+Table 3: Iterations and solver runtimes for lossless convexification problems
+Iterations Solver Runtime (s)
+Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
+lcvx_N_15_i_0 644 8 7 - 9 9 9 0.00074 0.00060 0.00438 0.00111 0.00049 0.00023
+lcvx_N_15_i_1 644 8 7 - 9 9 9 0.00072 0.00060 0.00428 0.00110 0.00049 0.00024
+lcvx_N_15_i_2 644 10 7 - 9 10 10 0.00085 0.00062 0.00458 0.00099 0.00055 0.00025
+lcvx_N_15_i_3 644 8 7 - 9 10 10 0.00073 0.00057 0.00442 0.00107 0.00052 0.00026
+lcvx_N_15_i_4 644 9 7 - 9 9 9 0.00078 0.00062 0.00472 0.00109 0.00049 0.00023
+lcvx_N_15_i_5 644 10 7 - 10 9 9 0.00084 0.00060 0.00446 0.00115 0.00049 0.00023
+lcvx_N_15_i_6 644 7 7 - 9 10 10 0.00068 0.00058 0.00417 0.00109 0.00053 0.00025
+lcvx_N_15_i_7 644 9 7 - 9 10 10 0.00078 0.00060 0.00450 0.00108 0.00053 0.00025
+lcvx_N_15_i_8 644 8 7 - 10 9 9 0.00074 0.00059 0.00444 0.00115 0.00048 0.00023
+lcvx_N_15_i_9 644 8 7 - 9 10 10 0.00073 0.00057 0.00458 0.00109 0.00053 0.00025
+lcvx_N_15_i_10 644 8 7 - 9 9 9 0.00072 0.00060 0.00441 0.00109 0.00049 0.00023
+lcvx_N_15_i_11 644 7 7 - 9 9 9 0.00067 0.00059 0.00413 0.00104 0.00049 0.00023
+lcvx_N_15_i_12 644 9 7 - 9 10 10 0.00087 0.00058 0.00439 0.00109 0.00054 0.00026
+lcvx_N_15_i_13 644 7 7 - 9 9 9 0.00069 0.00060 0.00426 0.00107 0.00050 0.00023
+lcvx_N_15_i_14 644 8 7 - 9 10 10 0.00074 0.00058 0.00441 0.00109 0.00054 0.00025
+lcvx_N_15_i_15 644 8 7 - 10 10 10 0.00080 0.00059 0.00453 0.00117 0.00054 0.00026
+lcvx_N_15_i_16 644 7 7 - 9 9 9 0.00066 0.00058 0.00415 0.00108 0.00049 0.00023
+lcvx_N_15_i_17 644 8 7 - 9 10 10 0.00072 0.00060 0.00455 0.00108 0.00054 0.00025
+lcvx_N_15_i_18 644 8 7 - 9 9 9 0.00073 0.00061 0.00441 0.00110 0.00050 0.00023
+lcvx_N_15_i_19 644 9 7 - 9 10 10 0.00079 0.00060 0.00462 0.00108 0.00053 0.00025
+lcvx_N_50_i_0 2114 9 8 - 13 10 10 0.00282 0.00210 0.01095 0.00445 0.00169 0.00091
+lcvx_N_50_i_1 2114 9 7 - 16 10 10 0.00266 0.00183 0.01104 0.00504 0.00171 0.00091
+lcvx_N_50_i_2 2114 9 7 - 15 10 10 0.00269 0.00190 0.01103 0.00486 0.00170 0.00092
+lcvx_N_50_i_3 2114 10 9 - 16 10 10 0.00298 0.00238 0.01180 0.00504 0.00170 0.00091
+lcvx_N_50_i_4 2114 8 7 - 15 10 10 0.00248 0.00185 0.01045 0.00484 0.00170 0.00092
+lcvx_N_50_i_5 2114 8 7 - 13 10 10 0.00246 0.00187 0.00975 0.00448 0.00169 0.00091
+lcvx_N_50_i_6 2114 9 7 - 16 10 10 0.00268 0.00186 0.01130 0.00502 0.00171 0.00091
+lcvx_N_50_i_7 2114 8 8 - 13 10 10 0.00247 0.00224 0.01052 0.00445 0.00171 0.00091
+lcvx_N_50_i_8 2114 10 8 - 16 10 10 0.00281 0.00218 0.01099 0.00502 0.00169 0.00090
+lcvx_N_50_i_9 2114 10 8 - 16 10 10 0.00284 0.00220 0.01120 0.00504 0.00168 0.00093
+lcvx_N_50_i_10 2114 10 8 - 16 10 10 0.00283 0.00219 0.01113 0.00502 0.00171 0.00092
+lcvx_N_50_i_11 2114 10 8 - 15 10 10 0.00298 0.00209 0.01104 0.00487 0.00170 0.00090
+lcvx_N_50_i_12 2114 9 7 - 15 10 10 0.00262 0.00187 0.01087 0.00487 0.00172 0.00092
+lcvx_N_50_i_13 2114 9 8 - 14 10 10 0.00280 0.00205 0.01061 0.00464 0.00170 0.00091
+lcvx_N_50_i_14 2114 10 7 - 16 10 10 0.00282 0.00186 0.01116 0.00506 0.00169 0.00091
+lcvx_N_50_i_15 2114 9 8 - 15 10 10 0.00263 0.00215 0.01134 0.00482 0.00173 0.00090
+lcvx_N_50_i_16 2114 9 7 - 15 10 10 0.00267 0.00188 0.01109 0.00485 0.00172 0.00089
+lcvx_N_50_i_17 2114 10 9 - 16 10 10 0.00298 0.00240 0.01185 0.00501 0.00169 0.00091
+lcvx_N_50_i_18 2114 9 8 - 13 10 10 0.00267 0.00211 0.01076 0.00447 0.00172 0.00090
+lcvx_N_50_i_19 2114 12 9 - 17 10 10 0.00335 0.00235 0.01167 0.00525 0.00168 0.00091
+lcvx_N_75_i_0 3164 10 8 - 16 10 10 0.00423 0.00324 0.01509 0.00753 0.00247 0.00141
+lcvx_N_75_i_1 3164 10 9 - 17 10 10 0.00441 0.00350 0.01635 0.00782 0.00248 0.00142
+lcvx_N_75_i_2 3164 10 8 - 15 10 10 0.00447 0.00309 0.01605 0.00729 0.00249 0.00145
+lcvx_N_75_i_3 3164 8 8 - 15 10 10 0.00371 0.00332 0.01549 0.00725 0.00249 0.00141
+lcvx_N_75_i_4 3164 10 9 - 17 10 10 0.00442 0.00351 0.01552 0.00779 0.00250 0.00145
+lcvx_N_75_i_5 3164 9 8 - 16 10 10 0.00396 0.00327 0.01546 0.00753 0.00250 0.00143
+lcvx_N_75_i_6 3164 8 8 - 12 10 10 0.00365 0.00323 0.01358 0.00646 0.00249 0.00141
+lcvx_N_75_i_7 3164 11 9 - 17 10 10 0.00471 0.00349 0.01546 0.00779 0.00251 0.00143
+lcvx_N_75_i_8 3164 10 8 - 15 10 10 0.00446 0.00310 0.01592 0.00723 0.00249 0.00142
+lcvx_N_75_i_9 3164 11 9 - 17 10 10 0.00474 0.00348 0.01640 0.00778 0.00252 0.00136
+lcvx_N_75_i_10 3164 10 8 - 16 10 10 0.00419 0.00322 0.01604 0.00751 0.00250 0.00142
+lcvx_N_75_i_11 3164 8 8 - 16 10 10 0.00366 0.00315 0.01535 0.00751 0.00247 0.00141
+lcvx_N_75_i_12 3164 11 9 - 17 10 10 0.00479 0.00350 0.01714 0.00779 0.00248 0.00141
+lcvx_N_75_i_13 3164 9 8 - 13 10 10 0.00421 0.00300 0.01504 0.00671 0.00248 0.00141
+lcvx_N_75_i_14 3164 10 9 - 17 10 10 0.00455 0.00353 0.01655 0.00779 0.00248 0.00140
+lcvx_N_75_i_15 3164 10 9 - 16 10 10 0.00442 0.00352 0.01661 0.00752 0.00250 0.00147
+lcvx_N_75_i_16 3164 10 8 - 16 10 10 0.00447 0.00308 0.01539 0.00752 0.00250 0.00142
+lcvx_N_75_i_17 3164 11 9 - 17 10 10 0.00473 0.00352 0.01640 0.00779 0.00251 0.00142
+lcvx_N_75_i_18 3164 10 8 - 17 10 10 0.00419 0.00326 0.01578 0.00779 0.00248 0.00141
+lcvx_N_75_i_19 3164 9 8 - 16 10 10 0.00384 0.00309 0.01508 0.00753 0.00261 0.00141
+lcvx_N_100_i_0 4214 9 8 - 16 10 10 0.00550 0.00424 0.01873 0.00996 0.00341 0.00200
+lcvx_N_100_i_1 4214 10 9 - 16 10 10 0.00612 0.00462 0.01938 0.01000 0.00336 0.00196
+lcvx_N_100_i_2 4214 12 9 - 17 11 11 0.00680 0.00450 0.02016 0.01029 0.00365 0.00214
+lcvx_N_100_i_3 4214 10 9 - 15 10 10 0.00618 0.00471 0.01943 0.01001 0.00344 0.00190
+lcvx_N_100_i_4 4214 8 8 - 15 10 10 0.00454 0.00416 0.01976 0.00965 0.00335 0.00195
+lcvx_N_100_i_5 4214 11 9 - 17 10 10 0.00655 0.00469 0.01987 0.01032 0.00338 0.00195
+lcvx_N_100_i_6 4214 10 9 - 16 10 10 0.00596 0.00461 0.01895 0.01004 0.00337 0.00198
+lcvx_N_100_i_7 4214 11 9 - 18 11 11 0.00625 0.00447 0.01985 0.01070 0.00365 0.00210
+lcvx_N_100_i_8 4214 9 8 - 15 10 10 0.00530 0.00422 0.01903 0.01003 0.00333 0.00195
+lcvx_N_100_i_9 4214 9 8 - 16 10 10 0.00495 0.00414 0.01832 0.00998 0.00339 0.00198
+lcvx_N_100_i_10 4214 10 8 - 16 10 10 0.00580 0.00405 0.01931 0.01000 0.00335 0.00196
+lcvx_N_100_i_11 4214 10 8 - 16 10 10 0.00579 0.00397 0.02012 0.00996 0.00346 0.00205
+lcvx_N_100_i_12 4214 10 9 - 17 10 10 0.00606 0.00464 0.01984 0.01033 0.00337 0.00194
+lcvx_N_100_i_13 4214 10 9 - 18 11 11 0.00593 0.00445 0.01944 0.01071 0.00366 0.00215
+lcvx_N_100_i_14 4214 11 9 - 17 11 11 0.00628 0.00446 0.02044 0.01027 0.00366 0.00211
+lcvx_N_100_i_15 4214 10 9 - 17 10 10 0.00579 0.00468 0.01965 0.01036 0.00340 0.00201
+lcvx_N_100_i_16 4214 10 8 - 17 10 10 0.00582 0.00421 0.02082 0.01032 0.00333 0.00197
+lcvx_N_100_i_17 4214 12 9 - 17 10 10 0.00682 0.00464 0.01991 0.01035 0.00343 0.00197
+lcvx_N_100_i_18 4214 10 8 - 18 10 10 0.00577 0.00422 0.01906 0.01075 0.00337 0.00192
+lcvx_N_100_i_19 4214 11 9 - 17 10 10 0.00662 0.00464 0.02034 0.01031 0.00339 0.00194
+lcvx_N_125_i_0 5264 11 9 - 18 11 11 0.00777 0.00565 0.02290 0.01312 0.00445 0.00273
+lcvx_N_125_i_1 5264 11 9 - 17 10 10 0.00765 0.00577 0.02467 0.01285 0.00417 0.00236
+lcvx_N_125_i_2 5264 10 9 - 18 10 10 0.00760 0.00562 0.02316 0.01321 0.00415 0.00242
+lcvx_N_125_i_3 5264 11 9 - 17 10 10 0.00762 0.00579 0.02365 0.01275 0.00418 0.00240
+lcvx_N_125_i_4 5264 9 8 - 15 10 10 0.00653 0.00504 0.02218 0.01198 0.00422 0.00239
+lcvx_N_125_i_5 5264 11 9 - 18 10 10 0.00807 0.00563 0.02226 0.01319 0.00417 0.00239
+lcvx_N_125_i_6 5264 11 9 - 16 10 10 0.00767 0.00584 0.02496 0.01285 0.00425 0.00245
+lcvx_N_125_i_7 5264 11 8 - 18 10 10 0.00750 0.00525 0.02348 0.01317 0.00415 0.00244
+lcvx_N_125_i_8 5264 9 8 - 16 10 10 0.00606 0.00519 0.02307 0.01232 0.00418 0.00238
+lcvx_N_125_i_9 5264 10 9 - 15 10 10 0.00779 0.00588 0.02401 0.01235 0.00415 0.00241
+lcvx_N_125_i_10 5264 9 8 - 17 10 10 0.00613 0.00523 0.02227 0.01273 0.00413 0.00240
+lcvx_N_125_i_11 5264 8 8 - 13 10 10 0.00577 0.00520 0.02197 0.01101 0.00413 0.00246
+lcvx_N_125_i_12 5264 9 8 - 15 10 10 0.00727 0.00507 0.02199 0.01193 0.00418 0.00241
+lcvx_N_125_i_13 5264 11 8 - 18 10 10 0.00708 0.00529 0.02405 0.01320 0.00408 0.00245
+lcvx_N_125_i_14 5264 9 8 - 18 10 10 0.00612 0.00515 0.02389 0.01320 0.00415 0.00238
+lcvx_N_125_i_15 5264 11 9 - 18 10 10 0.00801 0.00569 0.02199 0.01310 0.00414 0.00238
+lcvx_N_125_i_16 5264 11 9 - 17 10 10 0.00773 0.00589 0.02420 0.01281 0.00416 0.00243
+lcvx_N_125_i_17 5264 11 9 - 17 11 11 0.00797 0.00561 0.02357 0.01277 0.00454 0.00263
+lcvx_N_125_i_18 5264 11 9 - 18 10 10 0.00804 0.00554 0.02353 0.01319 0.00417 0.00238
+lcvx_N_125_i_19 5264 11 9 - 17 10 10 0.00829 0.00588 0.02513 0.01278 0.00413 0.00238
+lcvx_N_150_i_0 6314 10 8 - 19 10 - 0.00803 0.00623 0.02961 0.01603 0.00504 -
+lcvx_N_150_i_1 6314 11 9 - 18 11 - 0.00992 0.00671 0.02842 0.01561 0.00542 -
+lcvx_N_150_i_2 6314 9 8 - 15 10 - 0.00822 0.00598 0.02509 0.01408 0.00498 -
+lcvx_N_150_i_3 6314 10 9 - 16 10 - 0.00937 0.00697 0.02804 0.01450 0.00513 -
+lcvx_N_150_i_4 6314 9 8 - 15 10 - 0.00835 0.00600 0.02466 0.01403 0.00505 -
+lcvx_N_150_i_5 6314 11 10 - 17 11 - 0.00955 0.00737 0.02743 0.01504 0.00544 -
+lcvx_N_150_i_6 6314 8 8 - 13 10 - 0.00682 0.00627 0.02362 0.01299 0.00504 -
+lcvx_N_150_i_7 6314 9 8 - 17 10 - 0.00737 0.00600 0.02651 0.01505 0.00503 -
+lcvx_N_150_i_8 6314 11 9 - 17 10 - 0.00936 0.00723 0.02915 0.01502 0.00526 -
+lcvx_N_150_i_9 6314 12 11 - 20 11 - 0.01045 0.00825 0.02870 0.01658 0.00561 -
+lcvx_N_150_i_10 6314 11 9 - 18 10 - 0.00950 0.00699 0.02729 0.01550 0.00506 -
+lcvx_N_150_i_11 6314 10 8 - 16 10 - 0.00888 0.00601 0.02707 0.01451 0.00507 -
+lcvx_N_150_i_12 6314 10 8 - 16 10 - 0.00899 0.00605 0.02689 0.01458 0.00507 -
+lcvx_N_150_i_13 6314 12 9 - 18 11 - 0.01040 0.00675 0.02879 0.01552 0.00547 -
+lcvx_N_150_i_14 6314 9 8 - 15 10 - 0.00795 0.00596 0.02549 0.01397 0.00506 -
+lcvx_N_150_i_15 6314 10 8 - 17 10 - 0.00881 0.00600 0.02704 0.01508 0.00508 -
+lcvx_N_150_i_16 6314 11 9 - 18 10 - 0.00982 0.00678 0.02729 0.01558 0.00507 -
+lcvx_N_150_i_17 6314 9 8 - 18 10 - 0.00799 0.00606 0.02714 0.01560 0.00505 -
+lcvx_N_150_i_18 6314 9 8 - 17 10 - 0.00740 0.00613 0.02876 0.01510 0.00520 -
+lcvx_N_150_i_19 6314 11 9 - 17 10 - 0.00956 0.00722 0.02700 0.01508 0.00503 -
+lcvx_N_200_i_0 8414 10 9 - 17 11 - 0.01181 0.00899 0.03711 0.02025 0.00735 -
+lcvx_N_200_i_1 8414 10 9 - 17 11 - 0.01206 0.00902 0.03977 0.02033 0.00738 -
+lcvx_N_200_i_2 8414 9 9 - 15 11 - 0.01093 0.00935 0.03452 0.01897 0.00721 -
+lcvx_N_200_i_3 8414 12 9 - 18 11 - 0.01347 0.00911 0.03906 0.02090 0.00734 -
+lcvx_N_200_i_4 8414 12 9 - 17 11 - 0.01426 0.00914 0.03792 0.02017 0.00733 -
+lcvx_N_200_i_5 8414 11 9 - 17 10 - 0.01247 0.00955 0.03944 0.02029 0.00674 -
+lcvx_N_200_i_6 8414 11 8 - 17 10 - 0.01161 0.00814 0.03592 0.02027 0.00672 -
+lcvx_N_200_i_7 8414 11 9 - 18 11 - 0.01345 0.00907 0.03841 0.02110 0.00722 -
+lcvx_N_200_i_8 8414 10 8 - 17 10 - 0.01160 0.00820 0.03577 0.02033 0.00677 -
+lcvx_N_200_i_9 8414 9 8 - 17 10 - 0.01001 0.00822 0.03583 0.02019 0.00668 -
+lcvx_N_200_i_10 8414 11 12 - 18 11 - 0.01301 0.01216 0.03920 0.02107 0.00734 -
+lcvx_N_200_i_11 8414 9 8 - 16 10 - 0.01107 0.00824 0.03434 0.01956 0.00672 -
+lcvx_N_200_i_12 8414 10 8 - 16 10 - 0.01063 0.00807 0.03451 0.01957 0.00675 -
+lcvx_N_200_i_13 8414 10 8 - 17 10 - 0.01078 0.00811 0.03646 0.02030 0.00678 -
+lcvx_N_200_i_14 8414 11 11 - 18 11 - 0.01295 0.01127 0.03857 0.02098 0.00729 -
+lcvx_N_200_i_15 8414 11 9 - 18 10 - 0.01246 0.00952 0.03749 0.02083 0.00690 -
+lcvx_N_200_i_16 8414 9 8 - 18 10 - 0.01008 0.00839 0.03665 0.02101 0.00669 -
+lcvx_N_200_i_17 8414 12 9 - 18 11 - 0.01344 0.00907 0.03711 0.02103 0.00728 -
+lcvx_N_200_i_18 8414 9 8 - 14 10 - 0.01000 0.00806 0.03299 0.01820 0.00669 -
+lcvx_N_200_i_19 8414 12 10 - 17 11 - 0.01413 0.01004 0.03575 0.02032 0.00730 -
+lcvx_N_250_i_0 10514 10 9 - 11 11 - 0.01406 0.01104 0.05053 0.01684 0.00924 -
+lcvx_N_250_i_1 10514 11 10 - 12 11 - 0.01524 0.01236 0.05496 0.01762 0.00913 -
+lcvx_N_250_i_2 10514 11 9 - 12 11 - 0.01541 0.01128 0.05248 0.01766 0.00915 -
+lcvx_N_250_i_3 10514 9 8 - 11 10 - 0.01237 0.01002 0.04312 0.01670 0.00848 -
+lcvx_N_250_i_4 10514 10 8 - 11 10 - 0.01348 0.00992 0.04709 0.01676 0.00847 -
+lcvx_N_250_i_5 10514 11 8 - 12 10 - 0.01450 0.01018 0.04506 0.01763 0.00902 -
+lcvx_N_250_i_6 10514 9 8 - 11 10 - 0.01251 0.01000 0.04508 0.01683 0.00844 -
+lcvx_N_250_i_7 10514 11 8 - 11 10 - 0.01454 0.01008 0.04595 0.01692 0.00840 -
+lcvx_N_250_i_8 10514 12 12 - 12 11 - 0.01624 0.01454 0.05261 0.01774 0.00911 -
+lcvx_N_250_i_9 10514 9 8 - 11 10 - 0.01245 0.01014 0.04380 0.01679 0.00844 -
+lcvx_N_250_i_10 10514 12 9 - 12 10 - 0.01566 0.01161 0.04630 0.01759 0.00843 -
+lcvx_N_250_i_11 10514 11 9 - 11 10 - 0.01453 0.01193 0.04860 0.01690 0.00836 -
+lcvx_N_250_i_12 10514 12 9 - 12 11 - 0.01674 0.01127 0.05007 0.01760 0.00929 -
+lcvx_N_250_i_13 10514 11 8 - 11 10 - 0.01452 0.01019 0.04785 0.01683 0.00843 -
+lcvx_N_250_i_14 10514 10 9 - 11 11 - 0.01442 0.01137 0.05002 0.01687 0.00919 -
+lcvx_N_250_i_15 10514 9 8 - 11 10 - 0.01243 0.01009 0.04669 0.01690 0.00848 -
+lcvx_N_250_i_16 10514 10 8 - 10 10 - 0.01351 0.01007 0.04663 0.01604 0.00848 -
+lcvx_N_250_i_17 10514 10 8 - 11 10 - 0.01347 0.01012 0.04441 0.01690 0.00842 -
+lcvx_N_250_i_18 10514 13 9 - 12 11 - 0.01775 0.01133 0.05219 0.01766 0.00908 -
+lcvx_N_250_i_19 10514 10 9 - 11 10 - 0.01341 0.01172 0.05027 0.01685 0.00841 -
+lcvx_N_300_i_0 12614 11 11 - 12 11 - 0.01898 0.01659 0.06157 0.02122 0.01100 -
+lcvx_N_300_i_1 12614 11 9 - 12 10 - 0.01757 0.01430 0.05777 0.02123 0.01015 -
+lcvx_N_300_i_2 12614 10 11 - 12 11 - 0.01742 0.01652 0.05705 0.02147 0.01101 -
+lcvx_N_300_i_3 12614 12 10 - 13 12 - 0.01939 0.01514 0.05686 0.02239 0.01189 -
+lcvx_N_300_i_4 12614 10 8 - 12 10 - 0.01617 0.01223 0.05519 0.02120 0.01002 -
+lcvx_N_300_i_5 12614 12 10 - 13 11 - 0.01980 0.01521 0.06007 0.02236 0.01102 -
+lcvx_N_300_i_6 12614 10 8 - 11 10 - 0.01609 0.01220 0.05390 0.02013 0.01013 -
+lcvx_N_300_i_7 12614 9 8 - 11 10 - 0.01491 0.01226 0.05744 0.02019 0.01013 -
+lcvx_N_300_i_8 12614 10 9 - 11 11 - 0.01631 0.01384 0.05907 0.02027 0.01105 -
+lcvx_N_300_i_9 12614 11 9 - 12 11 - 0.01744 0.01359 0.05865 0.02138 0.01101 -
+lcvx_N_300_i_10 12614 13 10 - 13 11 - 0.02084 0.01507 0.05715 0.02225 0.01098 -
+lcvx_N_300_i_11 12614 10 9 - 11 11 - 0.01641 0.01386 0.05330 0.02026 0.01103 -
+lcvx_N_300_i_12 12614 11 9 - 12 11 - 0.01768 0.01376 0.05674 0.02117 0.01097 -
+lcvx_N_300_i_13 12614 9 9 - 11 11 - 0.01500 0.01384 0.05338 0.02027 0.01108 -
+lcvx_N_300_i_14 12614 10 9 - 12 10 - 0.01625 0.01423 0.05383 0.02125 0.01011 -
+lcvx_N_300_i_15 12614 11 8 - 11 10 - 0.01745 0.01226 0.05440 0.02019 0.01017 -
+lcvx_N_300_i_16 12614 9 8 - 11 10 - 0.01497 0.01204 0.05560 0.02032 0.01035 -
+lcvx_N_300_i_17 12614 10 8 - 12 10 - 0.01610 0.01207 0.05475 0.02112 0.01015 -
+lcvx_N_300_i_18 12614 9 8 - 11 10 - 0.01488 0.01205 0.05439 0.02027 0.01017 -
+lcvx_N_300_i_19 12614 10 10 - 11 11 - 0.01721 0.01514 0.05864 0.02016 0.01100 -
+lcvx_N_350_i_0 14714 9 9 - 12 11 - 0.01761 0.01590 0.06407 0.02445 0.01288 -
+lcvx_N_350_i_1 14714 11 10 - 12 11 - 0.02119 0.01799 0.06574 0.02439 0.01290 -
+lcvx_N_350_i_2 14714 11 9 - 12 11 - 0.02053 0.01606 0.06838 0.02432 0.01327 -
+lcvx_N_350_i_3 14714 10 8 - 12 10 - 0.01901 0.01428 0.06508 0.02446 0.01181 -
+lcvx_N_350_i_4 14714 9 8 - 11 10 - 0.01757 0.01428 0.06324 0.02319 0.01181 -
+lcvx_N_350_i_5 14714 11 9 - 12 11 - 0.02034 0.01593 0.06593 0.02441 0.01277 -
+lcvx_N_350_i_6 14714 10 10 - 11 11 - 0.01905 0.01783 0.06727 0.02322 0.01285 -
+lcvx_N_350_i_7 14714 11 9 - 12 11 - 0.02053 0.01591 0.06480 0.02435 0.01284 -
+lcvx_N_350_i_8 14714 10 8 - 12 10 - 0.01895 0.01422 0.06198 0.02463 0.01183 -
+lcvx_N_350_i_9 14714 11 9 - 12 11 - 0.02035 0.01588 0.06730 0.02448 0.01286 -
+lcvx_N_350_i_10 14714 9 8 - 11 11 - 0.01758 0.01444 0.06150 0.02327 0.01288 -
+lcvx_N_350_i_11 14714 12 10 - 13 12 - 0.02271 0.01792 0.06807 0.02549 0.01384 -
+lcvx_N_350_i_12 14714 12 10 - 13 12 - 0.02284 0.01772 0.06639 0.02558 0.01382 -
+lcvx_N_350_i_13 14714 11 10 - 13 11 - 0.02159 0.01788 0.06887 0.02561 0.01290 -
+lcvx_N_350_i_14 14714 10 11 - 11 11 - 0.02021 0.01889 0.06600 0.02322 0.01288 -
+lcvx_N_350_i_15 14714 11 9 - 12 11 - 0.02037 0.01601 0.06846 0.02443 0.01286 -
+lcvx_N_350_i_16 14714 11 9 - 12 11 - 0.02056 0.01593 0.06633 0.02416 0.01288 -
+lcvx_N_350_i_17 14714 8 8 - 11 10 - 0.01600 0.01423 0.06098 0.02317 0.01178 -
+lcvx_N_350_i_18 14714 9 11 - 11 11 - 0.01766 0.01929 0.06235 0.02325 0.01283 -
+lcvx_N_350_i_19 14714 10 8 - 12 10 - 0.01902 0.01435 0.06182 0.02450 0.01180 -
+Table 4: Iterations and solver runtimes for group lasso problems
+Iterations Solver Runtime (s)
+Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
+group_lasso_N_1_i_0 761 4 11 - 4 6 6 0.00025 0.00111 0.00564 0.00047 0.00020 0.00007
+group_lasso_N_1_i_1 761 4 11 - 5 6 6 0.00027 0.00109 0.00580 0.00049 0.00020 0.00007
+group_lasso_N_1_i_2 761 4 10 - 5 6 6 0.00027 0.00098 0.00597 0.00049 0.00021 0.00007
+group_lasso_N_1_i_3 761 4 12 - 5 6 6 0.00027 0.00122 0.00577 0.00050 0.00020 0.00007
+group_lasso_N_1_i_4 761 4 11 - 5 6 6 0.00027 0.00110 0.00545 0.00050 0.00020 0.00007
+group_lasso_N_1_i_5 761 4 11 - 5 6 6 0.00026 0.00113 0.00610 0.00048 0.00020 0.00007
+group_lasso_N_1_i_6 761 4 11 - 5 6 6 0.00027 0.00109 0.00588 0.00049 0.00022 0.00006
+group_lasso_N_1_i_7 761 5 10 - 5 6 6 0.00030 0.00099 0.00549 0.00049 0.00021 0.00007
+group_lasso_N_1_i_8 761 4 12 - 5 6 6 0.00026 0.00115 0.00552 0.00051 0.00020 0.00006
+group_lasso_N_1_i_9 761 5 10 - 5 6 6 0.00030 0.00098 0.00571 0.00048 0.00021 0.00007
+group_lasso_N_1_i_10 761 4 9 - 5 6 6 0.00027 0.00091 0.00631 0.00050 0.00020 0.00007
+group_lasso_N_1_i_11 761 4 11 - 4 6 6 0.00027 0.00112 0.00654 0.00047 0.00021 0.00007
+group_lasso_N_1_i_12 761 5 9 - 5 6 6 0.00030 0.00090 0.00634 0.00050 0.00021 0.00006
+group_lasso_N_1_i_13 761 4 11 - 5 6 6 0.00026 0.00108 0.00587 0.00049 0.00020 0.00006
+group_lasso_N_1_i_14 761 4 12 - 5 6 6 0.00026 0.00116 0.00571 0.00049 0.00020 0.00007
+group_lasso_N_1_i_15 761 4 12 - 5 6 6 0.00027 0.00109 0.00592 0.00049 0.00020 0.00006
+group_lasso_N_1_i_16 761 4 11 - 5 6 6 0.00027 0.00107 0.00588 0.00049 0.00020 0.00007
+group_lasso_N_1_i_17 761 4 12 - 5 6 6 0.00027 0.00121 0.00575 0.00050 0.00020 0.00007
+group_lasso_N_1_i_18 761 5 11 - 5 6 6 0.00030 0.00111 0.00571 0.00050 0.00020 0.00006
+group_lasso_N_1_i_19 761 5 10 - 4 6 6 0.00030 0.00098 0.00601 0.00047 0.00020 0.00006
+group_lasso_N_2_i_0 2022 6 11 - 7 7 7 0.00082 0.00257 0.03673 0.00107 0.00057 0.00021
+group_lasso_N_2_i_1 2022 7 11 - 7 7 7 0.00088 0.00254 0.03584 0.00108 0.00056 0.00020
+group_lasso_N_2_i_2 2022 5 11 - 7 7 7 0.00074 0.00260 0.03842 0.00111 0.00057 0.00020
+group_lasso_N_2_i_3 2022 7 12 - 7 8 8 0.00090 0.00278 0.04030 0.00108 0.00062 0.00023
+group_lasso_N_2_i_4 2022 6 11 - 7 7 7 0.00081 0.00254 0.03843 0.00109 0.00059 0.00021
+group_lasso_N_2_i_5 2022 7 11 - 8 8 8 0.00089 0.00249 0.03584 0.00112 0.00061 0.00023
+group_lasso_N_2_i_6 2022 6 11 - 7 7 7 0.00082 0.00256 0.03675 0.00109 0.00057 0.00021
+group_lasso_N_2_i_7 2022 6 10 - 7 7 7 0.00081 0.00239 0.03766 0.00107 0.00062 0.00020
+group_lasso_N_2_i_8 2022 6 10 - 7 7 7 0.00082 0.00239 0.03624 0.00107 0.00058 0.00021
+group_lasso_N_2_i_9 2022 6 11 - 7 7 7 0.00082 0.00241 0.05126 0.00109 0.00056 0.00020
+group_lasso_N_2_i_10 2022 6 11 - 7 6 6 0.00084 0.00250 0.03732 0.00107 0.00059 0.00017
+group_lasso_N_2_i_11 2022 7 11 - 7 7 7 0.00089 0.00242 0.03560 0.00106 0.00071 0.00020
+group_lasso_N_2_i_12 2022 6 10 - 7 7 7 0.00081 0.00236 0.04354 0.00108 0.00057 0.00020
+group_lasso_N_2_i_13 2022 7 13 - 7 7 7 0.00090 0.00305 0.03834 0.00109 0.00058 0.00021
+group_lasso_N_2_i_14 2022 6 12 - 7 7 7 0.00082 0.00271 0.03596 0.00109 0.00058 0.00020
+group_lasso_N_2_i_15 2022 6 11 - 7 7 7 0.00082 0.00249 0.03711 0.00108 0.00056 0.00020
+group_lasso_N_2_i_16 2022 6 11 - 7 7 7 0.00082 0.00256 0.03639 0.00108 0.00060 0.00020
+group_lasso_N_2_i_17 2022 6 11 - 7 7 7 0.00082 0.00251 0.03742 0.00109 0.00061 0.00021
+group_lasso_N_2_i_18 2022 7 11 - 7 7 7 0.00090 0.00249 0.04392 0.00109 0.00055 0.00021
+group_lasso_N_2_i_19 2022 6 10 - 7 7 7 0.00082 0.00236 0.03721 0.00109 0.00056 0.00020
+group_lasso_N_3_i_0 3783 7 12 - 8 8 8 0.00177 0.00642 0.08580 0.00233 0.00135 0.00041
+group_lasso_N_3_i_1 3783 7 12 - 8 8 8 0.00180 0.00665 0.09602 0.00229 0.00138 0.00042
+group_lasso_N_3_i_2 3783 6 11 - 7 7 7 0.00164 0.00624 0.05644 0.00218 0.00133 0.00037
+group_lasso_N_3_i_3 3783 6 11 - 6 7 7 0.00166 0.00655 0.09255 0.00223 0.00130 0.00037
+group_lasso_N_3_i_4 3783 7 12 - 7 8 8 0.00183 0.00650 0.09112 0.00219 0.00136 0.00042
+group_lasso_N_3_i_5 3783 6 12 - 7 7 7 0.00166 0.00618 0.08814 0.00218 0.00125 0.00037
+group_lasso_N_3_i_6 3783 6 12 - 7 7 7 0.00166 0.00679 0.09434 0.00219 0.00130 0.00037
+group_lasso_N_3_i_7 3783 7 12 - 8 8 8 0.00179 0.00663 0.08940 0.00209 0.00136 0.00042
+group_lasso_N_3_i_8 3783 8 12 - 8 8 8 0.00193 0.00672 0.09324 0.00204 0.00141 0.00042
+group_lasso_N_3_i_9 3783 7 12 - 8 7 7 0.00179 0.00674 0.08115 0.00227 0.00129 0.00037
+group_lasso_N_3_i_10 3783 7 12 - 7 8 8 0.00179 0.00644 0.08538 0.00194 0.00141 0.00042
+group_lasso_N_3_i_11 3783 8 12 - 9 8 8 0.00207 0.00663 0.08846 0.00241 0.00138 0.00042
+group_lasso_N_3_i_12 3783 8 12 - 9 8 8 0.00197 0.00664 0.09559 0.00238 0.00136 0.00042
+group_lasso_N_3_i_13 3783 7 12 - 7 7 7 0.00183 0.00669 0.08609 0.00218 0.00130 0.00037
+group_lasso_N_3_i_14 3783 8 12 - 8 8 8 0.00194 0.00646 0.05551 0.00227 0.00135 0.00042
+group_lasso_N_3_i_15 3783 8 12 - 8 8 8 0.00195 0.00625 0.08523 0.00226 0.00143 0.00042
+group_lasso_N_3_i_16 3783 8 12 - 7 8 8 0.00194 0.00618 0.05406 0.00218 0.00135 0.00042
+group_lasso_N_3_i_17 3783 7 11 - 7 7 7 0.00179 0.00600 0.08719 0.00220 0.00127 0.00037
+group_lasso_N_3_i_18 3783 7 12 - 7 7 7 0.00179 0.00651 0.08314 0.00234 0.00128 0.00037
+group_lasso_N_3_i_19 3783 7 12 - 7 8 8 0.00182 0.00658 0.08637 0.00226 0.00133 0.00042
+group_lasso_N_4_i_0 6044 7 11 - 8 8 8 0.00300 0.00908 0.05446 0.00337 0.00234 0.00070
+group_lasso_N_4_i_1 6044 7 11 - 7 7 7 0.00297 0.00955 0.05064 0.00300 0.00221 0.00061
+group_lasso_N_4_i_2 6044 7 11 - 8 8 8 0.00303 0.00940 0.04993 0.00319 0.00207 0.00069
+group_lasso_N_4_i_3 6044 6 11 - 7 7 7 0.00282 0.00981 0.05742 0.00330 0.00219 0.00060
+group_lasso_N_4_i_4 6044 6 10 - 7 7 7 0.00278 0.00832 0.06131 0.00313 0.00219 0.00063
+group_lasso_N_4_i_5 6044 7 11 - 8 8 8 0.00303 0.00939 0.05877 0.00314 0.00232 0.00069
+group_lasso_N_4_i_6 6044 8 11 - 8 9 9 0.00316 0.00966 0.05485 0.00325 0.00246 0.00078
+group_lasso_N_4_i_7 6044 7 11 - 7 8 8 0.00307 0.00955 0.05600 0.00312 0.00230 0.00069
+group_lasso_N_4_i_8 6044 7 12 - 8 7 7 0.00310 0.01008 0.05623 0.00344 0.00220 0.00061
+group_lasso_N_4_i_9 6044 8 12 - 8 8 8 0.00319 0.01056 0.05563 0.00322 0.00231 0.00068
+group_lasso_N_4_i_10 6044 7 12 - 7 7 7 0.00299 0.01032 0.06346 0.00307 0.00221 0.00061
+group_lasso_N_4_i_11 6044 7 11 - 7 7 7 0.00299 0.00932 0.05491 0.00310 0.00218 0.00062
+group_lasso_N_4_i_12 6044 8 12 - 10 8 8 0.00323 0.01024 0.05190 0.00378 0.00230 0.00068
+group_lasso_N_4_i_13 6044 8 12 - 9 8 8 0.00319 0.01032 0.05939 0.00336 0.00233 0.00070
+group_lasso_N_4_i_14 6044 7 11 - 8 8 8 0.00301 0.00938 0.05642 0.00331 0.00231 0.00068
+group_lasso_N_4_i_15 6044 7 11 - 7 7 7 0.00298 0.00909 0.05730 0.00309 0.00218 0.00061
+group_lasso_N_4_i_16 6044 7 12 - 6 8 8 0.00298 0.01051 0.05445 0.00311 0.00231 0.00068
+group_lasso_N_4_i_17 6044 7 12 - 7 7 7 0.00300 0.01066 0.05494 0.00304 0.00219 0.00062
+group_lasso_N_4_i_18 6044 7 12 - 8 8 8 0.00315 0.01057 0.04919 0.00324 0.00233 0.00068
+group_lasso_N_4_i_19 6044 7 11 - 8 8 8 0.00299 0.00936 0.05176 0.00323 0.00230 0.00071
+group_lasso_N_5_i_0 8805 7 12 - 8 8 8 0.00469 0.01341 0.08319 0.00421 0.00384 0.00104
+group_lasso_N_5_i_1 8805 7 12 - 7 8 8 0.00470 0.01325 0.10216 0.00416 0.00379 0.00106
+group_lasso_N_5_i_2 8805 7 12 - 8 8 8 0.00476 0.01535 0.08730 0.00363 0.00379 0.00104
+group_lasso_N_5_i_3 8805 7 12 - 9 8 8 0.00485 0.01440 0.07491 0.00442 0.00378 0.00104
+group_lasso_N_5_i_4 8805 7 12 - 8 8 8 0.00493 0.01482 0.08799 0.00419 0.00376 0.00104
+group_lasso_N_5_i_5 8805 8 12 - 8 8 8 0.00504 0.01447 0.08363 0.00418 0.00378 0.00104
+group_lasso_N_5_i_6 8805 8 13 - 10 9 9 0.00498 0.01613 0.10808 0.00407 0.00419 0.00116
+group_lasso_N_5_i_7 8805 8 13 - 10 9 9 0.00500 0.01470 0.08271 0.00449 0.00397 0.00119
+group_lasso_N_5_i_8 8805 7 12 - 8 8 8 0.00473 0.01473 0.09001 0.00415 0.00381 0.00105
+group_lasso_N_5_i_9 8805 8 12 - 9 9 9 0.00499 0.01269 0.07729 0.00438 0.00396 0.00115
+group_lasso_N_5_i_10 8805 8 12 - 9 8 8 0.00496 0.01467 0.08892 0.00432 0.00383 0.00103
+group_lasso_N_5_i_11 8805 7 12 - 8 8 8 0.00476 0.01321 0.09225 0.00420 0.00383 0.00105
+group_lasso_N_5_i_12 8805 7 12 - 8 8 8 0.00463 0.01325 0.09461 0.00419 0.00390 0.00103
+group_lasso_N_5_i_13 8805 7 12 - 8 8 8 0.00468 0.01545 0.08324 0.00420 0.00380 0.00102
+group_lasso_N_5_i_14 8805 8 13 - 8 8 8 0.00505 0.01502 0.08284 0.00418 0.00384 0.00107
+group_lasso_N_5_i_15 8805 6 11 - 8 7 7 0.00439 0.01318 0.11825 0.00425 0.00364 0.00092
+group_lasso_N_5_i_16 8805 8 13 - 9 8 8 0.00503 0.01559 0.09084 0.00435 0.00376 0.00104
+group_lasso_N_5_i_17 8805 7 12 - 7 8 8 0.00493 0.01263 0.08314 0.00426 0.00381 0.00105
+group_lasso_N_5_i_18 8805 8 12 - 8 8 8 0.00500 0.01342 0.09374 0.00423 0.00381 0.00104
+group_lasso_N_5_i_19 8805 7 12 - 8 8 8 0.00469 0.01493 0.08515 0.00421 0.00381 0.00106
+group_lasso_N_8_i_0 20088 8 12 - 9 8 - 0.01510 0.03081 0.23046 0.00891 0.01268 -
+group_lasso_N_8_i_1 20088 8 12 - 10 9 - 0.01562 0.03244 0.22238 0.00864 0.01326 -
+group_lasso_N_8_i_2 20088 8 11 - 9 8 - 0.01589 0.03120 0.25512 0.00829 0.01277 -
+group_lasso_N_8_i_3 20088 8 12 - 9 9 - 0.01565 0.03183 0.25449 0.00827 0.01317 -
+group_lasso_N_8_i_4 20088 8 13 - 9 9 - 0.01555 0.03495 0.24709 0.00835 0.01317 -
+group_lasso_N_8_i_5 20088 8 12 - 10 9 - 0.01568 0.03265 0.24470 0.00763 0.01316 -
+group_lasso_N_8_i_6 20088 7 11 - 9 8 - 0.01558 0.02985 0.23378 0.00829 0.01273 -
+group_lasso_N_8_i_7 20088 8 12 - 9 8 - 0.01546 0.03117 0.24632 0.00847 0.01270 -
+group_lasso_N_8_i_8 20088 8 13 - 10 9 - 0.01594 0.03331 0.24449 0.00891 0.01317 -
+group_lasso_N_8_i_9 20088 8 12 - 10 8 - 0.01574 0.03230 0.24570 0.00866 0.01274 -
+group_lasso_N_8_i_10 20088 8 12 - 10 8 - 0.01627 0.03130 0.25514 0.00867 0.01271 -
+group_lasso_N_8_i_11 20088 8 12 - 9 8 - 0.01566 0.03195 0.24611 0.00844 0.01276 -
+group_lasso_N_8_i_12 20088 7 11 - 9 8 - 0.01504 0.03068 0.22589 0.00866 0.01274 -
+group_lasso_N_8_i_13 20088 8 12 - 8 8 - 0.01627 0.03083 0.25352 0.00849 0.01269 -
+group_lasso_N_8_i_14 20088 7 12 - 9 8 - 0.01506 0.03242 0.24988 0.00830 0.01274 -
+group_lasso_N_8_i_15 20088 8 12 - 10 8 - 0.01586 0.03222 0.28708 0.00933 0.01270 -
+group_lasso_N_8_i_16 20088 8 12 - 10 8 - 0.01544 0.03166 0.25367 0.00892 0.01281 -
+group_lasso_N_8_i_17 20088 7 12 - 9 8 - 0.01509 0.03256 0.26745 0.00820 0.01271 -
+group_lasso_N_8_i_18 20088 7 12 - 8 8 - 0.01508 0.03457 0.22807 0.00801 0.01269 -
+group_lasso_N_8_i_19 20088 8 11 - 9 8 - 0.01641 0.03025 0.24183 0.00822 0.01267 -
+group_lasso_N_10_i_0 30110 8 12 - 10 8 - 0.02756 0.04259 0.42095 0.01052 0.02289 -
+group_lasso_N_10_i_1 30110 8 13 - 10 8 - 0.02695 0.04466 0.49732 0.01039 0.02288 -
+group_lasso_N_10_i_2 30110 8 13 - 10 8 - 0.02863 0.04405 0.38538 0.01061 0.02286 -
+group_lasso_N_10_i_3 30110 8 12 - 9 8 - 0.02738 0.04291 0.44772 0.01012 0.02288 -
+group_lasso_N_10_i_4 30110 8 12 - 10 9 - 0.02784 0.04340 0.39140 0.01026 0.02376 -
+group_lasso_N_10_i_5 30110 8 12 - 10 8 - 0.02721 0.04283 0.44021 0.01029 0.02308 -
+group_lasso_N_10_i_6 30110 8 13 - 11 9 - 0.02714 0.04510 0.38498 0.01081 0.02469 -
+group_lasso_N_10_i_7 30110 8 12 - 10 8 - 0.02728 0.03974 0.43310 0.01084 0.02287 -
+group_lasso_N_10_i_8 30110 8 12 - 9 8 - 0.02705 0.04384 0.46175 0.00990 0.02397 -
+group_lasso_N_10_i_9 30110 7 12 - 9 8 - 0.02639 0.04041 0.47321 0.00964 0.02276 -
+group_lasso_N_10_i_10 30110 8 13 - 10 9 - 0.02693 0.04484 0.38864 0.01007 0.02359 -
+group_lasso_N_10_i_11 30110 7 12 - 8 8 - 0.02591 0.04340 0.40515 0.00960 0.02277 -
+group_lasso_N_10_i_12 30110 8 12 - 9 8 - 0.02906 0.03989 0.39042 0.00981 0.02378 -
+group_lasso_N_10_i_13 30110 8 12 - 9 8 - 0.02754 0.04088 0.38960 0.00992 0.02323 -
+group_lasso_N_10_i_14 30110 8 12 - 10 8 - 0.02794 0.04135 0.47359 0.01084 0.02279 -
+group_lasso_N_10_i_15 30110 8 13 - 10 8 - 0.02995 0.04443 0.45111 0.01077 0.02382 -
+group_lasso_N_10_i_16 30110 8 12 - 9 8 - 0.02747 0.04256 0.45569 0.01025 0.02287 -
+group_lasso_N_10_i_17 30110 8 12 - 10 8 - 0.02945 0.04246 0.40807 0.01069 0.02267 -
+group_lasso_N_10_i_18 30110 8 12 - 9 8 - 0.02665 0.04161 0.40308 0.01116 0.02278 -
+group_lasso_N_10_i_19 30110 7 11 - 8 8 - 0.02629 0.03774 0.43661 0.00941 0.02381 -
+group_lasso_N_12_i_0 42132 8 12 - 9 8 - 0.04325 0.05675 0.72594 0.01384 0.03507 -
+group_lasso_N_12_i_1 42132 8 12 - 8 8 - 0.04418 0.05847 0.72517 0.01283 0.03495 -
+group_lasso_N_12_i_2 42132 7 12 - 8 7 - 0.04241 0.05995 0.72523 0.01339 0.03310 -
+group_lasso_N_12_i_3 42132 8 13 - 8 8 - 0.04294 0.06362 0.71941 0.01290 0.03478 -
+group_lasso_N_12_i_4 42132 8 12 - 8 9 - 0.04232 0.05920 0.80056 0.01300 0.03589 -
+group_lasso_N_12_i_5 42132 8 12 - 9 8 - 0.04281 0.05680 0.73030 0.01386 0.03468 -
+group_lasso_N_12_i_6 42132 8 12 - 8 8 - 0.04788 0.06034 0.65356 0.01307 0.03501 -
+group_lasso_N_12_i_7 42132 8 12 - 8 8 - 0.04578 0.06083 0.67306 0.01335 0.03442 -
+group_lasso_N_12_i_8 42132 8 12 - 8 8 - 0.04329 0.06219 0.66407 0.01339 0.03478 -
+group_lasso_N_12_i_9 42132 8 12 - 8 8 - 0.04575 0.05707 0.72486 0.01280 0.03471 -
+group_lasso_N_12_i_10 42132 8 12 - 8 8 - 0.04532 0.05731 0.64343 0.01296 0.03480 -
+group_lasso_N_12_i_11 42132 8 12 - 8 8 - 0.04423 0.05731 0.58129 0.01348 0.03428 -
+group_lasso_N_12_i_12 42132 8 12 - 8 8 - 0.04310 0.06025 0.68214 0.01284 0.03404 -
+group_lasso_N_12_i_13 42132 8 12 - 8 8 - 0.04308 0.06026 0.80281 0.01309 0.03423 -
+group_lasso_N_12_i_14 42132 8 12 - 8 8 - 0.04317 0.05748 0.69917 0.01355 0.03446 -
+group_lasso_N_12_i_15 42132 8 12 - 9 8 - 0.04401 0.05937 0.65879 0.01412 0.03489 -
+group_lasso_N_12_i_16 42132 8 12 - 8 8 - 0.04318 0.05796 0.73036 0.01336 0.03458 -
+group_lasso_N_12_i_17 42132 8 12 - 8 8 - 0.04254 0.06040 0.66171 0.01297 0.03405 -
+group_lasso_N_12_i_18 42132 8 12 - 9 8 - 0.04516 0.06173 0.86410 0.01419 0.03433 -
+group_lasso_N_12_i_19 42132 8 12 - 8 8 - 0.04368 0.05717 0.72283 0.01353 0.03530 -
+group_lasso_N_14_i_0 56154 7 11 - 8 8 - 0.06224 0.07874 1.23566 0.01759 0.04544 -
+group_lasso_N_14_i_1 56154 8 12 - 8 8 - 0.06245 0.08160 0.96851 0.01818 0.04671 -
+group_lasso_N_14_i_2 56154 8 12 - 9 8 - 0.06498 0.08289 0.97849 0.01848 0.04671 -
+group_lasso_N_14_i_3 56154 8 12 - 8 8 - 0.06662 0.08547 1.17808 0.01798 0.04580 -
+group_lasso_N_14_i_4 56154 8 12 - 8 8 - 0.06497 0.08176 1.22698 0.01762 0.04588 -
+group_lasso_N_14_i_5 56154 7 11 - 7 8 - 0.06280 0.07908 1.09989 0.01686 0.04782 -
+group_lasso_N_14_i_6 56154 8 13 - 8 9 - 0.06354 0.08527 1.09452 0.01801 0.04769 -
+group_lasso_N_14_i_7 56154 8 13 - 8 9 - 0.06345 0.08543 1.11042 0.01851 0.04802 -
+group_lasso_N_14_i_8 56154 8 12 - 8 8 - 0.06339 0.08132 0.99903 0.01885 0.04614 -
+group_lasso_N_14_i_9 56154 7 11 - 8 8 - 0.06288 0.07812 1.26237 0.01756 0.04631 -
+group_lasso_N_14_i_10 56154 7 11 - 8 8 - 0.06023 0.07890 1.10284 0.01826 0.04638 -
+group_lasso_N_14_i_11 56154 8 12 - 8 8 - 0.06586 0.08050 0.99469 0.01772 0.04630 -
+group_lasso_N_14_i_12 56154 7 12 - 8 8 - 0.06166 0.08194 1.15584 0.01847 0.04567 -
+group_lasso_N_14_i_13 56154 8 12 - 8 8 - 0.06373 0.08154 1.07342 0.02029 0.04599 -
+group_lasso_N_14_i_14 56154 7 12 - 7 8 - 0.06258 0.08137 1.00195 0.01702 0.04789 -
+group_lasso_N_14_i_15 56154 7 12 - 7 8 - 0.06407 0.08213 1.11372 0.01982 0.04643 -
+group_lasso_N_14_i_16 56154 8 12 - 8 8 - 0.06230 0.08161 1.06719 0.01917 0.04592 -
+group_lasso_N_14_i_17 56154 8 13 - 9 8 - 0.06332 0.08538 1.08016 0.01906 0.04597 -
+group_lasso_N_14_i_18 56154 7 12 - 8 8 - 0.06121 0.08230 0.97765 0.01952 0.04903 -
+group_lasso_N_14_i_19 56154 8 12 - 8 8 - 0.06337 0.08094 1.21798 0.01775 0.04637 -
+group_lasso_N_16_i_0 72176 7 12 - 8 8 - 0.08818 0.10965 1.53172 0.02212 0.06506 -
+group_lasso_N_16_i_1 72176 8 12 - 10 8 - 0.09073 0.11368 1.73110 0.02489 0.06656 -
+group_lasso_N_16_i_2 72176 7 13 - 8 8 - 0.08744 0.11966 1.39775 0.02242 0.06552 -
+group_lasso_N_16_i_3 72176 7 12 - 8 8 - 0.08782 0.11273 1.61282 0.02303 0.06462 -
+group_lasso_N_16_i_4 72176 7 12 - 8 8 - 0.09297 0.11399 1.57281 0.02254 0.06636 -
+group_lasso_N_16_i_5 72176 7 12 - 8 8 - 0.08922 0.11064 1.39139 0.02248 0.06505 -
+group_lasso_N_16_i_6 72176 8 12 - 8 8 - 0.09090 0.11306 1.59569 0.02503 0.06585 -
+group_lasso_N_16_i_7 72176 7 12 - 8 8 - 0.08715 0.11313 1.59741 0.02194 0.06630 -
+group_lasso_N_16_i_8 72176 7 12 - 8 8 - 0.08660 0.10653 1.56979 0.02195 0.06473 -
+group_lasso_N_16_i_9 72176 8 12 - 8 8 - 0.09179 0.11243 1.58186 0.02176 0.06512 -
+group_lasso_N_16_i_10 72176 8 12 - 8 8 - 0.09056 0.11126 1.44625 0.02328 0.06416 -
+group_lasso_N_16_i_11 72176 8 12 - 9 8 - 0.09036 0.11116 1.60883 0.02410 0.06535 -
+group_lasso_N_16_i_12 72176 8 12 - 8 8 - 0.09242 0.11283 1.61328 0.02212 0.06555 -
+group_lasso_N_16_i_13 72176 8 12 - 9 8 - 0.09120 0.11232 1.57203 0.02442 0.06505 -
+group_lasso_N_16_i_14 72176 8 12 - 8 8 - 0.09242 0.11370 1.44119 0.02223 0.06555 -
+group_lasso_N_16_i_15 72176 8 12 - 8 8 - 0.09026 0.11228 1.41785 0.02290 0.06527 -
+group_lasso_N_16_i_16 72176 8 12 - 8 8 - 0.09227 0.11340 1.45983 0.02243 0.06665 -
+group_lasso_N_16_i_17 72176 7 12 - 8 8 - 0.08756 0.11138 1.42790 0.02237 0.06617 -
+group_lasso_N_16_i_18 72176 8 12 - 8 8 - 0.09079 0.11191 1.61307 0.02328 0.06484 -
+group_lasso_N_16_i_19 72176 7 12 - 8 8 - 0.09151 0.11219 1.59887 0.02206 0.06481 -
+Table 5: Iterations and solver runtimes for portfolio optimization problems
+Iterations Solver Runtime (s)
+Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
+portfolio_N_2_i_0 804 9 12 14 11 9 9 0.00041 0.00087 0.00205 0.00100 0.00025 0.00012
+portfolio_N_2_i_1 804 9 14 13 11 8 8 0.00041 0.00096 0.00200 0.00101 0.00024 0.00010
+portfolio_N_2_i_2 804 8 12 12 11 8 8 0.00038 0.00089 0.00204 0.00100 0.00024 0.00010
+portfolio_N_2_i_3 804 8 15 16 11 8 8 0.00038 0.00110 0.00203 0.00101 0.00024 0.00011
+portfolio_N_2_i_4 804 9 14 15 9 8 8 0.00041 0.00099 0.00207 0.00092 0.00024 0.00011
+portfolio_N_2_i_5 804 10 14 14 13 9 9 0.00046 0.00097 0.00203 0.00110 0.00026 0.00011
+portfolio_N_2_i_6 804 10 18 11 10 10 10 0.00045 0.00133 0.00198 0.00096 0.00028 0.00013
+portfolio_N_2_i_7 804 8 10 11 9 7 7 0.00037 0.00078 0.00202 0.00091 0.00021 0.00009
+portfolio_N_2_i_8 804 9 17 13 9 9 9 0.00041 0.00134 0.00203 0.00091 0.00025 0.00011
+portfolio_N_2_i_9 804 16 13 15 12 9 9 0.00064 0.00090 0.00203 0.00104 0.00026 0.00011
+portfolio_N_2_i_10 804 10 13 17 13 10 10 0.00045 0.00089 0.00203 0.00110 0.00027 0.00013
+portfolio_N_2_i_11 804 10 12 14 13 12 12 0.00045 0.00087 0.00205 0.00109 0.00031 0.00015
+portfolio_N_2_i_12 804 10 11 15 11 10 10 0.00044 0.00084 0.00206 0.00101 0.00027 0.00013
+portfolio_N_2_i_13 804 9 12 13 11 8 8 0.00041 0.00088 0.00202 0.00100 0.00023 0.00010
+portfolio_N_2_i_14 804 9 14 14 11 9 9 0.00040 0.00101 0.00201 0.00103 0.00029 0.00011
+portfolio_N_2_i_15 804 9 11 14 11 8 8 0.00041 0.00081 0.00204 0.00103 0.00023 0.00010
+portfolio_N_2_i_16 804 10 14 12 10 8 8 0.00044 0.00101 0.00204 0.00095 0.00024 0.00011
+portfolio_N_2_i_17 804 8 11 11 8 8 8 0.00038 0.00084 0.00198 0.00085 0.00024 0.00010
+portfolio_N_2_i_18 804 9 13 12 11 9 9 0.00041 0.00097 0.00202 0.00102 0.00025 0.00011
+portfolio_N_2_i_19 804 10 11 12 9 8 8 0.00045 0.00083 0.00203 0.00092 0.00023 0.00010
+portfolio_N_4_i_0 2008 10 13 14 11 9 9 0.00110 0.00194 0.00230 0.00184 0.00056 0.00028
+portfolio_N_4_i_1 2008 10 14 14 12 11 11 0.00109 0.00212 0.00231 0.00191 0.00062 0.00034
+portfolio_N_4_i_2 2008 11 15 14 12 10 10 0.00116 0.00268 0.00232 0.00189 0.00059 0.00031
+portfolio_N_4_i_3 2008 11 16 15 13 9 9 0.00116 0.00227 0.00230 0.00198 0.00055 0.00027
+portfolio_N_4_i_4 2008 9 17 13 11 8 8 0.00100 0.00249 0.00229 0.00182 0.00049 0.00025
+portfolio_N_4_i_5 2008 11 13 14 12 9 9 0.00117 0.00202 0.00230 0.00193 0.00057 0.00027
+portfolio_N_4_i_6 2008 10 14 13 11 9 9 0.00106 0.00210 0.00226 0.00186 0.00055 0.00028
+portfolio_N_4_i_7 2008 11 15 13 11 9 9 0.00112 0.00221 0.00228 0.00185 0.00057 0.00028
+portfolio_N_4_i_8 2008 11 13 12 10 9 9 0.00113 0.00192 0.00227 0.00176 0.00055 0.00028
+portfolio_N_4_i_9 2008 11 16 14 13 10 10 0.00116 0.00245 0.00230 0.00200 0.00065 0.00030
+portfolio_N_4_i_10 2008 11 18 16 14 10 10 0.00118 0.00256 0.00231 0.00209 0.00058 0.00031
+portfolio_N_4_i_11 2008 9 15 13 11 9 9 0.00099 0.00230 0.00228 0.00183 0.00059 0.00028
+portfolio_N_4_i_12 2008 12 14 12 10 9 9 0.00122 0.00209 0.00227 0.00176 0.00056 0.00027
+portfolio_N_4_i_13 2008 10 15 16 12 9 9 0.00106 0.00239 0.00231 0.00192 0.00054 0.00027
+portfolio_N_4_i_14 2008 10 17 15 14 9 9 0.00108 0.00240 0.00231 0.00208 0.00055 0.00028
+portfolio_N_4_i_15 2008 11 12 13 11 10 10 0.00114 0.00190 0.00226 0.00183 0.00059 0.00031
+portfolio_N_4_i_16 2008 11 17 14 12 10 10 0.00116 0.00277 0.00231 0.00192 0.00059 0.00031
+portfolio_N_4_i_17 2008 9 13 12 10 9 9 0.00101 0.00194 0.00225 0.00174 0.00054 0.00028
+portfolio_N_4_i_18 2008 11 14 14 13 9 9 0.00117 0.00204 0.00231 0.00197 0.00055 0.00027
+portfolio_N_4_i_19 2008 9 19 14 11 8 8 0.00099 0.00298 0.00231 0.00181 0.00052 0.00025
+portfolio_N_6_i_0 3612 13 16 14 11 11 11 0.00223 0.00405 0.00261 0.00277 0.00112 0.00058
+portfolio_N_6_i_1 3612 14 16 14 10 8 8 0.00233 0.00418 0.00267 0.00270 0.00093 0.00043
+portfolio_N_6_i_2 3612 12 14 15 12 10 10 0.00211 0.00345 0.00269 0.00294 0.00111 0.00053
+portfolio_N_6_i_3 3612 15 18 15 13 10 10 0.00247 0.00431 0.00268 0.00302 0.00105 0.00052
+portfolio_N_6_i_4 3612 12 14 14 13 11 11 0.00208 0.00370 0.00264 0.00304 0.00112 0.00058
+portfolio_N_6_i_5 3612 11 13 12 10 9 9 0.00198 0.00332 0.00257 0.00266 0.00098 0.00048
+portfolio_N_6_i_6 3612 10 15 13 11 9 9 0.00186 0.00390 0.00261 0.00278 0.00100 0.00048
+portfolio_N_6_i_7 3612 12 14 14 12 9 9 0.00212 0.00340 0.00262 0.00292 0.00100 0.00047
+portfolio_N_6_i_8 3612 12 16 15 13 9 9 0.00210 0.00398 0.00269 0.00303 0.00098 0.00048
+portfolio_N_6_i_9 3612 13 15 15 14 10 10 0.00222 0.00392 0.00271 0.00317 0.00109 0.00053
+portfolio_N_6_i_10 3612 11 15 14 12 10 10 0.00195 0.00369 0.00266 0.00287 0.00105 0.00052
+portfolio_N_6_i_11 3612 11 14 13 11 9 9 0.00196 0.00359 0.00261 0.00279 0.00100 0.00047
+portfolio_N_6_i_12 3612 12 15 16 11 11 11 0.00211 0.00410 0.00265 0.00277 0.00115 0.00058
+portfolio_N_6_i_13 3612 11 15 14 12 10 10 0.00197 0.00369 0.00266 0.00288 0.00105 0.00053
+portfolio_N_6_i_14 3612 13 16 14 13 10 10 0.00226 0.00394 0.00262 0.00301 0.00107 0.00054
+portfolio_N_6_i_15 3612 12 17 14 12 10 10 0.00216 0.00417 0.00265 0.00289 0.00105 0.00053
+portfolio_N_6_i_16 3612 11 17 14 12 9 9 0.00200 0.00421 0.00265 0.00289 0.00097 0.00048
+portfolio_N_6_i_17 3612 13 15 16 15 10 10 0.00222 0.00358 0.00274 0.00323 0.00109 0.00052
+portfolio_N_6_i_18 3612 12 20 15 14 10 10 0.00210 0.00528 0.00264 0.00313 0.00106 0.00053
+portfolio_N_6_i_19 3612 13 13 14 11 9 9 0.00217 0.00344 0.00264 0.00278 0.00100 0.00047
+portfolio_N_8_i_0 5616 11 15 14 11 12 12 0.00332 0.00572 0.00302 0.00373 0.00166 0.00098
+portfolio_N_8_i_1 5616 12 15 14 13 11 11 0.00353 0.00585 0.00306 0.00405 0.00166 0.00089
+portfolio_N_8_i_2 5616 12 16 14 12 12 12 0.00346 0.00602 0.00308 0.00391 0.00163 0.00098
+portfolio_N_8_i_3 5616 13 15 16 14 11 11 0.00371 0.00564 0.00317 0.00423 0.00146 0.00095
+portfolio_N_8_i_4 5616 13 16 15 12 11 11 0.00368 0.00639 0.00312 0.00389 0.00156 0.00093
+portfolio_N_8_i_5 5616 12 15 13 13 10 10 0.00345 0.00587 0.00306 0.00405 0.00141 0.00084
+portfolio_N_8_i_6 5616 14 17 13 12 10 10 0.00386 0.00634 0.00303 0.00390 0.00136 0.00086
+portfolio_N_8_i_7 5616 12 14 16 15 9 9 0.00340 0.00550 0.00324 0.00435 0.00131 0.00077
+portfolio_N_8_i_8 5616 13 17 15 14 11 11 0.00370 0.00649 0.00310 0.00423 0.00156 0.00089
+portfolio_N_8_i_9 5616 14 18 14 13 10 10 0.00393 0.00621 0.00308 0.00405 0.00147 0.00085
+portfolio_N_8_i_10 5616 13 18 16 15 10 10 0.00374 0.00629 0.00325 0.00439 0.00136 0.00087
+portfolio_N_8_i_11 5616 12 18 15 15 10 10 0.00345 0.00668 0.00315 0.00436 0.00142 0.00085
+portfolio_N_8_i_12 5616 11 15 15 12 9 9 0.00325 0.00565 0.00313 0.00390 0.00136 0.00073
+portfolio_N_8_i_13 5616 14 17 15 13 10 10 0.00389 0.00647 0.00316 0.00406 0.00147 0.00083
+portfolio_N_8_i_14 5616 12 17 15 14 10 10 0.00349 0.00627 0.00313 0.00423 0.00138 0.00085
+portfolio_N_8_i_15 5616 14 18 15 15 10 10 0.00389 0.00672 0.00305 0.00437 0.00147 0.00084
+portfolio_N_8_i_16 5616 14 17 16 14 12 12 0.00387 0.00617 0.00315 0.00422 0.00168 0.00099
+portfolio_N_8_i_17 5616 12 15 15 13 9 9 0.00347 0.00575 0.00308 0.00404 0.00134 0.00075
+portfolio_N_8_i_18 5616 12 14 14 13 10 10 0.00352 0.00551 0.00308 0.00404 0.00142 0.00083
+portfolio_N_8_i_19 5616 12 16 16 14 10 10 0.00350 0.00603 0.00318 0.00420 0.00141 0.00083
+portfolio_N_10_i_0 8020 13 18 14 13 9 9 0.00581 0.01020 0.00366 0.00525 0.00185 0.00149
+portfolio_N_10_i_1 8020 12 19 13 13 10 10 0.00556 0.01172 0.00357 0.00521 0.00200 0.00166
+portfolio_N_10_i_2 8020 14 18 14 12 10 10 0.00613 0.00946 0.00364 0.00502 0.00198 0.00169
+portfolio_N_10_i_3 8020 14 16 15 14 10 10 0.00606 0.00823 0.00372 0.00544 0.00199 0.00165
+portfolio_N_10_i_4 8020 12 16 15 13 9 9 0.00546 0.00860 0.00371 0.00523 0.00186 0.00149
+portfolio_N_10_i_5 8020 11 15 14 12 11 11 0.00518 0.00977 0.00360 0.00504 0.00218 0.00180
+portfolio_N_10_i_6 8020 14 17 16 14 11 11 0.00605 0.00913 0.00381 0.00543 0.00216 0.00181
+portfolio_N_10_i_7 8020 14 15 15 13 12 12 0.00602 0.00865 0.00370 0.00524 0.00235 0.00194
+portfolio_N_10_i_8 8020 13 16 17 14 11 11 0.00580 0.00873 0.00385 0.00544 0.00217 0.00179
+portfolio_N_10_i_9 8020 12 15 15 12 11 11 0.00557 0.00826 0.00372 0.00501 0.00217 0.00177
+portfolio_N_10_i_10 8020 13 19 14 14 11 11 0.00582 0.00975 0.00364 0.00540 0.00215 0.00182
+portfolio_N_10_i_11 8020 14 18 16 16 10 10 0.00606 0.00962 0.00376 0.00583 0.00198 0.00167
+portfolio_N_10_i_12 8020 13 18 14 13 12 12 0.00575 0.00940 0.00364 0.00523 0.00235 0.00200
+portfolio_N_10_i_13 8020 13 13 13 13 9 9 0.00570 0.00734 0.00357 0.00523 0.00187 0.00150
+portfolio_N_10_i_14 8020 13 18 14 13 10 10 0.00587 0.00992 0.00366 0.00527 0.00199 0.00163
+portfolio_N_10_i_15 8020 12 16 15 12 11 11 0.00554 0.00864 0.00368 0.00502 0.00218 0.00178
+portfolio_N_10_i_16 8020 13 18 16 13 12 12 0.00580 0.00995 0.00379 0.00522 0.00236 0.00192
+portfolio_N_10_i_17 8020 13 17 15 14 10 10 0.00585 0.00932 0.00372 0.00542 0.00200 0.00168
+portfolio_N_10_i_18 8020 13 18 15 13 10 10 0.00576 0.00967 0.00370 0.00525 0.00198 0.00164
+portfolio_N_10_i_19 8020 13 15 15 15 12 12 0.00570 0.00853 0.00370 0.00564 0.00236 0.00194
+portfolio_N_15_i_0 15780 13 18 14 13 11 - 0.01637 0.01628 0.00539 0.00874 0.00454 -
+portfolio_N_15_i_1 15780 13 18 16 14 11 - 0.01608 0.01650 0.00567 0.00908 0.00456 -
+portfolio_N_15_i_2 15780 13 16 16 16 11 - 0.01628 0.01516 0.00572 0.00959 0.00456 -
+portfolio_N_15_i_3 15780 12 18 16 15 12 - 0.01570 0.02050 0.00563 0.00941 0.00490 -
+portfolio_N_15_i_4 15780 13 18 16 13 10 - 0.01600 0.01653 0.00561 0.00872 0.00422 -
+portfolio_N_15_i_5 15780 13 15 15 13 10 - 0.01613 0.01388 0.00554 0.00884 0.00418 -
+portfolio_N_15_i_6 15780 14 15 17 14 11 - 0.01664 0.01377 0.00577 0.00905 0.00460 -
+portfolio_N_15_i_7 15780 13 16 16 15 10 - 0.01615 0.01428 0.00562 0.00936 0.00424 -
+portfolio_N_15_i_8 15780 14 19 18 16 11 - 0.01667 0.01589 0.00614 0.00980 0.00455 -
+portfolio_N_15_i_9 15780 13 14 15 14 10 - 0.01606 0.01471 0.00551 0.00904 0.00421 -
+portfolio_N_15_i_10 15780 14 17 15 14 10 - 0.01686 0.01465 0.00549 0.00904 0.00422 -
+portfolio_N_15_i_11 15780 13 17 15 13 11 - 0.01626 0.01497 0.00552 0.00880 0.00455 -
+portfolio_N_15_i_12 15780 14 21 16 17 11 - 0.01670 0.01898 0.00574 0.00994 0.00458 -
+portfolio_N_15_i_13 15780 13 16 15 14 12 - 0.01601 0.01464 0.00550 0.00911 0.00501 -
+portfolio_N_15_i_14 15780 12 17 16 14 12 - 0.01563 0.01455 0.00574 0.00902 0.00497 -
+portfolio_N_15_i_15 15780 15 17 16 15 13 - 0.01715 0.01484 0.00562 0.00931 0.00534 -
+portfolio_N_15_i_16 15780 12 16 16 14 11 - 0.01557 0.01435 0.00568 0.00906 0.00459 -
+portfolio_N_15_i_17 15780 13 15 16 13 11 - 0.01612 0.01491 0.00569 0.00872 0.00458 -
+portfolio_N_15_i_18 15780 13 17 15 13 10 - 0.01614 0.01608 0.00555 0.00880 0.00418 -
+portfolio_N_15_i_19 15780 12 17 15 15 10 - 0.01582 0.01613 0.00559 0.00937 0.00421 -
+portfolio_N_20_i_0 26040 13 15 15 14 11 - 0.01608 0.02261 0.00864 0.01355 0.00848 -
+portfolio_N_20_i_1 26040 13 19 15 15 10 - 0.01652 0.02820 0.00860 0.01399 0.00780 -
+portfolio_N_20_i_2 26040 13 17 18 14 13 - 0.01644 0.02987 0.00915 0.01361 0.00985 -
+portfolio_N_20_i_3 26040 13 16 15 14 11 - 0.01514 0.02325 0.00849 0.01360 0.00855 -
+portfolio_N_20_i_4 26040 15 19 14 15 13 - 0.01828 0.03341 0.00854 0.01400 0.00992 -
+portfolio_N_20_i_5 26040 13 16 16 17 10 - 0.01646 0.02098 0.00871 0.01484 0.00782 -
+portfolio_N_20_i_6 26040 14 20 15 15 11 - 0.01747 0.02746 0.00835 0.01398 0.00852 -
+portfolio_N_20_i_7 26040 15 17 16 15 11 - 0.01858 0.02532 0.00888 0.01407 0.00862 -
+portfolio_N_20_i_8 26040 13 18 14 13 11 - 0.01628 0.02561 0.00837 0.01308 0.00845 -
+portfolio_N_20_i_9 26040 14 17 15 15 11 - 0.01726 0.02707 0.00880 0.01410 0.00843 -
+portfolio_N_20_i_10 26040 14 18 16 14 11 - 0.01711 0.02505 0.00872 0.01356 0.00849 -
+portfolio_N_20_i_11 26040 13 16 17 14 13 - 0.01621 0.02322 0.00890 0.01350 0.00990 -
+portfolio_N_20_i_12 26040 15 17 17 16 11 - 0.01829 0.02644 0.00895 0.01436 0.00851 -
+portfolio_N_20_i_13 26040 13 15 15 13 10 - 0.01599 0.02926 0.00906 0.01311 0.00777 -
+portfolio_N_20_i_14 26040 13 18 18 15 10 - 0.01616 0.02533 0.00919 0.01403 0.00775 -
+portfolio_N_20_i_15 26040 13 16 16 14 11 - 0.01636 0.02529 0.00865 0.01381 0.00854 -
+portfolio_N_20_i_16 26040 13 17 16 14 12 - 0.01634 0.02396 0.00867 0.01358 0.00933 -
+portfolio_N_20_i_17 26040 13 18 16 16 12 - 0.01627 0.02555 0.00881 0.01447 0.00929 -
+portfolio_N_20_i_18 26040 13 18 16 14 10 - 0.01645 0.02772 0.00881 0.01352 0.00777 -
+portfolio_N_20_i_19 26040 14 16 16 15 11 - 0.01738 0.02570 0.00892 0.01392 0.00846 -
+portfolio_N_25_i_0 38800 15 18 17 17 12 - 0.02906 0.03832 0.01357 0.02078 0.01460 -
+portfolio_N_25_i_1 38800 13 19 17 15 12 - 0.02572 0.04047 0.01364 0.01964 0.01463 -
+portfolio_N_25_i_2 38800 14 15 15 15 11 - 0.02703 0.03219 0.01285 0.01948 0.01357 -
+portfolio_N_25_i_3 38800 15 20 15 16 12 - 0.02894 0.03989 0.01286 0.02016 0.01476 -
+portfolio_N_25_i_4 38800 13 16 15 15 11 - 0.02520 0.03309 0.01289 0.01982 0.01353 -
+portfolio_N_25_i_5 38800 14 15 15 15 12 - 0.02751 0.03296 0.01305 0.01962 0.01489 -
+portfolio_N_25_i_6 38800 14 18 14 14 11 - 0.02750 0.03763 0.01266 0.01904 0.01364 -
+portfolio_N_25_i_7 38800 14 19 15 16 13 - 0.02732 0.03533 0.01281 0.02014 0.01590 -
+portfolio_N_25_i_8 38800 14 17 15 15 11 - 0.02728 0.03293 0.01274 0.01960 0.01364 -
+portfolio_N_25_i_9 38800 14 19 15 15 12 - 0.02739 0.03600 0.01290 0.01954 0.01477 -
+portfolio_N_25_i_10 38800 14 18 14 14 11 - 0.02736 0.03734 0.01236 0.01896 0.01347 -
+portfolio_N_25_i_11 38800 15 21 15 16 11 - 0.02910 0.04072 0.01281 0.02027 0.01354 -
+portfolio_N_25_i_12 38800 15 18 17 14 12 - 0.02923 0.03431 0.01344 0.01905 0.01463 -
+portfolio_N_25_i_13 38800 16 19 15 16 11 - 0.03070 0.03879 0.01290 0.02019 0.01364 -
+portfolio_N_25_i_14 38800 14 16 13 13 10 - 0.02719 0.03314 0.01209 0.01832 0.01244 -
+portfolio_N_25_i_15 38800 13 16 17 15 11 - 0.02563 0.03328 0.01350 0.01982 0.01352 -
+portfolio_N_25_i_16 38800 14 19 14 15 12 - 0.02742 0.05195 0.01246 0.01948 0.01469 -
+portfolio_N_25_i_17 38800 14 15 14 13 12 - 0.02716 0.03272 0.01235 0.01838 0.01469 -
+portfolio_N_25_i_18 38800 15 18 16 16 12 - 0.02915 0.03600 0.01317 0.02032 0.01499 -
+portfolio_N_25_i_19 38800 14 17 15 15 12 - 0.02670 0.03615 0.01267 0.01962 0.01478 -
+portfolio_N_30_i_0 54060 15 17 18 16 11 - 0.04222 0.04377 0.01902 0.02735 0.02026 -
+portfolio_N_30_i_1 54060 15 18 18 17 11 - 0.04154 0.04821 0.01899 0.02813 0.02007 -
+portfolio_N_30_i_2 54060 14 20 17 15 14 - 0.03917 0.05034 0.01822 0.02646 0.02502 -
+portfolio_N_30_i_3 54060 14 20 17 15 12 - 0.03929 0.05799 0.01819 0.02625 0.02192 -
+portfolio_N_30_i_4 54060 14 17 18 15 13 - 0.03964 0.04580 0.01906 0.02649 0.02396 -
+portfolio_N_30_i_5 54060 14 18 15 16 11 - 0.03956 0.04614 0.01716 0.02728 0.02006 -
+portfolio_N_30_i_6 54060 12 19 16 16 11 - 0.03474 0.04898 0.01781 0.02729 0.02011 -
+portfolio_N_30_i_7 54060 14 17 16 16 11 - 0.03936 0.04394 0.01761 0.02717 0.02021 -
+portfolio_N_30_i_8 54060 15 17 17 15 12 - 0.04215 0.04182 0.01836 0.02652 0.02186 -
+portfolio_N_30_i_9 54060 14 18 17 17 12 - 0.03966 0.04980 0.01842 0.02831 0.02199 -
+portfolio_N_30_i_10 54060 13 18 16 15 11 - 0.03722 0.04368 0.01767 0.02632 0.02027 -
+portfolio_N_30_i_11 54060 14 18 16 16 12 - 0.03977 0.04644 0.01754 0.02713 0.02191 -
+portfolio_N_30_i_12 54060 14 18 16 15 11 - 0.03923 0.04913 0.01778 0.02649 0.02034 -
+portfolio_N_30_i_13 54060 15 21 17 16 12 - 0.04164 0.05595 0.01808 0.02713 0.02175 -
+portfolio_N_30_i_14 54060 16 19 17 15 11 - 0.04419 0.05431 0.01838 0.02643 0.02015 -
+portfolio_N_30_i_15 54060 14 19 17 15 12 - 0.03961 0.04777 0.01842 0.02656 0.02187 -
+portfolio_N_30_i_16 54060 15 17 18 15 14 - 0.04202 0.04420 0.01883 0.02649 0.02490 -
+portfolio_N_30_i_17 54060 16 16 17 14 11 - 0.04433 0.04140 0.01831 0.02563 0.02025 -
+portfolio_N_30_i_18 54060 15 19 16 14 12 - 0.04180 0.05000 0.01772 0.02550 0.02201 -
+portfolio_N_30_i_19 54060 16 18 16 15 11 - 0.04432 0.04734 0.01777 0.02642 0.02010 -
+portfolio_N_35_i_0 71820 15 19 18 19 12 - 0.05816 0.05931 0.02139 0.03880 0.03173 -
+portfolio_N_35_i_1 71820 15 20 17 16 12 - 0.05906 0.06253 0.02106 0.03483 0.03089 -
+portfolio_N_35_i_2 71820 14 18 15 14 12 - 0.05513 0.06219 0.01980 0.03284 0.03091 -
+portfolio_N_35_i_3 71820 15 20 16 17 13 - 0.05894 0.06915 0.02002 0.03579 0.03306 -
+portfolio_N_35_i_4 71820 16 19 16 15 13 - 0.06250 0.06503 0.02030 0.03380 0.03310 -
+portfolio_N_35_i_5 71820 14 19 17 15 13 - 0.05460 0.06351 0.02085 0.03382 0.03323 -
+portfolio_N_35_i_6 71820 15 18 18 15 12 - 0.05876 0.05876 0.02086 0.03374 0.03071 -
+portfolio_N_35_i_7 71820 17 17 17 15 12 - 0.06557 0.05745 0.02104 0.03375 0.03093 -
+portfolio_N_35_i_8 71820 15 20 17 17 11 - 0.05911 0.06865 0.02056 0.03577 0.02855 -
+portfolio_N_35_i_9 71820 14 17 19 17 12 - 0.05430 0.05969 0.02247 0.03597 0.03085 -
+portfolio_N_35_i_10 71820 15 18 17 16 12 - 0.05799 0.05773 0.02076 0.03484 0.03083 -
+portfolio_N_35_i_11 71820 15 20 16 16 12 - 0.05800 0.06867 0.02054 0.03480 0.03071 -
+portfolio_N_35_i_12 71820 15 18 16 15 11 - 0.05866 0.06290 0.01991 0.03373 0.02864 -
+portfolio_N_35_i_13 71820 15 20 16 17 12 - 0.05746 0.06915 0.02026 0.03583 0.03091 -
+portfolio_N_35_i_14 71820 16 20 18 18 13 - 0.06144 0.06523 0.02127 0.03676 0.03324 -
+portfolio_N_35_i_15 71820 15 19 16 15 11 - 0.05826 0.06382 0.02012 0.03400 0.02846 -
+portfolio_N_35_i_16 71820 14 18 19 16 13 - 0.05486 0.06249 0.02178 0.03498 0.03310 -
+portfolio_N_35_i_17 71820 13 17 16 14 12 - 0.05057 0.05973 0.02016 0.03373 0.03074 -
+portfolio_N_35_i_18 71820 16 19 18 17 13 - 0.06121 0.06841 0.02141 0.03582 0.03322 -
+portfolio_N_35_i_19 71820 16 17 18 16 12 - 0.06150 0.05831 0.02125 0.03476 0.03118 -
+Table 6: Iterations and solver runtimes for oscillating masses problems
+Iterations Solver Runtime (s)
+Problem Size clarabel ecos gurobi mosek qoco qoco_custom cvxgen clarabel ecos gurobi mosek qoco qoco_custom cvxgen
+oscillating_masses_N_8_i_0 1344 5 13 11 11 5 5 7 0.00049 0.00108 0.00222 0.00187 0.00027 0.00011 0.00014
+oscillating_masses_N_8_i_1 1344 5 14 8 10 5 5 6 0.00048 0.00117 0.00218 0.00181 0.00028 0.00012 0.00013
+oscillating_masses_N_8_i_2 1344 5 17 9 11 5 5 6 0.00048 0.00139 0.00216 0.00189 0.00026 0.00011 0.00013
+oscillating_masses_N_8_i_3 1344 5 14 8 11 5 5 6 0.00047 0.00115 0.00217 0.00190 0.00026 0.00011 0.00012
+oscillating_masses_N_8_i_4 1344 5 12 10 11 5 5 7 0.00049 0.00099 0.00219 0.00189 0.00029 0.00011 0.00015
+oscillating_masses_N_8_i_5 1344 5 12 8 11 5 5 6 0.00047 0.00100 0.00216 0.00190 0.00026 0.00011 0.00013
+oscillating_masses_N_8_i_6 1344 10 13 12 11 7 7 10 0.00083 0.00110 0.00223 0.00188 0.00036 0.00015 0.00020
+oscillating_masses_N_8_i_7 1344 5 13 9 10 5 5 6 0.00048 0.00104 0.00219 0.00178 0.00027 0.00011 0.00012
+oscillating_masses_N_8_i_8 1344 5 14 9 11 4 4 6 0.00048 0.00112 0.00217 0.00188 0.00023 0.00009 0.00012
+oscillating_masses_N_8_i_9 1344 5 15 8 11 5 5 6 0.00048 0.00121 0.00219 0.00153 0.00027 0.00011 0.00013
+oscillating_masses_N_8_i_10 1344 5 14 9 11 5 5 6 0.00048 0.00111 0.00221 0.00190 0.00026 0.00011 0.00013
+oscillating_masses_N_8_i_11 1344 5 13 8 13 5 5 6 0.00048 0.00105 0.00217 0.00210 0.00028 0.00011 0.00013
+oscillating_masses_N_8_i_12 1344 5 14 9 12 5 5 7 0.00048 0.00113 0.00220 0.00203 0.00027 0.00011 0.00014
+oscillating_masses_N_8_i_13 1344 5 12 8 12 5 5 6 0.00048 0.00097 0.00216 0.00160 0.00027 0.00011 0.00012
+oscillating_masses_N_8_i_14 1344 5 12 8 9 5 5 6 0.00048 0.00098 0.00215 0.00138 0.00026 0.00011 0.00013
+oscillating_masses_N_8_i_15 1344 5 13 8 11 5 5 7 0.00048 0.00109 0.00218 0.00156 0.00026 0.00011 0.00015
+oscillating_masses_N_8_i_16 1344 6 14 10 11 6 6 8 0.00055 0.00117 0.00220 0.00190 0.00031 0.00013 0.00016
+oscillating_masses_N_8_i_17 1344 5 13 10 12 5 5 6 0.00047 0.00105 0.00221 0.00199 0.00027 0.00011 0.00014
+oscillating_masses_N_8_i_18 1344 5 12 9 11 4 4 6 0.00048 0.00100 0.00218 0.00190 0.00024 0.00009 0.00013
+oscillating_masses_N_8_i_19 1344 5 12 10 11 5 5 6 0.00048 0.00100 0.00221 0.00187 0.00026 0.00011 0.00013
+oscillating_masses_N_20_i_0 3312 5 15 9 12 5 5 7 0.00110 0.00300 0.00265 0.00541 0.00062 0.00029 0.00039
+oscillating_masses_N_20_i_1 3312 5 14 9 12 5 5 7 0.00117 0.00275 0.00266 0.00523 0.00061 0.00029 0.00041
+oscillating_masses_N_20_i_2 3312 5 15 9 12 5 5 7 0.00115 0.00290 0.00266 0.00538 0.00065 0.00029 0.00042
+oscillating_masses_N_20_i_3 3312 7 14 10 13 7 7 8 0.00145 0.00274 0.00270 0.00561 0.00080 0.00039 0.00046
+oscillating_masses_N_20_i_4 3312 7 14 10 14 7 7 9 0.00151 0.00275 0.00271 0.00579 0.00077 0.00039 0.00051
+oscillating_masses_N_20_i_5 3312 5 14 8 13 5 5 6 0.00113 0.00272 0.00262 0.00553 0.00062 0.00029 0.00035
+oscillating_masses_N_20_i_6 3312 5 16 8 11 5 5 7 0.00114 0.00308 0.00259 0.00510 0.00062 0.00029 0.00040
+oscillating_masses_N_20_i_7 3312 5 14 9 12 5 5 6 0.00115 0.00278 0.00265 0.00517 0.00062 0.00029 0.00035
+oscillating_masses_N_20_i_8 3312 5 13 8 12 5 5 7 0.00116 0.00259 0.00260 0.00530 0.00064 0.00029 0.00041
+oscillating_masses_N_20_i_9 3312 5 16 9 13 5 5 6 0.00113 0.00306 0.00266 0.00562 0.00061 0.00029 0.00035
+oscillating_masses_N_20_i_10 3312 5 14 8 12 5 5 6 0.00114 0.00274 0.00262 0.00547 0.00062 0.00029 0.00034
+oscillating_masses_N_20_i_11 3312 5 14 10 12 5 5 7 0.00114 0.00276 0.00270 0.00532 0.00064 0.00029 0.00041
+oscillating_masses_N_20_i_12 3312 5 15 9 14 5 5 7 0.00118 0.00287 0.00263 0.00580 0.00062 0.00029 0.00041
+oscillating_masses_N_20_i_13 3312 5 14 8 12 5 5 6 0.00112 0.00268 0.00262 0.00486 0.00061 0.00029 0.00036
+oscillating_masses_N_20_i_14 3312 5 15 9 12 5 5 6 0.00111 0.00289 0.00265 0.00552 0.00063 0.00029 0.00035
+oscillating_masses_N_20_i_15 3312 5 13 8 11 5 5 6 0.00113 0.00260 0.00259 0.00510 0.00062 0.00029 0.00036
+oscillating_masses_N_20_i_16 3312 5 14 8 12 5 5 6 0.00113 0.00276 0.00259 0.00510 0.00063 0.00029 0.00034
+oscillating_masses_N_20_i_17 3312 5 15 10 13 5 5 7 0.00115 0.00294 0.00266 0.00533 0.00062 0.00029 0.00040
+oscillating_masses_N_20_i_18 3312 5 15 8 13 5 5 6 0.00113 0.00280 0.00261 0.00560 0.00063 0.00029 0.00035
+oscillating_masses_N_20_i_19 3312 5 15 10 14 5 5 7 0.00109 0.00298 0.00267 0.00543 0.00061 0.00029 0.00039
+oscillating_masses_N_32_i_0 5280 5 14 9 13 5 5 - 0.00179 0.00420 0.00309 0.00885 0.00097 0.00048 -
+oscillating_masses_N_32_i_1 5280 5 14 8 13 5 5 - 0.00181 0.00434 0.00306 0.00881 0.00096 0.00049 -
+oscillating_masses_N_32_i_2 5280 5 15 8 13 5 5 - 0.00178 0.00456 0.00303 0.00882 0.00095 0.00048 -
+oscillating_masses_N_32_i_3 5280 7 13 9 13 6 6 - 0.00226 0.00398 0.00308 0.00925 0.00110 0.00057 -
+oscillating_masses_N_32_i_4 5280 9 15 10 15 8 8 - 0.00285 0.00470 0.00318 0.00958 0.00137 0.00078 -
+oscillating_masses_N_32_i_5 5280 5 15 8 10 5 5 - 0.00177 0.00457 0.00306 0.00761 0.00097 0.00049 -
+oscillating_masses_N_32_i_6 5280 6 14 7 13 6 6 - 0.00197 0.00419 0.00299 0.00929 0.00110 0.00058 -
+oscillating_masses_N_32_i_7 5280 6 15 9 13 6 6 - 0.00202 0.00450 0.00307 0.00865 0.00113 0.00056 -
+oscillating_masses_N_32_i_8 5280 5 16 9 13 5 5 - 0.00182 0.00486 0.00311 0.00885 0.00097 0.00048 -
+oscillating_masses_N_32_i_9 5280 5 14 8 12 5 5 - 0.00178 0.00429 0.00306 0.00839 0.00097 0.00048 -
+oscillating_masses_N_32_i_10 5280 6 15 10 13 5 5 - 0.00204 0.00462 0.00323 0.00882 0.00098 0.00049 -
+oscillating_masses_N_32_i_11 5280 5 13 8 12 5 5 - 0.00179 0.00395 0.00307 0.00842 0.00098 0.00049 -
+oscillating_masses_N_32_i_12 5280 5 15 8 11 5 5 - 0.00177 0.00441 0.00305 0.00849 0.00098 0.00049 -
+oscillating_masses_N_32_i_13 5280 5 15 9 13 5 5 - 0.00179 0.00464 0.00314 0.00887 0.00115 0.00050 -
+oscillating_masses_N_32_i_14 5280 6 16 9 13 6 6 - 0.00199 0.00482 0.00312 0.00895 0.00110 0.00058 -
+oscillating_masses_N_32_i_15 5280 5 16 8 12 5 5 - 0.00179 0.00486 0.00305 0.00842 0.00096 0.00049 -
+oscillating_masses_N_32_i_16 5280 5 15 9 12 5 5 - 0.00179 0.00459 0.00311 0.00846 0.00097 0.00049 -
+oscillating_masses_N_32_i_17 5280 5 15 8 12 5 5 - 0.00175 0.00460 0.00302 0.00850 0.00102 0.00049 -
+oscillating_masses_N_32_i_18 5280 5 15 8 12 5 5 - 0.00179 0.00471 0.00298 0.00845 0.00100 0.00049 -
+oscillating_masses_N_32_i_19 5280 5 15 8 13 5 5 - 0.00181 0.00465 0.00304 0.00916 0.00098 0.00049 -
+oscillating_masses_N_44_i_0 7248 5 15 9 11 5 5 - 0.00237 0.00623 0.00356 0.01125 0.00132 0.00069 -
+oscillating_masses_N_44_i_1 7248 6 16 10 12 6 6 - 0.00277 0.00660 0.00365 0.01121 0.00148 0.00083 -
+oscillating_masses_N_44_i_2 7248 6 15 11 14 6 6 - 0.00282 0.00623 0.00376 0.01247 0.00156 0.00081 -
+oscillating_masses_N_44_i_3 7248 5 14 8 12 5 5 - 0.00247 0.00604 0.00350 0.01175 0.00134 0.00070 -
+oscillating_masses_N_44_i_4 7248 5 16 8 13 5 5 - 0.00241 0.00676 0.00348 0.01229 0.00138 0.00070 -
+oscillating_masses_N_44_i_5 7248 5 14 8 13 5 5 - 0.00245 0.00587 0.00349 0.01222 0.00133 0.00070 -
+oscillating_masses_N_44_i_6 7248 6 15 9 14 6 6 - 0.00287 0.00642 0.00361 0.01279 0.00150 0.00082 -
+oscillating_masses_N_44_i_7 7248 5 15 9 12 5 5 - 0.00243 0.00638 0.00355 0.01170 0.00146 0.00070 -
+oscillating_masses_N_44_i_8 7248 5 15 9 12 5 5 - 0.00245 0.00643 0.00359 0.01171 0.00134 0.00069 -
+oscillating_masses_N_44_i_9 7248 5 15 9 12 5 5 - 0.00245 0.00624 0.00359 0.01178 0.00130 0.00069 -
+oscillating_masses_N_44_i_10 7248 5 15 9 13 5 5 - 0.00252 0.00618 0.00358 0.01230 0.00133 0.00069 -
+oscillating_masses_N_44_i_11 7248 5 14 9 13 5 5 - 0.00240 0.00600 0.00358 0.01233 0.00136 0.00068 -
+oscillating_masses_N_44_i_12 7248 5 16 9 12 5 5 - 0.00241 0.00651 0.00356 0.01173 0.00133 0.00070 -
+oscillating_masses_N_44_i_13 7248 5 14 8 12 5 5 - 0.00246 0.00607 0.00352 0.01169 0.00133 0.00069 -
+oscillating_masses_N_44_i_14 7248 5 16 8 13 5 5 - 0.00249 0.00677 0.00352 0.01231 0.00133 0.00069 -
+oscillating_masses_N_44_i_15 7248 5 14 10 13 5 5 - 0.00241 0.00592 0.00366 0.01204 0.00133 0.00071 -
+oscillating_masses_N_44_i_16 7248 5 16 9 13 5 5 - 0.00245 0.00671 0.00357 0.01214 0.00132 0.00068 -
+oscillating_masses_N_44_i_17 7248 5 15 9 13 5 5 - 0.00244 0.00638 0.00362 0.01167 0.00138 0.00069 -
+oscillating_masses_N_44_i_18 7248 5 14 9 13 5 5 - 0.00245 0.00584 0.00352 0.01216 0.00134 0.00071 -
+oscillating_masses_N_44_i_19 7248 5 14 9 12 5 5 - 0.00248 0.00594 0.00366 0.01185 0.00137 0.00068 -
+oscillating_masses_N_56_i_0 9216 5 15 9 13 5 5 - 0.00312 0.00795 0.00405 0.01711 0.00171 0.00092 -
+oscillating_masses_N_56_i_1 9216 5 16 9 12 5 5 - 0.00303 0.00856 0.00404 0.01631 0.00167 0.00092 -
+oscillating_masses_N_56_i_2 9216 5 16 10 11 5 5 - 0.00303 0.00849 0.00416 0.01550 0.00189 0.00093 -
+oscillating_masses_N_56_i_3 9216 5 15 8 12 5 5 - 0.00319 0.00797 0.00396 0.01628 0.00172 0.00092 -
+oscillating_masses_N_56_i_4 9216 5 15 9 13 5 5 - 0.00315 0.00786 0.00409 0.01698 0.00167 0.00093 -
+oscillating_masses_N_56_i_5 9216 7 16 8 13 7 7 - 0.00407 0.00848 0.00392 0.01688 0.00217 0.00125 -
+oscillating_masses_N_56_i_6 9216 5 14 9 12 5 5 - 0.00313 0.00751 0.00410 0.01632 0.00189 0.00092 -
+oscillating_masses_N_56_i_7 9216 5 19 8 13 5 5 - 0.00312 0.01001 0.00397 0.01723 0.00169 0.00096 -
+oscillating_masses_N_56_i_8 9216 5 14 9 12 5 5 - 0.00312 0.00765 0.00404 0.01606 0.00170 0.00093 -
+oscillating_masses_N_56_i_9 9216 6 16 9 13 6 6 - 0.00354 0.00851 0.00406 0.01712 0.00192 0.00110 -
+oscillating_masses_N_56_i_10 9216 5 19 9 13 5 5 - 0.00313 0.01006 0.00401 0.01689 0.00169 0.00093 -
+oscillating_masses_N_56_i_11 9216 5 14 10 12 5 5 - 0.00311 0.00758 0.00416 0.01613 0.00168 0.00095 -
+oscillating_masses_N_56_i_12 9216 5 15 8 12 5 5 - 0.00311 0.00801 0.00397 0.01629 0.00171 0.00092 -
+oscillating_masses_N_56_i_13 9216 5 14 8 12 5 5 - 0.00311 0.00760 0.00394 0.01614 0.00169 0.00092 -
+oscillating_masses_N_56_i_14 9216 5 15 8 12 5 5 - 0.00311 0.00783 0.00399 0.01622 0.00171 0.00095 -
+oscillating_masses_N_56_i_15 9216 6 14 9 13 6 6 - 0.00357 0.00749 0.00407 0.01692 0.00193 0.00109 -
+oscillating_masses_N_56_i_16 9216 5 16 8 13 5 5 - 0.00311 0.00849 0.00394 0.01687 0.00167 0.00093 -
+oscillating_masses_N_56_i_17 9216 6 15 9 14 6 6 - 0.00353 0.00801 0.00404 0.01801 0.00191 0.00110 -
+oscillating_masses_N_56_i_18 9216 5 16 8 13 5 5 - 0.00312 0.00879 0.00395 0.01696 0.00168 0.00091 -
+oscillating_masses_N_56_i_19 9216 7 16 10 16 7 7 - 0.00422 0.00866 0.00424 0.01921 0.00216 0.00128 -
+oscillating_masses_N_76_i_0 12496 5 16 9 12 5 - - 0.00417 0.01169 0.00509 0.02337 0.00229 - -
+oscillating_masses_N_76_i_1 12496 6 15 10 12 5 - - 0.00491 0.01090 0.00521 0.02242 0.00231 - -
+oscillating_masses_N_76_i_2 12496 5 15 9 12 5 - - 0.00425 0.01106 0.00518 0.02213 0.00238 - -
+oscillating_masses_N_76_i_3 12496 5 14 8 12 6 - - 0.00430 0.01017 0.00494 0.02243 0.00265 - -
+oscillating_masses_N_76_i_4 12496 5 14 8 12 5 - - 0.00418 0.01023 0.00488 0.02254 0.00231 - -
+oscillating_masses_N_76_i_5 12496 5 15 9 14 5 - - 0.00427 0.01097 0.00511 0.02418 0.00245 - -
+oscillating_masses_N_76_i_6 12496 5 15 8 13 5 - - 0.00424 0.01141 0.00492 0.02355 0.00258 - -
+oscillating_masses_N_76_i_7 12496 5 15 9 14 5 - - 0.00419 0.01078 0.00506 0.02419 0.00257 - -
+oscillating_masses_N_76_i_8 12496 6 14 9 12 6 - - 0.00488 0.01022 0.00507 0.02243 0.00262 - -
+oscillating_masses_N_76_i_9 12496 5 14 9 13 5 - - 0.00423 0.01017 0.00509 0.02297 0.00228 - -
+oscillating_masses_N_76_i_10 12496 6 15 9 11 6 - - 0.00466 0.01093 0.00511 0.02109 0.00264 - -
+oscillating_masses_N_76_i_11 12496 5 14 8 12 5 - - 0.00423 0.01031 0.00494 0.02259 0.00233 - -
+oscillating_masses_N_76_i_12 12496 7 15 10 13 7 - - 0.00543 0.01121 0.00517 0.02368 0.00294 - -
+oscillating_masses_N_76_i_13 12496 7 17 11 14 6 - - 0.00550 0.01210 0.00542 0.02445 0.00259 - -
+oscillating_masses_N_76_i_14 12496 5 15 10 14 5 - - 0.00424 0.01108 0.00518 0.02432 0.00224 - -
+oscillating_masses_N_76_i_15 12496 5 16 9 13 5 - - 0.00425 0.01168 0.00506 0.02324 0.00231 - -
+oscillating_masses_N_76_i_16 12496 5 15 9 11 5 - - 0.00427 0.01089 0.00507 0.02136 0.00233 - -
+oscillating_masses_N_76_i_17 12496 5 15 9 13 5 - - 0.00429 0.01105 0.00502 0.02345 0.00230 - -
+oscillating_masses_N_76_i_18 12496 5 16 9 12 5 - - 0.00421 0.01152 0.00504 0.02265 0.00228 - -
+oscillating_masses_N_76_i_19 12496 5 15 9 12 5 - - 0.00420 0.01093 0.00503 0.02258 0.00228 - -
+oscillating_masses_N_96_i_0 15776 5 15 8 13 5 - - 0.00537 0.01378 0.00572 0.02968 0.00287 - -
+oscillating_masses_N_96_i_1 15776 6 16 9 13 6 - - 0.00618 0.01448 0.00589 0.02978 0.00333 - -
+oscillating_masses_N_96_i_2 15776 5 16 8 12 5 - - 0.00528 0.01466 0.00580 0.02827 0.00288 - -
+oscillating_masses_N_96_i_3 15776 5 17 8 13 5 - - 0.00530 0.01562 0.00575 0.02965 0.00287 - -
+oscillating_masses_N_96_i_4 15776 5 13 10 11 5 - - 0.00526 0.01187 0.00609 0.02489 0.00291 - -
+oscillating_masses_N_96_i_5 15776 5 15 8 14 5 - - 0.00535 0.01372 0.00577 0.03095 0.00290 - -
+oscillating_masses_N_96_i_6 15776 5 14 9 12 5 - - 0.00536 0.01284 0.00593 0.02815 0.00332 - -
+oscillating_masses_N_96_i_7 15776 5 15 9 13 5 - - 0.00543 0.01424 0.00603 0.02947 0.00290 - -
+oscillating_masses_N_96_i_8 15776 5 16 8 13 5 - - 0.00552 0.01457 0.00577 0.02981 0.00289 - -
+oscillating_masses_N_96_i_9 15776 5 15 8 11 5 - - 0.00536 0.01356 0.00574 0.02750 0.00293 - -
+oscillating_masses_N_96_i_10 15776 6 14 10 13 6 - - 0.00589 0.01277 0.00609 0.02968 0.00327 - -
+oscillating_masses_N_96_i_11 15776 6 15 9 12 6 - - 0.00608 0.01380 0.00583 0.02834 0.00331 - -
+oscillating_masses_N_96_i_12 15776 5 17 9 13 5 - - 0.00527 0.01555 0.00592 0.02948 0.00289 - -
+oscillating_masses_N_96_i_13 15776 5 15 8 13 5 - - 0.00531 0.01377 0.00574 0.02977 0.00293 - -
+oscillating_masses_N_96_i_14 15776 5 17 10 13 5 - - 0.00534 0.01547 0.00612 0.02916 0.00288 - -
+oscillating_masses_N_96_i_15 15776 6 15 9 12 6 - - 0.00603 0.01406 0.00591 0.02810 0.00328 - -
+oscillating_masses_N_96_i_16 15776 5 14 8 13 5 - - 0.00543 0.01295 0.00574 0.02947 0.00290 - -
+oscillating_masses_N_96_i_17 15776 5 16 8 14 5 - - 0.00529 0.01427 0.00579 0.03088 0.00290 - -
+oscillating_masses_N_96_i_18 15776 5 14 9 13 5 - - 0.00517 0.01306 0.00603 0.02977 0.00290 - -
+oscillating_masses_N_96_i_19 15776 5 16 8 13 5 - - 0.00527 0.01481 0.00577 0.02987 0.00295 - -
+oscillating_masses_N_116_i_0 19056 5 15 9 13 5 - - 0.00645 0.01652 0.00678 0.03539 0.00354 - -
+oscillating_masses_N_116_i_1 19056 6 14 9 12 6 - - 0.00723 0.01589 0.00676 0.03201 0.00404 - -
+oscillating_masses_N_116_i_2 19056 5 15 10 12 5 - - 0.00653 0.01675 0.00712 0.03463 0.00351 - -
+oscillating_masses_N_116_i_3 19056 5 16 8 12 5 - - 0.00637 0.01775 0.00654 0.03461 0.00347 - -
+oscillating_masses_N_116_i_4 19056 5 15 8 12 5 - - 0.00637 0.01692 0.00664 0.03405 0.00351 - -
+oscillating_masses_N_116_i_5 19056 6 16 8 13 6 - - 0.00724 0.01791 0.00661 0.03601 0.00395 - -
+oscillating_masses_N_116_i_6 19056 6 14 9 11 6 - - 0.00670 0.01542 0.00674 0.03227 0.00398 - -
+oscillating_masses_N_116_i_7 19056 5 15 9 14 5 - - 0.00663 0.01652 0.00677 0.03753 0.00347 - -
+oscillating_masses_N_116_i_8 19056 7 17 9 14 7 - - 0.00864 0.01920 0.00681 0.03718 0.00454 - -
+oscillating_masses_N_116_i_9 19056 5 15 10 13 5 - - 0.00646 0.01671 0.00698 0.03599 0.00350 - -
+oscillating_masses_N_116_i_10 19056 5 14 9 12 5 - - 0.00640 0.01547 0.00678 0.03438 0.00354 - -
+oscillating_masses_N_116_i_11 19056 5 16 9 13 5 - - 0.00628 0.01727 0.00673 0.03629 0.00349 - -
+oscillating_masses_N_116_i_12 19056 5 16 9 12 5 - - 0.00635 0.01787 0.00674 0.03427 0.00355 - -
+oscillating_masses_N_116_i_13 19056 5 15 11 12 6 - - 0.00641 0.01696 0.00725 0.03417 0.00401 - -
+oscillating_masses_N_116_i_14 19056 6 15 10 13 6 - - 0.00744 0.01671 0.00700 0.03591 0.00397 - -
+oscillating_masses_N_116_i_15 19056 5 16 9 13 5 - - 0.00639 0.01772 0.00676 0.03526 0.00353 - -
+oscillating_masses_N_116_i_16 19056 5 14 9 13 5 - - 0.00639 0.01565 0.00678 0.03580 0.00362 - -
+oscillating_masses_N_116_i_17 19056 5 16 8 13 5 - - 0.00646 0.01823 0.00662 0.03592 0.00346 - -
+oscillating_masses_N_116_i_18 19056 5 15 8 11 6 - - 0.00638 0.01739 0.00661 0.03253 0.00402 - -
+oscillating_masses_N_116_i_19 19056 5 16 9 13 5 - - 0.00655 0.01813 0.00678 0.03584 0.00350 - -
+oscillating_masses_N_136_i_0 22336 5 16 9 13 5 - - 0.00746 0.02089 0.00775 0.04277 0.00412 - -
+oscillating_masses_N_136_i_1 22336 5 15 9 12 5 - - 0.00758 0.01966 0.00779 0.04064 0.00407 - -
+oscillating_masses_N_136_i_2 22336 5 14 10 12 5 - - 0.00765 0.01879 0.00805 0.04135 0.00411 - -
+oscillating_masses_N_136_i_3 22336 5 16 9 12 5 - - 0.00758 0.02064 0.00774 0.04057 0.00413 - -
+oscillating_masses_N_136_i_4 22336 5 15 8 13 5 - - 0.00747 0.02013 0.00759 0.04242 0.00409 - -
+oscillating_masses_N_136_i_5 22336 5 16 9 13 5 - - 0.00752 0.02073 0.00777 0.03960 0.00425 - -
+oscillating_masses_N_136_i_6 22336 5 15 12 14 5 - - 0.00764 0.01978 0.00859 0.04500 0.00409 - -
+oscillating_masses_N_136_i_7 22336 5 15 10 13 5 - - 0.00744 0.01972 0.00795 0.04260 0.00411 - -
+oscillating_masses_N_136_i_8 22336 5 15 8 13 5 - - 0.00744 0.01993 0.00757 0.04299 0.00404 - -
+oscillating_masses_N_136_i_9 22336 5 14 9 12 5 - - 0.00746 0.01840 0.00776 0.04081 0.00414 - -
+oscillating_masses_N_136_i_10 22336 6 14 10 13 6 - - 0.00844 0.01862 0.00804 0.04146 0.00469 - -
+oscillating_masses_N_136_i_11 22336 6 15 8 13 6 - - 0.00843 0.01963 0.00749 0.04219 0.00471 - -
+oscillating_masses_N_136_i_12 22336 5 15 8 12 5 - - 0.00752 0.01976 0.00757 0.04099 0.00449 - -
+oscillating_masses_N_136_i_13 22336 5 15 8 14 5 - - 0.00773 0.02007 0.00764 0.04463 0.00408 - -
+oscillating_masses_N_136_i_14 22336 5 16 8 13 5 - - 0.00747 0.02072 0.00768 0.04340 0.00417 - -
+oscillating_masses_N_136_i_15 22336 6 15 9 12 6 - - 0.00835 0.01984 0.00774 0.04062 0.00481 - -
+oscillating_masses_N_136_i_16 22336 6 16 9 13 6 - - 0.00833 0.02041 0.00778 0.04251 0.00472 - -
+oscillating_masses_N_136_i_17 22336 6 16 9 11 5 - - 0.00830 0.02048 0.00773 0.03879 0.00411 - -
+oscillating_masses_N_136_i_18 22336 5 17 9 12 5 - - 0.00749 0.02231 0.00781 0.04124 0.00415 - -
+oscillating_masses_N_136_i_19 22336 6 16 8 12 6 - - 0.00835 0.02071 0.00741 0.04067 0.00473 - -
+oscillating_masses_N_156_i_0 25616 6 15 10 13 7 - - 0.01005 0.02309 0.00922 0.04915 0.00601 - -
+oscillating_masses_N_156_i_1 25616 6 14 8 13 6 - - 0.00959 0.02071 0.00842 0.04941 0.00539 - -
+oscillating_masses_N_156_i_2 25616 7 20 11 14 5 - - 0.01111 0.02926 0.00921 0.05136 0.00469 - -
+oscillating_masses_N_156_i_3 25616 5 16 10 14 5 - - 0.00860 0.02390 0.00897 0.05125 0.00467 - -
+oscillating_masses_N_156_i_4 25616 6 16 10 12 6 - - 0.00953 0.02348 0.00909 0.04725 0.00544 - -
+oscillating_masses_N_156_i_5 25616 5 16 9 12 5 - - 0.00858 0.02381 0.00855 0.04722 0.00475 - -
+oscillating_masses_N_156_i_6 25616 5 15 10 13 5 - - 0.00871 0.02267 0.00897 0.04866 0.00466 - -
+oscillating_masses_N_156_i_7 25616 5 16 8 13 5 - - 0.00842 0.02388 0.00832 0.04869 0.00468 - -
+oscillating_masses_N_156_i_8 25616 5 15 9 12 5 - - 0.00860 0.02269 0.00877 0.04664 0.00473 - -
+oscillating_masses_N_156_i_9 25616 5 15 9 14 5 - - 0.00869 0.02238 0.00865 0.05132 0.00481 - -
+oscillating_masses_N_156_i_10 25616 6 15 9 13 6 - - 0.00988 0.02311 0.00865 0.04906 0.00544 - -
+oscillating_masses_N_156_i_11 25616 5 17 9 12 5 - - 0.00842 0.02520 0.00880 0.04729 0.00471 - -
+oscillating_masses_N_156_i_12 25616 5 15 9 12 5 - - 0.00859 0.02292 0.00881 0.04701 0.00471 - -
+oscillating_masses_N_156_i_13 25616 5 14 10 12 5 - - 0.00872 0.02105 0.00898 0.04681 0.00473 - -
+oscillating_masses_N_156_i_14 25616 6 16 8 12 6 - - 0.00954 0.02290 0.00847 0.04685 0.00539 - -
+oscillating_masses_N_156_i_15 25616 5 14 8 13 5 - - 0.00860 0.02131 0.00842 0.04894 0.00474 - -
+oscillating_masses_N_156_i_16 25616 6 13 7 11 6 - - 0.00952 0.01997 0.00811 0.04467 0.00543 - -
+oscillating_masses_N_156_i_17 25616 5 17 10 13 5 - - 0.00869 0.02538 0.00900 0.04941 0.00476 - -
+oscillating_masses_N_156_i_18 25616 7 16 11 13 7 - - 0.01143 0.02377 0.00915 0.04932 0.00598 - -
+oscillating_masses_N_156_i_19 25616 5 15 9 13 5 - - 0.00878 0.02218 0.00868 0.04924 0.00471 - -
+Table 7: Iterations and solver runtimes for mpc problems
+Iterations Solver Runtime (s)
+Problem Size qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek
+forcesExample_3 189 9 9 9 17 10 9 0.00009 0.00003 0.00014 0.00029 0.00194 0.00050
+spacecraft_1 1137 9 9 9 16 10 13 0.00039 0.00018 0.00074 0.00110 0.00217 0.00167
+dcMotor_4 5514 8 8 8 17 13 17 0.00170 0.00084 0.00370 0.00677 0.00459 0.00823
+pendulum_3 155 5 5 5 10 7 9 0.00005 0.00001 0.00008 0.00014 0.00191 0.00039
+forcesExample_2 170 7 7 7 13 12 9 0.00006 0.00002 0.00010 0.00021 0.00188 0.00046
+binaryDistillationColumn_2 2936 9 9 9 17 12 8 0.00085 0.00043 0.00164 0.00269 0.00288 0.00251
+quadcopter_2 3172 7 7 7 13 14 11 0.00105 0.00050 0.00180 0.00274 0.00353 0.00469
+pendulum_1 435 6 6 5 12 10 14 0.00012 0.00004 0.00020 0.00041 0.00209 0.00091
+shell_3 739 10 10 12 18 13 14 0.00045 0.00018 0.00085 0.00139 0.00218 0.00113
+fiordosExample_3 119 10 10 8 - 10 12 0.00006 0.00002 0.00009 - 0.00191 0.00039
+fiordosExample_1 119 8 8 7 12 13 10 0.00005 0.00001 0.00008 0.00014 0.00193 0.00037
+aircraft_11 636 11 11 12 23 9 14 0.00026 0.00010 0.00054 0.00097 0.00203 0.00099
+tripleInvertedPendulum_3 2919 8 8 8 19 14 12 0.00079 0.00038 0.00299 0.00406 0.00317 0.00515
+robotArm_1 276 5 5 5 - 9 14 0.00008 0.00002 0.00018 - 0.00196 0.00060
+aircraft_12 636 11 11 12 23 9 14 0.00026 0.00011 0.00055 0.00093 0.00202 0.00100
+quadcopter_5 3072 - - 26 - 40 20 - - 0.00705 - 0.00557 0.00596
+dcMotor_2 1114 8 8 9 17 11 15 0.00036 0.00015 0.00099 0.00135 0.00235 0.00140
+springMass_1 21239 10 10 12 24 17 40 0.00706 0.00452 0.03067 0.03561 0.01192 0.10089
+aircraft_10 584 8 8 9 15 10 13 0.00020 0.00008 0.00041 0.00055 0.00204 0.00081
+doubleInvertedPendulum_1 502 5 5 6 13 10 16 0.00012 0.00004 0.00031 0.00054 0.00211 0.00095
+springMass_3 2144 5 5 7 21 10 21 0.00045 0.00019 0.00107 0.00307 0.00247 0.00548
+robotArm_2 1056 10 10 10 - 13 21 0.00042 0.00019 0.00142 - 0.00230 0.00194
+ballOnPlate_1 398 7 7 6 10 12 13 0.00014 0.00005 0.00023 0.00038 0.00200 0.00081
+aircraft_13 3116 14 14 15 25 11 19 0.00141 0.00073 0.00334 0.00523 0.00276 0.00965
+aircraft_3 464 7 7 7 14 16 12 0.00015 0.00005 0.00025 0.00048 0.00209 0.00070
+nonlinearCstr_2 1164 9 9 10 17 11 11 0.00037 0.00016 0.00085 0.00112 0.00221 0.00164
+quadcopter_6 1678 7 7 7 15 13 12 0.00055 0.00025 0.00103 0.00207 0.00265 0.00176
+quadcopter_3 7912 7 7 7 13 14 12 0.00266 0.00138 0.00446 0.00676 0.00612 0.01181
+aircraft_4 504 8 8 9 15 10 11 0.00018 0.00007 0.00035 0.00052 0.00203 0.00076
+toyExample_1 199 6 6 7 18 9 11 0.00006 0.00002 0.00013 0.00031 0.00194 0.00054
+shell_1 739 6 6 7 15 11 10 0.00028 0.00011 0.00056 0.00116 0.00216 0.00096
+helicopter_2 1051 6 6 5 23 12 18 0.00032 0.00012 0.00068 0.00213 0.00256 0.00230
+aircraft_2 524 8 8 9 17 10 13 0.00017 0.00007 0.00035 0.00053 0.00202 0.00085
+ballOnPlate_2 398 7 7 6 12 10 14 0.00014 0.00005 0.00023 0.00044 0.00201 0.00083
+tripleInvertedPendulum_2 2919 7 7 7 19 13 14 0.00073 0.00034 0.00265 0.00393 0.00312 0.00589
+forcesExample_1 171 7 7 7 13 12 9 0.00006 0.00002 0.00010 0.00022 0.00198 0.00046
+fiordosExample_2 139 7 7 8 15 11 10 0.00006 0.00002 0.00012 0.00019 0.00193 0.00040
+dcMotor_6 1114 6 6 7 17 11 11 0.00030 0.00012 0.00110 0.00134 0.00236 0.00119
+binaryDistillationColumn_1 2936 9 9 9 17 12 8 0.00085 0.00042 0.00162 0.00262 0.00285 0.00252
+dcMotor_1 564 7 7 8 15 10 14 0.00018 0.00006 0.00048 0.00062 0.00207 0.00084
+ballOnPlate_4 658 8 8 9 15 11 16 0.00024 0.00010 0.00053 0.00091 0.00208 0.00126
+doubleInvertedPendulum_3 542 5 5 6 15 11 15 0.00013 0.00005 0.00033 0.00065 0.00208 0.00094
+quadcopter_1 1592 7 7 7 12 13 9 0.00054 0.00023 0.00092 0.00136 0.00256 0.00137
+dcMotor_5 1114 8 8 12 18 15 15 0.00036 0.00015 0.00161 0.00143 0.00241 0.00140
+shell_2 1469 7 7 7 16 11 10 0.00062 0.00027 0.00113 0.00342 0.00269 0.00174
+dcMotor_3 5514 8 8 10 21 15 20 0.00168 0.00083 0.00532 0.00803 0.00478 0.00868
+helicopter_3 536 5 5 5 22 10 18 0.00015 0.00004 0.00034 0.00105 0.00210 0.00125
+quadcopter_4 3172 10 10 10 15 15 16 0.00139 0.00069 0.00258 0.00383 0.00359 0.00554
+doubleInvertedPendulum_2 502 8 8 9 16 12 18 0.00016 0.00007 0.00041 0.00063 0.00210 0.00101
+tripleInvertedPendulum_1 2739 7 7 7 19 13 12 0.00067 0.00029 0.00250 0.00369 0.00305 0.00489
+springMass_2 2159 6 6 7 21 10 21 0.00050 0.00022 0.00110 0.00303 0.00266 0.00665
+toyExample_5 1909 9 9 11 - 17 16 0.00062 0.00029 0.00151 - 0.00312 0.00473
+toyExample_3 199 6 6 12 - 8 15 0.00007 0.00002 0.00019 - 0.00193 0.00065
+helicopter_1 201 4 4 4 21 8 13 0.00006 0.00001 0.00011 0.00034 0.00198 0.00049
+spacecraft_2 1137 9 9 9 16 10 13 0.00042 0.00018 0.00073 0.00110 0.00214 0.00168
+forcesExample_4 369 7 7 7 15 12 10 0.00012 0.00004 0.00020 0.00051 0.00208 0.00076
+toyExample_2 389 6 6 6 18 10 11 0.00012 0.00004 0.00020 0.00061 0.00207 0.00080
+ballOnPlate_3 398 7 7 7 14 9 14 0.00014 0.00005 0.00026 0.00052 0.00198 0.00084
+toyExample_4 959 9 9 10 - 14 18 0.00034 0.00014 0.00071 - 0.00242 0.00217
+nonlinearCstr_3 294 7 7 8 16 9 10 0.00009 0.00003 0.00019 0.00027 0.00190 0.00050
+nonlinearCstr_1 1164 9 9 10 17 11 11 0.00037 0.00016 0.00085 0.00112 0.00222 0.00163
+springMass_4 4279 7 7 7 22 10 23 0.00111 0.00055 0.00240 0.00637 0.00332 0.01068
+aircraft_1 504 8 8 9 15 10 11 0.00018 0.00007 0.00036 0.00053 0.00202 0.00074
+pendulum_2 432 8 8 8 13 14 15 0.00014 0.00005 0.00027 0.00045 0.00206 0.00095
 Table 8: Iterations and solver runtimes for Maros–Mészáros problems
@@
 Problem Size qoco clarabel ecos gurobi mosek qoco clarabel ecos gurobi mosek
-LASER 9216 10 10 - 21 24 0.00258 0.00414 - 0.02421 0.03150
-QPCBOEI1 4743 21 15 - 21 33 0.00613 0.00632 - 0.01534 0.03697
-HUESTIS 39968 7 8 - 14 - 0.01261 0.02198 - 0.03369 -
-QBANDM 3007 21 20 32 29 22 0.00345 0.00502 0.01198 0.01520 0.02451
-QSC205 775 15 18 17 30 15 0.00111 0.00192 0.00183 0.00950 0.00823
-LISWET11 40002 - - - 40 31 - - - 0.15115 0.26106
-LISWET8 40002 - - - 417 70 - - - 1.83018 0.90131
-AUG2DC 60200 0 0 - 2 15 0.01392 0.02377 - 0.02266 0.46025
-QSEBA 6548 30 29 - 27 30 0.00717 0.01452 - 0.02000 0.05188
-AUG2D 59800 0 0 - 2 11 0.01088 0.02133 - 0.01970 0.30137
-QGROW15 7365 18 20 17 20 28 0.00616 0.01114 0.01225 0.02356 0.06942
-CONT-200 319196 13 10 30 12 28 3.41491 3.91965 8.99180 0.75313 2.27627
-LISWET1 40002 - - - 27 52 - - - 0.09510 0.55497
-ZECEVIC2 9 6 7 8 8 7 0.00002 0.00005 0.00004 0.00275 0.00102
-QFFFFF80 8997 95 25 45 48 22 0.08216 0.03295 0.49989 0.06666 0.10850
-AUG2DCQP 80400 11 12 - 17 - 0.08341 0.14402 - 0.09956 -
-QPILOTNO 16258 39 31 43 56 27 0.09448 0.11113 0.21933 0.12265 0.15589
-CVXQP2_M 6733 13 9 30 15 17 0.03255 0.02544 11.03589 0.06452 1.07815
-POWELL20 30000 38 - - 60 - 0.05451 - - 0.21773 -
-STADAT2 15997 17 44 - 18 21 0.00939 0.04388 - 0.02200 0.04019
-QPCBOEI2 1623 40 19 - 26 48 0.00298 0.00274 - 0.00741 0.01278
-PRIMAL3 22292 11 8 18 17 8 0.01240 0.01680 0.03449 0.01265 0.04159
-QSHIP08S 21178 14 15 26 24 18 0.02043 0.02884 1.58040 0.04072 0.17699
-DUALC2 1645 18 10 13 11 16 0.00054 0.00061 0.00091 0.00860 0.00116
-LISWET4 40002 10 22 - 13 19 0.01575 0.05955 - 0.05271 0.13555
-QPCSTAIR 4872 25 22 - 23 39 0.00897 0.01026 - 0.02300 0.05212
-PRIMALC2 2076 24 15 30 11 16 0.00130 0.00142 0.00353 0.00385 0.00828
-LISWET6 40002 9 17 - 13 20 0.01906 0.05229 - 0.05445 0.14808
-QSHIP04L 8506 14 15 26 21 28 0.00769 0.00918 0.02240 0.02517 0.06894
-QE226 3824 18 24 22 25 19 0.00466 0.00709 0.02897 0.01216 0.01492
-PRIMAL4 17520 9 9 16 16 11 0.01266 0.01348 0.02799 0.00947 0.04455
-QSCSD8 13844 11 10 15 15 12 0.00735 0.01020 0.05332 0.02534 0.06593
-QSCORPIO 1824 15 11 13 24 15 0.00128 0.00169 0.00258 0.00776 0.03019
-CONT-201 290393 12 10 - 14 - 4.17722 5.71965 - 0.82300 -
-GOULDQP2 3142 9 12 20 13 10 0.00188 0.00313 0.58865 0.00962 0.05237
-LISWET10 40002 - - - 125 20 - - - 0.44295 0.14301
-HS268 40 5 10 15 12 15 0.00003 0.00007 0.00009 0.00356 0.00057
-QBORE3D 1834 18 28 24 26 20 0.00202 0.00442 0.00487 0.00950 0.01198
-QSHIP12L 83825 18 15 34 23 39 0.21193 0.19337 79.39230 0.10314 4.08685
-QSHIP04S 5866 14 14 24 19 21 0.00373 0.00592 0.01427 0.01438 0.04390
-QSHARE1B 1415 28 28 46 30 31 0.00200 0.00343 0.00512 0.00929 0.02247
-QGFRDXPN 3889 24 20 - 43 31 0.00589 0.00682 - 0.03540 0.08180
-CONT-050 19796 10 9 24 10 18 0.03964 0.04369 0.10372 0.02808 0.08417
-MOSARQP1 8467 9 10 24 15 11 0.00520 0.00843 0.03596 0.01658 0.02250
-BOYD2 517049 68 - - 78 - 2.99925 - - 6.10176 -
-QSCAGR25 2182 17 19 34 24 - 0.00160 0.00345 0.00844 0.01023 -
-HUES-MOD 39968 13 12 - 13 23 0.02296 0.02834 - 0.02918 0.36701
-YAO 8007 - - - 25 36 - - - 0.01411 0.05955
-CONT-101 72693 10 10 - 13 25 0.30755 0.44878 - 0.09657 0.47646
-HS53 24 5 6 9 8 7 0.00003 0.00004 0.00005 0.00273 0.00030
-CVXQP1_M 7482 13 10 - 16 21 0.04223 0.03895 - 0.09931 0.43868
-STCQP2 48135 7 8 25 12 14 0.08184 0.10365 0.53351 0.21152 0.16697
-QGROW22 10930 22 25 20 25 31 0.00986 0.01891 0.02126 0.03386 0.10940
-HS35 11 5 7 10 9 9 0.00002 0.00005 0.00004 0.00366 0.00077
-STCQP1 48135 7 7 28 - 24 0.03555 0.04303 0.32154 - 0.24021
-STADAT1 15997 20 15 - 185 - 0.01086 0.01243 - 0.17253 -
-PRIMAL1 6140 12 9 17 18 8 0.00325 0.00402 0.00986 0.00567 0.00878
-DUAL3 6441 9 11 13 11 7 0.00313 0.00378 0.00485 0.01875 0.00723
-HS21 8 9 8 8 8 10 0.00002 0.00003 0.00003 0.00255 0.00027
-QBEACONF 3664 12 - 29 18 22 0.00182 - 0.00717 0.00796 0.01373
-QISRAEL 3109 26 25 31 23 23 0.00366 0.00576 0.01015 0.01732 0.00747
-LISWET2 40002 9 17 - 15 20 0.01800 0.04961 - 0.05521 0.14195
-DTOC3 49994 4 6 - 8 26 0.01081 0.02869 - 0.02627 0.38133
-GOULDQP3 3840 6 7 25 9 17 0.00109 0.00237 3.86470 0.01007 0.17719
-PRIMAL2 8691 11 8 16 17 8 0.00431 0.00527 0.01403 0.00586 0.01458
-TAME 7 4 5 5 6 4 0.00002 0.00004 0.00003 0.00368 0.00031
-PRIMALC5 2860 20 13 14 11 14 0.00141 0.00149 0.00241 0.00348 0.00938
-AUG3DC 10419 0 0 14 2 10 0.00227 0.00270 0.03786 0.00950 0.04388
-AUG3DCQP 14292 9 10 29 16 22 0.01343 0.01937 0.08650 0.02952 0.16977
-QSCFXM2 7228 32 30 36 56 - 0.01022 0.01737 0.04520 0.05266 -
-VALUES 4428 12 13 - - - 0.00149 0.00242 - - -
-QAFIRO 121 15 13 14 15 13 0.00017 0.00023 0.00020 0.00347 0.00100
-QETAMACR 7761 26 19 25 28 27 0.05831 0.04680 0.91543 0.04582 0.10767
-QSCSD6 7070 14 13 16 17 13 0.00407 0.00613 0.02588 0.01711 0.05499
-QSTAIR 5423 23 33 32 41 40 0.01108 0.01966 0.03227 0.03192 0.05891
-PRIMALC1 2514 25 16 27 10 20 0.00237 0.00153 0.00347 0.00359 0.01178
-QSHELL 40488 28 36 - 33 27 0.10774 0.12657 - 0.07681 0.18787
-CVXQP2_L 67483 22 9 - 16 31 15.34297 9.37783 - 2.75816 750.19697
-QSCFXM1 3779 26 25 35 49 31 0.00448 0.00762 0.02215 0.02710 0.08131
-DUAL1 3813 10 12 14 12 12 0.00182 0.00219 0.00283 0.01026 0.00219
-UBH1 66033 7 17 - 16 8 0.08430 0.08050 - 0.05857 0.15917
-DUALC5 2276 10 10 18 14 12 0.00043 0.00084 0.00126 0.01239 0.00120
-LISWET9 40002 - - - - 62 - - - - 0.75226
-Q25FV47 71470 27 25 34 37 53 0.18642 0.18053 2.34299 0.16627 0.52865
-PRIMALC8 5182 20 12 14 10 22 0.00247 0.00322 0.00401 0.00440 0.02708
-QPCBLEND 657 16 16 17 20 17 0.00060 0.00095 0.00097 0.00530 0.00192
-LOTSCHD 72 7 8 18 11 15 0.00004 0.00008 0.00011 0.00273 0.00067
-QSCAGR7 585 15 15 29 23 32 0.00044 0.00084 0.00174 0.00461 0.00756
-QSCRS8 4472 23 31 33 35 30 0.00722 0.01372 0.01325 0.02854 0.06120
-CVXQP1_S 734 8 8 12 13 12 0.00054 0.00085 0.01330 0.00830 0.00374
-QADLITTL 567 11 13 17 16 14 0.00041 0.00076 0.00140 0.00485 0.00326
-CVXQP2_S 660 8 9 12 14 11 0.00045 0.00073 0.01244 0.00789 0.00310
-QRECIPE 988 18 16 19 20 15 0.00110 0.00153 0.00195 0.00703 0.00469
-QSCSD1 3893 11 10 12 16 12 0.00267 0.00363 0.00896 0.01347 0.01324
-GENHS28 43 0 0 9 1 5 0.00002 0.00004 0.00008 0.00576 0.00039
-QSCTAP1 2325 15 10 10 17 13 0.00176 0.00185 0.00272 0.00821 0.02044
-QBRANDY 2462 19 18 28 40 28 0.00330 0.00420 0.00662 0.01947 0.02632
-HS52 14 0 0 7 2 8 0.00002 0.00003 0.00005 0.00333 0.00039
-HS51 14 0 0 7 2 7 0.00001 0.00002 0.00004 0.00242 0.00025
-CVXQP3_S 808 9 10 13 14 14 0.00050 0.00089 0.01110 0.00681 0.00338
-CVXQP3_L 82481 - 10 - 14 42 - 21.86824 - 6.19038 187.95385
-STADAT3 31997 17 54 28 27 22 0.02096 0.10802 0.08724 0.06329 0.07766
-S268 40 5 10 15 12 15 0.00002 0.00006 0.00007 0.00266 0.00044
-QSTANDAT 5030 24 20 11 47 22 0.00536 0.00792 0.04844 0.03088 0.05748
-EXDATA 1137750 16 21 19 14 13 4.82029 6.01972 36.06845 1.83909 0.83910
-QGROW7 3550 19 21 16 20 27 0.00310 0.00556 0.00557 0.01416 0.03009
-DUAL4 3024 8 11 13 10 13 0.00134 0.00224 0.00236 0.00856 0.00247
-QFORPLAN 5591 28 21 - 56 - 0.00617 0.00799 - 0.02855 -
-QSCTAP3 12401 16 11 11 19 12 0.01999 0.01361 0.07613 0.03330 0.08330
-QSHARE2B 828 19 15 24 21 20 0.00065 0.00092 0.00141 0.00396 0.00258
-HS35MOD 12 9 11 11 15 13 0.00002 0.00003 0.00005 0.00227 0.00032
-DUAL2 4796 8 10 13 11 7 0.00168 0.00263 0.00429 0.00878 0.00594
-MOSARQP2 4775 8 9 27 16 13 0.00445 0.00614 0.02766 0.01492 0.01722
-QPTEST 10 5 7 9 10 10 0.00001 0.00003 0.00003 0.00225 0.00021
-AUG3D 9219 0 0 13 2 6 0.00191 0.00252 0.03937 0.00711 0.04820
-QSIERRA 11557 22 26 38 24 21 0.01371 0.02585 0.05065 0.03655 0.12325
-LISWET5 40002 8 9 - 13 39 0.01421 0.03237 - 0.05394 0.40437
-LISWET3 40002 9 16 - 13 20 0.01809 0.05045 - 0.05186 0.14603
-LISWET12 40002 - - - - 60 - - - - 0.72638
-CVXQP3_M 8231 23 11 - 18 21 0.08985 0.05016 - 0.14954 0.46277
-HS118 108 11 11 9 13 10 0.00007 0.00012 0.00011 0.00229 0.00048
-QSHIP12S 28344 18 15 31 24 26 0.02525 0.03026 7.90644 0.04829 0.79477
-DUALC1 1998 23 11 29 15 14 0.00076 0.00074 0.00162 0.01125 0.00095
-CVXQP1_L 74982 - 10 - 15 37 - 19.33309 - 4.44083 157.56906
-LISWET7 40002 11 - - 58 44 0.01843 - - 0.26488 0.49234
-CONT-100 79596 11 9 22 11 19 0.32731 0.38091 0.80000 0.08902 0.27225
-QSCFXM3 10369 31 32 34 58 31 0.02114 0.02997 0.07188 0.08477 0.09320
-DPKLO1 1652 0 0 13 2 4 0.00018 0.00027 0.00204 0.00283 0.00136
-KSIP 18871 14 14 21 17 11 0.00666 0.00982 0.01018 0.01295 0.00380
-DUALC8 4076 14 9 27 14 22 0.00111 0.00136 0.00493 0.02466 0.00201
-QSHIP08L 52050 14 15 28 22 20 0.11742 0.14559 13.08404 0.08743 0.45644
-AUG3DQP 13092 11 11 31 17 13 0.01592 0.02040 0.08446 0.03143 0.07889
-AUG2DQP 80000 11 13 - 17 - 0.07453 0.14683 - 0.10080 -
-HS76 20 6 6 10 10 7 0.00002 0.00004 0.00005 0.00300 0.00038
-QSCTAP2 9371 14 10 10 18 11 0.00958 0.00923 0.03774 0.02440 0.07304
-BOYD1 745507 37 - - 30 - 0.90575 - - 1.08528 -
-QCAPRI 3147 40 31 - 33 39 0.00557 0.00889 - 0.02008 0.06086
-CONT-300 653093 12 10 - 14 - 15.35224 15.73776 - 2.06562 -
-Table 3: Iterations and solver runtimes for robust Kalman filter problems
+LASER 9216 10 10 - 21 24 0.00295 0.00418 - 0.02075 0.03384
+QPCBOEI1 4743 21 15 - 22 44 0.00557 0.00623 - 0.00675 0.04344
+HUESTIS 39968 7 8 - 14 - 0.01247 0.02049 - 0.01075 -
+QBANDM 3007 21 20 30 21 17 0.00283 0.00447 0.00984 0.00504 0.00617
+QSC205 775 15 18 17 17 22 0.00085 0.00158 0.00146 0.00349 0.00221
+LISWET11 40002 - - - 40 31 - - - 0.13571 0.27070
+LISWET8 40002 - - - 417 57 - - - 1.72564 0.69070
+AUG2DC 60200 0 0 - 2 12 0.01347 0.01958 - 0.02400 0.23210
+QSEBA 6548 30 29 - 20 34 0.00622 0.01452 - 0.00404 0.01449
+AUG2D 59800 0 0 - 2 14 0.00958 0.01984 - 0.02289 0.31375
+QGROW15 7365 18 20 17 17 22 0.00529 0.00997 0.00961 0.00758 0.01993
+CONT-200 319196 13 10 30 12 24 3.08655 4.03513 8.48776 0.39624 1.77756
+LISWET1 40002 - - - 27 45 - - - 0.09196 0.50075
+ZECEVIC2 9 6 7 8 8 7 0.00003 0.00005 0.00004 0.00255 0.00106
+QFFFFF80 8997 93 25 45 26 26 0.06986 0.02945 0.42187 0.01826 0.09688
+AUG2DCQP 80400 11 12 - 17 - 0.08231 0.13210 - 0.05561 -
+QPILOTNO 16258 38 31 39 26 29 0.07808 0.09577 0.32104 0.03651 0.11626
+CVXQP2_M 6733 13 9 30 12 16 0.02790 0.02295 10.49264 0.02690 0.70517
+POWELL20 30000 38 - - 60 - 0.05336 - - 0.20098 -
+STADAT2 15997 17 44 - 18 21 0.00891 0.04415 - 0.01993 0.04982
+QPCBOEI2 1623 40 19 - 27 35 0.00236 0.00273 - 0.00413 0.00854
+PRIMAL3 22292 11 8 18 18 8 0.01258 0.01602 0.03150 0.01333 0.01050
+QSHIP08S 21178 14 15 27 20 17 0.01986 0.02778 1.55641 0.02175 0.15395
+DUALC2 1645 18 10 13 9 15 0.00053 0.00056 0.00085 0.00207 0.00080
+LISWET4 40002 10 22 - 13 20 0.01496 0.05909 - 0.05490 0.15414
+QPCSTAIR 4872 25 22 - 21 34 0.00806 0.01108 - 0.01309 0.02036
+PRIMALC2 2076 24 15 30 10 12 0.00120 0.00119 0.00306 0.00213 0.00196
+LISWET6 40002 9 17 - 13 21 0.01726 0.05086 - 0.04968 0.17387
+QSHIP04L 8506 14 15 25 17 22 0.00711 0.00944 0.01949 0.00936 0.01942
+QE226 3824 18 24 22 19 18 0.00422 0.00677 0.02736 0.00802 0.01922
+PRIMAL4 17520 9 9 16 16 9 0.01190 0.01330 0.02602 0.00949 0.01691
+QSCSD8 13844 11 10 15 14 10 0.00644 0.00955 0.05031 0.01281 0.03718
+QSCORPIO 1824 15 11 13 15 14 0.00122 0.00165 0.00236 0.00320 0.00257
+CONT-201 290393 12 10 - 12 - 3.68995 5.09101 - 0.37022 -
+GOULDQP2 3142 9 12 20 11 7 0.00177 0.00315 0.56048 0.00697 0.04653
+LISWET10 40002 - - - 125 20 - - - 0.40801 0.15746
+HS268 40 5 10 15 12 15 0.00003 0.00007 0.00009 0.00301 0.00053
+QBORE3D 1834 18 28 24 11 15 0.00198 0.00440 0.00724 0.00253 0.00180
+QSHIP12L 83825 18 15 32 18 29 0.20070 0.19269 72.93811 0.05007 2.43260
+QSHIP04S 5866 14 14 25 16 20 0.00358 0.00600 0.01307 0.00542 0.01263
+QSHARE1B 1415 28 28 43 21 26 0.00190 0.00319 0.00469 0.00467 0.02870
+QGFRDXPN 3889 24 20 - 19 - 0.00537 0.00846 - 0.00708 -
+CONT-050 19796 10 9 24 10 18 0.03774 0.04240 0.10226 0.01759 0.08744
+MOSARQP1 8467 9 10 24 16 12 0.00509 0.00884 0.03554 0.01028 0.01954
+BOYD2 517049 68 - - 90 - 1.79598 - - 0.87993 -
+QSCAGR25 2182 17 19 34 18 31 0.00173 0.00351 0.00796 0.00450 0.01722
+HUES-MOD 39968 13 12 - 13 18 0.02211 0.02751 - 0.01053 0.07958
+YAO 8007 - - - 28 34 - - - 0.01528 0.05277
+CONT-101 72693 10 10 - 11 80 0.30443 0.43120 - 0.08288 1.72033
+HS53 24 5 6 9 8 6 0.00003 0.00005 0.00007 0.00261 0.00047
+CVXQP1_M 7482 13 10 - 14 22 0.04193 0.03792 - 0.05341 0.45150
+STCQP2 48135 7 8 25 10 11 0.07536 0.10185 0.51387 0.03917 0.04279
+QGROW22 10930 22 25 20 19 24 0.00963 0.01915 0.01957 0.01047 0.04234
+HS35 11 5 7 10 9 9 0.00002 0.00005 0.00003 0.00266 0.00999
+STCQP1 48135 7 7 28 9 11 0.03430 0.04184 0.31809 0.00664 0.02806
+STADAT1 15997 20 15 - - - 0.01204 0.01224 - - -
+PRIMAL1 6140 12 9 17 18 11 0.00330 0.00377 0.00911 0.00539 0.00803
+DUAL3 6441 9 11 13 15 7 0.00287 0.00433 0.00484 0.00864 0.00370
+HS21 8 9 8 8 0 11 0.00002 0.00003 0.00003 0.00205 0.00034
+QBEACONF 3664 12 - 29 11 14 0.00182 - 0.00648 0.00268 0.00234
+QISRAEL 3109 26 25 30 21 22 0.00340 0.00542 0.00873 0.01089 0.01127
+LISWET2 40002 9 17 - 15 20 0.01736 0.04994 - 0.05463 0.15266
+DTOC3 49994 4 6 - 2 14 0.01034 0.02675 - 0.01440 0.13182
+GOULDQP3 3840 6 7 25 8 16 0.00103 0.00196 3.92135 0.00670 0.17622
+PRIMAL2 8691 11 8 16 17 7 0.00437 0.00506 0.01278 0.00645 0.00473
+TAME 7 4 5 5 6 4 0.00002 0.00004 0.00003 0.00259 0.00034
+PRIMALC5 2860 20 13 14 11 11 0.00141 0.00145 0.00216 0.00210 0.00236
+AUG3DC 10419 0 0 14 2 7 0.00201 0.00302 0.03647 0.00928 0.01731
+AUG3DCQP 14292 9 10 29 16 10 0.01278 0.01802 0.08718 0.01828 0.02658
+QSCFXM2 7228 32 30 37 31 37 0.01037 0.01937 0.04069 0.01564 0.06499
+VALUES 4428 11 13 - - - 0.00165 0.00226 - - -
+QAFIRO 121 15 13 15 12 14 0.00015 0.00021 0.00018 0.00265 0.00661
+QETAMACR 7761 26 19 26 21 21 0.05323 0.04463 0.94189 0.01268 0.06093
+QSCSD6 $70$ 14 13 16 15 13 0.00372 0.00569 0.02306 0.00688 0.04069
+QSTAIR 5423 23 33 32 20 38 0.01015 0.01912 0.02920 0.00891 0.03350
+PRIMALC1 2514 25 16 27 11 14 0.00216 0.00148 0.00321 0.00279 0.00321
+QSHELL 40488 28 36 - 38 26 0.09449 0.12115 - 0.07993 0.18460
+CVXQP2_L 67483 22 9 - 15 23 14.21961 8.60789 - 0.47604 390.09870
+QSCFXM1 3779 26 25 36 23 35 0.00385 0.00677 0.02216 0.00790 0.03538
+DUAL1 3813 10 12 14 15 12 0.00173 0.00259 0.00275 0.00672 0.00236
+UBH1 66033 7 17 - 6 11 0.02002 0.07995 - 0.01160 0.05267
+DUALC5 2276 10 10 18 11 12 0.00047 0.00071 0.00116 0.00244 0.00067
+LISWET9 40002 - - - - 53 - - - - 0.63997
+Q25FV47 71470 27 25 33 27 41 0.16469 0.17414 4.81007 0.08488 0.33967
+PRIMALC8 5182 20 12 14 9 13 0.00247 0.00326 0.00375 0.00220 0.00417
+QPCBLEND 657 16 16 17 19 20 0.00055 0.00088 0.00115 0.00355 0.00547
+LOTSCHD 72 7 8 18 10 13 0.00004 0.00006 0.00012 0.00201 0.00053
+QSCAGR7 585 15 15 29 18 49 0.00042 0.00081 0.00128 0.00274 0.00534
+QSCRS8 4472 23 31 33 26 24 0.00487 0.01325 0.01523 0.00647 0.01403
+CVXQP1_S 734 8 8 12 11 12 0.00050 0.00076 0.01201 0.00416 0.00402
+QADLITTL 567 11 13 17 12 16 0.00037 0.00070 0.00128 0.00330 0.00193
+CVXQP2_S 660 8 9 12 12 10 0.00035 0.00058 0.01285 0.00294 0.00394
+QRECIPE 988 18 16 19 19 18 0.00111 0.00149 0.00185 0.00347 0.00222
+QSCSD1 3893 11 10 12 13 9 0.00251 0.00348 0.00695 0.00614 0.00966
+GENHS28 43 0 0 9 1 5 0.00001 0.00002 0.00006 0.00262 0.01276
+QSCTAP1 2325 15 10 10 0 12 0.00173 0.00191 0.00247 0.00246 0.01201
+QBRANDY 2462 19 18 27 17 14 0.00299 0.00430 0.00576 0.00514 0.00510
+HS52 14 0 0 7 2 6 0.00001 0.00003 0.00004 0.00266 0.00030
+HS51 14 0 0 7 2 7 0.00001 0.00002 0.00004 0.00201 0.00025
+CVXQP3_S 808 9 10 13 12 14 0.00052 0.00086 0.01384 0.00406 0.00455
+CVXQP3_L 82481 - 10 - 18 24 - 21.84589 - 2.47151 106.68357
+STADAT3 31997 16 54 28 26 22 0.01792 0.10851 0.07749 0.04597 0.07562
+S268 40 5 10 15 12 15 0.00002 0.00005 0.00007 0.00239 0.00047
+QSTANDAT 5030 24 20 11 18 18 0.00456 0.00757 0.03674 0.00550 0.05588
+EXDATA 1137750 16 21 19 11 13 4.14068 5.91543 41.02600 0.93205 0.88089
+QGROW7 3550 19 21 16 17 21 0.00292 0.00498 0.00441 0.00494 0.01086
+DUAL4 3024 8 11 13 14 13 0.00105 0.00190 0.00228 0.00526 0.00334
+QFORPLAN 5591 28 21 - 23 - 0.00554 0.00757 - 0.00635 -
+QSCTAP3 12401 16 11 11 0 10 0.01539 0.01249 0.06928 0.00262 0.06155
+QSHARE2B 828 19 15 24 18 21 0.00060 0.00086 0.00130 0.00313 0.00202
+HS35MOD 12 9 11 11 9 13 0.00002 0.00004 0.00003 0.00203 0.00031
+DUAL2 4796 8 10 13 13 8 0.00144 0.00237 0.00345 0.00495 0.00281
+MOSARQP2 4775 8 9 27 16 13 0.00384 0.00567 0.02281 0.00931 0.01393
+QPTEST 10 5 7 9 11 10 0.00002 0.00003 0.00003 0.00204 0.00030
+AUG3D 9219 0 0 13 2 8 0.00148 0.00230 0.03595 0.00389 0.03126
+QSIERRA 11557 22 26 38 18 30 0.01250 0.02299 0.04104 0.01296 0.05868
+LISWET5 40002 8 9 - 13 24 0.01327 0.03012 - 0.05090 0.20424
+LISWET3 40002 9 16 - 13 20 0.01723 0.05153 - 0.04864 0.15546
+LISWET12 40002 - - - - 51 - - - - 0.61876
+CVXQP3_M 8231 23 11 - 14 24 0.09018 0.05447 - 0.08596 0.55480
+HS118 108 11 11 9 12 10 0.00008 0.00012 0.00010 0.00250 0.00051
+QSHIP12S 28344 18 15 32 17 19 0.02393 0.02954 7.88479 0.01928 0.33531
+DUALC1 1998 23 11 29 12 14 0.00088 0.00079 0.00184 0.00230 0.00084
+CVXQP1_L 74982 - 10 - 13 34 - 20.36505 - 1.39803 127.34169
+LISWET7 40002 11 - - 58 46 0.01622 - - 0.23969 0.53076
+CONT-100 79596 11 9 22 12 20 0.27128 0.33981 0.67966 0.08552 0.35265
+QSCFXM3 10369 31 32 35 29 43 0.01834 0.02726 0.06242 0.01993 0.10596
+DPKLO1 1652 0 0 13 2 5 0.00016 0.00025 0.00173 0.00276 0.00111
+KSIP 18871 14 14 21 17 11 0.00483 0.00848 0.00887 0.01248 0.00428
+DUALC8 4076 14 9 27 10 17 0.00104 0.00143 0.00446 0.00241 0.00098
+QSHIP08L 52050 14 15 28 19 17 0.09698 0.12749 11.94011 0.04549 0.47082
+AUG3DQP 13092 11 11 31 18 17 0.01532 0.01932 0.08252 0.01974 0.10652
+AUG2DQP 80000 11 13 - 19 - 0.07272 0.13726 - 0.05924 -
+HS76 20 6 6 10 9 7 0.00002 0.00004 0.00004 0.00232 0.00054
+QSCTAP2 9371 14 10 10 0 9 0.00771 0.00859 0.02937 0.00252 0.06080
+BOYD1 745507 37 - - 30 - 0.93186 - - 0.31081 -
+QCAPRI 3147 40 31 - 29 21 0.00544 0.00852 - 0.00698 0.01174
+CONT-300 653093 12 10 - 12 - 13.65596 15.11321 - 0.84948 -
+Table 9: Iterations and solver runtimes for SuiteSparse problems
 Iterations Solver Runtime (s)
-Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
-robust_kalman_filter_N_25_i_0 775 12 15 - 12 9 9 0.00091 0.00167 0.00801 0.00420 0.00051 0.00020
-robust_kalman_filter_N_25_i_1 775 10 15 - 11 9 9 0.00081 0.00170 0.00895 0.00415 0.00051 0.00020
-robust_kalman_filter_N_25_i_2 775 11 18 - 12 10 10 0.00087 0.00197 0.00998 0.00416 0.00056 0.00021
-robust_kalman_filter_N_25_i_3 775 10 17 - 12 9 9 0.00082 0.00170 0.00776 0.00398 0.00044 0.00020
-robust_kalman_filter_N_25_i_4 775 11 15 - 12 9 9 0.00087 0.00166 0.00933 0.00478 0.00051 0.00019
-robust_kalman_filter_N_25_i_5 775 10 14 - 10 8 8 0.00080 0.00160 0.00887 0.00400 0.00047 0.00018
-robust_kalman_filter_N_25_i_6 775 11 15 - 12 9 9 0.00086 0.00169 0.01004 0.00413 0.00052 0.00020
-robust_kalman_filter_N_25_i_7 775 10 17 - 11 9 9 0.00080 0.00193 0.00975 0.00389 0.00051 0.00020
-robust_kalman_filter_N_25_i_8 775 10 15 - 11 9 9 0.00083 0.00163 0.00885 0.00390 0.00052 0.00020
-robust_kalman_filter_N_25_i_9 775 9 18 - 11 9 9 0.00075 0.00176 0.00724 0.00356 0.00046 0.00018
-robust_kalman_filter_N_25_i_10 775 11 16 - 13 10 10 0.00075 0.00153 0.00809 0.00419 0.00048 0.00020
-robust_kalman_filter_N_25_i_11 775 12 19 - 12 10 10 0.00080 0.00177 0.00831 0.00378 0.00049 0.00020
-robust_kalman_filter_N_25_i_12 775 10 17 - 12 9 9 0.00068 0.00159 0.00728 0.00372 0.00044 0.00018
-robust_kalman_filter_N_25_i_13 775 8 18 - 13 8 8 0.00059 0.00175 0.00799 0.00413 0.00041 0.00016
-robust_kalman_filter_N_25_i_14 775 10 17 - 12 9 9 0.00068 0.00162 0.00760 0.00381 0.00045 0.00019
-robust_kalman_filter_N_25_i_15 775 10 16 - 11 8 8 0.00070 0.00157 0.00805 0.00351 0.00042 0.00016
-robust_kalman_filter_N_25_i_16 775 10 17 - 12 9 9 0.00070 0.00161 0.00762 0.00386 0.00044 0.00018
-robust_kalman_filter_N_25_i_17 775 9 17 - 11 9 9 0.00064 0.00180 0.01127 0.00367 0.00045 0.00018
-robust_kalman_filter_N_25_i_18 775 11 16 - 12 9 9 0.00080 0.00156 0.00795 0.00396 0.00048 0.00018
-robust_kalman_filter_N_25_i_19 775 11 19 - 11 10 10 0.00075 0.00196 0.00793 0.00370 0.00051 0.00021
-robust_kalman_filter_N_50_i_0 1550 13 18 - 13 12 12 0.00167 0.00415 0.01567 0.00701 0.00127 0.00053
-robust_kalman_filter_N_50_i_1 1550 14 19 - 14 12 12 0.00209 0.00410 0.01851 0.00789 0.00126 0.00054
-robust_kalman_filter_N_50_i_2 1550 11 20 - 12 10 10 0.00167 0.00441 0.01895 0.00712 0.00108 0.00044
-robust_kalman_filter_N_50_i_3 1550 11 19 - 12 10 10 0.00169 0.00426 0.02149 0.00714 0.00108 0.00040
-robust_kalman_filter_N_50_i_4 1550 11 20 - 11 10 10 0.00166 0.00426 0.01637 0.00651 0.00108 0.00046
-robust_kalman_filter_N_50_i_5 1550 11 19 - 11 10 10 0.00168 0.00386 0.01841 0.00680 0.00101 0.00045
-robust_kalman_filter_N_50_i_6 1550 12 19 - 12 11 11 0.00157 0.00363 0.01517 0.00638 0.00103 0.00044
-robust_kalman_filter_N_50_i_7 1550 10 17 - 10 9 9 0.00135 0.00338 0.01494 0.00546 0.00089 0.00036
-robust_kalman_filter_N_50_i_8 1550 11 20 - 10 10 10 0.00145 0.00374 0.01444 0.00548 0.00096 0.00041
-robust_kalman_filter_N_50_i_9 1550 12 24 - 12 10 10 0.00152 0.00454 0.01599 0.00647 0.00097 0.00041
-robust_kalman_filter_N_50_i_10 1550 11 21 - 11 9 9 0.00145 0.00406 0.01574 0.00576 0.00093 0.00037
-robust_kalman_filter_N_50_i_11 1550 11 19 - 12 10 10 0.00146 0.00374 0.01475 0.00612 0.00099 0.00040
-robust_kalman_filter_N_50_i_12 1550 12 18 - 11 11 11 0.00155 0.00402 0.01724 0.00593 0.00117 0.00050
-robust_kalman_filter_N_50_i_13 1550 13 21 - 14 12 12 0.00199 0.00425 0.01969 0.00778 0.00127 0.00048
-robust_kalman_filter_N_50_i_14 1550 13 19 - 13 12 12 0.00167 0.00421 0.01539 0.00650 0.00109 0.00053
-robust_kalman_filter_N_50_i_15 1550 11 19 - 12 10 10 0.00172 0.00364 0.01478 0.00624 0.00096 0.00039
-robust_kalman_filter_N_50_i_16 1550 12 21 - 11 10 10 0.00155 0.00401 0.01620 0.00600 0.00094 0.00044
-robust_kalman_filter_N_50_i_17 1550 13 21 - 12 11 11 0.00165 0.00394 0.01572 0.00610 0.00102 0.00043
-robust_kalman_filter_N_50_i_18 1550 11 18 - 11 10 10 0.00144 0.00345 0.01480 0.00610 0.00102 0.00040
-robust_kalman_filter_N_50_i_19 1550 13 19 - 13 12 12 0.00170 0.00380 0.01551 0.00675 0.00116 0.00053
-robust_kalman_filter_N_75_i_0 2325 13 18 - 12 11 11 0.00250 0.00548 0.02104 0.00899 0.00153 0.00067
-robust_kalman_filter_N_75_i_1 2325 12 19 - 11 10 10 0.00229 0.00552 0.02102 0.00838 0.00147 0.00062
-robust_kalman_filter_N_75_i_2 2325 11 21 - 11 10 10 0.00221 0.00629 0.02039 0.00836 0.00145 0.00062
-robust_kalman_filter_N_75_i_3 2325 13 21 - 14 11 11 0.00247 0.00600 0.02210 0.01016 0.00161 0.00067
-robust_kalman_filter_N_75_i_4 2325 13 22 - 14 12 12 0.00251 0.00646 0.02187 0.01024 0.00163 0.00073
-robust_kalman_filter_N_75_i_5 2325 12 23 - 13 11 11 0.00232 0.00661 0.02040 0.00925 0.00154 0.00068
-robust_kalman_filter_N_75_i_6 2325 12 21 - 12 10 10 0.00230 0.00634 0.02142 0.00912 0.00143 0.00063
-robust_kalman_filter_N_75_i_7 2325 12 24 - 12 10 10 0.00230 0.00727 0.02007 0.00882 0.00146 0.00069
-robust_kalman_filter_N_75_i_8 2325 12 21 - 12 10 10 0.00229 0.00603 0.01992 0.00915 0.00142 0.00062
-robust_kalman_filter_N_75_i_9 2325 12 20 - 11 10 10 0.00232 0.00575 0.02182 0.00839 0.00143 0.00063
-robust_kalman_filter_N_75_i_10 2325 14 23 - 14 12 12 0.00264 0.00648 0.02124 0.01016 0.00168 0.00072
-robust_kalman_filter_N_75_i_11 2325 13 21 - 12 11 11 0.00246 0.00622 0.01754 0.00891 0.00154 0.00067
-robust_kalman_filter_N_75_i_12 2325 12 22 - 12 11 11 0.00234 0.00624 0.02128 0.00904 0.00154 0.00069
-robust_kalman_filter_N_75_i_13 2325 13 21 - 12 11 11 0.00245 0.00611 0.02185 0.00887 0.00154 0.00068
-robust_kalman_filter_N_75_i_14 2325 13 19 - 12 11 11 0.00244 0.00558 0.02053 0.00882 0.00160 0.00065
-robust_kalman_filter_N_75_i_15 2325 14 22 - 15 12 12 0.00267 0.00634 0.02140 0.01082 0.00165 0.00075
-robust_kalman_filter_N_75_i_16 2325 13 22 - 12 11 11 0.00246 0.00630 0.01927 0.00899 0.00154 0.00069
-robust_kalman_filter_N_75_i_17 2325 12 21 - 12 10 10 0.00231 0.00611 0.02065 0.00906 0.00142 0.00062
-robust_kalman_filter_N_75_i_18 2325 13 20 - 14 11 11 0.00250 0.00578 0.02298 0.01035 0.00154 0.00069
-robust_kalman_filter_N_75_i_19 2325 12 20 - 13 10 10 0.00229 0.00568 0.02107 0.00992 0.00144 0.00062
-robust_kalman_filter_N_125_i_0 3875 13 22 - 14 11 11 0.00418 0.01056 0.03251 0.01433 0.00257 0.00117
-robust_kalman_filter_N_125_i_1 3875 14 23 - 13 12 12 0.00447 0.01152 0.03139 0.01300 0.00282 0.00140
-robust_kalman_filter_N_125_i_2 3875 13 20 - 12 11 11 0.00420 0.01122 0.03640 0.01222 0.00300 0.00117
-robust_kalman_filter_N_125_i_3 3875 13 21 - 13 10 10 0.00488 0.01034 0.03417 0.01317 0.00240 0.00120
-robust_kalman_filter_N_125_i_4 3875 13 20 - 12 10 10 0.00416 0.00991 0.03691 0.01212 0.00247 0.00106
-robust_kalman_filter_N_125_i_5 3875 13 26 - 14 12 12 0.00418 0.01285 0.03163 0.01427 0.00275 0.00125
-robust_kalman_filter_N_125_i_6 3875 14 21 - 13 12 12 0.00443 0.01017 0.03108 0.01283 0.00276 0.00126
-robust_kalman_filter_N_125_i_7 3875 13 21 - 14 12 12 0.00416 0.01198 0.03990 0.01367 0.00319 0.00141
-robust_kalman_filter_N_125_i_8 3875 11 23 - 12 10 10 0.00421 0.01136 0.03093 0.01260 0.00252 0.00119
-robust_kalman_filter_N_125_i_9 3875 12 23 - 12 10 10 0.00454 0.01325 0.04026 0.01384 0.00275 0.00120
-robust_kalman_filter_N_125_i_10 3875 15 24 - 15 12 12 0.00554 0.01366 0.04077 0.01713 0.00317 0.00142
-robust_kalman_filter_N_125_i_11 3875 13 24 - 13 11 11 0.00496 0.01335 0.04015 0.01551 0.00300 0.00130
-robust_kalman_filter_N_125_i_12 3875 13 23 - 15 11 11 0.00493 0.01106 0.03885 0.01553 0.00258 0.00115
-robust_kalman_filter_N_125_i_13 3875 11 21 - 11 10 10 0.00365 0.01220 0.03033 0.01155 0.00277 0.00105
-robust_kalman_filter_N_125_i_14 3875 14 23 - 13 11 11 0.00519 0.01256 0.03178 0.01334 0.00298 0.00129
-robust_kalman_filter_N_125_i_15 3875 13 21 - 13 11 11 0.00485 0.01181 0.03807 0.01472 0.00300 0.00116
-robust_kalman_filter_N_125_i_16 3875 12 21 - 13 11 11 0.00459 0.01058 0.03455 0.01461 0.00260 0.00129
-robust_kalman_filter_N_125_i_17 3875 13 23 - 13 12 12 0.00419 0.01117 0.03539 0.01349 0.00280 0.00124
-robust_kalman_filter_N_125_i_18 3875 13 21 - 12 11 11 0.00416 0.01051 0.03393 0.01216 0.00258 0.00119
-robust_kalman_filter_N_125_i_19 3875 14 21 - 13 12 12 0.00440 0.01028 0.03846 0.01306 0.00281 0.00122
-robust_kalman_filter_N_175_i_0 5425 13 23 - 14 11 11 0.00576 0.01609 0.04946 0.02003 0.00372 0.00182
-robust_kalman_filter_N_175_i_1 5425 13 24 - 14 11 11 0.00586 0.01695 0.04905 0.01895 0.00369 0.00167
-robust_kalman_filter_N_175_i_2 5425 13 24 - 14 11 11 0.00581 0.01677 0.04803 0.01918 0.00385 0.00162
-robust_kalman_filter_N_175_i_3 5425 13 22 - 13 11 11 0.00583 0.01531 0.04494 0.01763 0.00383 0.00159
-robust_kalman_filter_N_175_i_4 5425 13 22 - 14 11 11 0.00675 0.01538 0.04717 0.01978 0.00374 0.00166
-robust_kalman_filter_N_175_i_5 5425 14 26 - 14 11 11 0.00622 0.01812 0.04802 0.01980 0.00373 0.00162
-robust_kalman_filter_N_175_i_6 5425 15 22 - 14 12 12 0.00657 0.01523 0.04868 0.01976 0.00405 0.00197
-robust_kalman_filter_N_175_i_7 5425 14 25 - 14 11 11 0.00651 0.01728 0.04697 0.01966 0.00377 0.00163
-robust_kalman_filter_N_175_i_8 5425 14 24 - 15 11 11 0.00620 0.01666 0.04773 0.02065 0.00381 0.00183
-robust_kalman_filter_N_175_i_9 5425 14 24 - 15 12 12 0.00727 0.01647 0.04878 0.02087 0.00401 0.00198
-robust_kalman_filter_N_175_i_10 5425 13 24 - 14 11 11 0.00584 0.01975 0.04389 0.01975 0.00431 0.00169
-robust_kalman_filter_N_175_i_11 5425 13 24 - 15 12 12 0.00679 0.01684 0.04417 0.02087 0.00403 0.00175
-robust_kalman_filter_N_175_i_12 5425 14 23 - 16 12 12 0.00621 0.01603 0.04882 0.02320 0.00398 0.00170
-robust_kalman_filter_N_175_i_13 5425 12 26 - 14 10 10 0.00543 0.02103 0.04248 0.01993 0.00403 0.00168
-robust_kalman_filter_N_175_i_14 5425 13 22 - 13 10 10 0.00674 0.01584 0.05128 0.01833 0.00350 0.00145
-robust_kalman_filter_N_175_i_15 5425 14 26 - 15 13 13 0.00618 0.01838 0.04897 0.02098 0.00427 0.00187
-robust_kalman_filter_N_175_i_16 5425 12 22 - 13 11 11 0.00553 0.01566 0.04850 0.01898 0.00372 0.00162
-robust_kalman_filter_N_175_i_17 5425 14 25 - 15 11 11 0.00623 0.01752 0.04675 0.02127 0.00372 0.00160
-robust_kalman_filter_N_175_i_18 5425 13 25 - 14 12 12 0.00594 0.01770 0.04420 0.01985 0.00397 0.00170
-robust_kalman_filter_N_175_i_19 5425 13 24 - 15 11 11 0.00584 0.01683 0.04848 0.02111 0.00374 0.00161
-robust_kalman_filter_N_225_i_0 6975 13 24 - 14 11 - 0.00757 0.02219 0.05807 0.02588 0.00496 -
-robust_kalman_filter_N_225_i_1 6975 13 24 - 13 11 - 0.00748 0.02199 0.05756 0.02292 0.00512 -
-robust_kalman_filter_N_225_i_2 6975 13 26 - 15 12 - 0.00754 0.02367 0.06344 0.02702 0.00524 -
-robust_kalman_filter_N_225_i_3 6975 12 23 - 13 10 - 0.00708 0.02084 0.06611 0.02305 0.00457 -
-robust_kalman_filter_N_225_i_4 6975 13 25 - 14 11 - 0.00757 0.02393 0.06104 0.02447 0.00496 -
-robust_kalman_filter_N_225_i_5 6975 13 25 - 13 11 - 0.00874 0.02696 0.05785 0.02351 0.00492 -
-robust_kalman_filter_N_225_i_6 6975 13 26 - 15 11 - 0.00880 0.02462 0.06758 0.02683 0.00490 -
-robust_kalman_filter_N_225_i_7 6975 13 24 - 16 11 - 0.00755 0.02537 0.05830 0.02795 0.00567 -
-robust_kalman_filter_N_225_i_8 6975 13 24 - 13 11 - 0.00877 0.02552 0.06356 0.02308 0.00569 -
-robust_kalman_filter_N_225_i_9 6975 12 25 - 14 11 - 0.00816 0.02334 0.05694 0.02564 0.00492 -
-robust_kalman_filter_N_225_i_10 6975 14 26 - 15 12 - 0.00796 0.02748 0.06755 0.02696 0.00605 -
-robust_kalman_filter_N_225_i_11 6975 14 25 - 15 12 - 0.00928 0.02284 0.06145 0.02675 0.00528 -
-robust_kalman_filter_N_225_i_12 6975 12 25 - 12 10 - 0.00697 0.02360 0.06125 0.02157 0.00459 -
-robust_kalman_filter_N_225_i_13 6975 13 24 - 14 11 - 0.00759 0.02204 0.08955 0.02526 0.00563 -
-robust_kalman_filter_N_225_i_14 6975 13 24 - 15 11 - 0.00754 0.02213 0.05645 0.02666 0.00491 -
-robust_kalman_filter_N_225_i_15 6975 14 24 - 13 10 - 0.00797 0.02213 0.06271 0.02311 0.00459 -
-robust_kalman_filter_N_225_i_16 6975 13 24 - 13 11 - 0.00750 0.02191 0.06950 0.02277 0.00486 -
-robust_kalman_filter_N_225_i_17 6975 14 23 - 13 11 - 0.00794 0.02353 0.06576 0.02390 0.00566 -
-robust_kalman_filter_N_225_i_18 6975 13 25 - 15 11 - 0.00877 0.02643 0.06683 0.03059 0.00569 -
-robust_kalman_filter_N_225_i_19 6975 13 25 - 15 12 - 0.00878 0.02639 0.06711 0.03064 0.00606 -
-robust_kalman_filter_N_300_i_0 9300 14 23 - 15 11 - 0.01230 0.02920 0.07393 0.03522 0.00675 -
-robust_kalman_filter_N_300_i_1 9300 13 25 - 16 11 - 0.00996 0.03078 0.07924 0.03700 0.00673 -
-robust_kalman_filter_N_300_i_2 9300 14 25 - 13 10 - 0.01053 0.03125 0.08683 0.03018 0.00626 -
-robust_kalman_filter_N_300_i_3 9300 13 25 - 14 11 - 0.01003 0.03218 0.08651 0.03377 0.00679 -
-robust_kalman_filter_N_300_i_4 9300 13 23 - 13 11 - 0.01006 0.02924 0.08816 0.03026 0.00672 -
-robust_kalman_filter_N_300_i_5 9300 13 25 - 15 11 - 0.01000 0.03138 0.08341 0.03518 0.00674 -
-robust_kalman_filter_N_300_i_6 9300 13 24 - 14 11 - 0.00996 0.03027 0.08081 0.03316 0.00685 -
-robust_kalman_filter_N_300_i_7 9300 14 25 - 16 11 - 0.01068 0.03071 0.08734 0.03687 0.00688 -
-robust_kalman_filter_N_300_i_8 9300 13 25 - 15 10 - 0.00999 0.03179 0.08249 0.03536 0.00639 -
-robust_kalman_filter_N_300_i_9 9300 14 26 - 13 11 - 0.01063 0.03282 0.08182 0.03022 0.00677 -
-robust_kalman_filter_N_300_i_10 9300 13 24 - 13 11 - 0.01002 0.03008 0.08660 0.03015 0.00678 -
-robust_kalman_filter_N_300_i_11 9300 13 25 - 16 11 - 0.01003 0.03607 0.08646 0.03671 0.00776 -
-robust_kalman_filter_N_300_i_12 9300 14 26 - 15 11 - 0.01245 0.03253 0.07619 0.04085 0.00691 -
-robust_kalman_filter_N_300_i_13 9300 13 25 - 15 11 - 0.01001 0.03161 0.08319 0.03523 0.00678 -
-robust_kalman_filter_N_300_i_14 9300 14 25 - 15 11 - 0.01066 0.03163 0.07509 0.03532 0.00786 -
-robust_kalman_filter_N_300_i_15 9300 13 27 - 15 11 - 0.00997 0.03292 0.08569 0.03499 0.00687 -
-robust_kalman_filter_N_300_i_16 9300 16 26 - 16 12 - 0.01194 0.03260 0.08105 0.03728 0.00725 -
-robust_kalman_filter_N_300_i_17 9300 13 23 - 13 10 - 0.00987 0.02901 0.09474 0.03012 0.00634 -
-robust_kalman_filter_N_300_i_18 9300 13 24 - 13 10 - 0.00986 0.03074 0.08284 0.03015 0.00637 -
-robust_kalman_filter_N_300_i_19 9300 13 25 - 19 11 - 0.01003 0.03526 0.08672 0.05072 0.00690 -
-robust_kalman_filter_N_375_i_0 11625 13 26 - 13 11 - 0.01465 0.04878 0.10385 0.03771 0.01025 -
-robust_kalman_filter_N_375_i_1 11625 13 26 - 16 10 - 0.01451 0.04146 0.10711 0.04625 0.00820 -
-robust_kalman_filter_N_375_i_2 11625 12 26 - 14 11 - 0.01180 0.04283 0.10340 0.04159 0.00882 -
-robust_kalman_filter_N_375_i_3 11625 13 26 - 14 11 - 0.01253 0.04203 0.09593 0.03980 0.00894 -
-robust_kalman_filter_N_375_i_4 11625 13 24 - 16 11 - 0.01258 0.03876 0.09791 0.04590 0.00876 -
-robust_kalman_filter_N_375_i_5 11625 12 23 - 13 10 - 0.01174 0.03728 0.11823 0.03755 0.00820 -
-robust_kalman_filter_N_375_i_6 11625 14 25 - 16 12 - 0.01331 0.04077 0.13839 0.05476 0.00940 -
-robust_kalman_filter_N_375_i_7 11625 14 25 - 15 11 - 0.01339 0.04071 0.11648 0.04368 0.00874 -
-robust_kalman_filter_N_375_i_8 11625 13 26 - 16 11 - 0.01259 0.04182 0.10793 0.04589 0.01014 -
-robust_kalman_filter_N_375_i_9 11625 14 27 - 15 11 - 0.01346 0.04374 0.13064 0.04391 0.00891 -
-robust_kalman_filter_N_375_i_10 11625 14 25 - 14 11 - 0.01348 0.04713 0.11812 0.04127 0.01033 -
-robust_kalman_filter_N_375_i_11 11625 12 26 - 15 10 - 0.01367 0.04176 0.10039 0.04392 0.00828 -
-robust_kalman_filter_N_375_i_12 11625 13 24 - 16 11 - 0.01250 0.04469 0.11300 0.04598 0.01010 -
-robust_kalman_filter_N_375_i_13 11625 13 27 - 14 11 - 0.01467 0.04935 0.11771 0.04580 0.01018 -
-robust_kalman_filter_N_375_i_14 11625 14 26 - 15 11 - 0.01557 0.04223 0.13709 0.04383 0.00888 -
-robust_kalman_filter_N_375_i_15 11625 14 26 - 15 11 - 0.01346 0.04199 0.10943 0.04412 0.00887 -
-robust_kalman_filter_N_375_i_16 11625 13 27 - 16 12 - 0.01264 0.04334 0.10027 0.04589 0.01077 -
-robust_kalman_filter_N_375_i_17 11625 13 26 - 16 11 - 0.01258 0.04814 0.10561 0.04790 0.01013 -
-robust_kalman_filter_N_375_i_18 11625 13 24 - 15 11 - 0.01459 0.03878 0.11506 0.05058 0.01005 -
-robust_kalman_filter_N_375_i_19 11625 14 25 - 16 11 - 0.01344 0.04601 0.11903 0.04579 0.01016 -
-robust_kalman_filter_N_450_i_0 13950 13 26 - 16 11 - 0.01755 0.05943 0.13787 0.06379 0.01251 -
-robust_kalman_filter_N_450_i_1 13950 13 27 - 15 11 - 0.01761 0.06068 0.13805 0.06043 0.01265 -
-robust_kalman_filter_N_450_i_2 13950 13 26 - 15 11 - 0.01755 0.05289 0.15038 0.06417 0.01084 -
-robust_kalman_filter_N_450_i_3 13950 13 28 - 15 10 - 0.01499 0.06268 0.13392 0.05238 0.01051 -
-robust_kalman_filter_N_450_i_4 13950 13 31 - 15 11 - 0.01753 0.05969 0.12347 0.06045 0.01090 -
-robust_kalman_filter_N_450_i_5 13950 14 26 - 15 11 - 0.01596 0.05177 0.13747 0.05258 0.01252 -
-robust_kalman_filter_N_450_i_6 13950 14 29 - 16 11 - 0.01600 0.05609 0.13698 0.05515 0.01088 -
-robust_kalman_filter_N_450_i_7 13950 13 27 - 16 11 - 0.01501 0.05145 0.14600 0.05486 0.01091 -
-robust_kalman_filter_N_450_i_8 13950 14 28 - 16 12 - 0.01609 0.05432 0.13131 0.05507 0.01152 -
-robust_kalman_filter_N_450_i_9 13950 13 29 - 14 10 - 0.01497 0.05663 0.14099 0.04749 0.01039 -
-robust_kalman_filter_N_450_i_10 13950 14 27 - 15 10 - 0.01595 0.06025 0.14367 0.05227 0.01175 -
-robust_kalman_filter_N_450_i_11 13950 12 25 - 15 10 - 0.01635 0.05020 0.13812 0.06005 0.01029 -
-robust_kalman_filter_N_450_i_12 13950 15 26 - 16 12 - 0.01697 0.05084 0.13032 0.05577 0.01160 -
-robust_kalman_filter_N_450_i_13 13950 13 26 - 16 11 - 0.01504 0.05123 0.12504 0.05499 0.01091 -
-robust_kalman_filter_N_450_i_14 13950 13 26 - 14 10 - 0.01502 0.05176 0.13912 0.04784 0.01034 -
-robust_kalman_filter_N_450_i_15 13950 14 28 - 16 10 - 0.01606 0.05606 0.12310 0.05692 0.01182 -
-robust_kalman_filter_N_450_i_16 13950 14 29 - 16 11 - 0.01590 0.05792 0.13368 0.05491 0.01091 -
-robust_kalman_filter_N_450_i_17 13950 13 27 - 14 11 - 0.01747 0.05376 0.15784 0.05465 0.01091 -
-robust_kalman_filter_N_450_i_18 13950 14 26 - 15 13 - 0.01610 0.05996 0.13933 0.05233 0.01228 -
-robust_kalman_filter_N_450_i_19 13950 13 25 - 17 11 - 0.01744 0.04958 0.14799 0.06067 0.01084 -
-robust_kalman_filter_N_500_i_0 15500 14 25 - 14 11 - 0.01781 0.05705 0.18759 0.05262 0.01240 -
-robust_kalman_filter_N_500_i_1 15500 13 25 - 16 11 - 0.01664 0.05575 0.14857 0.06107 0.01249 -
-robust_kalman_filter_N_500_i_2 15500 13 25 - 15 10 - 0.01664 0.05711 0.15082 0.05883 0.01156 -
-robust_kalman_filter_N_500_i_3 15500 13 26 - 14 11 - 0.01672 0.06066 0.16043 0.05296 0.01227 -
-robust_kalman_filter_N_500_i_4 15500 13 29 - 16 11 - 0.01941 0.07490 0.15324 0.06134 0.01428 -
-robust_kalman_filter_N_500_i_5 15500 13 25 - 15 10 - 0.01652 0.06453 0.14032 0.05833 0.01341 -
-robust_kalman_filter_N_500_i_6 15500 14 29 - 16 11 - 0.02072 0.06508 0.17977 0.06136 0.01437 -
-robust_kalman_filter_N_500_i_7 15500 13 25 - 14 11 - 0.01673 0.06524 0.16417 0.05285 0.01429 -
-robust_kalman_filter_N_500_i_8 15500 14 27 - 16 11 - 0.02074 0.07022 0.17136 0.06185 0.01437 -
-robust_kalman_filter_N_500_i_9 15500 13 30 - 16 10 - 0.01941 0.06671 0.17077 0.06119 0.01201 -
-robust_kalman_filter_N_500_i_10 15500 14 26 - 15 11 - 0.01779 0.05815 0.14614 0.05873 0.01234 -
-robust_kalman_filter_N_500_i_11 15500 13 28 - 15 11 - 0.01657 0.06165 0.14662 0.05840 0.01236 -
-robust_kalman_filter_N_500_i_12 15500 13 26 - 16 10 - 0.01666 0.05863 0.15932 0.06092 0.01163 -
-robust_kalman_filter_N_500_i_13 15500 12 25 - 15 10 - 0.01554 0.05651 0.14555 0.05786 0.01232 -
-robust_kalman_filter_N_500_i_14 15500 13 26 - 16 11 - 0.01957 0.05692 0.16304 0.06113 0.01237 -
-robust_kalman_filter_N_500_i_15 15500 14 27 - 16 11 - 0.01773 0.05933 0.19075 0.06103 0.01254 -
-robust_kalman_filter_N_500_i_16 15500 13 26 - 16 11 - 0.01670 0.05790 0.16259 0.06289 0.01237 -
-robust_kalman_filter_N_500_i_17 15500 13 27 - 16 11 - 0.01671 0.06048 0.15740 0.06095 0.01250 -
-robust_kalman_filter_N_500_i_18 15500 13 26 - 15 11 - 0.01664 0.05763 0.16208 0.05853 0.01254 -
-robust_kalman_filter_N_500_i_19 15500 13 28 - 16 11 - 0.01677 0.07028 0.16353 0.06141 0.01430 -
-Table 4: Iterations and solver runtimes for lossless convexification problems
-Iterations Solver Runtime (s)
-Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
-lcvx_N_15_i_0 644 8 7 - 7 9 9 0.00084 0.00067 0.00505 0.00198 0.00055 0.00026
-lcvx_N_15_i_1 644 8 7 - 7 9 9 0.00082 0.00065 0.00515 0.00193 0.00056 0.00024
-lcvx_N_15_i_2 644 10 7 - 7 10 10 0.00097 0.00068 0.00535 0.00194 0.00060 0.00028
-lcvx_N_15_i_3 644 8 7 - 8 10 10 0.00085 0.00064 0.00505 0.00240 0.00060 0.00028
-lcvx_N_15_i_4 644 9 7 - 7 9 9 0.00090 0.00059 0.00443 0.00195 0.00047 0.00023
-lcvx_N_15_i_5 644 10 7 - 7 9 9 0.00082 0.00059 0.00432 0.00162 0.00051 0.00023
-lcvx_N_15_i_6 644 7 7 - 7 10 10 0.00065 0.00058 0.00440 0.00162 0.00051 0.00026
-lcvx_N_15_i_7 644 9 7 - 7 10 10 0.00076 0.00058 0.00454 0.00161 0.00051 0.00026
-lcvx_N_15_i_8 644 8 7 - 7 9 9 0.00071 0.00058 0.00430 0.00168 0.00047 0.00023
-lcvx_N_15_i_9 644 8 7 - 8 10 10 0.00072 0.00056 0.00457 0.00215 0.00051 0.00025
-lcvx_N_15_i_10 644 8 7 - 7 9 9 0.00071 0.00059 0.00456 0.00173 0.00048 0.00023
-lcvx_N_15_i_11 644 7 7 - 7 9 9 0.00065 0.00058 0.00430 0.00163 0.00047 0.00023
-lcvx_N_15_i_12 644 9 7 - 8 10 10 0.00083 0.00055 0.00444 0.00195 0.00051 0.00025
-lcvx_N_15_i_13 644 7 7 - 7 9 9 0.00064 0.00058 0.00446 0.00163 0.00047 0.00023
-lcvx_N_15_i_14 644 8 7 - 7 10 10 0.00071 0.00056 0.00443 0.00174 0.00053 0.00026
-lcvx_N_15_i_15 644 8 7 - 7 10 10 0.00077 0.00056 0.00445 0.00163 0.00051 0.00025
-lcvx_N_15_i_16 644 7 7 - 7 9 9 0.00065 0.00058 0.00448 0.00160 0.00047 0.00023
-lcvx_N_15_i_17 644 8 7 - 7 10 10 0.00075 0.00059 0.00459 0.00163 0.00052 0.00025
-lcvx_N_15_i_18 644 8 7 - 7 9 9 0.00071 0.00059 0.00444 0.00161 0.00048 0.00023
-lcvx_N_15_i_19 644 9 7 - 8 10 10 0.00078 0.00059 0.00461 0.00197 0.00051 0.00025
-lcvx_N_50_i_0 2114 9 8 - 10 10 10 0.00278 0.00205 0.01024 0.00657 0.00166 0.00093
-lcvx_N_50_i_1 2114 9 7 - 11 10 10 0.00261 0.00185 0.01066 0.00742 0.00167 0.00094
-lcvx_N_50_i_2 2114 9 7 - 10 10 10 0.00260 0.00212 0.01193 0.00662 0.00192 0.00101
-lcvx_N_50_i_3 2114 10 9 - 13 10 10 0.00339 0.00269 0.01270 0.00951 0.00194 0.00100
-lcvx_N_50_i_4 2114 8 7 - 10 10 10 0.00281 0.00209 0.01227 0.00737 0.00191 0.00103
-lcvx_N_50_i_5 2114 8 7 - 11 10 10 0.00244 0.00185 0.01040 0.00765 0.00166 0.00092
-lcvx_N_50_i_6 2114 9 7 - 11 10 10 0.00259 0.00182 0.01134 0.00764 0.00165 0.00093
-lcvx_N_50_i_7 2114 8 8 - 10 10 10 0.00239 0.00218 0.01072 0.00672 0.00167 0.00099
-lcvx_N_50_i_8 2114 10 8 - 10 10 10 0.00274 0.00213 0.01097 0.00670 0.00165 0.00090
-lcvx_N_50_i_9 2114 10 8 - 10 10 10 0.00277 0.00214 0.01102 0.00672 0.00167 0.00087
-lcvx_N_50_i_10 2114 10 8 - 11 10 10 0.00278 0.00212 0.01095 0.00784 0.00168 0.00089
-lcvx_N_50_i_11 2114 10 8 - 11 10 10 0.00291 0.00203 0.01069 0.00713 0.00166 0.00092
-lcvx_N_50_i_12 2114 9 7 - 10 10 10 0.00257 0.00183 0.01087 0.00660 0.00166 0.00092
-lcvx_N_50_i_13 2114 9 8 - 11 10 10 0.00272 0.00204 0.01022 0.00728 0.00167 0.00100
-lcvx_N_50_i_14 2114 10 7 - 10 10 10 0.00289 0.00210 0.01576 0.00823 0.00189 0.00102
-lcvx_N_50_i_15 2114 9 8 - 11 10 10 0.00300 0.00207 0.01070 0.00796 0.00167 0.00092
-lcvx_N_50_i_16 2114 9 7 - 10 10 10 0.00255 0.00176 0.01094 0.00739 0.00167 0.00091
-lcvx_N_50_i_17 2114 10 9 - 11 10 10 0.00290 0.00269 0.01292 0.00770 0.00192 0.00101
-lcvx_N_50_i_18 2114 9 8 - 11 10 10 0.00310 0.00238 0.01204 0.00853 0.00192 0.00102
-lcvx_N_50_i_19 2114 12 9 - 12 10 10 0.00382 0.00267 0.01244 0.00827 0.00191 0.00101
-lcvx_N_75_i_0 3164 10 8 - 10 10 10 0.00483 0.00367 0.01726 0.01255 0.00281 0.00157
-lcvx_N_75_i_1 3164 10 9 - 12 10 10 0.00503 0.00395 0.01880 0.01296 0.00284 0.00157
-lcvx_N_75_i_2 3164 10 8 - 11 10 10 0.00520 0.00350 0.01749 0.01170 0.00284 0.00160
-lcvx_N_75_i_3 3164 8 8 - 10 10 10 0.00420 0.00326 0.01437 0.01028 0.00242 0.00139
-lcvx_N_75_i_4 3164 10 9 - 12 10 10 0.00426 0.00398 0.01629 0.01286 0.00282 0.00159
-lcvx_N_75_i_5 3164 9 8 - 10 10 10 0.00456 0.00373 0.01735 0.01230 0.00288 0.00159
-lcvx_N_75_i_6 3164 8 8 - 10 10 10 0.00421 0.00314 0.01460 0.01023 0.00243 0.00137
-lcvx_N_75_i_7 3164 11 9 - 11 10 10 0.00464 0.00343 0.01502 0.01178 0.00242 0.00141
-lcvx_N_75_i_8 3164 10 8 - 11 10 10 0.00437 0.00302 0.01486 0.01073 0.00244 0.00142
-lcvx_N_75_i_9 3164 11 9 - 11 10 10 0.00460 0.00341 0.01494 0.01078 0.00244 0.00139
-lcvx_N_75_i_10 3164 10 8 - 10 10 10 0.00404 0.00311 0.01478 0.01152 0.00243 0.00141
-lcvx_N_75_i_11 3164 8 8 - 10 10 10 0.00363 0.00314 0.01447 0.01010 0.00244 0.00141
-lcvx_N_75_i_12 3164 11 9 - 11 10 10 0.00470 0.00344 0.01473 0.01054 0.00243 0.00142
-lcvx_N_75_i_13 3164 9 8 - 11 10 10 0.00410 0.00300 0.01479 0.01059 0.00244 0.00158
-lcvx_N_75_i_14 3164 10 9 - 11 10 10 0.00442 0.00348 0.01499 0.01232 0.00246 0.00144
-lcvx_N_75_i_15 3164 10 9 - 11 10 10 0.00438 0.00346 0.01438 0.01071 0.00243 0.00140
-lcvx_N_75_i_16 3164 10 8 - 11 10 10 0.00440 0.00300 0.01516 0.01072 0.00247 0.00145
-lcvx_N_75_i_17 3164 11 9 - 11 10 10 0.00467 0.00345 0.01504 0.01073 0.00247 0.00142
-lcvx_N_75_i_18 3164 10 8 - 11 10 10 0.00407 0.00326 0.01829 0.01182 0.00244 0.00140
-lcvx_N_75_i_19 3164 9 8 - 10 10 10 0.00383 0.00301 0.01468 0.01010 0.00243 0.00145
-lcvx_N_100_i_0 4214 9 8 - 10 10 10 0.00531 0.00413 0.01846 0.01415 0.00332 0.00202
-lcvx_N_100_i_1 4214 10 9 - 11 10 10 0.00587 0.00528 0.02145 0.01364 0.00380 0.00219
-lcvx_N_100_i_2 4214 12 9 - 12 11 11 0.00772 0.00505 0.02538 0.01594 0.00419 0.00245
-lcvx_N_100_i_3 4214 10 9 - 11 10 10 0.00713 0.00530 0.02058 0.01506 0.00382 0.00222
-lcvx_N_100_i_4 4214 8 8 - 10 10 10 0.00522 0.00469 0.02155 0.01427 0.00382 0.00224
-lcvx_N_100_i_5 4214 11 9 - 11 10 10 0.00756 0.00530 0.02170 0.01503 0.00379 0.00221
-lcvx_N_100_i_6 4214 10 9 - 11 10 10 0.00692 0.00461 0.01840 0.01704 0.00329 0.00188
-lcvx_N_100_i_7 4214 11 9 - 12 11 11 0.00616 0.00435 0.01749 0.01617 0.00362 0.00216
-lcvx_N_100_i_8 4214 9 8 - 10 10 10 0.00523 0.00412 0.01727 0.01262 0.00332 0.00196
-lcvx_N_100_i_9 4214 9 8 - 10 10 10 0.00474 0.00404 0.01726 0.01276 0.00331 0.00191
-lcvx_N_100_i_10 4214 10 8 - 11 10 10 0.00560 0.00395 0.01779 0.01347 0.00331 0.00189
-lcvx_N_100_i_11 4214 10 8 - 11 10 10 0.00574 0.00388 0.01737 0.01600 0.00331 0.00190
-lcvx_N_100_i_12 4214 10 9 - 12 10 10 0.00592 0.00456 0.01834 0.01565 0.00331 0.00194
-lcvx_N_100_i_13 4214 10 9 - 12 11 11 0.00578 0.00507 0.01823 0.01437 0.00418 0.00242
-lcvx_N_100_i_14 4214 11 9 - 12 11 11 0.00722 0.00435 0.02542 0.01744 0.00388 0.00205
-lcvx_N_100_i_15 4214 10 9 - 11 10 10 0.00565 0.00536 0.02161 0.01337 0.00387 0.00223
-lcvx_N_100_i_16 4214 10 8 - 10 10 10 0.00666 0.00479 0.02177 0.01598 0.00380 0.00220
-lcvx_N_100_i_17 4214 12 9 - 12 10 10 0.00783 0.00533 0.02160 0.01689 0.00381 0.00219
-lcvx_N_100_i_18 4214 10 8 - 10 10 10 0.00668 0.00410 0.01870 0.01297 0.00334 0.00221
-lcvx_N_100_i_19 4214 11 9 - 11 10 10 0.00760 0.00529 0.02164 0.01657 0.00380 0.00222
-lcvx_N_125_i_0 5264 11 9 - 12 11 11 0.00889 0.00639 0.03142 0.02272 0.00506 0.00297
-lcvx_N_125_i_1 5264 11 9 - 11 10 10 0.00876 0.00562 0.03142 0.01875 0.00406 0.00235
-lcvx_N_125_i_2 5264 10 9 - 12 10 10 0.00742 0.00549 0.02604 0.02137 0.00407 0.00233
-lcvx_N_125_i_3 5264 11 9 - 11 10 10 0.00759 0.00566 0.02148 0.01894 0.00401 0.00237
-lcvx_N_125_i_4 5264 9 8 - 10 10 10 0.00647 0.00494 0.02034 0.01619 0.00407 0.00232
-lcvx_N_125_i_5 5264 11 9 - 12 10 10 0.00770 0.00540 0.02805 0.01804 0.00406 0.00233
-lcvx_N_125_i_6 5264 11 9 - 11 10 10 0.00745 0.00566 0.02949 0.01920 0.00406 0.00234
-lcvx_N_125_i_7 5264 11 8 - 11 10 10 0.00727 0.00513 0.02713 0.01875 0.00404 0.00275
-lcvx_N_125_i_8 5264 9 8 - 10 10 10 0.00589 0.00582 0.02469 0.01816 0.00489 0.00276
-lcvx_N_125_i_9 5264 10 9 - 11 10 10 0.00749 0.00572 0.02529 0.01747 0.00409 0.00272
-lcvx_N_125_i_10 5264 9 8 - 11 10 10 0.00597 0.00513 0.01997 0.02011 0.00405 0.00234
-lcvx_N_125_i_11 5264 8 8 - 11 10 10 0.00561 0.00499 0.02039 0.01770 0.00411 0.00237
-lcvx_N_125_i_12 5264 9 8 - 11 10 10 0.00714 0.00491 0.02246 0.01729 0.00410 0.00233
-lcvx_N_125_i_13 5264 11 8 - 10 10 10 0.00701 0.00595 0.03027 0.01666 0.00471 0.00275
-lcvx_N_125_i_14 5264 9 8 - 10 10 10 0.00709 0.00590 0.03029 0.01800 0.00471 0.00278
-lcvx_N_125_i_15 5264 11 9 - 12 10 10 0.00911 0.00544 0.02933 0.02052 0.00405 0.00234
-lcvx_N_125_i_16 5264 11 9 - 11 10 10 0.00753 0.00575 0.02619 0.01939 0.00402 0.00245
-lcvx_N_125_i_17 5264 11 9 - 12 11 11 0.00774 0.00546 0.02632 0.02054 0.00441 0.00256
-lcvx_N_125_i_18 5264 11 9 - 12 10 10 0.00783 0.00549 0.02611 0.02030 0.00404 0.00274
-lcvx_N_125_i_19 5264 11 9 - 11 10 10 0.00806 0.00575 0.02910 0.01742 0.00408 0.00232
-lcvx_N_150_i_0 6314 10 8 - 11 10 - 0.00786 0.00612 0.02926 0.02194 $0$
-lcvx_N_150_i_1 6314 11 9 - 12 11 - 0.00957 0.00654 0.03172 0.02258 0.00538 -
-lcvx_N_150_i_2 6314 9 8 - 11 10 - 0.00803 0.00585 0.02330 0.02031 0.00490 -
-lcvx_N_150_i_3 6314 10 9 - 12 10 - 0.00922 0.00777 0.02501 0.02377 0.00494 -
-lcvx_N_150_i_4 6314 9 8 - 11 10 - 0.00948 0.00680 0.02824 0.02464 0.00571 -
-lcvx_N_150_i_5 6314 11 10 - 12 11 - 0.00982 0.00724 0.02544 0.02362 0.00539 -
-lcvx_N_150_i_6 6314 8 8 - 11 10 - 0.00673 0.00608 0.02354 0.02240 0.00495 -
-lcvx_N_150_i_7 6314 9 8 - 11 10 - 0.00718 0.00586 0.02917 0.02252 0.00494 -
-lcvx_N_150_i_8 6314 11 9 - 11 10 - 0.00915 0.00702 0.03529 0.02027 0.00497 -
-lcvx_N_150_i_9 6314 12 11 - 12 11 - 0.01013 0.00809 0.03704 0.02204 0.00533 -
-lcvx_N_150_i_10 6314 11 9 - 11 10 - 0.00926 0.00695 0.03020 0.01989 0.00495 -
-lcvx_N_150_i_11 6314 10 8 - 11 10 - 0.00868 0.00591 0.03679 0.02231 0.00496 -
-lcvx_N_150_i_12 6314 10 8 - 11 10 - 0.00866 0.00586 0.02538 0.02202 0.00493 -
-lcvx_N_150_i_13 6314 12 9 - 13 11 - 0.01014 0.00660 0.02530 0.02384 0.00540 -
-lcvx_N_150_i_14 6314 9 8 - 11 10 - 0.00776 0.00591 0.02464 0.02160 0.00496 -
-lcvx_N_150_i_15 6314 10 8 - 11 10 - 0.00869 0.00580 0.03040 0.02198 0.00492 -
-lcvx_N_150_i_16 6314 11 9 - 12 10 - 0.00948 0.00658 0.03178 0.02267 0.00492 -
-lcvx_N_150_i_17 6314 9 8 - 11 10 - 0.00766 0.00680 0.03032 0.02280 0.00571 -
-lcvx_N_150_i_18 6314 9 8 - 11 10 - 0.00852 0.00605 0.02975 0.02419 0.00494 -
-lcvx_N_150_i_19 6314 11 9 - 11 10 - 0.00920 0.00709 0.03412 0.01958 0.00491 -
-lcvx_N_200_i_0 8414 10 9 - 12 11 - 0.01143 0.00892 0.03671 0.02733 0.00717 -
-lcvx_N_200_i_1 8414 10 9 - 12 11 - 0.01170 0.00886 0.03590 0.02720 0.00720 -
-lcvx_N_200_i_2 8414 9 9 - 12 11 - 0.01059 0.00905 0.03362 0.02715 0.00720 -
-lcvx_N_200_i_3 8414 12 9 - 12 11 - 0.01327 0.00889 0.04336 0.02782 0.00710 -
-lcvx_N_200_i_4 8414 12 9 - 12 11 - 0.01386 0.00874 0.03456 0.03067 0.00713 -
-lcvx_N_200_i_5 8414 11 9 - 11 10 - 0.01195 0.00924 0.04326 0.02603 0.00664 -
-lcvx_N_200_i_6 8414 11 8 - 11 10 - 0.01121 0.00793 0.03630 0.02847 0.00658 -
-lcvx_N_200_i_7 8414 11 9 - 12 11 - 0.01286 0.00881 0.05245 0.03026 0.00714 -
-lcvx_N_200_i_8 8414 10 8 - 12 10 - 0.01117 0.00798 0.03576 0.03125 0.00659 -
-lcvx_N_200_i_9 8414 9 8 - 11 10 - 0.00971 0.00796 0.03288 0.02907 0.00663 -
-lcvx_N_200_i_10 8414 11 12 - 12 11 - 0.01259 0.01169 0.04709 0.02880 0.00715 -
-lcvx_N_200_i_11 8414 9 8 - 13 10 - 0.01061 0.00797 0.03540 0.02946 0.00659 -
-lcvx_N_200_i_12 8414 10 8 - 11 10 - 0.01031 0.00781 0.03495 0.02987 0.00659 -
-lcvx_N_200_i_13 8414 10 8 - 11 10 - 0.01042 0.00799 0.03464 0.02826 0.00668 -
-lcvx_N_200_i_14 8414 11 11 - 12 11 - 0.01269 0.01097 0.04928 0.03008 0.00718 -
-lcvx_N_200_i_15 8414 11 9 - 12 10 - 0.01204 0.00929 0.04334 0.03131 0.00658 -
-lcvx_N_200_i_16 8414 9 8 - 11 10 - 0.00971 0.00813 0.04864 0.02568 0.00668 -
-lcvx_N_200_i_17 8414 12 9 - 12 11 - 0.01312 0.00888 0.04630 0.03020 0.00717 -
-lcvx_N_200_i_18 8414 9 8 - 11 10 - 0.00966 0.00914 0.03348 0.02590 0.00764 -
-lcvx_N_200_i_19 8414 12 10 - 12 11 - 0.01595 0.01134 0.04035 0.03376 0.00833 -
-lcvx_N_250_i_0 10514 10 9 - 12 11 - 0.01613 0.01238 0.04357 0.03788 0.01035 -
-lcvx_N_250_i_1 10514 11 10 - 12 11 - 0.01734 0.01417 0.04455 0.03854 0.01039 -
-lcvx_N_250_i_2 10514 11 9 - 12 11 - 0.01491 0.01096 0.06412 0.03377 0.00894 -
-lcvx_N_250_i_3 10514 9 8 - 11 10 - 0.01209 0.00976 0.04371 0.03196 0.00824 -
-lcvx_N_250_i_4 10514 10 8 - 12 10 - 0.01297 0.00965 0.05525 0.03753 0.00829 -
-lcvx_N_250_i_5 10514 11 8 - 11 10 - 0.01413 0.00983 0.05159 0.03588 0.00827 -
-lcvx_N_250_i_6 10514 9 8 - 12 10 - 0.01215 0.00974 0.04114 0.03851 0.00829 -
-lcvx_N_250_i_7 10514 11 8 - 11 10 - 0.01415 0.00984 0.07292 0.03271 0.00830 -
-lcvx_N_250_i_8 10514 12 12 - 12 11 - 0.01583 0.01423 0.07551 0.03456 $0$
-lcvx_N_250_i_9 10514 9 8 - 11 10 - 0.01214 0.01133 0.04417 0.03677 0.00958 -
-lcvx_N_250_i_10 10514 12 9 - 11 10 - 0.01802 0.01307 0.06299 0.03559 0.00956 -
-lcvx_N_250_i_11 10514 11 9 - 11 10 - 0.01673 0.01162 0.06861 0.03580 0.00821 -
-lcvx_N_250_i_12 10514 12 9 - 13 11 - 0.01621 0.01095 0.06591 0.03763 0.00890 -
-lcvx_N_250_i_13 10514 11 8 - 11 10 - 0.01414 0.01006 0.06107 0.03165 0.00960 -
-lcvx_N_250_i_14 10514 10 9 - 12 11 - 0.01417 0.01097 0.05955 0.03464 0.00902 -
-lcvx_N_250_i_15 10514 9 8 - 11 10 - 0.01203 0.00981 0.06903 0.03250 0.00830 -
-lcvx_N_250_i_16 10514 10 8 - 11 10 - 0.01306 0.00984 0.05888 0.03269 0.00824 -
-lcvx_N_250_i_17 10514 10 8 - 11 10 - 0.01316 0.01139 0.04364 0.03279 0.00822 -
-lcvx_N_250_i_18 10514 13 9 - 12 11 - 0.02033 0.01271 0.06754 0.03774 0.01034 -
-lcvx_N_250_i_19 10514 10 9 - 11 10 - 0.01540 0.01307 0.06816 0.03144 0.00956 -
-lcvx_N_300_i_0 12614 11 11 - 13 11 - 0.01840 0.01611 0.06711 0.04499 0.01068 -
-lcvx_N_300_i_1 12614 11 9 - 12 10 - 0.01711 0.01391 0.07451 0.04436 0.00989 -
-lcvx_N_300_i_2 12614 10 11 - 13 11 - 0.01689 0.01585 0.05271 0.04656 0.01069 -
-lcvx_N_300_i_3 12614 12 10 - 13 12 - 0.01890 0.01467 0.07362 0.04415 0.01154 -
-lcvx_N_300_i_4 12614 10 8 - 12 10 - 0.01585 0.01191 0.07343 0.04025 0.00994 -
-lcvx_N_300_i_5 12614 12 10 - 14 11 - 0.01901 0.01476 0.08295 0.04839 0.01077 -
-lcvx_N_300_i_6 12614 10 8 - 13 10 - 0.01555 0.01262 0.05116 0.04782 0.01023 -
-lcvx_N_300_i_7 12614 9 8 - 12 10 - 0.01461 0.01368 0.07382 0.04529 0.01144 -
-lcvx_N_300_i_8 12614 10 9 - 13 11 - 0.01857 0.01335 0.05193 0.04822 0.01078 -
-lcvx_N_300_i_9 12614 11 9 - 13 11 - 0.01695 0.01322 0.06888 0.04389 0.01076 -
-lcvx_N_300_i_10 12614 13 10 - 13 11 - 0.02362 0.01457 0.08371 0.04780 0.01241 -
-lcvx_N_300_i_11 12614 10 9 - 13 11 - 0.01585 0.01350 0.05985 0.04982 0.01250 -
-lcvx_N_300_i_12 12614 11 9 - 13 11 - 0.01714 0.01542 0.07645 0.04790 0.01251 -
-lcvx_N_300_i_13 12614 9 9 - 13 11 - 0.01711 0.01361 0.04904 0.04500 0.01081 -
-lcvx_N_300_i_14 12614 10 9 - 12 10 - 0.01572 0.01389 0.08113 0.04145 0.00987 -
-lcvx_N_300_i_15 12614 11 8 - 12 10 - 0.01689 0.01197 0.06695 0.04455 0.00989 -
-lcvx_N_300_i_16 12614 9 8 - 12 10 - 0.01464 0.01368 0.05260 0.04061 0.01152 -
-lcvx_N_300_i_17 12614 10 8 - 12 10 - 0.01862 0.01196 0.08553 0.04934 0.00985 -
-lcvx_N_300_i_18 12614 9 8 - 12 10 - 0.01446 0.01191 0.04953 0.04100 0.00994 -
-lcvx_N_300_i_19 12614 10 10 - 13 11 - 0.01679 0.01469 0.05266 0.04568 0.01084 -
-lcvx_N_350_i_0 14714 9 9 - 13 11 - 0.02004 0.01545 0.07680 0.05598 0.01249 -
-lcvx_N_350_i_1 14714 11 10 - 13 11 - 0.02061 0.01728 0.06032 0.05012 0.01260 -
-lcvx_N_350_i_2 14714 11 9 - 12 11 - 0.01994 0.01560 0.06840 0.05095 0.01260 -
-lcvx_N_350_i_3 14714 10 8 - 12 10 - 0.01839 0.01380 0.07929 0.05342 0.01161 -
-lcvx_N_350_i_4 14714 9 8 - 12 10 - 0.01700 0.01389 0.05833 0.04700 0.01158 -
-lcvx_N_350_i_5 14714 11 9 - 12 11 - 0.01983 0.01550 0.10310 0.05331 0.01261 -
-lcvx_N_350_i_6 14714 10 10 - 12 11 - 0.01853 0.01738 0.06233 0.04658 0.01265 -
-lcvx_N_350_i_7 14714 11 9 - 12 11 - 0.01990 0.01556 0.10463 0.05172 0.01256 -
-lcvx_N_350_i_8 14714 10 8 - 12 10 - 0.01841 0.01379 0.05820 0.04862 0.01168 -
-lcvx_N_350_i_9 14714 11 9 - 12 11 - 0.01967 0.01550 0.07686 0.04768 0.01260 -
-lcvx_N_350_i_10 14714 9 8 - 12 11 - 0.01696 0.01403 0.06066 0.05296 0.01266 -
-lcvx_N_350_i_11 14714 12 10 - 14 12 - 0.02212 0.01970 0.08369 0.05846 0.01574 -
-lcvx_N_350_i_12 14714 12 10 - 14 12 - 0.02615 0.01714 0.07824 0.05940 0.01350 -
-lcvx_N_350_i_13 14714 11 10 - 14 11 - 0.02099 0.01721 0.08116 0.05812 0.01254 -
-lcvx_N_350_i_14 14714 10 11 - 13 11 - 0.01949 0.01833 0.07396 0.05615 0.01257 -
-lcvx_N_350_i_15 14714 11 9 - 14 11 - 0.01968 0.01787 0.07639 0.06019 0.01456 -
-lcvx_N_350_i_16 14714 11 9 - 12 11 - 0.02357 0.01550 0.07644 0.05193 0.01264 -
-lcvx_N_350_i_17 14714 8 8 - 12 10 - 0.01551 0.01374 0.05710 0.05051 0.01343 -
-lcvx_N_350_i_18 14714 9 11 - 13 11 - 0.01701 0.01860 0.06107 0.05173 0.01263 -
-lcvx_N_350_i_19 14714 10 8 - 12 10 - 0.01837 0.01601 0.05973 0.05333 0.01342 -
-Table 5: Iterations and solver runtimes for group lasso problems
-Iterations Solver Runtime (s)
-Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
-group_lasso_N_1_i_0 761 4 11 - 5 6 6 0.00025 0.00109 0.01260 0.00126 0.00019 0.00006
-group_lasso_N_1_i_1 761 4 11 - 7 6 6 0.00026 0.00108 0.01325 0.00153 0.00020 0.00006
-group_lasso_N_1_i_2 761 4 10 - 7 6 6 0.00027 0.00097 0.01365 0.00152 0.00019 0.00006
-group_lasso_N_1_i_3 761 4 12 - 7 6 6 0.00026 0.00119 0.01350 0.00156 0.00020 0.00007
-group_lasso_N_1_i_4 761 4 11 - 7 6 6 0.00026 0.00110 0.01364 0.00151 0.00020 0.00007
-group_lasso_N_1_i_5 761 4 11 - 7 6 6 0.00026 0.00112 0.01389 0.00151 0.00020 0.00006
-group_lasso_N_1_i_6 761 4 11 - 6 6 6 0.00026 0.00107 0.01362 0.00139 0.00019 0.00006
-group_lasso_N_1_i_7 761 5 10 - 7 6 6 0.00030 0.00097 0.01332 0.00150 0.00019 0.00006
-group_lasso_N_1_i_8 761 4 12 - 7 6 6 0.00026 0.00114 0.01329 0.00159 0.00019 0.00006
-group_lasso_N_1_i_9 761 5 10 - 6 6 6 0.00030 0.00109 0.01349 0.00148 0.00023 0.00007
-group_lasso_N_1_i_10 761 4 9 - 7 6 6 0.00029 0.00101 0.01752 0.00167 0.00022 0.00007
-group_lasso_N_1_i_11 761 4 11 - 5 6 6 0.00030 0.00123 0.01574 0.00139 0.00022 0.00007
-group_lasso_N_1_i_12 761 5 9 - 6 6 6 0.00033 0.00100 0.01571 0.00153 0.00022 0.00007
-group_lasso_N_1_i_13 761 4 11 - 7 6 6 0.00029 0.00122 0.01684 0.00166 0.00022 0.00007
-group_lasso_N_1_i_14 761 4 12 - 6 6 6 0.00029 0.00129 0.01562 0.00154 0.00022 0.00007
-group_lasso_N_1_i_15 761 4 12 - 7 6 6 0.00030 0.00121 0.01704 0.00169 0.00022 0.00007
-group_lasso_N_1_i_16 761 4 11 - 7 6 6 0.00030 0.00121 0.01793 0.00166 0.00022 0.00007
-group_lasso_N_1_i_17 761 4 12 - 7 6 6 0.00029 0.00120 0.01522 0.00164 0.00019 0.00006
-group_lasso_N_1_i_18 761 5 11 - 7 6 6 0.00029 0.00110 0.01455 0.00152 0.00019 0.00007
-group_lasso_N_1_i_19 761 5 10 - 5 6 6 0.00030 0.00098 0.01329 0.00126 0.00019 0.00007
-group_lasso_N_2_i_0 2022 6 11 - 7 7 7 0.00080 0.00246 0.03452 0.00311 0.00054 0.00021
-group_lasso_N_2_i_1 2022 7 11 - 7 7 7 0.00087 0.00248 0.04259 0.00309 0.00053 0.00022
-group_lasso_N_2_i_2 2022 5 11 - 7 7 7 0.00073 0.00292 0.03949 0.00314 0.00057 0.00022
-group_lasso_N_2_i_3 2022 7 12 - 8 8 8 0.00102 0.00316 0.04699 0.00371 0.00072 0.00024
-group_lasso_N_2_i_4 2022 6 11 - 7 7 7 0.00092 0.00293 0.04221 0.00346 0.00062 0.00022
-group_lasso_N_2_i_5 2022 7 11 - 9 8 8 0.00103 0.00243 0.04056 0.00357 0.00057 0.00021
-group_lasso_N_2_i_6 2022 6 11 - 7 7 7 0.00079 0.00245 0.03730 0.00311 0.00052 0.00020
-group_lasso_N_2_i_7 2022 6 10 - 7 7 7 0.00078 0.00230 0.03744 0.00310 0.00054 0.00020
-group_lasso_N_2_i_8 2022 6 10 - 8 7 7 0.00079 0.00234 0.03938 0.00333 0.00053 0.00020
-group_lasso_N_2_i_9 2022 6 11 - 8 7 7 0.00079 0.00235 0.03837 0.00336 0.00053 0.00022
-group_lasso_N_2_i_10 2022 6 11 - 7 6 6 0.00080 0.00289 0.04676 0.00311 0.00056 0.00019
-group_lasso_N_2_i_11 2022 7 11 - 7 7 7 0.00102 0.00287 0.04623 0.00345 0.00061 0.00019
-group_lasso_N_2_i_12 2022 6 10 - 7 7 7 0.00091 0.00270 0.05127 0.00349 0.00061 0.00022
-group_lasso_N_2_i_13 2022 7 13 - 7 7 7 0.00101 0.00347 0.04492 0.00356 0.00062 0.00022
-group_lasso_N_2_i_14 2022 6 12 - 7 7 7 0.00094 0.00259 0.03859 0.00344 0.00055 0.00020
-group_lasso_N_2_i_15 2022 6 11 - 8 7 7 0.00079 0.00243 0.03694 0.00336 0.00052 0.00022
-group_lasso_N_2_i_16 2022 6 11 - 7 7 7 0.00081 0.00240 0.04211 0.00311 0.00052 0.00022
-group_lasso_N_2_i_17 2022 6 11 - 7 7 7 0.00080 0.00251 0.03627 0.00307 0.00052 0.00020
-group_lasso_N_2_i_18 2022 7 11 - 8 7 7 0.00087 0.00286 0.04580 0.00335 0.00062 0.00022
-group_lasso_N_2_i_19 2022 6 10 - 8 7 7 0.00092 0.00235 0.03820 0.00342 0.00052 0.00022
-group_lasso_N_3_i_0 3783 7 12 - 7 8 8 0.00179 0.00632 0.09303 0.00640 0.00136 0.00045
-group_lasso_N_3_i_1 3783 7 12 - 8 8 8 0.00181 0.00663 0.10263 0.00699 0.00131 0.00045
-group_lasso_N_3_i_2 3783 6 11 - 8 7 7 0.00167 0.00665 0.06413 0.00738 0.00134 0.00036
-group_lasso_N_3_i_3 3783 6 11 - 8 7 7 0.00184 0.00611 0.09396 0.00745 0.00122 0.00040
-group_lasso_N_3_i_4 3783 7 12 - 9 8 8 0.00179 0.00652 0.09673 0.00747 0.00129 0.00045
-group_lasso_N_3_i_5 3783 6 12 - 8 7 7 0.00169 0.00627 0.09973 0.00699 0.00123 0.00040
-group_lasso_N_3_i_6 3783 6 12 - 8 7 7 0.00167 0.00674 0.09808 0.00601 0.00123 0.00040
-group_lasso_N_3_i_7 3783 7 12 - 9 8 8 0.00178 0.00665 0.10586 0.00756 0.00131 0.00045
-group_lasso_N_3_i_8 3783 8 12 - 9 8 8 0.00195 0.00654 0.10374 0.00751 0.00131 0.00045
-group_lasso_N_3_i_9 3783 7 12 - 9 7 7 0.00178 0.00671 0.09236 0.00744 0.00124 0.00040
-group_lasso_N_3_i_10 3783 7 12 - 8 8 8 0.00179 0.00694 0.10241 0.00586 0.00145 0.00041
-group_lasso_N_3_i_11 3783 8 12 - 10 8 8 0.00221 0.00650 0.09390 0.00712 0.00134 0.00046
-group_lasso_N_3_i_12 3783 8 12 - 10 8 8 0.00194 0.00719 0.09329 0.00809 0.00146 0.00041
-group_lasso_N_3_i_13 3783 7 12 - 8 7 7 0.00203 0.00664 0.09344 0.00617 0.00122 0.00041
-group_lasso_N_3_i_14 3783 8 12 - 9 8 8 0.00198 0.00643 0.05164 0.00744 0.00136 0.00045
-group_lasso_N_3_i_15 3783 8 12 - 9 8 8 0.00192 0.00626 0.09667 0.00624 0.00132 0.00045
-group_lasso_N_3_i_16 3783 8 12 - 9 8 8 0.00198 0.00616 0.06490 0.00748 0.00137 0.00045
-group_lasso_N_3_i_17 3783 7 11 - 9 7 7 0.00184 0.00646 0.09337 0.00752 0.00123 0.00036
-group_lasso_N_3_i_18 3783 7 12 - 8 7 7 0.00200 0.00671 0.08989 0.00693 0.00124 0.00040
-group_lasso_N_3_i_19 3783 7 12 - 8 8 8 0.00179 0.00707 0.09407 0.00697 0.00140 0.00039
-group_lasso_N_4_i_0 6044 7 11 - 9 8 8 0.00372 0.00825 0.08675 0.00929 0.00254 0.00068
-group_lasso_N_4_i_1 6044 7 11 - 9 7 7 0.00331 0.01034 0.07636 0.01163 0.00241 0.00060
-group_lasso_N_4_i_2 6044 7 11 - 9 8 8 0.00332 0.00936 0.09203 0.01130 0.00233 0.00074
-group_lasso_N_4_i_3 6044 6 11 - 7 7 7 0.00279 0.01065 0.07408 0.00972 0.00239 0.00060
-group_lasso_N_4_i_4 6044 6 10 - 8 7 7 0.00309 0.00910 0.07713 0.01098 0.00240 0.00060
-group_lasso_N_4_i_5 6044 7 11 - 9 8 8 0.00332 0.00932 0.07537 0.01131 0.00234 0.00075
-group_lasso_N_4_i_6 6044 8 11 - 9 9 9 0.00319 0.01044 0.07465 0.01045 0.00272 0.00075
-group_lasso_N_4_i_7 6044 7 11 - 9 8 8 0.00332 0.01038 0.08235 0.01195 0.00257 0.00067
-group_lasso_N_4_i_8 6044 7 12 - 8 7 7 0.00332 0.01096 0.09033 0.01048 0.00239 0.00060
-group_lasso_N_4_i_9 6044 8 12 - 9 8 8 0.00364 0.01153 0.09164 0.01165 0.00254 0.00067
-group_lasso_N_4_i_10 6044 7 12 - 9 7 7 0.00338 0.01116 0.08548 0.00989 0.00240 0.00060
-group_lasso_N_4_i_11 6044 7 11 - 8 7 7 0.00350 0.01017 0.08882 0.01088 0.00240 0.00061
-group_lasso_N_4_i_12 6044 8 12 - 10 8 8 0.00373 0.01022 0.07332 0.01159 0.00231 0.00074
-group_lasso_N_4_i_13 6044 8 12 - 10 8 8 0.00325 0.01025 0.08315 0.00932 0.00229 0.00074
-group_lasso_N_4_i_14 6044 7 11 - 9 8 8 0.00303 0.01005 0.08752 0.01085 0.00255 0.00067
-group_lasso_N_4_i_15 6044 7 11 - 8 7 7 0.00332 0.00827 0.08699 0.00823 0.00240 0.00060
-group_lasso_N_4_i_16 6044 7 12 - 8 8 8 0.00328 0.01145 0.08421 0.01026 0.00248 0.00067
-group_lasso_N_4_i_17 6044 7 12 - 8 7 7 0.00326 0.01063 0.07738 0.00827 0.00220 0.00066
-group_lasso_N_4_i_18 6044 7 12 - 9 8 8 0.00307 0.01160 0.07562 0.01173 0.00255 0.00068
-group_lasso_N_4_i_19 6044 7 11 - 9 8 8 0.00336 0.01011 0.08114 0.01054 0.00258 0.00067
-group_lasso_N_5_i_0 8805 7 12 - 9 8 8 0.00539 0.01336 0.15490 0.01425 0.00376 0.00113
-group_lasso_N_5_i_1 8805 7 12 - 8 8 8 0.00478 0.01432 0.12789 0.01328 0.00416 0.00099
-group_lasso_N_5_i_2 8805 7 12 - 9 8 8 0.00527 0.01689 0.13607 0.01258 0.00412 0.00103
-group_lasso_N_5_i_3 8805 7 12 - 9 8 8 0.00520 0.01569 0.12261 0.01514 0.00412 0.00101
-group_lasso_N_5_i_4 8805 7 12 - 8 8 8 0.00522 0.01640 0.13366 0.01215 0.00409 0.00102
-group_lasso_N_5_i_5 8805 8 12 - 9 8 8 0.00557 0.01594 0.13463 0.01448 0.00412 0.00103
-group_lasso_N_5_i_6 8805 8 13 - 11 9 9 0.00557 0.01799 0.14506 0.01703 0.00430 0.00111
-group_lasso_N_5_i_7 8805 8 13 - 11 9 9 0.00557 0.01593 0.13196 0.01626 0.00433 0.00114
-group_lasso_N_5_i_8 8805 7 12 - 9 8 8 0.00578 0.01632 0.14206 0.01525 0.00414 0.00100
-group_lasso_N_5_i_9 8805 8 12 - 10 9 9 0.00563 0.01265 0.10661 0.01594 0.00396 0.00125
-group_lasso_N_5_i_10 8805 8 12 - 10 8 8 0.00507 0.01444 0.12481 0.01522 0.00386 0.00112
-group_lasso_N_5_i_11 8805 7 12 - 9 8 8 0.00474 0.01315 0.13159 0.01476 0.00375 0.00113
-group_lasso_N_5_i_12 8805 7 12 - 9 8 8 0.00471 0.01321 0.11506 0.01415 0.00376 0.00104
-group_lasso_N_5_i_13 8805 7 12 - 9 8 8 0.00480 0.01528 0.12522 0.01420 0.00379 0.00113
-group_lasso_N_5_i_14 8805 8 13 - 10 8 8 0.00504 0.01657 0.13945 0.01488 0.00412 0.00103
-group_lasso_N_5_i_15 8805 6 11 - 9 7 7 0.00489 0.01440 0.15750 0.01530 0.00392 0.00091
-group_lasso_N_5_i_16 8805 8 13 - 10 8 8 0.00556 0.01707 0.15121 0.01627 0.00411 0.00103
-group_lasso_N_5_i_17 8805 7 12 - 8 8 8 0.00530 0.01363 0.12424 0.01428 0.00412 0.00098
-group_lasso_N_5_i_18 8805 8 12 - 9 8 8 0.00552 0.01342 0.13743 0.01483 0.00376 0.00114
-group_lasso_N_5_i_19 8805 7 12 - 8 8 8 0.00477 0.01616 0.15250 0.01364 0.00410 0.00103
-group_lasso_N_8_i_0 20088 8 12 - 11 8 - 0.01663 0.03164 0.42595 0.03193 0.01264 -
-group_lasso_N_8_i_1 20088 8 12 - 11 9 - 0.01533 0.03202 0.43969 0.03056 0.01308 -
-group_lasso_N_8_i_2 20088 8 11 - 10 8 - 0.01536 0.03429 0.52114 0.03041 0.01263 -
-group_lasso_N_8_i_3 20088 8 12 - 11 9 - 0.01680 0.03238 0.53397 0.03206 0.01312 -
-group_lasso_N_8_i_4 20088 8 13 - 11 9 - 0.01536 0.03464 0.43105 0.03163 0.01310 -
-group_lasso_N_8_i_5 20088 8 12 - 11 9 - 0.01529 0.03548 0.44647 0.03194 0.01441 -
-group_lasso_N_8_i_6 20088 7 11 - 9 8 - 0.01607 0.03011 0.43955 0.03064 0.01260 -
-group_lasso_N_8_i_7 20088 8 12 - 10 8 - 0.01534 0.03510 0.37856 0.03001 0.01385 -
-group_lasso_N_8_i_8 20088 8 13 - 11 9 - 0.01667 0.03326 0.44272 0.03138 0.01322 -
-group_lasso_N_8_i_9 20088 8 12 - 10 8 - 0.01531 0.03289 0.36489 0.02823 0.01272 -
-group_lasso_N_8_i_10 20088 8 12 - 10 8 - 0.01546 0.02967 0.44515 0.03007 0.01382 -
-group_lasso_N_8_i_11 20088 8 12 - 10 8 - 0.01682 0.03171 0.36262 0.02546 0.01268 -
-group_lasso_N_8_i_12 20088 7 11 - 9 8 - 0.01448 0.03253 0.39282 0.02447 0.01382 -
-group_lasso_N_8_i_13 20088 8 12 - 10 8 - 0.01684 0.03112 0.38244 0.03053 0.01271 -
-group_lasso_N_8_i_14 20088 7 12 - 10 8 - 0.01587 0.03158 0.43929 0.02986 0.01263 -
-group_lasso_N_8_i_15 20088 8 12 - 10 8 - 0.01534 0.03216 0.43672 0.02979 0.01272 -
-group_lasso_N_8_i_16 20088 8 12 - 10 8 - 0.01541 0.03232 0.45250 0.03021 0.01268 -
-group_lasso_N_8_i_17 20088 7 12 - 9 8 - 0.01487 0.03449 0.39412 0.02670 0.01387 -
-group_lasso_N_8_i_18 20088 7 12 - 9 8 - 0.01594 0.03084 0.45118 0.02881 0.01258 -
-group_lasso_N_8_i_19 20088 8 11 - 10 8 - 0.01543 0.03438 0.48799 0.03068 0.01379 -
-group_lasso_N_10_i_0 30110 8 12 - 10 8 - 0.02990 0.04260 0.81705 0.03441 0.02297 -
-group_lasso_N_10_i_1 30110 8 13 - 10 8 - 0.02640 0.04393 0.71839 0.03325 0.02295 -
-group_lasso_N_10_i_2 30110 8 13 - 10 8 - 0.02634 0.04325 0.71089 0.03299 0.02309 -
-group_lasso_N_10_i_3 30110 8 12 - 10 8 - 0.02629 0.04314 0.82863 0.03320 0.02288 -
-group_lasso_N_10_i_4 30110 8 12 - 10 9 - 0.02683 0.04325 0.74742 0.03280 0.02406 -
-group_lasso_N_10_i_5 30110 8 12 - 10 8 - 0.02697 0.04647 0.71730 0.03281 0.02506 -
-group_lasso_N_10_i_6 30110 8 13 - 11 9 - 0.02953 0.04871 0.72293 0.03796 0.02569 -
-group_lasso_N_10_i_7 30110 8 12 - 11 8 - 0.02942 0.04018 0.78965 0.03761 0.02297 -
-group_lasso_N_10_i_8 30110 8 12 - 10 8 - 0.02650 0.04055 0.73732 0.03352 0.02290 -
-group_lasso_N_10_i_9 30110 7 12 - 10 8 - 0.02551 0.04065 0.70643 0.03675 $0$
-group_lasso_N_10_i_10 30110 8 13 - 10 9 - 0.02617 0.04443 0.71020 0.03381 0.02367 -
-group_lasso_N_10_i_11 30110 7 12 - 10 8 - 0.02521 0.04700 0.85205 0.03292 0.02512 -
-group_lasso_N_10_i_12 30110 8 12 - 11 8 - 0.02932 0.03969 0.84706 0.03805 0.02333 -
-group_lasso_N_10_i_13 30110 8 12 - 10 8 - 0.02654 0.04085 0.71497 0.03355 0.02305 -
-group_lasso_N_10_i_14 30110 8 12 - 11 8 - 0.02646 0.04156 0.71650 0.03545 0.02291 -
-group_lasso_N_10_i_15 30110 8 13 - 11 8 - 0.02719 0.04412 0.70731 0.03440 0.02288 -
-group_lasso_N_10_i_16 30110 8 12 - 10 8 - 0.02660 0.04692 0.72722 0.03345 0.02508 -
-group_lasso_N_10_i_17 30110 8 12 - 10 8 - 0.02956 0.04225 0.81926 0.03716 0.02288 -
-group_lasso_N_10_i_18 30110 8 12 - 10 8 - 0.02757 0.04193 0.71638 0.03325 0.02331 -
-group_lasso_N_10_i_19 30110 7 11 - 9 8 - 0.02621 0.03705 0.57778 0.03114 0.02293 -
-group_lasso_N_12_i_0 42132 8 12 - 10 8 - 0.04176 0.05646 1.13886 0.04270 0.03342 -
-group_lasso_N_12_i_1 42132 8 12 - 11 8 - 0.04718 0.06379 1.21747 0.04359 0.03531 -
-group_lasso_N_12_i_2 42132 7 12 - 10 7 - 0.04523 0.06090 1.29220 0.04656 0.03292 -
-group_lasso_N_12_i_3 42132 8 13 - 11 8 - 0.04103 0.06254 1.14120 0.04422 0.03333 -
-group_lasso_N_12_i_4 42132 8 12 - 11 9 - 0.04334 0.05760 1.33111 0.04989 0.03406 -
-group_lasso_N_12_i_5 42132 8 12 - 10 8 - 0.04179 0.05618 1.13627 0.04234 0.03340 -
-group_lasso_N_12_i_6 42132 8 12 - 10 8 - 0.04743 0.06710 1.14156 0.04154 0.03561 -
-group_lasso_N_12_i_7 42132 8 12 - 10 8 - 0.04925 0.06160 1.40540 0.04672 0.03352 -
-group_lasso_N_12_i_8 42132 8 12 - 10 8 - 0.04718 0.05998 1.35485 0.04598 0.03311 -
-group_lasso_N_12_i_9 42132 8 12 - 10 8 - 0.04905 0.05656 1.21177 0.04653 0.03330 -
-group_lasso_N_12_i_10 42132 8 12 - 10 8 - 0.04734 0.06416 1.16632 0.04149 0.03473 -
-group_lasso_N_12_i_11 42132 8 12 - 10 8 - 0.04929 0.05646 1.16314 0.05081 0.03312 -
-group_lasso_N_12_i_12 42132 8 12 - 11 8 - 0.04759 0.06398 1.17661 0.05339 0.03473 -
-group_lasso_N_12_i_13 42132 8 12 - 10 8 - 0.04926 0.05970 1.15317 0.05015 0.03347 -
-group_lasso_N_12_i_14 42132 8 12 - 10 8 - 0.04236 0.05699 1.15607 0.04171 0.03296 -
-group_lasso_N_12_i_15 42132 8 12 - 11 8 - 0.04935 0.06750 1.37147 0.04819 0.03532 -
-group_lasso_N_12_i_16 42132 8 12 - 11 8 - 0.05005 0.06500 1.13286 0.04832 0.03555 -
-group_lasso_N_12_i_17 42132 8 12 - 10 8 - 0.04763 0.06239 1.19049 0.04723 0.03542 -
-group_lasso_N_12_i_18 42132 8 12 - 11 8 - 0.04341 0.06100 1.56013 0.04476 0.03346 -
-group_lasso_N_12_i_19 42132 8 12 - 11 8 - 0.04741 0.06446 1.36957 0.05266 $0$
-group_lasso_N_14_i_0 56154 7 11 - 10 8 - 0.06732 0.08825 1.84934 0.06011 0.04971 -
-group_lasso_N_14_i_1 56154 8 12 - 10 8 - 0.07072 0.09272 1.74199 0.06349 0.04997 -
-group_lasso_N_14_i_2 56154 8 12 - 11 8 - 0.07052 0.08224 1.75298 0.06603 0.04582 -
-group_lasso_N_14_i_3 56154 8 12 - 10 8 - 0.06288 0.08092 1.81541 0.05672 0.04587 -
-group_lasso_N_14_i_4 56154 8 12 - 11 8 - 0.06408 0.08039 1.77375 0.06373 0.04576 -
-group_lasso_N_14_i_5 56154 7 11 - 10 8 - 0.05957 0.08723 1.73803 0.05546 0.05051 -
-group_lasso_N_14_i_6 56154 8 13 - 11 9 - 0.07036 0.08742 1.78430 0.06126 0.05179 -
-group_lasso_N_14_i_7 56154 8 13 - 11 9 - 0.06455 0.08512 1.75818 0.05494 0.04697 -
-group_lasso_N_14_i_8 56154 8 12 - 11 8 - 0.07086 0.08050 1.71973 0.06242 0.04588 -
-group_lasso_N_14_i_9 56154 7 11 - 10 8 - 0.06089 0.07734 1.76336 0.05817 0.04569 -
-group_lasso_N_14_i_10 56154 7 11 - 10 8 - 0.06136 0.07749 2.37906 0.06483 0.04582 -
-group_lasso_N_14_i_11 56154 8 12 - 11 8 - 0.06271 0.08033 1.74188 0.05966 0.04585 -
-group_lasso_N_14_i_12 56154 7 12 - 11 8 - 0.05962 0.09189 1.82984 0.05886 0.05023 -
-group_lasso_N_14_i_13 56154 8 12 - 10 8 - 0.07108 0.08138 1.73734 0.05703 0.04582 -
-group_lasso_N_14_i_14 56154 7 12 - 10 8 - 0.06805 0.09148 1.83138 0.05970 0.05013 -
-group_lasso_N_14_i_15 56154 7 12 - 9 8 - 0.06833 0.08114 1.76118 0.06040 0.04593 -
-group_lasso_N_14_i_16 56154 8 12 - 11 8 - 0.07149 0.08162 1.74090 0.05902 0.04518 -
-group_lasso_N_14_i_17 56154 8 13 - 11 8 - 0.06544 0.08407 1.72602 0.05831 0.04578 -
-group_lasso_N_14_i_18 56154 7 12 - 10 8 - 0.05882 0.08192 1.85183 0.06374 0.04648 -
-group_lasso_N_14_i_19 56154 8 12 - 10 8 - 0.07138 0.08001 1.72838 0.05553 0.04590 -
-group_lasso_N_16_i_0 72176 7 12 - 9 8 - 0.08672 0.10810 2.73494 0.06387 0.06441 -
-group_lasso_N_16_i_1 72176 8 12 - 9 8 - 0.09175 0.10956 2.57650 0.06641 0.06458 -
-group_lasso_N_16_i_2 72176 7 13 - 8 8 - 0.08420 0.11815 2.47541 0.06230 0.06381 -
-group_lasso_N_16_i_3 72176 7 12 - 8 8 - 0.09670 0.11035 2.61814 0.06587 0.06468 -
-group_lasso_N_16_i_4 72176 7 12 - 9 8 - 0.08805 0.11655 2.53132 0.06833 0.06994 -
-group_lasso_N_16_i_5 72176 7 12 - 8 8 - 0.08591 0.11024 2.56714 0.06273 0.06479 -
-group_lasso_N_16_i_6 72176 8 12 - 9 8 - 0.09397 0.12189 2.57184 0.06543 0.07089 -
-group_lasso_N_16_i_7 72176 7 12 - 8 8 - 0.09662 0.11715 2.62896 0.07091 0.07108 -
-group_lasso_N_16_i_8 72176 7 12 - 8 8 - 0.09697 0.10666 2.52536 0.06546 0.06463 -
-group_lasso_N_16_i_9 72176 8 12 - 9 8 - 0.10105 0.11045 2.59268 0.07594 0.06850 -
-group_lasso_N_16_i_10 72176 8 12 - 8 8 - 0.08739 0.12383 2.55628 0.06493 0.07137 -
-group_lasso_N_16_i_11 72176 8 12 - 8 8 - 0.09985 0.11001 2.27946 0.07167 0.06446 -
-group_lasso_N_16_i_12 72176 8 12 - 9 8 - 0.09288 0.12516 2.65303 0.06511 0.07138 -
-group_lasso_N_16_i_13 72176 8 12 - 9 8 - 0.10170 0.12524 2.70107 0.06619 0.07118 -
-group_lasso_N_16_i_14 72176 8 12 - 8 8 - 0.10075 0.11424 2.69967 0.06278 0.06494 -
-group_lasso_N_16_i_15 72176 8 12 - 9 8 - 0.10100 0.12544 2.59960 0.06730 0.07098 -
-group_lasso_N_16_i_16 72176 8 12 - 9 8 - 0.10077 0.11113 2.91541 0.07759 0.06459 -
-group_lasso_N_16_i_17 72176 7 12 - 8 8 - 0.08758 0.12504 2.53327 0.07146 0.07131 -
-group_lasso_N_16_i_18 72176 8 12 - 8 8 - 0.10206 0.11034 2.64789 0.07121 0.06484 -
-group_lasso_N_16_i_19 72176 7 12 - 8 8 - 0.08468 0.11729 2.65421 0.07299 0.07011 -
-Table 6: Iterations and solver runtimes for portfolio optimization problems
-Iterations Solver Runtime (s)
-Problem Size clarabel ecos gurobi mosek qoco qoco_custom clarabel ecos gurobi mosek qoco qoco_custom
-portfolio_N_2_i_0 804 9 12 14 12 9 9 0.00046 0.00099 0.00263 0.00485 0.00028 0.00013
-portfolio_N_2_i_1 804 9 14 13 12 8 8 0.00045 0.00108 0.00255 0.00486 0.00027 0.00011
-portfolio_N_2_i_2 804 8 12 11 11 8 8 0.00041 0.00102 0.00251 0.00461 0.00026 0.00010
-portfolio_N_2_i_3 804 8 15 12 12 8 8 0.00041 0.00120 0.00252 0.00486 0.00025 0.00011
-portfolio_N_2_i_4 804 9 14 13 11 8 8 0.00044 0.00113 0.00259 0.00461 0.00026 0.00010
-portfolio_N_2_i_5 804 10 14 13 14 9 9 0.00049 0.00112 0.00256 0.00535 0.00027 0.00012
-portfolio_N_2_i_6 804 10 18 12 12 10 10 0.00049 0.00151 0.00255 0.00485 0.00030 0.00014
-portfolio_N_2_i_7 804 8 10 11 11 7 7 0.00041 0.00090 0.00248 0.00447 0.00024 0.00010
-portfolio_N_2_i_8 804 9 17 11 10 9 9 0.00045 0.00150 0.00251 0.00423 0.00027 0.00012
-portfolio_N_2_i_9 804 16 13 14 13 9 9 0.00070 0.00101 0.00258 0.00493 0.00028 0.00012
-portfolio_N_2_i_10 804 10 13 15 13 9 9 0.00050 0.00100 0.00262 0.00490 0.00027 0.00011
-portfolio_N_2_i_11 804 10 12 13 13 10 10 0.00044 0.00087 0.00229 0.00451 0.00028 0.00013
-portfolio_N_2_i_12 804 10 11 13 12 10 10 0.00043 0.00079 0.00225 0.00426 0.00026 0.00012
-portfolio_N_2_i_13 804 9 12 11 11 8 8 0.00040 0.00085 0.00220 0.00402 0.00023 0.00010
-portfolio_N_2_i_14 804 9 14 13 13 9 9 0.00040 0.00095 0.00228 0.00444 0.00025 0.00011
-portfolio_N_2_i_15 804 9 11 12 12 8 8 0.00040 0.00077 0.00224 0.00424 0.00022 0.00010
-portfolio_N_2_i_16 804 10 14 11 12 8 8 0.00043 0.00095 0.00224 0.00425 0.00022 0.00010
-portfolio_N_2_i_17 804 8 11 10 9 8 8 0.00037 0.00081 0.00219 0.00359 0.00023 0.00010
-portfolio_N_2_i_18 804 9 13 11 11 9 9 0.00040 0.00092 0.00223 0.00400 0.00024 0.00011
-portfolio_N_2_i_19 804 10 11 12 11 7 7 0.00044 0.00081 0.00224 0.00401 0.00021 0.00009
-portfolio_N_4_i_0 2008 10 13 13 12 9 9 0.00107 0.00191 0.00293 0.00703 0.00054 0.00026
-portfolio_N_4_i_1 2008 10 14 15 12 11 11 0.00107 0.00243 0.00310 0.00709 0.00062 0.00033
-portfolio_N_4_i_2 2008 11 15 12 12 10 10 0.00131 0.00256 0.00287 0.00670 0.00058 0.00030
-portfolio_N_4_i_3 2008 11 16 14 14 9 9 0.00113 0.00220 0.00299 0.00807 0.00056 0.00027
-portfolio_N_4_i_4 2008 9 17 13 12 8 8 0.00098 0.00239 0.00284 0.00705 0.00049 0.00024
-portfolio_N_4_i_5 2008 11 13 12 14 9 9 0.00113 0.00193 0.00279 0.00805 0.00053 0.00028
-portfolio_N_4_i_6 2008 10 14 13 11 9 9 0.00107 0.00204 0.00290 0.00650 0.00054 0.00030
-portfolio_N_4_i_7 2008 11 15 13 12 9 9 0.00110 0.00211 0.00293 0.00689 0.00053 0.00028
-portfolio_N_4_i_8 2008 11 13 12 12 9 9 0.00109 0.00189 0.00292 0.00708 0.00053 0.00027
-portfolio_N_4_i_9 2008 11 16 13 13 10 10 0.00111 0.00240 0.00295 0.00742 0.00057 0.00030
-portfolio_N_4_i_10 2008 11 18 15 14 10 10 0.00112 0.00246 0.00311 0.00797 0.00057 0.00031
-portfolio_N_4_i_11 2008 9 15 14 12 9 9 0.00097 0.00223 0.00291 0.00702 0.00056 0.00027
-portfolio_N_4_i_12 2008 12 14 12 12 9 9 0.00120 0.00207 0.00278 0.00698 0.00054 0.00027
-portfolio_N_4_i_13 2008 10 15 14 11 9 9 0.00102 0.00230 0.00303 0.00666 0.00053 0.00027
-portfolio_N_4_i_14 2008 10 17 15 14 9 9 0.00106 0.00233 0.00305 0.00750 0.00054 0.00027
-portfolio_N_4_i_15 2008 11 12 12 13 10 10 0.00115 0.00185 0.00289 0.00746 0.00058 0.00030
-portfolio_N_4_i_16 2008 11 17 14 12 10 10 0.00114 0.00270 0.00292 0.00698 0.00058 0.00030
-portfolio_N_4_i_17 2008 9 13 14 13 9 9 0.00098 0.00187 0.00292 0.00748 0.00052 0.00028
-portfolio_N_4_i_18 2008 11 14 13 14 9 9 0.00112 0.00195 0.00292 0.00776 0.00052 0.00027
-portfolio_N_4_i_19 2008 9 19 12 12 8 8 0.00096 0.00286 0.00285 0.00692 0.00049 0.00027
-portfolio_N_6_i_0 3612 13 16 14 14 11 11 0.00215 0.00403 0.00395 0.01172 0.00112 0.00056
-portfolio_N_6_i_1 3612 14 16 14 12 8 8 0.00224 0.00401 0.00387 0.01036 0.00090 0.00041
-portfolio_N_6_i_2 3612 12 14 14 13 10 10 0.00210 0.00408 0.00452 0.01138 0.00124 0.00057
-portfolio_N_6_i_3 3612 15 18 13 14 10 10 0.00279 0.00505 0.00437 0.01193 0.00121 0.00057
-portfolio_N_6_i_4 3612 12 14 13 15 11 11 0.00235 0.00424 0.00442 0.01279 0.00129 0.00063
-portfolio_N_6_i_5 3612 11 13 12 12 9 9 0.00221 0.00385 0.00422 0.01174 0.00112 0.00052
-portfolio_N_6_i_6 3612 10 15 14 14 9 9 0.00208 0.00457 0.00453 0.01331 0.00118 0.00052
-portfolio_N_6_i_7 3612 12 14 14 13 9 9 0.00237 0.00393 0.00451 0.01255 0.00114 0.00051
-portfolio_N_6_i_8 3612 12 16 14 14 10 10 0.00235 0.00405 0.00454 0.01340 0.00124 0.00053
-portfolio_N_6_i_9 3612 13 15 15 14 10 10 0.00217 0.00389 0.00398 0.01178 0.00102 0.00057
-portfolio_N_6_i_10 3612 11 15 14 13 10 10 0.00198 0.00365 0.00387 0.01253 0.00104 0.00050
-portfolio_N_6_i_11 3612 11 14 13 11 9 9 0.00194 0.00415 0.00434 0.00978 0.00114 0.00051
-portfolio_N_6_i_12 3612 12 15 16 13 11 11 0.00236 0.00388 0.00410 0.01128 0.00117 0.00056
-portfolio_N_6_i_13 3612 11 15 13 13 10 10 0.00197 0.00428 0.00435 0.01116 0.00122 0.00053
-portfolio_N_6_i_14 3612 13 16 14 14 10 10 0.00252 0.00461 0.00452 0.01329 0.00122 0.00058
-portfolio_N_6_i_15 3612 12 17 13 14 10 10 0.00238 0.00484 0.00438 0.01335 0.00120 0.00058
-portfolio_N_6_i_16 3612 11 17 14 12 9 9 0.00225 0.00431 0.00386 0.01178 0.00103 0.00047
-portfolio_N_6_i_17 3612 13 15 15 13 10 10 0.00250 0.00414 0.00485 0.01259 0.00121 0.00057
-portfolio_N_6_i_18 3612 12 20 15 14 10 10 0.00244 0.00614 0.00461 0.01338 0.00123 0.00057
-portfolio_N_6_i_19 3612 13 13 12 14 9 9 0.00248 0.00333 0.00375 0.01202 0.00096 0.00046
-portfolio_N_8_i_0 5616 11 15 14 14 11 11 0.00327 0.00665 0.00574 0.01784 0.00183 0.00101
-portfolio_N_8_i_1 5616 12 15 13 13 11 11 0.00396 0.00685 0.00584 0.01858 0.00189 0.00097
-portfolio_N_8_i_2 5616 12 16 13 13 11 11 0.00392 0.00691 0.00569 0.01836 0.00181 0.00090
-portfolio_N_8_i_3 5616 13 15 16 15 10 10 0.00418 0.00551 0.00537 0.02087 0.00136 0.00087
-portfolio_N_8_i_4 5616 13 16 13 12 11 11 0.00357 0.00635 0.00478 0.01551 0.00159 0.00100
-portfolio_N_8_i_5 5616 12 15 13 15 10 10 0.00340 0.00580 0.00476 0.01890 0.00138 0.00083
-portfolio_N_8_i_6 5616 14 17 14 14 10 10 0.00377 0.00628 0.00500 0.01778 0.00137 0.00085
-portfolio_N_8_i_7 5616 12 14 15 15 9 9 0.00340 0.00544 0.00526 0.01896 0.00133 0.00084
-portfolio_N_8_i_8 5616 13 17 14 15 11 11 0.00363 0.00637 0.00525 0.01838 0.00150 0.00088
-portfolio_N_8_i_9 5616 14 18 16 15 10 10 0.00379 0.00621 0.00547 0.01805 0.00144 0.00084
-portfolio_N_8_i_10 5616 13 18 15 16 10 10 0.00360 0.00625 0.00516 0.01937 0.00135 0.00087
-portfolio_N_8_i_11 5616 12 18 15 14 10 10 0.00338 0.00660 0.00522 0.01747 0.00138 0.00084
-portfolio_N_8_i_12 5616 11 15 13 13 9 9 0.00322 0.00561 0.00483 0.01652 0.00138 0.00074
-portfolio_N_8_i_13 5616 14 17 15 14 10 10 0.00376 0.00652 0.00520 0.01770 0.00143 0.00083
-portfolio_N_8_i_14 5616 12 17 15 15 10 10 0.00342 0.00622 0.00521 0.01825 0.00136 0.00086
-portfolio_N_8_i_15 5616 14 18 15 15 10 10 0.00379 0.00653 0.00515 0.01854 0.00143 0.00084
-portfolio_N_8_i_16 5616 14 17 15 15 12 12 0.00379 0.00609 0.00514 0.01866 0.00158 0.00098
-portfolio_N_8_i_17 5616 12 15 15 13 9 9 0.00337 0.00571 0.00525 0.01621 0.00133 0.00075
-portfolio_N_8_i_18 5616 12 14 14 15 10 10 0.00345 0.00547 0.00507 0.01890 0.00136 0.00083
-portfolio_N_8_i_19 5616 12 16 15 15 10 10 0.00342 0.00583 0.00514 0.01860 0.00137 0.00083
-portfolio_N_10_i_0 8020 13 18 13 14 9 9 0.00567 0.01012 0.00598 0.02258 0.00187 0.00144
-portfolio_N_10_i_1 8020 12 19 14 13 10 10 0.00543 0.01160 0.00641 0.02101 0.00202 0.00157
-portfolio_N_10_i_2 8020 14 18 15 14 10 10 0.00595 0.00946 0.00654 0.02224 0.00202 0.00158
-portfolio_N_10_i_3 8020 14 16 15 15 10 10 0.00590 0.00945 0.00747 0.02371 0.00237 0.00178
-portfolio_N_10_i_4 8020 12 16 15 14 9 9 0.00619 0.00860 0.00655 0.02196 0.00188 0.00143
-portfolio_N_10_i_5 8020 11 15 15 14 11 11 0.00517 0.00979 0.00648 0.02214 0.00222 0.00195
-portfolio_N_10_i_6 8020 14 17 17 15 11 11 0.00587 0.01039 0.00793 0.02337 0.00256 0.00198
-portfolio_N_10_i_7 8020 14 15 14 14 11 11 0.00677 0.01008 0.00753 0.02500 0.00258 0.00195
-portfolio_N_10_i_8 8020 13 16 16 13 11 11 0.00647 0.00867 0.00691 0.02054 0.00224 0.00172
-portfolio_N_10_i_9 8020 12 15 15 13 11 11 0.00541 0.00941 0.00759 0.02041 0.00249 0.00195
-portfolio_N_10_i_10 8020 13 19 15 14 11 11 0.00653 0.01125 0.00770 0.02481 0.00257 0.00196
-portfolio_N_10_i_11 8020 14 18 17 15 10 10 0.00681 0.01118 0.00832 0.02592 0.00240 0.00161
-portfolio_N_10_i_12 8020 13 18 14 13 12 12 0.00645 0.00939 0.00615 0.02070 0.00242 0.00185
-portfolio_N_10_i_13 8020 13 13 14 14 9 9 0.00562 0.00730 0.00627 0.02257 0.00187 0.00143
-portfolio_N_10_i_14 8020 13 18 16 15 10 10 0.00577 0.00979 0.00696 0.02354 0.00206 0.00161
-portfolio_N_10_i_15 8020 12 16 15 13 11 11 0.00540 0.00855 0.00646 0.02049 0.00221 0.00174
-portfolio_N_10_i_16 8020 13 18 15 14 12 12 0.00570 0.00988 0.00650 0.02208 0.00235 0.00188
-portfolio_N_10_i_17 8020 13 17 16 15 10 10 0.00572 0.00930 0.00695 0.02366 0.00207 0.00157
-portfolio_N_10_i_18 8020 13 18 15 13 10 10 0.00565 0.01098 0.00751 0.02122 0.00236 0.00179
-portfolio_N_10_i_19 8020 13 15 14 14 12 12 0.00643 0.00842 0.00650 0.02185 0.00245 0.00186
-portfolio_N_15_i_0 15780 13 18 14 15 11 - 0.01573 0.01625 0.01104 0.03435 0.00476 -
-portfolio_N_15_i_1 15780 13 18 15 16 11 - 0.01555 0.01698 0.01102 0.03537 0.00480 -
-portfolio_N_15_i_2 15780 13 16 17 16 11 - 0.01575 0.01525 0.01227 0.03390 0.00488 -
-portfolio_N_15_i_3 15780 12 18 15 15 12 - 0.01520 0.02108 0.01112 0.03209 0.00529 -
-portfolio_N_15_i_4 15780 13 18 15 14 10 - 0.01562 0.01648 0.01098 0.03003 0.00447 -
-portfolio_N_15_i_5 15780 13 15 15 15 10 - 0.01560 0.01402 0.01105 0.03340 0.00443 -
-portfolio_N_15_i_6 15780 14 15 17 15 11 - 0.01616 0.01385 0.01165 0.03171 0.00479 -
-portfolio_N_15_i_7 15780 13 16 16 15 10 - 0.01560 0.01443 0.01140 0.03200 0.00441 -
-portfolio_N_15_i_8 15780 14 19 17 15 11 - 0.01606 0.01592 0.01228 0.03191 0.00484 -
-portfolio_N_15_i_9 15780 13 14 16 15 10 - 0.01554 0.01470 0.01153 0.03193 0.00446 -
-portfolio_N_15_i_10 15780 14 17 15 16 10 - 0.01614 0.01479 0.01128 0.03533 0.00447 -
-portfolio_N_15_i_11 15780 13 17 14 15 11 - 0.01566 0.01509 0.01094 0.03390 0.00480 -
-portfolio_N_15_i_12 15780 14 21 17 16 11 - 0.01614 0.01952 0.01213 0.03332 0.00481 -
-portfolio_N_15_i_13 15780 13 16 15 14 12 - 0.01543 0.01474 0.01108 0.02999 0.00519 -
-portfolio_N_15_i_14 15780 12 17 16 16 12 - 0.01511 0.01508 0.01151 0.03332 0.00521 -
-portfolio_N_15_i_15 15780 15 17 18 16 13 - 0.01665 0.01487 0.01277 0.03564 0.00569 -
-portfolio_N_15_i_16 15780 12 16 15 15 11 - 0.01519 0.01448 0.01143 0.03405 0.00487 -
-portfolio_N_15_i_17 15780 13 15 15 14 11 - 0.01557 0.01546 0.01118 0.03211 0.00489 -
-portfolio_N_15_i_18 15780 13 17 15 15 10 - 0.01562 0.01611 0.01145 0.03200 0.00438 -
-portfolio_N_15_i_19 15780 12 17 16 14 10 - 0.01538 0.01628 0.01147 0.03246 0.00449 -
-portfolio_N_20_i_0 26040 13 15 16 14 11 - 0.01579 0.02319 0.01691 0.04306 0.00867 -
-portfolio_N_20_i_1 26040 13 19 16 15 10 - 0.01596 0.02887 0.01702 0.04652 $0$
-portfolio_N_20_i_2 26040 13 17 16 15 13 - 0.01584 0.02901 0.01678 0.04624 0.00982 -
-portfolio_N_20_i_3 26040 13 16 16 14 11 - 0.01462 0.02260 0.01738 0.04299 0.00871 -
-portfolio_N_20_i_4 26040 15 19 16 16 13 - 0.01763 0.03265 0.01731 0.04858 0.01017 -
-portfolio_N_20_i_5 26040 13 16 16 15 10 - 0.01851 0.02140 0.01698 0.04604 0.00802 -
-portfolio_N_20_i_6 26040 14 20 16 15 11 - 0.01711 0.02815 0.01691 0.04646 0.00873 -
-portfolio_N_20_i_7 26040 15 17 17 15 11 - 0.01803 0.03054 0.02014 0.04598 0.00995 -
-portfolio_N_20_i_8 26040 13 18 15 15 11 - 0.01829 0.02487 0.01659 0.04978 0.00834 -
-portfolio_N_20_i_9 26040 14 17 15 15 11 - 0.01674 0.02765 0.01654 0.04574 0.00868 -
-portfolio_N_20_i_10 26040 14 18 17 14 11 - 0.01678 0.02426 0.01734 0.04360 0.00856 -
-portfolio_N_20_i_11 26040 13 16 17 15 13 - 0.01566 0.02243 0.01722 0.04618 0.01007 -
-portfolio_N_20_i_12 26040 15 17 16 16 11 - 0.01770 0.02716 0.01693 0.04894 0.00866 -
-portfolio_N_20_i_13 26040 13 15 14 13 10 - 0.01550 0.02817 0.01549 0.04090 0.00795 -
-portfolio_N_20_i_14 26040 13 18 17 14 10 - 0.01559 0.02457 0.01755 0.04368 0.00788 -
-portfolio_N_20_i_15 26040 13 16 16 14 11 - 0.01849 0.03036 0.01962 0.05036 0.01012 -
-portfolio_N_20_i_16 26040 13 17 15 15 12 - 0.01848 0.02328 0.01900 0.05309 0.01077 -
-portfolio_N_20_i_17 26040 13 18 17 16 12 - 0.01567 0.02475 0.01766 0.04937 0.00933 -
-portfolio_N_20_i_18 26040 13 18 17 15 10 - 0.01576 0.02845 0.01770 0.04576 0.00794 -
-portfolio_N_20_i_19 26040 14 16 16 16 11 - 0.01681 0.02650 0.01691 0.04792 0.00869 -
-portfolio_N_25_i_0 38800 15 18 18 15 12 - 0.02811 0.03697 0.02796 0.06077 0.01479 -
-portfolio_N_25_i_1 38800 13 19 17 15 12 - 0.02489 0.03911 0.02666 0.06077 0.01491 -
-portfolio_N_25_i_2 38800 14 15 17 16 11 - 0.02627 0.03189 0.02655 0.06440 0.01396 -
-portfolio_N_25_i_3 38800 15 20 16 15 12 - 0.02832 0.03838 0.02565 0.06078 0.01489 -
-portfolio_N_25_i_4 38800 13 16 16 15 11 - 0.02451 0.03297 0.02543 0.06389 0.01387 -
-portfolio_N_25_i_5 38800 14 15 16 15 12 - 0.02645 0.03208 0.02487 0.06108 0.01474 -
-portfolio_N_25_i_6 38800 14 18 16 15 11 - 0.02653 0.03622 0.02487 0.06103 0.01371 -
-portfolio_N_25_i_7 38800 14 19 17 14 13 - 0.02668 0.04121 0.03103 0.05905 0.01871 -
-portfolio_N_25_i_8 38800 14 17 16 15 11 - 0.03083 0.03273 0.02523 0.06138 0.01385 -
-portfolio_N_25_i_9 38800 14 19 17 16 12 - 0.02653 0.03570 0.02635 0.06449 0.01502 -
-portfolio_N_25_i_10 38800 14 18 15 15 11 - 0.03111 0.03593 0.02820 0.07083 0.01615 -
-portfolio_N_25_i_11 38800 15 21 17 16 11 - 0.02816 0.03953 0.02618 0.06469 0.01383 -
-portfolio_N_25_i_12 38800 15 18 16 15 12 - 0.02831 0.03316 0.02426 0.06071 0.01493 -
-portfolio_N_25_i_13 38800 16 19 17 16 11 - 0.02977 0.03784 0.02601 0.06431 0.01380 -
-portfolio_N_25_i_14 38800 14 16 17 15 10 - 0.02643 0.03204 0.02684 0.06072 0.01276 -
-portfolio_N_25_i_15 38800 13 16 17 14 11 - 0.02484 0.03209 0.02575 0.05748 0.01383 -
-portfolio_N_25_i_16 38800 14 19 16 14 12 - 0.02663 0.06186 0.02918 0.05802 0.01725 -
-portfolio_N_25_i_17 38800 14 15 17 14 11 - 0.03080 0.03863 0.03059 0.06712 0.01616 -
-portfolio_N_25_i_18 38800 15 18 16 16 12 - 0.03304 0.03557 0.02505 0.06457 0.01495 -
-portfolio_N_25_i_19 38800 14 17 17 14 12 - 0.02603 0.03617 0.02648 0.05771 0.01506 -
-portfolio_N_30_i_0 54060 15 17 17 15 11 - 0.04078 0.04232 0.03562 0.07782 0.01979 -
-portfolio_N_30_i_1 54060 15 18 19 14 11 - 0.04038 0.04706 0.03848 0.07431 0.01976 -
-portfolio_N_30_i_2 54060 14 20 16 14 14 - 0.03791 0.04927 0.03356 0.07436 0.02453 -
-portfolio_N_30_i_3 54060 14 20 17 16 12 - 0.03790 0.05631 0.03541 0.08273 0.02143 -
-portfolio_N_30_i_4 54060 14 17 17 15 13 - 0.03847 0.04522 0.03540 0.07840 0.02313 -
-portfolio_N_30_i_5 54060 14 18 18 15 11 - 0.03849 0.04513 0.03699 0.08154 0.01995 -
-portfolio_N_30_i_6 54060 12 19 16 14 11 - 0.03361 0.04782 0.03416 0.07452 0.01964 -
-portfolio_N_30_i_7 54060 14 17 17 16 11 - 0.03834 0.04254 0.03533 0.08245 0.01989 -
-portfolio_N_30_i_8 54060 15 17 16 15 12 - 0.04104 0.04122 0.03423 0.07857 0.02149 -
-portfolio_N_30_i_9 54060 14 18 17 14 11 - 0.03878 0.05691 0.03565 0.07451 0.02303 -
-portfolio_N_30_i_10 54060 13 18 16 14 11 - 0.04241 0.04240 0.03351 0.07457 0.02012 -
-portfolio_N_30_i_11 54060 14 18 17 15 12 - 0.03828 0.05314 0.03513 0.07874 0.02485 -
-portfolio_N_30_i_12 54060 14 18 16 16 11 - 0.04465 0.04759 0.03386 0.08244 0.01961 -
-portfolio_N_30_i_13 54060 15 21 16 15 12 - 0.04048 0.05437 0.03384 0.07851 0.02141 -
-portfolio_N_30_i_14 54060 16 19 17 15 11 - 0.04310 0.05265 0.03532 0.07844 0.01982 -
-portfolio_N_30_i_15 54060 14 19 17 15 12 - 0.03834 0.04615 0.03562 0.07836 0.02172 -
-portfolio_N_30_i_16 54060 15 17 17 16 14 - 0.04096 0.05073 0.04097 0.08245 0.02876 -
-portfolio_N_30_i_17 54060 16 16 15 14 11 - 0.05060 0.04024 0.03210 0.07422 0.02012 -
-portfolio_N_30_i_18 54060 15 19 16 15 12 - 0.04045 0.04823 0.03355 0.07831 0.02184 -
-portfolio_N_30_i_19 54060 16 18 16 15 11 - 0.04294 0.04589 0.03384 0.07850 0.02013 -
-portfolio_N_35_i_0 71820 15 19 18 16 12 - 0.05584 0.05739 0.04426 0.10356 0.02967 -
-portfolio_N_35_i_1 71820 15 20 16 15 12 - 0.05662 0.06042 0.04109 0.09836 0.02957 -
-portfolio_N_35_i_2 71820 14 18 16 14 11 - 0.05282 0.07135 0.04395 0.10664 0.03179 -
-portfolio_N_35_i_3 71820 15 20 17 16 13 - 0.06657 0.07905 0.04376 0.10612 0.03325 -
-portfolio_N_35_i_4 71820 16 19 16 15 13 - 0.07068 0.07492 0.04398 0.11254 0.03697 -
-portfolio_N_35_i_5 71820 14 19 17 16 13 - 0.06208 0.06215 0.04377 0.11941 0.03192 -
-portfolio_N_35_i_6 71820 15 18 16 15 12 - 0.05657 0.05702 0.03915 0.09702 0.02968 -
-portfolio_N_35_i_7 71820 17 17 17 16 12 - 0.06331 0.05526 0.04299 0.10258 0.02950 -
-portfolio_N_35_i_8 71820 15 20 16 16 13 - 0.05627 0.06575 0.04163 0.10279 0.03188 -
-portfolio_N_35_i_9 71820 14 17 18 16 12 - 0.05318 0.06848 0.04473 0.10844 0.03426 -
-portfolio_N_35_i_10 71820 15 18 17 17 11 - 0.06658 0.06619 0.04656 0.12444 0.03171 -
-portfolio_N_35_i_11 71820 15 20 15 17 12 - 0.06651 0.06644 0.04067 0.10984 0.02958 -
-portfolio_N_35_i_12 71820 15 18 16 15 11 - 0.05663 0.06125 0.04135 0.09925 0.02731 -
-portfolio_N_35_i_13 71820 15 20 17 15 12 - 0.05644 0.06690 0.04293 0.09958 0.02954 -
-portfolio_N_35_i_14 71820 16 20 19 16 13 - 0.05986 0.07504 0.04919 0.10278 0.03688 -
-portfolio_N_35_i_15 71820 15 19 16 15 11 - 0.06674 0.06148 0.04025 0.11378 0.02719 -
-portfolio_N_35_i_16 71820 14 18 18 15 13 - 0.05262 0.06579 0.04489 0.09922 0.03192 -
-portfolio_N_35_i_17 71820 13 17 16 16 13 - 0.04921 0.05822 0.04275 0.10370 0.03170 -
-portfolio_N_35_i_18 71820 16 19 18 16 13 - 0.05986 0.06595 0.04472 0.10778 0.03192 -
-portfolio_N_35_i_19 71820 16 17 17 15 12 - 0.05994 0.05646 0.04273 0.09732 0.02933 -
-Table 7: Iterations and solver runtimes for oscillating masses problems
-Iterations Solver Runtime (s)
-Problem Size clarabel ecos gurobi mosek qoco qoco_custom cvxgen clarabel ecos gurobi mosek qoco qoco_custom cvxgen
-oscillating_masses_N_8_i_0 1344 5 13 9 10 5 5 7 0.00055 0.00128 0.00327 0.00375 0.00029 0.00012 0.00016
-oscillating_masses_N_8_i_1 1344 5 14 9 10 5 5 6 0.00052 0.00132 0.00326 0.00377 0.00029 0.00011 0.00013
-oscillating_masses_N_8_i_2 1344 5 17 9 11 5 5 6 0.00052 0.00156 0.00329 0.00427 0.00030 0.00012 0.00014
-oscillating_masses_N_8_i_3 1344 5 14 9 11 5 5 6 0.00052 0.00130 0.00327 0.00422 0.00030 0.00011 0.00013
-oscillating_masses_N_8_i_4 1344 5 12 10 11 5 5 7 0.00053 0.00114 0.00338 0.00434 0.00030 0.00011 0.00015
-oscillating_masses_N_8_i_5 1344 5 12 9 10 5 5 6 0.00052 0.00113 0.00326 0.00381 0.00030 0.00012 0.00014
-oscillating_masses_N_8_i_6 1344 10 13 10 12 7 7 10 0.00093 0.00105 0.00304 0.00426 0.00034 0.00016 0.00022
-oscillating_masses_N_8_i_7 1344 5 13 12 11 5 5 6 0.00047 0.00119 0.00357 0.00400 0.00030 0.00012 0.00014
-oscillating_masses_N_8_i_8 1344 5 14 8 11 4 4 6 0.00052 0.00125 0.00316 0.00404 0.00026 0.00010 0.00013
-oscillating_masses_N_8_i_9 1344 5 15 9 11 5 5 6 0.00052 0.00133 0.00326 0.00407 0.00030 0.00011 0.00013
-oscillating_masses_N_8_i_10 1344 5 14 8 13 5 5 6 0.00052 0.00125 0.00317 0.00472 0.00030 0.00012 0.00014
-oscillating_masses_N_8_i_11 1344 5 13 10 12 5 5 6 0.00052 0.00119 0.00337 0.00426 0.00030 0.00011 0.00013
-oscillating_masses_N_8_i_12 1344 5 14 9 12 5 5 7 0.00052 0.00130 0.00332 0.00453 0.00030 0.00012 0.00016
-oscillating_masses_N_8_i_13 1344 5 12 9 10 5 5 6 0.00052 0.00113 0.00330 0.00375 0.00030 0.00012 0.00013
-oscillating_masses_N_8_i_14 1344 5 12 9 10 5 5 6 0.00052 0.00095 0.00284 0.00375 0.00026 0.00011 0.00012
-oscillating_masses_N_8_i_15 1344 5 13 9 10 5 5 7 0.00047 0.00105 0.00284 0.00353 0.00025 0.00011 0.00015
-oscillating_masses_N_8_i_16 1344 6 14 10 11 6 6 8 0.00055 0.00115 0.00293 0.00379 0.00029 0.00013 0.00017
-oscillating_masses_N_8_i_17 1344 5 13 8 11 5 5 6 0.00047 0.00102 0.00276 0.00365 0.00026 0.00012 0.00013
-oscillating_masses_N_8_i_18 1344 5 12 9 10 4 4 6 0.00047 0.00096 0.00286 0.00334 0.00023 0.00009 0.00012
-oscillating_masses_N_8_i_19 1344 5 12 9 11 5 5 6 0.00048 0.00097 0.00288 0.00379 0.00026 0.00011 0.00012
-oscillating_masses_N_20_i_0 3312 5 15 9 12 5 5 7 0.00111 0.00287 0.00466 0.01414 0.00060 0.00028 0.00040
-oscillating_masses_N_20_i_1 3312 5 14 10 13 5 5 7 0.00113 0.00263 0.00488 0.01447 0.00060 0.00028 0.00038
-oscillating_masses_N_20_i_2 3312 5 15 10 13 5 5 7 0.00112 0.00280 0.00497 0.01484 0.00061 0.00029 0.00040
-oscillating_masses_N_20_i_3 3312 7 14 12 13 7 7 8 0.00141 0.00265 0.00529 0.01475 0.00077 0.00038 0.00045
-oscillating_masses_N_20_i_4 3312 7 14 9 14 7 7 9 0.00146 0.00266 0.00462 0.01572 0.00077 0.00039 0.00051
-oscillating_masses_N_20_i_5 3312 5 14 9 12 5 5 6 0.00109 0.00262 0.00462 0.01343 0.00061 0.00028 0.00034
-oscillating_masses_N_20_i_6 3312 5 16 10 12 5 5 7 0.00110 0.00296 0.00492 0.01387 0.00061 0.00028 0.00040
-oscillating_masses_N_20_i_7 3312 5 14 9 13 5 5 6 0.00109 0.00270 0.00463 0.01458 0.00060 0.00029 0.00034
-oscillating_masses_N_20_i_8 3312 5 13 9 13 5 5 7 0.00111 0.00255 0.00467 0.01429 0.00062 0.00031 0.00039
-oscillating_masses_N_20_i_9 3312 5 16 10 13 5 5 6 0.00109 0.00297 0.00492 0.01376 0.00064 0.00027 0.00034
-oscillating_masses_N_20_i_10 3312 5 14 9 13 5 5 6 0.00110 0.00269 0.00463 0.01413 0.00062 0.00028 0.00035
-oscillating_masses_N_20_i_11 3312 5 14 9 13 5 5 7 0.00111 0.00268 0.00467 0.01373 0.00061 0.00031 0.00040
-oscillating_masses_N_20_i_12 3312 5 15 9 13 5 5 7 0.00112 0.00276 0.00465 0.01485 0.00062 0.00031 0.00040
-oscillating_masses_N_20_i_13 3312 5 14 9 12 5 5 6 0.00108 0.00262 0.00464 0.01416 0.00062 0.00027 0.00035
-oscillating_masses_N_20_i_14 3312 5 15 9 13 5 5 6 0.00109 0.00280 0.00465 0.01435 0.00061 0.00029 0.00038
-oscillating_masses_N_20_i_15 3312 5 13 9 12 5 5 6 0.00110 0.00253 0.00469 0.01373 0.00062 0.00029 0.00034
-oscillating_masses_N_20_i_16 3312 5 14 9 14 5 5 6 0.00109 0.00270 0.00460 0.01496 0.00061 0.00029 0.00035
-oscillating_masses_N_20_i_17 3312 5 15 10 13 5 5 7 0.00116 0.00286 0.00488 0.01485 0.00061 0.00029 0.00040
-oscillating_masses_N_20_i_18 3312 5 15 9 14 5 5 6 0.00111 0.00279 0.00467 0.01549 0.00061 0.00029 0.00035
-oscillating_masses_N_20_i_19 3312 5 15 9 13 5 5 7 0.00110 0.00288 0.00464 0.01489 0.00063 0.00028 0.00040
-oscillating_masses_N_32_i_0 5280 5 14 10 14 5 5 - 0.00172 0.00422 0.00671 0.02045 0.00097 0.00053 -
-oscillating_masses_N_32_i_1 5280 5 14 9 13 5 5 - 0.00181 0.00420 0.00629 0.01951 0.00096 0.00048 -
-oscillating_masses_N_32_i_2 5280 5 15 9 12 5 5 - 0.00173 0.00442 0.00628 0.01808 0.00101 0.00047 -
-oscillating_masses_N_32_i_3 5280 7 13 11 12 6 6 - 0.00219 0.00385 0.00695 0.01738 0.00109 0.00056 -
-oscillating_masses_N_32_i_4 5280 9 15 10 14 8 8 - 0.00281 0.00461 0.00681 0.02077 0.00137 0.00073 -
-oscillating_masses_N_32_i_5 5280 5 15 9 11 5 5 - 0.00172 0.00446 0.00631 0.01756 0.00098 0.00048 -
-oscillating_masses_N_32_i_6 5280 6 14 10 11 6 6 - 0.00194 0.00415 0.00677 0.01658 0.00110 0.00056 -
-oscillating_masses_N_32_i_7 5280 6 15 11 12 6 6 - 0.00196 0.00443 0.00702 0.01703 0.00110 0.00058 -
-oscillating_masses_N_32_i_8 5280 5 16 9 13 5 5 - 0.00176 0.00477 0.00631 0.01762 0.00099 0.00048 -
-oscillating_masses_N_32_i_9 5280 5 14 9 12 5 5 - 0.00171 0.00414 0.00630 0.01706 0.00099 0.00048 -
-oscillating_masses_N_32_i_10 5280 6 15 10 12 5 5 - 0.00193 0.00448 0.00678 0.01870 0.00097 0.00048 -
-oscillating_masses_N_32_i_11 5280 5 13 9 12 5 5 - 0.00172 0.00391 0.00622 0.01769 0.00102 0.00048 -
-oscillating_masses_N_32_i_12 5280 5 15 9 11 5 5 - 0.00170 0.00438 0.00631 0.01586 0.00099 0.00049 -
-oscillating_masses_N_32_i_13 5280 5 15 9 14 5 5 - 0.00173 0.00454 0.00632 0.01974 0.00097 0.00048 -
-oscillating_masses_N_32_i_14 5280 6 16 10 11 6 6 - 0.00196 0.00470 0.00678 0.01605 0.00111 0.00057 -
-oscillating_masses_N_32_i_15 5280 5 16 9 13 5 5 - 0.00173 0.00480 0.00629 0.01822 0.00098 0.00053 -
-oscillating_masses_N_32_i_16 5280 5 15 9 13 5 5 - 0.00174 0.00444 0.00622 0.01941 0.00096 0.00048 -
-oscillating_masses_N_32_i_17 5280 5 15 10 12 5 5 - 0.00171 0.00446 0.00665 0.01877 0.00097 0.00048 -
-oscillating_masses_N_32_i_18 5280 5 15 9 13 5 5 - 0.00170 0.00459 0.00731 0.01839 0.00115 0.00048 -
-oscillating_masses_N_32_i_19 5280 5 15 10 13 5 5 - 0.00173 0.00454 0.00670 0.01873 0.00100 0.00048 -
-oscillating_masses_N_44_i_0 7248 5 15 9 12 5 5 - 0.00230 0.00611 0.00789 0.02372 0.00134 0.00065 -
-oscillating_masses_N_44_i_1 7248 6 16 10 12 6 6 - 0.00266 0.00640 0.00840 0.02427 0.00153 0.00080 -
-oscillating_masses_N_44_i_2 7248 6 15 11 13 6 6 - 0.00272 0.00604 0.00884 0.02677 0.00156 0.00080 -
-oscillating_masses_N_44_i_3 7248 5 14 9 13 5 5 - 0.00240 0.00599 0.00783 0.02612 0.00134 0.00068 -
-oscillating_masses_N_44_i_4 7248 5 16 9 13 5 5 - 0.00235 0.00668 0.00783 0.02650 0.00137 0.00069 -
-oscillating_masses_N_44_i_5 7248 5 14 9 12 5 5 - 0.00235 0.00573 0.00780 0.02321 0.00136 0.00068 -
-oscillating_masses_N_44_i_6 7248 6 15 11 13 6 6 - 0.00278 0.00623 0.00880 0.02700 0.00154 0.00080 -
-oscillating_masses_N_44_i_7 7248 5 15 9 12 5 5 - 0.00235 0.00624 0.00781 0.02492 0.00132 0.00066 -
-oscillating_masses_N_44_i_8 7248 5 15 9 13 5 5 - 0.00238 0.00622 0.00785 0.02579 0.00133 0.00069 -
-oscillating_masses_N_44_i_9 7248 5 15 9 12 5 5 - 0.00239 0.00604 0.00778 0.02401 0.00134 0.00069 -
-oscillating_masses_N_44_i_10 7248 5 15 9 13 5 5 - 0.00246 0.00608 0.00775 0.02662 0.00134 0.00069 -
-oscillating_masses_N_44_i_11 7248 5 14 9 12 5 5 - 0.00236 0.00586 0.00777 0.02545 0.00133 0.00068 -
-oscillating_masses_N_44_i_12 7248 5 16 9 13 5 5 - 0.00234 0.00642 0.00786 0.02667 0.00134 0.00068 -
-oscillating_masses_N_44_i_13 7248 5 14 9 12 5 5 - 0.00235 0.00589 0.00779 0.02546 0.00133 0.00076 -
-oscillating_masses_N_44_i_14 7248 5 16 9 13 5 5 - 0.00240 0.00668 0.00790 0.02632 0.00134 0.00068 -
-oscillating_masses_N_44_i_15 7248 5 14 9 14 5 5 - 0.00237 0.00578 0.00781 0.02633 0.00135 0.00069 -
-oscillating_masses_N_44_i_16 7248 5 16 9 13 5 5 - 0.00237 0.00647 0.00781 0.02624 0.00135 0.00068 -
-oscillating_masses_N_44_i_17 7248 5 15 9 13 5 5 - 0.00237 0.00618 0.00781 0.02594 0.00135 0.00070 -
-oscillating_masses_N_44_i_18 7248 5 14 9 13 5 5 - 0.00238 0.00576 0.00776 0.02651 0.00133 0.00069 -
-oscillating_masses_N_44_i_19 7248 5 14 9 13 5 5 - 0.00238 0.00584 0.00783 0.02656 0.00134 0.00068 -
-oscillating_masses_N_56_i_0 9216 5 15 9 14 5 5 - 0.00304 0.00893 0.00976 0.03563 0.00197 0.00101 -
-oscillating_masses_N_56_i_1 9216 5 16 9 12 5 5 - 0.00340 0.00968 0.00980 0.02983 0.00201 0.00101 -
-oscillating_masses_N_56_i_2 9216 5 16 9 12 5 5 - 0.00341 0.00834 0.00839 0.03087 0.00171 0.00092 -
-oscillating_masses_N_56_i_3 9216 5 15 9 13 5 5 - 0.00306 0.00785 0.00836 0.03244 0.00178 0.00092 -
-oscillating_masses_N_56_i_4 9216 5 15 9 13 5 5 - 0.00351 0.00900 0.00832 0.03240 0.00198 0.00101 -
-oscillating_masses_N_56_i_5 9216 7 16 11 13 6 6 - 0.00401 0.00826 0.00938 0.03206 0.00194 0.00107 -
-oscillating_masses_N_56_i_6 9216 5 14 9 12 5 5 - 0.00304 0.00734 0.00846 0.03137 0.00173 0.00090 -
-oscillating_masses_N_56_i_7 9216 5 19 10 14 5 5 - 0.00298 0.00972 0.00894 0.03386 0.00170 0.00091 -
-oscillating_masses_N_56_i_8 9216 5 14 9 12 5 5 - 0.00300 0.00740 0.00838 0.02979 0.00175 0.00091 -
-oscillating_masses_N_56_i_9 9216 6 16 9 13 6 6 - 0.00340 0.00822 0.00859 0.03321 0.00194 0.00105 -
-oscillating_masses_N_56_i_10 9216 5 19 9 12 5 5 - 0.00301 0.00978 0.00830 0.03279 0.00171 0.00091 -
-oscillating_masses_N_56_i_11 9216 5 14 9 12 5 5 - 0.00301 0.00738 0.00836 0.03012 0.00172 0.00089 -
-oscillating_masses_N_56_i_12 9216 5 15 9 12 5 5 - 0.00299 0.00775 0.00846 0.03024 0.00172 0.00092 -
-oscillating_masses_N_56_i_13 9216 5 14 9 12 5 5 - 0.00299 0.00730 0.00832 0.03158 0.00170 0.00090 -
-oscillating_masses_N_56_i_14 9216 5 15 9 13 5 5 - 0.00298 0.00762 0.00838 0.03245 0.00172 0.00090 -
-oscillating_masses_N_56_i_15 9216 6 14 11 13 6 6 - 0.00351 0.00844 0.01098 0.03534 0.00228 0.00119 -
-oscillating_masses_N_56_i_16 9216 5 16 9 13 5 5 - 0.00353 0.00958 0.00973 0.03549 0.00200 0.00101 -
-oscillating_masses_N_56_i_17 9216 6 15 10 14 6 6 - 0.00393 0.00903 0.01064 0.03770 0.00225 0.00120 -
-oscillating_masses_N_56_i_18 9216 5 16 9 13 5 5 - 0.00351 0.00846 0.00840 0.03598 0.00170 0.00090 -
-oscillating_masses_N_56_i_19 9216 7 16 10 16 7 7 - 0.00404 0.00836 0.00929 0.03900 0.00224 0.00121 -
-oscillating_masses_N_76_i_0 12496 5 16 9 12 5 - - 0.00404 0.01138 0.01050 0.04013 0.00236 - -
-oscillating_masses_N_76_i_1 12496 6 15 10 13 5 - - 0.00479 0.01055 0.01141 0.04462 0.00233 - -
-oscillating_masses_N_76_i_2 12496 5 15 9 13 5 - - 0.00409 0.01073 0.01062 0.04632 0.00236 - -
-oscillating_masses_N_76_i_3 12496 5 14 9 13 6 - - 0.00409 0.00973 0.01075 0.04351 0.00268 - -
-oscillating_masses_N_76_i_4 12496 5 14 9 12 5 - - 0.00406 0.01169 0.01070 0.04310 0.00235 - -
-oscillating_masses_N_76_i_5 12496 5 15 9 13 5 - - 0.00476 0.01237 0.01246 0.04689 0.00272 - -
-oscillating_masses_N_76_i_6 12496 5 15 9 13 5 - - 0.00472 0.01268 0.01276 0.04681 0.00270 - -
-oscillating_masses_N_76_i_7 12496 5 15 9 14 5 - - 0.00471 0.01057 0.01069 0.04750 0.00236 - -
-oscillating_masses_N_76_i_8 12496 6 14 10 12 6 - - 0.00469 0.00984 0.01144 0.04088 0.00266 - -
-oscillating_masses_N_76_i_9 12496 5 14 10 14 5 - - 0.00415 0.00987 0.01140 0.04779 0.00235 - -
-oscillating_masses_N_76_i_10 12496 6 15 10 12 6 - - 0.00454 0.01066 0.01137 0.04373 0.00266 - -
-oscillating_masses_N_76_i_11 12496 5 14 9 12 5 - - 0.00412 0.00997 0.01064 0.04083 0.00234 - -
-oscillating_masses_N_76_i_12 12496 7 15 10 14 7 - - 0.00533 0.01259 0.01348 0.04971 0.00346 - -
-oscillating_masses_N_76_i_13 12496 7 17 12 14 6 - - 0.00618 0.01371 0.01473 0.04861 0.00310 - -
-oscillating_masses_N_76_i_14 12496 5 15 9 14 5 - - 0.00477 0.01248 0.01232 0.05148 0.00271 - -
-oscillating_masses_N_76_i_15 12496 5 16 9 13 5 - - 0.00476 0.01134 0.01066 0.04689 0.00232 - -
-oscillating_masses_N_76_i_16 12496 5 15 9 11 5 - - 0.00433 0.01051 0.01066 0.03930 0.00235 - -
-oscillating_masses_N_76_i_17 12496 5 15 9 13 5 - - 0.00411 0.01072 0.01072 0.04533 0.00236 - -
-oscillating_masses_N_76_i_18 12496 5 16 9 13 5 - - 0.00406 0.01292 0.01258 0.04323 0.00273 - -
-oscillating_masses_N_76_i_19 12496 5 15 9 12 5 - - 0.00471 0.01225 0.01249 0.04638 0.00272 - -
-oscillating_masses_N_96_i_0 15776 5 15 9 13 5 - - 0.00602 0.01550 0.01450 0.05848 0.00348 - -
-oscillating_masses_N_96_i_1 15776 6 16 11 13 6 - - 0.00698 0.01420 0.01402 0.05978 0.00339 - -
-oscillating_masses_N_96_i_2 15776 5 16 9 13 5 - - 0.00511 0.01426 0.01238 0.05680 $0$
-oscillating_masses_N_96_i_3 15776 5 17 9 13 5 - - 0.00512 0.01756 0.01263 0.05612 0.00347 - -
-oscillating_masses_N_96_i_4 15776 5 13 13 12 5 - - 0.00581 0.01338 0.01845 0.05332 0.00346 - -
-oscillating_masses_N_96_i_5 15776 5 15 9 15 5 - - 0.00603 0.01337 0.01244 0.07083 0.00305 - -
-oscillating_masses_N_96_i_6 15776 5 14 9 13 5 - - 0.00512 0.01271 0.01272 0.05773 0.00301 - -
-oscillating_masses_N_96_i_7 15776 5 15 9 13 5 - - 0.00520 0.01367 0.01246 0.05746 0.00308 - -
-oscillating_masses_N_96_i_8 15776 5 16 9 13 5 - - 0.00523 0.01456 0.01244 0.05736 $0$
-oscillating_masses_N_96_i_9 15776 5 15 9 11 5 - - 0.00513 0.01334 0.01273 0.05200 $0$
-oscillating_masses_N_96_i_10 15776 6 14 10 14 6 - - 0.00571 0.01237 0.01338 0.05674 0.00340 - -
-oscillating_masses_N_96_i_11 15776 6 15 10 12 6 - - 0.00590 0.01362 0.01320 0.05358 0.00341 - -
-oscillating_masses_N_96_i_12 15776 5 17 9 13 5 - - 0.00512 0.01509 0.01248 0.05466 0.00298 - -
-oscillating_masses_N_96_i_13 15776 5 15 9 14 5 - - 0.00520 0.01339 0.01243 0.05965 $0$
-oscillating_masses_N_96_i_14 15776 5 17 9 13 5 - - 0.00518 0.01539 0.01246 0.05609 0.00302 - -
-oscillating_masses_N_96_i_15 15776 6 15 11 12 6 - - 0.00584 0.01363 0.01427 0.05236 0.00339 - -
-oscillating_masses_N_96_i_16 15776 5 14 10 14 5 - - 0.00528 0.01269 0.01350 0.05957 0.00299 - -
-oscillating_masses_N_96_i_17 15776 5 16 9 15 5 - - 0.00510 0.01405 0.01259 0.06245 0.00307 - -
-oscillating_masses_N_96_i_18 15776 5 14 9 12 5 - - 0.00504 0.01257 0.01469 0.05114 0.00347 - -
-oscillating_masses_N_96_i_19 15776 5 16 9 13 5 - - 0.00514 0.01434 0.01248 0.05771 0.00298 - -
-oscillating_masses_N_116_i_0 19056 5 15 9 13 5 - - 0.00627 0.01601 0.01484 0.06574 0.00368 - -
-oscillating_masses_N_116_i_1 19056 6 14 10 12 6 - - 0.00695 0.01539 0.01599 0.06194 0.00419 - -
-oscillating_masses_N_116_i_2 19056 5 15 9 12 5 - - 0.00637 0.01643 0.01469 0.06273 0.00367 - -
-oscillating_masses_N_116_i_3 19056 5 16 9 13 5 - - 0.00622 0.01964 0.01696 0.06735 0.00422 - -
-oscillating_masses_N_116_i_4 19056 5 15 10 12 5 - - 0.00715 0.01912 0.01840 0.06859 0.00423 - -
-oscillating_masses_N_116_i_5 19056 6 16 10 12 6 - - 0.00811 0.01714 0.01554 0.07024 0.00412 - -
-oscillating_masses_N_116_i_6 19056 6 14 10 10 6 - - 0.00646 0.01515 0.01561 0.05609 0.00416 - -
-oscillating_masses_N_116_i_7 19056 5 15 9 13 5 - - 0.00638 0.01621 0.01461 0.06767 0.00369 - -
-oscillating_masses_N_116_i_8 19056 7 17 11 14 7 - - 0.00841 0.01864 0.01665 0.07346 0.00465 - -
-oscillating_masses_N_116_i_9 19056 5 15 9 14 5 - - 0.00626 0.01642 0.01470 0.07328 0.00373 - -
-oscillating_masses_N_116_i_10 19056 5 14 10 13 5 - - 0.00623 0.01727 0.01818 0.06714 0.00425 - -
-oscillating_masses_N_116_i_11 19056 5 16 9 13 5 - - 0.00700 0.01690 0.01750 0.06795 0.00426 - -
-oscillating_masses_N_116_i_12 19056 5 16 9 12 5 - - 0.00608 0.01737 0.01494 0.06339 0.00366 - -
-oscillating_masses_N_116_i_13 19056 5 15 9 13 6 - - 0.00619 0.01623 0.01471 0.07116 0.00419 - -
-oscillating_masses_N_116_i_14 19056 6 15 10 13 6 - - 0.00718 0.01647 0.01562 0.07111 0.00415 - -
-oscillating_masses_N_116_i_15 19056 5 16 9 13 5 - - 0.00619 0.01733 0.01465 0.07095 0.00366 - -
-oscillating_masses_N_116_i_16 19056 5 14 9 14 5 - - 0.00620 0.01531 0.01466 0.07134 0.00363 - -
-oscillating_masses_N_116_i_17 19056 5 16 9 14 5 - - 0.00627 0.02028 0.01707 0.07243 0.00424 - -
-oscillating_masses_N_116_i_18 19056 5 15 10 12 6 - - 0.00719 0.01678 0.01573 0.07083 0.00417 - -
-oscillating_masses_N_116_i_19 19056 5 16 9 14 5 - - 0.00635 0.01737 0.01461 0.06908 0.00362 - -
-oscillating_masses_N_136_i_0 22336 5 16 9 13 5 - - 0.00721 0.02032 0.01688 0.08071 0.00432 - -
-oscillating_masses_N_136_i_1 22336 5 15 10 12 5 - - 0.00733 0.01903 0.01791 0.07587 0.00430 - -
-oscillating_masses_N_136_i_2 22336 5 14 9 12 5 - - 0.00743 0.01827 0.01725 0.07667 0.00436 - -
-oscillating_masses_N_136_i_3 22336 5 16 10 13 5 - - 0.00734 0.02020 0.01796 0.08116 0.00437 - -
-oscillating_masses_N_136_i_4 22336 5 15 9 13 5 - - 0.00729 0.01963 0.01699 0.07803 0.00438 - -
-oscillating_masses_N_136_i_5 22336 5 16 10 12 5 - - 0.00737 0.02324 0.02100 0.07585 0.00503 - -
-oscillating_masses_N_136_i_6 22336 5 15 10 14 5 - - 0.00860 0.02233 0.02066 0.09022 0.00504 - -
-oscillating_masses_N_136_i_7 22336 5 15 9 14 5 - - 0.00839 0.01910 0.01967 0.09066 0.00509 - -
-oscillating_masses_N_136_i_8 22336 5 15 9 13 5 - - 0.00723 0.01913 0.01691 0.07890 0.00433 - -
-oscillating_masses_N_136_i_9 22336 5 14 9 12 5 - - 0.00725 0.01794 0.01715 0.07905 0.00436 - -
-oscillating_masses_N_136_i_10 22336 6 14 10 13 6 - - 0.00821 0.01806 0.01861 0.07904 $0$
-oscillating_masses_N_136_i_11 22336 6 15 10 13 6 - - 0.00817 0.01932 0.01792 0.07561 0.00504 - -
-oscillating_masses_N_136_i_12 22336 5 15 9 13 5 - - 0.00727 0.01911 0.01712 0.07888 0.00435 - -
-oscillating_masses_N_136_i_13 22336 5 15 9 14 5 - - 0.00751 0.01968 0.01682 0.08226 0.00433 - -
-oscillating_masses_N_136_i_14 22336 5 16 9 14 5 - - 0.00726 0.02032 0.01678 0.08191 0.00440 - -
-oscillating_masses_N_136_i_15 22336 6 15 9 12 6 - - 0.00809 0.01898 0.01684 0.07451 0.00493 - -
-oscillating_masses_N_136_i_16 22336 6 16 10 12 6 - - 0.00810 0.01970 0.01790 0.07366 0.00493 - -
-oscillating_masses_N_136_i_17 22336 6 16 10 12 5 - - 0.00806 0.01983 0.01801 0.07859 0.00436 - -
-oscillating_masses_N_136_i_18 22336 5 17 9 13 5 - - 0.00729 0.02159 0.01689 0.08000 0.00432 - -
-oscillating_masses_N_136_i_19 22336 6 16 9 12 6 - - 0.00810 0.01987 0.01685 0.07215 0.00495 - -
-oscillating_masses_N_156_i_0 25616 6 15 10 14 7 - - 0.00975 0.02223 0.02161 0.09490 0.00636 - -
-oscillating_masses_N_156_i_1 25616 6 14 10 13 6 - - 0.00928 0.02040 0.02191 0.09083 0.00581 - -
-oscillating_masses_N_156_i_2 25616 7 20 12 14 5 - - 0.01080 0.02879 0.02462 0.09866 0.00515 - -
-oscillating_masses_N_156_i_3 25616 5 16 9 14 5 - - 0.00833 0.02689 0.02182 0.09493 0.00589 - -
-oscillating_masses_N_156_i_4 25616 6 16 10 13 6 - - 0.01073 0.02285 0.02305 0.09158 0.00571 - -
-oscillating_masses_N_156_i_5 25616 5 16 9 12 5 - - 0.00830 0.02309 0.02070 0.08317 0.00504 - -
-oscillating_masses_N_156_i_6 25616 5 15 9 14 5 - - 0.00851 0.02211 0.02133 0.09482 $0$
-oscillating_masses_N_156_i_7 25616 5 16 9 13 5 - - 0.00815 0.02310 0.02083 0.08544 0.00506 - -
-oscillating_masses_N_156_i_8 25616 5 15 9 13 5 - - 0.00841 0.02230 0.02131 0.09227 0.00514 - -
-oscillating_masses_N_156_i_9 25616 5 15 9 14 5 - - 0.00839 0.02164 0.02064 0.09737 0.00506 - -
-oscillating_masses_N_156_i_10 25616 6 15 10 13 6 - - 0.00960 0.02565 0.02176 0.09410 0.00671 - -
-oscillating_masses_N_156_i_11 25616 5 17 9 13 5 - - 0.00818 0.02474 0.02085 0.08873 0.00504 - -
-oscillating_masses_N_156_i_12 25616 5 15 9 13 5 - - 0.00840 0.02249 0.02040 0.09105 0.00503 - -
-oscillating_masses_N_156_i_13 25616 5 14 9 12 5 - - 0.00841 0.02069 0.02099 0.08307 0.00514 - -
-oscillating_masses_N_156_i_14 25616 6 16 9 12 6 - - 0.00932 0.02232 0.02112 0.08811 0.00664 - -
-oscillating_masses_N_156_i_15 25616 5 14 9 14 5 - - 0.00832 0.02064 0.02225 0.09329 0.00589 - -
-oscillating_masses_N_156_i_16 25616 6 13 11 11 6 - - 0.00927 0.01946 0.02327 0.07696 0.00576 - -
-oscillating_masses_N_156_i_17 25616 5 17 9 13 5 - - 0.00838 0.02463 0.02070 0.09187 0.00505 - -
-oscillating_masses_N_156_i_18 25616 7 16 13 13 7 - - 0.01111 0.02653 0.02615 0.08826 0.00740 - -
-oscillating_masses_N_156_i_19 25616 5 15 9 13 5 - - 0.00985 0.02505 0.02235 0.09991 0.00586 - -
+Problem Size qoco clarabel ecos gurobi mosek qoco clarabel ecos gurobi mosek
+NYPA_Maragal_6_huber 686479 6 6 13 - 65 18.85022 26.29638 47.81607 - 30.89082
+NYPA_Maragal_6_lasso 620812 7 7 - 13 13 20.08794 27.58199 - 1.69537 3.33290
+Pereyra_landmark_huber 1650512 8 8 16 9 11 0.75585 1.39359 3.07564 1.48089 4.40789
+Pereyra_landmark_lasso 1301568 7 7 - 12 24 0.50615 0.80743 - 1.37256 0.99517
+NYPA_Maragal_1_huber 458 5 5 10 7 8 0.00012 0.00018 0.00045 0.00237 0.00102
+NYPA_Maragal_1_lasso 354 8 8 12 10 10 0.00011 0.00018 0.00020 0.00217 0.00051
+HB_ash608_huber 5472 8 8 14 11 12 0.00192 0.00315 0.00993 0.00690 0.01230
+HB_ash608_lasso 3184 7 7 21 9 11 0.00102 0.00163 0.00458 0.00759 0.00376
+ANSYS_Delor64K_huber 1105173 6 7 7 0 3 1.30160 2.34456 2.86475 0.41221 1.00721
+ANSYS_Delor64K_lasso 7922958 7 7 - 17 20 6.12221 9.98326 - 4.06191 13.62744
+HB_ash85_huber 1118 5 5 6 6 3 0.00038 0.00057 0.00084 0.00264 0.00222
+HB_ash85_lasso 1033 7 7 16 9 10 0.00042 0.00061 0.00120 0.00295 0.00153
+NYPA_Maragal_3_huber 30221 6 7 12 9 10 0.02718 0.03827 0.07005 0.04743 0.12791
+NYPA_Maragal_3_lasso 25211 7 6 15 8 10 0.03499 0.03233 0.06423 0.04115 0.05463
+HB_well1033_huber 11963 6 6 11 7 9 0.00307 0.00518 0.01379 0.01955 0.02565
+HB_well1033_lasso 8078 6 13 - 8 27 0.00185 0.00573 - 0.00885 0.01698
+NYPA_Maragal_7_huber 1528452 6 9 37 - - 36.98824 55.93122 236.53158 - -
+NYPA_Maragal_7_lasso 1400483 7 6 - 13 11 38.76392 41.96500 - 4.47169 5.46443
+HB_illc1033_huber 11950 6 6 11 6 9 0.00307 0.00676 0.01472 0.01855 0.02622
+HB_illc1033_lasso 8065 6 19 32 9 43 0.00231 0.00883 0.01289 0.00981 0.03614
+NYPA_Maragal_4_huber 40467 6 6 12 8 55 0.06713 0.08642 0.14732 0.05997 0.74862
+NYPA_Maragal_4_lasso 34783 6 6 17 7 9 0.06302 0.07162 0.17837 0.05356 0.07373
+HB_ash219_huber 1971 7 7 12 10 11 0.00064 0.00105 0.00312 0.00369 0.00461
+HB_ash219_lasso 1216 8 8 18 9 11 0.00043 0.00068 0.00142 0.00359 0.00127
+NYPA_Maragal_2_huber 8242 6 7 12 9 11 0.00304 0.00456 0.01130 0.01764 0.01897
+NYPA_Maragal_2_lasso 6867 6 6 14 8 11 0.00243 0.00328 0.00613 0.01660 0.00786
+HB_ash331_huber 2979 8 8 12 10 12 0.00104 0.00169 0.00456 0.00451 0.00679
+HB_ash331_lasso 1740 7 7 20 10 10 0.00055 0.00089 0.00226 0.00467 0.00174
+HB_illc1850_huber 21586 6 6 14 8 9 0.00669 0.01189 0.03405 0.02046 0.05253
+HB_illc1850_lasso 15184 6 13 - 7 42 0.00443 0.01323 - 0.01773 0.07218
+HB_ash958_huber 8622 8 8 14 10 12 0.00302 0.00499 0.01556 0.01024 0.01923
+HB_ash958_lasso 5000 7 7 23 9 12 0.00154 0.00250 0.00853 0.01077 0.00546
+ANSYS_Delor295K_huber 4471461 7 7 8 - 4 3.12931 5.19758 10.03436 - 3.17687
+ANSYS_Delor295K_lasso 10288503 7 7 - 12 33 7.50801 12.16215 - 12.36872 24.15621
+ANSYS_Delor338K_huber 6614251 7 7 - 0 4 4.05127 6.12029 - 0.92594 3.42096
+ANSYS_Delor338K_lasso 8446303 7 6 - 15 23 5.27505 7.54937 - 13.85264 13.07142
+NYPA_Maragal_8_huber 1540899 6 6 13 - 16 22.18652 25.05274 45.98680 - 10.81397
+NYPA_Maragal_8_lasso 1675147 7 7 - 12 14 24.59839 29.16281 - 5.67463 7.79855
+HB_abb313_huber 3748 8 8 14 11 12 0.00124 0.00205 0.00550 0.00456 0.00744
+HB_abb313_lasso 2887 6 6 22 8 11 0.00085 0.00129 0.00467 0.00434 0.00264
+NYPA_Maragal_5_huber 125669 6 7 14 14 82 0.47005 0.60018 1.05444 0.17352 3.24096
+NYPA_Maragal_5_lasso 115679 7 7 22 14 11 0.51054 0.54681 1.35915 0.19943 0.27376
+HB_ash292_huber 4252 6 6 13 8 8 0.00171 0.00263 0.00664 0.00551 0.01193
+HB_ash292_lasso 3960 8 8 24 9 15 0.00195 0.00273 0.00825 0.00678 0.00735
+HB_well1850_huber 21705 6 6 14 7 9 0.00622 0.00968 0.03297 0.01690 0.04740
+HB_well1850_lasso 15303 6 12 - 8 40 0.00432 0.01136 - 0.01770 0.07075